A Complete Bibliography of Publications in Canadian Journal of Statistics = Revue

canadienne de statistique

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

12 October 2023 Version 1.23

Title word cross-reference

1 [855, 54, 959]. $10.50 [148]. $10.95 [103]. $100.00 [151]. $105.00 [134]. $12.00 [221]. $12.40 [311]. $12.50 [169]. $12.80 [310]. $120.00 [462]. $13.85 [202]. $13.95 [124, 465]. $15.00 [203]. $16.20 [148]. $16.50 [217]. $16.80 [168]. $16.95 [135]. $17.10 [218]. $18.40 [122]. $18.95 [170, 123]. $19.50 [287]. 2 [727, 99]. $20.00 [222]. $21.20 [204]. $21.85 [121]. $22.00 [205]. $22.50 [219, 165]. $22.95 [288]. $23.50 [167]. $24.00 [239]. $25.00 [149, 165]. $26.95 [220, 309]. $27.95 [266]. $28.50 [147, 308]. $28.95 [237]. $29.75 [265]. $29.80 [464]. 2^{m−p} [1149]. 2 × 2 [693]. 2 × k [1135]. 305 [411]. $31.50 [238]. $31.95 [368]. $33.95 [306]. $34.20 [149]. $34.50 [150, 466]. $35.00 [185, 369]. $36.95 [267]. $37.00 [201]. $38.95 [459]. $39.00 [461]. 3 × 3 [1303]. $4.50 [204, 284]. $40.00 [222]. $40.35 [238]. $41.50 [289]. $49.75 [189]. 5 [33]. $5.70 [290]. $50.00 [222]. $50.95 [267]. $8.00 [186, 152, 184]. $8.50 [166]. $8.75 [105]. $9.00 [104]. $9.95 [187, 200, 102]. < ∞ [264]. _{2} [56]. _{3} [56]. A [1163]. AR(1) [1021]. AR(p) [1104]. Ē^{2} [496]. x̄ [793]. C [632, 10]. D [1137, 1616, 1411, 1523, 1163]. D([0, ∞)^{q}, E) [192]. D(IR) [529]. E [905]. ℓ_{1} [1776]. ϵ [455]. F [547, 1849, 1, 22, 989]. γ [178]. H [2105, 2]. K [1473, 1954, 95, 1867, 452, 1550]. L [557, 244, 235, 399, 988, 575, 929]. L^{1} [529, 1942]. L^{2} [1300]. L^{p} [2072, 1942]. L^{q} [1925]. L_{0} [1596, 1619]. L_{1} [769, 1282, 1136, 999, 515]. L_{2} [769, 2041]. LR [21]. M [969, 717, 1104, 800, 372, 762, 1942, 1201, 835, 687, 1169, 406, 931, 32, 1]. M/G/1 [1447]. Z [1963]. Es^{2} [877]. n [1007]. n < p [1605]. P [538, 1426, 753]. P A [99]. Q [1589]. R [935, 399, 575, 757, 727]. r × c [751]. S [1128, 935]. s^{2} [935]. σ [935]. T [49, 94, 1067, 426]. τ [642]. U [357]. Vec [143]. vech [143].

-confidence [178]. -consistency [1282]. -contamination [455]. -content [727]. -day [411]. -dependence [32]. -distance [1925]. -distribution [547, 426]. -errors [769]. -estimate [244]. -estimates [235, 835, 687]. -estimateurs [557]. -estimation [1169, 929]. -estimators [557, 969, 717, 1128, 1104, 800, 372, 762, 399, 999, 1201, 642, 988, 406, 575, 931]. -factor [99]. -Function [2105]. -learning [1589]. -optimal [905, 1137, 1616, 1411, 1163]. -penalized [1776]. -quantile [1942]. -quantile-based [2072]. -sample [1867, 95, 1954]. -sequences [632]. -simplex [727]. -statistic [1]. -statistics [357, 2]. -test [94]. -tests [1849, 989]. -Tightness [529]. -tuples [452]. -values [1426]. -variates [22]. -vine [1523].

/Commentaries [571]. /Réplique [572].

0 [149, 166, 148, 170, 237, 238, 239, 266, 265, 519, 122, 188, 205, 169, 104, 221, 222, 186, 103, 147, 167, 168, 189, 202, 201, 307, 217, 464, 285, 187, 204, 218, 287, 290, 150, 123, 289, 368, 185, 135, 459, 152, 522, 284, 286, 105, 203, 520, 466, 460, 121, 369, 463, 184]. 0-02-424110-5 [135]. 0-02-917960-2 [151]. 0-12-285650-3 [149]. 0-12-426280-5 [369]. 0-12-473750-1 [286]. 0-12-504050-4 [239]. 0-12-618850-5 [219]. 0-201-07809-0 [368]. 0-205-05982-1 [167]. 0-205-05983-X [167]. 0-205-06038-2 [167]. 0-262-06063-9 [103]. 0-333-19056-4 [202]. 0-333-90115-0 [202]. 0-387-9008-0 [520]. 0-387-90443-3 [221]. 0-387-90804-8 [520]. 0-387-90805-6 [520]. 0-387-96155-0 [463]. 0-387-96186-0 [464]. 0-393-09041-8 [124]. 0-393-09076-0 [124]. 0-442-25369-9 [308]. 0-444-00302-9 [287]. 0-444-00449-1 [266]. 0-444-00621-4 [466]. 0-444-00756-3 [460]. 0-444-86019-3 [289]. 0-470-26380-6 [187]. 0-470-26861-1 [284]. 0-470-26911-1 [204]. 0-470-99322-7 [166]. 0-471-01412-5 [288]. 0-471-02563-1 [121]. 0-471-02772-3 [122]. 0-471-03036-8 [309]. 0-471-03210-7 [218]. 0-471-05519-0 [238]. 0-471-07859-X [306]. 0-471-86764-0 [459]. 0-471-86769-1 [519].

0-471-95974-X [220]. 0-471-95982-0 [170]. 0-520-03569-0 [148]. 0-660-90202-8 [200]. 0-669-01310-2 [205]. 0-691-04222-5 [217]. 0-7201-0930-2 [201]. 0-8020-5405-6 [203]. 0-8020-5489-7 [240]. 0-8162-1150-7 [237]. 0-8218-1905-4 [168]. 0-8218-1906-2 [310]. 0-8218-1907-0 [311]. 0-8247-6515-X [105]. 0-8247-6653-9 [265]. 0-8247-6774-8 [185]. 0-8247-6800-0 [150]. 0-8247-6836-1 [169]. 0-8247-6944-9 [189]. 0-8403-2287-9 [290]. 0-8405-0412-8 [186]. 0-8405-0413-6 [184]. 0-8405-0425-X [152]. 0-85226-158-6 [284]. 0-85274-330-0 [188]. 0-87872-099-5 [123]. 0-88275-477-7 [147]. 0-88864-017-X [307]. 0-88864-074-9 [285]. 0-89116-573-8 [522]. 0-904425-03-7 [104]. 0-920788-04-1 [222]. 0076 [240]. 0076-5333 [240]. 0094 [310, 311]. 0094-8837 [310, 311]. 0471 [267]. 0471-08544-8. [267].

1 [266, 519, 169, 222, 204, 284, 286, 121]. 11-511E [200]. 12th [165]. 1870s [1533]. 19 [721, 778, 2004, 745, 1947, 1949, 1988]. 1957-1972 [67]. 1973 [134]. 1979 [2164, 289]. 1981 [462, 521]. 1994 [1245].

2 [205, 201, 461, 2005]. 2-88124-013-5 [461]. 2016 [1740, 1723, 1732]. 22 [1245]. 23 [941]. 24 [897, 1247]. 26 [1013]. 27 [1063]. 28 [1140].

3 [149, 122, 221, 289, 460]. 30 [1202]. 31 [1210]. 32 [1235]. 3G1 [165].

4 [239, 168, 466].

5 [217, 123, 135, 461, 369]. 50th [2022]. 511E [200]. 5333 [240].

6 [187, 203, 520, 184]. 65th [2159].

7 [166, 237, 104, 147, 218]. 77 [2089, 148].

8 [186, 185, 522]. 8. [267]. ’81 [462]. 8837 [310, 311].

9 [265, 103, 189, 285, 287, 290]. 90 [134]. 90-277-0614-X [134]. 92k [778]. 93e [721]. 93g [745]. 97f [897]. 99f [1013].

A- [2051]. A83 [124]. Abraham [459]. absence [153]. absolute [906]. absolute-error [906]. abstracted [67]. absurd [648, 746]. abundance [1322, 1132, 130, 304]. abundances [254]. Academic [219, 239, 286, 369, 521, 149]. Accelerated [1277, 900, 2060]. acceleration [752]. acceptance [622]. access [1340, 1314, 1329, 1364]. accommodating [1743]. account [854]. accuracy [979, 1762, 1643]. accurate [1097, 870, 1537, 1641, 545]. acid [490, 732, 491]. acidity [268, 492]. Acknowledgement [1258, 1343, 1434, 1477, 1575, 1609, 1656, 1683, 1715, 1748, 1783, 1821, 1863, 1912, 2056, 1986]. Acknowledgements [1258]. active [1779, 1988]. activities [486]. Acute [609, 610, 487]. Adam [188]. Adaptive [1972, 1072, 1943, 1191, 1628, 1911, 1319, 700, 232, 1989, 1914, 1118, 1419, 604, 1100, 882, 1127, 1414, 1760, 1799, 1475, 1797, 1702]. Addendum [523, 722, 744, 940, 778, 897, 1106, 1013, 721, 777, 896, 941, 1063, 745, 1012]. Addendums [759, 952]. Addison [368]. Addison-Wesley [368]. additional [621]. Additive [1687, 1487, 1994, 953, 930, 677, 1591, 2049, 1652, 1369, 1362, 1882, 1940, 1996, 142, 1917, 1697]. additivity [1547, 1737]. adequacy [83]. adjusted [1695, 1584, 1212, 1677, 1952, 1415, 1702]. Adjusting [613, 1331, 1721]. Adjustment [2138, 449, 484, 1376, 483, 308, 1469]. Adjustments [747, 922, 869]. Adler [200]. administrative [1123, 488]. administratives [488]. admissibility [686, 625, 120, 180]. admissible [415]. admission [781]. admissions [431]. admitting [277]. ADNI [1910]. Adolf [102]. Advanced [368]. aerosols [491]. affected [1152, 1268, 1367]. affected-sib-pair [1367]. affecting [296]. affine [1207]. affine-invariant [1207]. Afriat [217]. after [1463]. against [640, 1054, 153, 796, 773, 791, 751, 1472]. Age [1533, 414, 354, 786, 262, 350]. age-period-cohort [786]. agent [1090]. agglomerative [138]. aggregate [330, 342]. aggregation [50, 1678]. aggressive [1453]. agreement [964, 813, 941, 1412]. agricoles [477]. agricultural [943, 477]. AIC [1518]. AIDS [626]. Air [781, 784, 609, 613, 782, 610, 780, 783]. Akademia [521]. Akaike [1791]. albatross [1392]. Albert [286]. Alberta [265, 307, 285]. aléatoire [155]. aléatoires [59, 110, 98]. Alexander [460]. Algebra [2144, 2106, 188, 2123, 460]. algorithm [1837, 1781, 1914, 956, 1922, 1010, 2043, 1751, 1999]. algorithms [1416, 1635, 138, 1836]. aliased [1474, 1629]. aliasing [292, 177]. aligned [172, 374]. aligned-rank [374]. allele [2034]. allele-based [2034]. allocation [700]. Allowance [472]. Allyn [167]. Alternative [903, 79, 1579, 438, 1222, 242, 1897, 159]. alternatives [1396, 711, 1054, 895, 1346, 751, 1472, 496]. always [685]. Alzheimer [1144, 1904]. ambient [491]. ambiguous [1969]. American [168, 310, 311]. Ames [740, 742]. among [1473, 1041, 63, 1992, 917, 673, 1855, 1040]. amplitude [1028, 2080]. Amsterdam [289]. anaerobic [518, 635, 634]. Analyse [582, 1325, 1042, 723, 1028]. Analyses [1046, 812, 331, 1116, 1539, 1661]. analysing [1440]. Analysis [1045, 2086, 611, 2133, 2088, 1043, 1152, 1867, 782, 1153, 2114, 2160, 103, 1044, 2090, 1390, 2164, 674, 2128, 2136, 314, 955, 846, 2138, 1324, 2130, 1154, 1674, 1835, 1786, 1256, 237, 238, 258, 1150, 1322, 914, 215, 612, 698, 1470, 1304, 250, 699, 613, 1375, 363, 1223, 1499, 794, 1437, 1879, 2029, 1124, 2032, 1128, 659, 1486, 983, 2157, 723, 1360, 1758, 2007, 1569, 1069, 371, 1041, 1151, 417, 1367, 2049, 1408, 708, 626, 313, 1521, 1387, 883, 524, 1880, 874, 1554, 1659, 1906, 1572, 285, 1767, 1082, 1978, 423, 1570, 1533, 72, 919, 1548, 1778, 1840]. analysis [1894, 1148, 1594, 1896, 2044, 2066, 1167, 1233, 1028, 289, 334, 1200, 1907, 1968, 1386, 1899, 2084, 55, 982, 1352, 1933, 918, 801, 81, 889, 789, 1626, 469, 1657, 783, 1727, 284, 619, 709, 1593, 1528, 748, 67, 936, 2050, 493, 1281, 1040, 908, 1381, 173, 1311, 1096, 467, 181, 496, 1700, 1952, 1011, 990, 1252, 1571, 2014, 1271, 1688, 2069, 1583, 954, 468, 582, 1325, 1326, 1042, 122, 147, 240, 308]. analytic [831, 2042]. Analyzing [137, 350, 1215, 1607, 1009, 1865]. Anatole [150]. ancillarity [349, 754]. ancillary [1951]. Anderson [147]. André [1446, 1445]. anniversary [2022]. annotated [832]. announcements [1686]. annual [165]. ANOVA [1237, 2014]. Anscombe [977]. answering [720]. Antarctic [944, 1670]. Anthony [185]. antibody [2005]. antifiducial [326]. any [2017, 118, 33]. Apercu [450]. Application [54, 2046, 482, 1817, 1747, 879, 257, 1558, 2079, 987, 844, 1486, 1909, 1143, 1752, 477, 1889, 1098, 229, 1484, 1791, 556, 1346, 251, 356, 1731, 1537, 1100, 1655, 1861, 11, 431, 1640, 1727, 586, 510, 1872, 1955, 1593, 2050, 483, 1323, 472, 1960, 1196, 16, 1465, 1475, 1549, 1629, 54, 547, 471, 626, 792, 483, 80]. Applications [2085, 1269, 692, 2119, 2105, 2121, 2122, 2130, 907, 732, 134, 24, 1810, 699, 949, 2070, 2024, 1957, 230, 212, 1082, 187, 452, 749, 1908, 1393, 1160, 1704, 292, 295, 1190, 1444, 1877, 420, 13, 1381, 541, 806, 239, 240, 123, 286, 203]. Applied [265, 2112, 2151, 2107, 185, 2149, 465, 218, 466, 462, 167, 460, 463]. appraisal [863]. Approach [372, 2110, 669, 561, 1546, 1470, 781, 205, 1223, 1499, 904, 1694, 1397, 1125, 371, 1318, 474, 1990, 1261, 1449, 874, 1915, 1136, 456, 2016, 1710, 947, 1370, 1894, 1108, 359, 1684, 928, 1012, 1013, 1209, 1931, 297, 1182, 1812, 1852, 1954, 48, 1507, 1565, 1215, 1606, 1357, 1562, 2071, 1895, 2031, 1249, 1287, 1171, 1556, 1519]. approaches [2027, 1401, 1542]. approche [1261]. Approximate [392, 730, 1212, 1782, 1147, 1560, 427, 1625, 761, 1161, 1938, 925]. approximately [1779, 704, 826, 1021]. Approximating [2145, 933, 1308, 520]. Approximation [855, 1654, 690, 956, 1180, 1005, 1828, 1483, 182, 282, 213, 533, 2055]. Approximations [441, 800, 1653, 508, 618, 1597, 852, 547, 208, 818, 1147, 890, 1745, 832, 514, 521, 915, 398, 599, 429]. April [462]. arbitrarily [1176]. ARCH [1198, 1497, 1576, 1312]. ARCH/GARCH [1576]. Archimedean [405, 1273]. archimédiennes [405]. area [1801, 869, 1893, 1735, 1585, 1789, 1791, 1204, 1489, 1512, 785, 1644, 1734, 1450, 1964, 1502, 1679, 1839, 1689, 1919, 1171, 1142]. areas [832]. argument [56]. arguments [408]. ARIMA [274, 457, 856, 1246, 1247]. arising [1332, 1513]. ARMA [1725, 880, 688, 1353]. armed [1371]. Aroian [311]. arrangements [177]. arrays [74, 987, 808, 1831, 1646]. Article [1757]. Articles [1259, 1327, 1279, 1363, 1313]. Askey [2159]. aspects [226, 758, 1433]. assays [48]. Assessing [1997, 585, 1643, 1547, 948, 1634, 1049]. assessment [1815, 1054, 741, 1388, 440, 1496, 962]. assignment [1035]. assisted [489]. associate [316, 1, 361]. associated [665, 610, 12, 1669]. Association [2117, 1930, 221, 329, 1827, 100, 1730, 1805, 1910, 1539, 1819, 1249, 2034, 1936]. associations [612]. assumption [1413]. asthma [782]. asthmatic [610]. asymmetric [1284, 1934, 835, 706]. asymmetry [406]. asymptomatic [1988]. Asymptotic [277, 817, 1003, 183, 214, 641, 1104, 513, 1367, 660, 212, 766, 435, 1850, 112, 594, 966, 336, 341, 850, 798, 600, 959, 1220, 620, 769, 1001, 299, 301, 250, 557, 904, 133, 331, 21, 651, 92, 567, 867, 378, 1809, 447, 988, 1, 857, 172, 374, 387, 337, 2055, 1799, 939, 566]. asymptotically [276]. asymptotically-independent [276]. Asymptotics [999, 799, 1599, 1294, 631, 2023, 842, 1810]. asymptotique [939, 1001, 557, 566, 387]. asymptotiquement [276]. Atlas [2161, 2155, 522]. attempts [2046]. attending [609]. attribute [774]. AUC [1468]. auteurs [1497, 1985]. Author [721, 722, 744, 759, 777, 896, 940, 952, 941, 1063, 1985, 745, 1012, 654, 624]. Authority [1949]. Authors [1497]. autocorrelated [1695, 71]. autocorrelations [601, 1001, 566, 1001, 566]. autocovariance [987, 1953]. autocovariances [601, 1001, 457, 348]. Automatic [1389, 1933, 545]. autorégressif [855]. autoregression [1577, 1087, 2071]. Autoregressive [207, 855, 50, 1074, 1138, 1490, 1351, 1072, 1764, 554, 1420, 1843, 1963, 1352, 1660, 1765, 1917]. autoregressive-parameter [554]. autoregressives [50]. aux [1258, 1434, 1477, 1575, 1609, 1656, 1683, 1715, 1748, 1783, 1821, 1863, 1912, 2056, 1986, 251, 572, 483]. auxiliaire [1261]. auxiliary [1238, 1261, 1204, 1263, 701, 199, 1114, 980]. available [151, 306, 679]. average [1998, 1352, 487, 1086, 128, 207]. averaged [1789]. averages [113, 844, 88, 843]. averaging [1834, 2047]. Avoiding [1993]. axiomatic [80].

B [148, 168, 189, 464, 368, 522, 105]. backcasting [1131]. backward [1877]. Bacon [167]. balance [86]. balanced [74, 1663, 1887, 1062, 53, 1281]. Ballot [116]. band [1390]. band-recovery [1390]. banding [1943]. bands [1808, 404, 537]. bandwidth [993]. Banks [2115, 201]. bar [577, 641]. barriers [47]. BART [1475]. Bartholomew [496]. Bartlett [1841]. base [947, 878, 15]. based [1795, 1625, 1837, 1035, 1202, 1648, 1747, 1668, 1802, 1304, 302, 300, 299, 301, 1088, 2041, 904, 1694, 1529, 1879, 1678, 715, 723, 1044, 1090, 1662, 2048, 1172, 1681, 1893, 859, 1027, 1048, 1649, 1380, 1828, 1514, 1916, 1850, 434, 1579, 112, 422, 1495, 947, 742, 1076, 1778, 1840, 1622, 1979, 1262, 453, 1476, 1602, 15, 1371, 1412, 1684, 332, 1394, 1807, 1686, 1745, 373, 1182, 1402, 1765, 1955, 1964, 638, 2072, 1174, 1481, 1689, 1195, 593, 841, 1547, 1836, 1453, 1509, 1511, 1249, 1582, 1556, 1703, 2034, 1936, 1583]. bases [1961]. basic [1533, 293, 80]. Basilevsky [460]. basis [1118, 322]. basketball [2046]. Bay [1816]. Bayes [616, 58, 776, 625, 1055, 903, 1125, 773, 869, 1093, 435, 981, 419, 972, 1056, 1033, 709, 1759, 1689, 806, 1231, 1571]. Bayes-frequentist [1093]. Bayesian [1444, 2019, 1625, 561, 608, 847, 1117, 247, 418, 658, 914, 1456, 1692, 2027, 1144, 1053, 649, 421, 1705, 823, 1591, 1240, 59, 1054, 2028, 1904, 983, 1079, 1360, 1758, 1137, 1264, 2004, 1408, 626, 819, 1387, 1796, 1989, 1449, 1410, 874, 1915, 1735, 818, 1016, 1145, 1288, 423, 1687, 1570, 1934, 900, 1908, 694, 1782, 1811, 2045, 1523, 1744, 1907, 1386, 1921, 281, 982, 801, 971, 1626, 1310, 431, 1190, 1401, 1938, 1402, 1017, 709, 1496, 748, 936, 1232, 1720, 682, 790, 1308, 495, 1661, 1721, 1895, 1002, 1556, 1969, 2014]. bayesienne [59]. be [37, 94, 8]. Beaufort [229]. Beer [581, 582, 583]. Behara [222]. behave [538]. behavior [1001, 1003, 337]. Behaviour [673, 1270, 1104, 674, 594, 988, 857, 798]. behaviours [1453]. Behrens [598, 1093]. Behrens-Fisher [598, 1093]. Belief [592]. beliefs [41]. Benchmarked [1801]. benefits [1842, 489]. Bent [1490]. Bent-cable [1490]. Berkeley [148]. Berkson [1348, 1718, 1357]. Berlin [520]. Bernoulli [844]. Bernstein [1103, 971, 1745, 1571]. Berry [244, 357]. Best [657, 2016, 1201, 497, 52, 1663, 810, 685, 1171]. beta [1625, 1187, 1886, 727, 1915, 1906, 2025, 899]. beta-process [899]. beta-Stacy [1625]. beta-type- [727]. Better [761, 2052, 354, 1762, 132]. between [247, 526, 278, 340, 497, 612, 782, 407, 1578, 898, 2001, 558, 69, 293, 1195, 1491, 1628, 360]. Beyond [964]. Bias [1413, 1858, 937, 969, 153, 1455, 1111, 1592, 2064, 735, 512, 918, 527, 2072, 1621, 931, 1518]. bias-corrected [1518]. bias-reduced [2072]. bias-robust [931]. Biased [1202, 1438, 1957, 1019, 1419, 1394, 1540, 1931, 1541, 1169, 1174, 1755, 525]. biases [1759]. Bibliographiques [221, 222, 290, 289]. Bibliography [2086, 233, 147, 832, 67]. bidimensionnelles [605, 405, 926]. bière [582]. Big [1777, 1778, 1894, 1780, 1773]. bilinéaires [1300]. bilinear [1300]. Bill [309]. Binary [771, 685, 1333, 659, 1761, 1695, 1229, 1349, 2052, 1711, 1412, 756, 1217, 1122, 1033, 876, 1890, 748, 1215, 1481, 1876, 1249, 1513, 1792]. Binary-image [771]. binomial [761, 686, 403, 40, 1061, 1007, 1170, 229, 977, 451, 709, 1571]. binomiales [977]. bioassay [740, 741, 853, 739]. bioassays [1090]. bioinformatics [1393]. biological [1407, 1745, 130, 254]. biology [1392]. biomarker [2048, 1651, 2050, 1549, 1628]. biomarkers [1527]. Biomedical [2089, 148]. biomolecular [1640]. Biostatistics [2127, 290]. biplot [676]. Birnbaum [105]. birth [277, 414, 363, 14]. birth-and-death [277, 414]. birth-death [363]. birthday [2159]. Bivariate [793, 1270, 653, 721, 1930, 1277, 1143, 1176, 1410, 622, 1767, 1288, 1881, 140, 1623, 1524, 1000, 1413, 1634, 12, 1614, 907, 325]. black [1392]. black-footed [1392]. Blake [218]. Blending [1592]. Blest [1184]. Blind [1107]. Block [1236, 1297, 1972, 1887, 427, 1727, 99, 728, 172, 374, 35, 65]. block-thresholding [1972]. blockchain [1965, 1885]. blocking [1216]. blood [626]. BLUEs [278, 340]. BLUS [1032]. Blyth [222]. BMDP [2089, 148]. BMDP-77 [2089, 148]. board [1716, 1707, 255, 1724, 1733]. bodies [644]. Bombay [202]. Book [149, 166, 148, 170, 219, 220, 237, 238, 239, 240, 266, 267, 265, 288, 309, 519, 122, 188, 205, 169, 104, 186, 103, 147, 165, 167, 168, 189, 124, 202, 201, 307, 217, 464, 285, 187, 204, 218, 287, 102, 310, 311, 150, 123, 368, 2095, 55, 185, 135, 459, 80, 152, 522, 43, 200, 30, 284, 286, 105, 203, 461, 520, 466, 151, 460, 462, 121, 369, 521, 134, 463, 308, 184, 306, 465]. Books [205]. Bootstrap [1293, 869, 478, 604, 1645, 527, 1439, 2042, 733, 715, 979, 1576, 690, 1064, 1605, 1579, 1294, 1676, 860, 1921, 1067, 538, 1727, 667, 706, 1709, 923, 1236]. bootstrap- [1067]. Bootstrapping [1418, 1032, 980]. born [1984]. borrowing [1584]. Boston [167]. both [849, 1754, 1743]. bound [78, 906, 357]. boundaries [76, 363]. boundary [1338, 2078, 1283]. Bounded [1563, 763, 804]. Bounded-influence [763]. Bounds [45, 68, 1149, 211, 405, 281, 806]. Bovas [459]. bowel [1150, 1153, 1151, 1154]. Box [904, 1156, 894, 1939, 1405]. brain [1909, 1910]. brain-wide [1910]. branching [260, 298, 1485, 2163, 150]. brand [36, 194]. Breach [461]. break [431]. break-point [431]. breaking [815]. breaks [2009]. breast [1465]. breeding [24]. Bregman [1850, 1433]. Brick [167]. bridge [19, 2031]. bridges [560]. Brillinger [237]. Bristol [188]. British [1323, 498, 1949]. broad [1927]. broken [1228, 1797]. broken-line [1228]. Brown [148]. Brownian [2146, 456, 509, 520]. Brusegard [200]. Budapest [521]. Byers [167].

C [149, 148, 238, 267, 1202, 188, 205, 222, 165, 1210, 287, 633, 522, 200, 284, 203, 462, 134, 465, 2051]. C$17.50 [307]. C$20.85 [202]. C$25.00 [240]. C$30.00 [285]. C$37.50 [309]. C$6.95 [290]. c- [2051]. cable [1490]. calcul [2098, 184]. calculation [1713]. calculations [600]. Calculator [2161]. calculus [1837, 125]. calibrated [1677]. calibration [1997, 1094, 1368]. California [148]. call [1044]. Calle [1235]. Cambridge [103]. camp [609]. Campbell [222]. camper [612]. Canad [721, 1202, 1140, 778, 941, 1063, 1210, 897, 745, 1235, 1245, 1013, 1247]. Canada [2125, 200, 462, 134, 312, 636, 5, 171, 391, 487, 638, 338, 489]. Canadian [1314, 1329, 1947, 1814, 2024, 499, 2036, 637, 379, 380, 381, 382, 194, 510, 483, 2037, 467, 16, 468, 1340, 1364]. canadiana [222]. Canadians [582, 583, 581]. canadiennede [1014]. canadiens [483, 582]. cancer [786, 1465, 1556]. candidate [1155]. canonical [145, 173]. capacity [697, 698]. Capture [1385, 1389, 1390, 1132, 1391, 1464, 1452]. Capture-recapture [1385, 1389, 1390, 1132, 1391]. caractéristiques [1325]. carcinogen [853]. care [637, 636, 638]. Carl [121]. Carl-Erik [121]. Carleton [309]. Carlo [1077, 914, 2028, 232, 1005, 1483, 1209, 910, 1606]. carpocapse [1325]. Carter [290, 466]. cas [1026, 932, 50, 335]. Case [258, 1150, 1322, 1041, 1151, 313, 1521, 1040, 1239, 1026, 932, 1574, 1742, 50, 1162, 499, 585, 637, 1098, 580, 1543, 1842, 1241, 1617, 297, 44, 440, 1657, 1984, 1507, 1392, 1816, 959, 697]. case-cohort [1543, 1241]. case-control [1239, 1742, 1842, 1617, 1657]. Case-study [697]. cases [1139, 471]. Cassel [121]. Catalogue [200]. Categorical [103, 174, 2090, 2130, 240, 1459, 1398, 1437, 875, 1263, 990]. categories [73, 157]. category [1337]. cattle [411]. Cauchy [444, 85]. Causal [2033, 2062, 1590, 1998, 2007, 1639]. cause [916]. CDF [1176]. cell [1817, 1027]. Cellular [908]. Cellwise [2015]. Censored [1352, 1765, 669, 1625, 757, 1251, 1785, 1613, 404, 1176, 354, 645, 1235, 1535, 1882, 1926, 333, 1493, 1580, 890, 1358, 889, 1738, 1643, 899, 1825, 231, 1253, 1755, 1786, 1712, 303, 1252, 1415, 1611]. Censoring [425, 1653, 1202, 1276, 661, 615, 1828, 1235, 505, 1234, 1931, 1657, 1487, 1691, 1406, 1870]. censorship [690, 537, 959]. Census [489]. central [552, 430, 1622]. centrality [433]. centre [1760]. centroid [1977]. century [2156]. Certain [92, 12, 497, 57, 979, 447, 448, 254]. Chaînes [2103, 186]. chain [1668, 2028, 342, 1209, 910]. Chains [2145, 2147, 2162, 257, 715, 1408, 111, 920, 520, 520]. challenges [1775, 1965, 1014]. Chambers [122]. Champs [110]. Change [1471, 86, 115, 1130, 976, 1244, 1860, 1879, 1670, 1883, 1719, 824, 1848, 1601, 1056, 857, 1727, 1877, 2068]. Change-over [86]. change-point [1130, 976, 1860, 1719, 1848, 1601, 857, 1877]. changepoint [1456, 718, 513]. changing [1074]. characteristic [588, 692, 816, 81, 426, 2002, 1469]. characteristics [612, 1325, 1989, 972]. Characterization [841, 6, 28, 705, 725, 114, 38, 377, 304, 144]. Characterizations [2093, 1235, 1234, 142, 280, 166]. Characterizing [1988]. charts [793, 935, 738]. Chaubey [462]. Chebyshev [1671]. Checking [1832, 1795, 1420, 1406]. chef [1181, 1248, 1292]. chemicals [741, 739]. chemistry [583, 581]. Chen [1245]. chi [1668, 830, 577, 641, 1135, 751, 397]. chi-bar-square [641]. chi-bar-squared [577]. chi-square [1668, 830, 751, 397]. child [637, 636, 638]. child-care [638]. children [609, 613, 610, 1453]. choice [1441, 905, 553, 562]. choix [723]. Cholesky [1916]. Cholesky-based [1916]. chromosome [1153]. chronologiques [482, 348]. cigarettes [194]. circle [1009, 207]. circles [618, 1597]. circuits [2080]. circular [577, 640, 983, 1735, 1770, 1178, 1472, 1854, 44]. circular-cone [640]. CJS [2022, 1531, 1587]. Claes [121]. Claes-Magnus [121]. class [653, 721, 277, 978, 949, 454, 552, 1305, 1603, 873, 1951, 1082, 71, 132, 1824, 1298, 792, 966, 988, 1580, 751, 336, 1927, 563, 109, 1190, 1444, 209, 579, 199, 1544, 172, 1669]. Classement [488]. Classes [276, 788, 657, 1233, 1]. classical [421, 1348]. Classification [1339, 1866, 2070, 59, 984, 657, 92, 10, 37, 1605, 31, 694, 1873, 1527, 1321, 1684, 191, 1433, 2014, 1044]. Classifications [2117, 1412, 221]. Classified [103, 2090, 2067, 1545]. classifier [332, 1977]. Clayton [1955]. clear [1149]. climate [1670]. clinical [655, 1240, 73, 157, 700, 2048, 1989, 1371, 1760, 1583, 1628]. Closed [1494, 1391]. Closed-form [1494]. closeness [805, 576]. Cluster [1978, 1344, 1567, 1025, 1771, 1207, 310, 1704, 2129, 2012, 834, 1606, 1729, 1876, 1482]. cluster-randomized [1482]. cluster-sample [1606]. clustered [1098, 2046, 1996, 1538, 1356, 1854, 1658, 1033, 1825, 954]. Clustering [62, 1747, 138, 2011, 1372, 2045, 1476, 1976, 1684, 1509]. clusters [1057, 1105, 875, 1513]. Co [205, 147, 124, 135, 466, 460]. coarse [1586]. Cocozza [1210]. Cocozza-Thivent [1210]. codling [1322, 1325, 1326, 1323, 1324]. coeditors [310, 311]. coefficient [653, 721, 1787, 1492, 1843, 378, 325, 790, 1465, 1317, 1944]. coefficients [1300, 1561, 516, 2049, 2046, 419, 1413, 548, 1835, 2039]. cognitive [1144]. coherence [2001]. cohort [1533, 1543, 786, 1241]. cohorts [1754]. Coin [612, 662, 611, 609, 613, 610, 673, 1419]. cointegrated [1353]. Collaborative [1814]. collection [489]. Collier [151]. collision [358, 360]. collisions [908]. Colloquium [134]. coloured [1782]. Columbia [498, 1949, 1323]. combination [1527, 397, 1872]. combinations [475, 512, 1060]. Combinatorial [2154, 2]. Combinatorics [2122, 206, 203]. Combined [1666, 790, 1853]. Combining [1615, 2013, 1790, 1213, 785, 501]. commentaires [572]. Commentaries [571]. Comments [743, 812, 584, 383, 639, 316, 361, 470]. Common [1859, 1858, 807, 376, 1295, 802]. Common-factor [1859]. commutation [295]. Compact [2106, 2123, 2158, 821, 188]. companies [488]. Company [202, 290, 289, 368, 134, 308]. comparative [1757, 1812]. Comparing [526, 667, 90, 614, 562, 1714]. Comparison [399, 100, 322, 1535, 2148, 579, 394, 1550, 321, 1192, 922, 138, 853, 524, 443, 228, 918, 159, 1195, 1542, 1618, 368]. comparisons [1546, 1866, 362, 49]. compatible [118]. compensateur [1075]. compensation [2054]. compensator [1075]. compensators [352]. competing [1696, 1935, 916, 1730, 334, 1614, 1643, 1461]. competing-risk [334, 1643]. competing-risks [916]. Compilation [102]. Compiled [105]. Complete [1364, 602, 1951, 86, 1114]. Complex [2035, 1459, 1155, 1869, 1838, 1595, 1890, 1675, 2080, 1315]. complexity [1347, 1195]. compliance [1482]. complications [1731]. complier [1998]. component [893, 1706, 498, 1594, 1483, 1474, 1933, 1640, 1959]. component-guided [1959]. Components [742, 442, 775, 893, 723, 945, 816, 1060, 11, 1182, 1439, 1068, 1206, 213, 1245, 119, 1562, 227]. comportement [1001]. composantes [723]. Composite [2072, 1271, 2041, 1305, 1676, 2082, 1666, 1618]. compound [1688]. compromise [915]. Comptes [219, 220, 265, 288, 309, 307, 217, 285, 218, 287, 284, 286, 308, 306]. computation [1530, 1751, 1002]. Computational [2088, 1470, 61, 122]. Computationally [788]. computations [1299]. Computer [2089, 148, 19, 1591, 165, 1894, 1148, 489]. computer-assisted [489]. computer-dealt [19]. Computers [2106, 188, 2123, 2158]. computing [1227]. concave [1164]. concentration [589]. concentrations [1610]. Concepts [831, 605, 2153, 464, 80, 1014]. concise [107]. concomitant [449]. concomitants [1311]. concordance [1620]. Concordia [462]. concurrent [274, 1974]. Condition [264, 595, 395, 1235, 802]. Conditional [840, 813, 1791, 650, 1023, 446, 939, 1276, 250, 1918, 1074, 1138, 1238, 1437, 393, 825, 392, 660, 331, 474, 730, 118, 1659, 1212, 1420, 1623, 1660, 798, 2050, 126, 841, 1953, 54, 941, 745]. conditionality [425]. conditionnel [939]. conditionnelle [54]. Conditions [756, 1104, 487]. cone [577, 640]. Conference [1269, 2164, 462]. Confidence [1808, 1085, 1140, 1501, 404, 894, 807, 973, 992, 2082, 1426, 1653, 1035, 761, 2042, 1830, 403, 1185, 1192, 261, 545, 585, 223, 178, 474, 443, 537, 424, 1099, 1789, 26, 1493, 1127, 648, 746, 1897, 1067, 336, 538, 934, 1102, 1621, 90, 704, 1424, 1315]. confidence-interval [648]. conflict [1993, 671]. confounding [1608]. conjecture [695]. Conjugate [1069, 845]. Connectedness [445]. Connectivity [1911]. Connectivity-informed [1911]. connectome [1905]. conquer [1894, 2060]. consequences [700, 595, 595]. conservation [1392]. Conservative [1316, 1552]. considerations [334]. Consistency [1187, 1603, 691, 1113, 1130, 1282, 645, 2071, 17, 802]. Consistent [550, 1719, 439, 1218, 1353, 1525, 1355, 118, 568, 756, 648, 746, 1823, 389, 515, 648, 746]. constant [955, 1235]. constant-sum [1235]. Constrained [1564, 1555, 958, 1253, 729, 1600, 1010]. constrained- [729]. constraint [1268]. constraints [1567, 796, 1208, 1059, 734, 797, 483]. constructed [1804]. constructing [53, 877, 1836]. construction [74, 1636, 329, 65]. constructions [1559, 1523]. Consul [633]. Consultant [612, 518, 662, 611, 609, 613, 622, 580, 635, 610, 634, 673]. consulting [622]. consumption [1818]. containing [1185, 1501]. Contaminated [1486, 2018, 1845, 1941]. contamination [455]. content [727, 1947]. contents [1269]. Contiguity [2150, 461]. continent [944]. continental [942]. contingency [608, 100, 751]. Continuous [1896, 2020, 1408, 1229, 1663, 47, 1711, 2045, 1738, 1621, 1442, 1590, 1378]. continuous-scale [1621]. continuous-time [1408, 2045]. Continuum [1977]. contoured [396]. contraintes [483]. contrast [830, 1319, 996]. contrastes [830]. contrasts [99]. Contre [633]. Contre-exemple [633]. Contribution [50]. contributions [2024, 2036]. control [1239, 1742, 793, 935, 2015, 1842, 1744, 1617, 362, 1657, 1507, 774, 738]. controlled [559]. controlling [1819]. controls [853, 650, 745]. Convergence [976, 932, 632, 98, 1010, 352, 126, 234, 530, 197, 960, 643, 127, 412, 1078, 389, 932]. convergent [939]. convergents [939]. conversation [2075, 356]. convex [291, 536, 644]. convexe [291]. convexity [1222, 139, 1516]. convolution [207]. convolutions [1633, 1402]. Copula [1893, 1840, 1815, 1559, 1276, 1274, 1359, 1678, 1600, 1805, 1523, 1861, 1413, 1278, 1825, 1691, 1272, 1553]. Copula-based [1893, 1840, 1678]. copula-graphic [1276]. copulas [1642, 405, 926, 1500, 1551, 1431, 1634, 1601, 1734, 1614, 1273]. Copules [405, 926]. copy [1877]. cores [944]. Corner [612, 518, 662, 611, 609, 613, 635, 610, 634, 673, 622]. Corner/Le [612, 662, 611, 609, 613, 610, 673]. Corners [580]. Cornwell [311]. Corporation [522]. Corrected [1317, 1698, 1518]. Correction [768, 1210, 1597, 1841, 1283, 1455, 1245]. Corrections [654, 1497, 1985, 1759]. corrector [657]. Correlated [619, 1890, 1567, 1267, 246, 1082, 824, 1514, 47, 2001, 1887, 1207, 1148, 834, 1377]. correlates [2050]. Correlation [1079, 1311, 1729, 653, 721, 814, 411, 879, 1561, 547, 1355, 1184, 131, 228, 112, 378, 447, 1413, 974, 548, 1626, 1303, 1353, 1728, 325, 69, 62, 540, 102]. correlations [141, 619, 394, 856, 1246, 1247]. corresponding [868]. Corrigenda [1140, 1105, 1497, 1985]. corrigenda/corrections [1497, 1985].

Corrigendum [339, 340]. cortical [1904]. cost [35]. couchiching [612, 611]. count [1480, 1210, 1193, 1649, 1372, 846, 1536, 1619, 1607]. Countable [2145, 520]. counterexample [695, 633]. Counting [2121, 239, 628, 1537, 564]. counting-process [628]. counts [861]. Coupling [398]. courbe [15]. Course [2139, 220, 30]. courses [248]. CoV [2005]. covariance [921, 449, 18, 259, 766, 1916, 1630, 82, 726, 1186, 1566, 448, 591, 693, 1943, 1728, 1641, 1874]. covariances [600]. Covariate [1702, 1757, 1584, 1081, 1791, 795, 1829, 1062, 1507, 1491, 1721, 1895, 1698, 1469, 1582, 1700, 1826]. Covariate-adjusted [1702, 1584]. covariate-environment [1826]. Covariates [1130, 1808, 1230, 1920, 1124, 1664, 1889, 1430, 1652, 873, 1488, 1145, 1978, 1111, 1882, 1809, 789, 1422, 1404, 1994, 2050, 1891, 1515, 1214, 1464, 1638, 1702]. covariation [1981]. Cover [1740, 1723, 1732]. coverage [1789, 704]. COVID [2004, 1949, 1988, 1947]. COVID-19 [2004, 1949, 1988]. Covid-19-related [1947]. Cox [985, 904, 1156, 690, 1889, 894, 1515, 1939, 1698, 1405, 1826]. Cox-type [1405]. Cramér [755, 1080, 140, 1350, 881, 455]. Cramér-von [755, 1080, 1350, 881, 455]. cricket [1435]. critère [723]. Criteria [497, 396, 562]. criterion [305, 723, 1335, 21, 576, 1076, 82, 1056, 128, 1146]. Critical [2033, 2027]. Critiques [219, 220, 265, 288, 309, 307, 217, 285, 218, 287, 284, 286, 308, 306, 1715, 1748, 1783, 1821, 1863, 1912, 1986]. Cross [103, 2090, 2117, 1117, 1470, 221, 1545, 785, 1033, 350, 1593, 1991, 856, 1246, 1247, 990]. Cross-Classified [103, 2090, 1545]. cross-correlations [856, 1246, 1247]. cross-national [1593]. cross-sectional [785, 350, 990]. cross-sectionally [1033]. cross-validation [1117, 1470, 1991]. crossings [116]. crossover [1281]. crude [1618]. crude-oil [1618]. cryptic [1819]. cryptocurrency [1885]. Csiszár [429]. Csörgo [222, 521]. cube [1294]. Cui [1202, 1202]. cumulant [556]. cumulants [556]. cumulative [1081, 1717, 1654, 2050]. Cure [1743, 1287, 1696, 1889, 1565, 1638, 1556]. cured [1672]. Current [1537, 655, 2032, 2049, 1847, 1985, 1767, 1543, 1254, 1461, 1381, 1515, 2165, 369]. curtailment [681]. Curvature [367]. Curve [1031, 658, 787, 1769, 1584, 1934, 854, 15, 1512]. curves [1085, 1140, 1818, 1192, 1167, 1790, 69, 1469]. cut [1061, 1945]. cut-off [1945]. cut-points [1061]. cutpoints [1081]. CVX [1836]. CVX-based [1836]. cyclic [1881]. cylinders [618, 1597].

D [238, 288, 205, 222, 168, 189, 310, 123, 368, 522, 200, 462, 369, 134]. D. [2023]. d’abnonnement [1290]. d’abonnement [1280, 1328, 1365]. Dabrowski [1446, 1445]. d’admissibilité [595]. DAG [1559]. d’agrégation [50]. daily [486, 2005]. Dalhousie [2164, 289]. dall/Hunt [290]. damage [142, 315]. Daniel [167]. Danièle [152]. dans [1026, 932, 1300, 59, 1261, 450]. d’Anscombe [977]. Dantzig [1599]. d’arrondi [646]. Data [237, 2133, 122, 1441, 2088, 1529, 2090, 2115, 2128, 2141, 2130, 669, 1625, 1480, 1045, 1230, 696, 814, 240, 258, 1150, 1322, 1265, 914, 757, 1648, 58, 831, 953, 1742, 1459, 817, 1304, 611, 1251, 1558, 776, 1375, 1830, 1223, 949, 793, 1561, 1152, 1437, 1613, 1841, 1930, 1972, 1879, 1277, 1672, 1777, 1486, 983, 740, 1567, 375, 1360, 1769, 1348, 1957, 1569, 1909, 1069, 2004, 2011, 1752, 1993, 2054, 1041, 1151, 1326, 201, 2049, 1418, 1407, 1581, 1670, 1133, 1761, 626, 1390, 729, 313, 1176, 1521, 480, 1116, 1301, 524, 956, 1847, 1985, 1210, 1193, 815, 1827]. data [1098, 674, 1659, 1260, 1568, 1735, 354, 1770, 1982, 1488, 1767, 1649, 1468, 1349, 1881, 1978, 2006, 749, 423, 851, 897, 1374, 824, 1372, 270, 330, 1484, 1687, 645, 1570, 1586, 324, 992, 314, 1207, 955, 1235, 1131, 846, 1535, 1548, 1543, 1536, 1730, 1778, 1805, 1840, 1882, 1873, 1894, 1630, 1639, 1701, 1926, 1996, 1204, 1838, 2044, 786, 1645, 333, 1537, 1306, 1476, 1907, 1493, 1589, 1995, 1766, 1845, 1052, 1968, 719, 1386, 1899, 756, 1538, 1356, 1973, 1580, 332, 967, 1807, 982, 1440, 1062, 473, 488, 1943, 889, 1540, 789, 752, 785, 1009, 1854, 1658, 1033, 1337, 1657, 2061]. data [1738, 510, 1872, 1983, 899, 1307, 1507, 1674, 667, 876, 1593, 1853, 394, 732, 748, 1825, 2060, 1604, 1254, 1461, 1544, 1835, 834, 1215, 1232, 493, 1040, 1774, 1381, 1253, 1096, 1432, 1729, 1755, 1786, 1780, 1891, 1960, 2010, 1712, 1515, 303, 525, 1453, 1464, 1753, 16, 1698, 1249, 1513, 1582, 1700, 1941, 1465, 1556, 1932, 990, 1252, 1442, 1571, 1619, 2009, 1271, 1415, 1607, 1631, 1797, 2069, 1977, 954, 1378, 1773, 103, 238, 267]. data-analytic [831]. Data-driven [1441, 817]. Dataset [811, 664]. David [219, 237, 124, 201, 520]. Dawson [222]. Day [237, 411, 1435]. days [487, 338]. Deal [376]. dealing [1507]. dealt [19]. death [277, 414, 363]. decay [726]. December [1740]. d’échelle [544]. Decision [2110, 58, 205, 1240, 1011]. Decision-Making [2110, 205]. decision-theoretic [1240]. deck [1645]. decline [1144]. decomposable [1641]. Decomposition [364, 792, 1874]. decompositions [1320]. Deconvolution [1224, 668, 1319, 550, 1603, 568, 656, 1243, 1884]. Dedicated [2159]. default [1125]. Deficiencies [70]. deficiency [1590]. defined [714, 1025]. defining [177]. définis [714]. definite [1908]. definition [193]. Definitions [305]. d’égalité [348]. degenerate [719]. degradation [1432]. Dehen [152]. Dekker [265, 169, 189, 150, 185, 105]. Delayed [88, 1419]. delays [747]. d’éléments [98]. deletion [1049, 440]. deletions [137]. Delhi [166, 187, 204, 284]. Demand [2126, 5, 217]. Demography [2151, 463]. demonstration [293, 293]. DeMoSTAFI [1269]. denoising [2043]. dense [1992]. densité [768, 1250, 646]. densités [892]. densities [530, 1224, 1633, 800, 892, 63, 556, 559, 325, 420]. Density [615, 768, 1183, 669, 769, 817, 1250, 1187, 78, 1251, 1296, 1003, 1282, 1441, 1867, 1319, 1758, 1348, 156, 550, 646, 645, 1283, 1399, 1058, 1106, 275, 568, 1718, 1023, 2025, 735, 1243, 594, 627, 1218, 971, 1008, 1086, 128, 647, 993, 1712, 17, 1945]. d’épandage [1325]. Département [152]. Department [165, 465]. dépendance [605]. Dependence [1269, 1038, 817, 1924, 1359, 1446, 1073, 1409, 1458, 32, 1500, 1600, 1410, 791, 1770, 452, 1804, 1551, 1400, 988, 1275, 1273, 1491, 605]. dependent [1757, 197, 930, 1276, 1375, 1806, 987, 1277, 454, 808, 825, 1569, 762, 1468, 672, 1536, 1894, 1651, 594, 1386, 1538, 1051, 1657, 579, 1691, 1311, 2055, 1697, 1607]. deposition [493]. Dept [219, 220, 222, 462]. Depth [1752, 1529, 1879, 2070]. depth-based [1529]. Depth-weighted [1752]. depths [1807]. derailment [315]. derivatives [932, 1525, 531]. derivées [932]. describing [1318]. description [697, 1321]. Design [2114, 2157, 2160, 1148, 1831, 1456, 321, 1333, 1591, 2029, 1240, 2154, 944, 1162, 73, 157, 950, 1989, 902, 862, 246, 772, 1800, 1951, 206, 2052, 1873, 1922, 1100, 1, 1194, 1964, 878, 1519, 1573, 1188, 284]. design-based [1964]. Designing [1452]. Designs [153, 2116, 704, 2051, 24, 1216, 1256, 445, 319, 1742, 305, 1636, 905, 1137, 1264, 146, 39, 1019, 1208, 1334, 1101, 228, 437, 1419, 1887, 1608, 1242, 1616, 2053, 1411, 1651, 966, 1186, 2084, 1163, 2062, 427, 185, 109, 1760, 562, 685, 53, 86, 1205, 1368, 99, 1179, 876, 728, 877, 1149, 172, 374, 1281, 65, 1876, 826, 1021, 1836, 1799, 1703, 1792, 1628, 1702, 55]. Dessaisonalisation [483]. d’estimateurs [939]. d’estimation [1022]. details [2058]. detect [1491, 1628]. Detecting [708, 1670, 1576, 1020, 198, 540]. Detection [1110, 1084, 1471, 1883, 1710, 1719, 498, 1840, 2015, 1413, 1601, 1310, 1640, 1877, 2010, 2068]. determinants [106]. Determination [518, 635, 740]. determining [1182]. Deterministic [2134, 265]. Deux [1022, 830, 348]. development [148]. Developments [2101, 863, 318, 2033, 307]. deviations [1447, 32, 563]. Df1.85.00 [289]. diabetic [1731]. Diagnosing [1030]. Diagnostic [1111, 1420, 1643, 603, 682, 1286, 1621]. diagnostics [1293, 1379, 1237]. diagonal [82, 1874]. Diagrams [1533]. diatom [494]. diatom-inferred [494]. dichotomous [1628]. Diegel [102]. dietary [949]. difference [247, 830, 320, 115, 830]. differences [1041, 1040]. différencier [1042]. different [228, 2013]. Differential [1047, 76, 264, 78, 56, 1533, 1971, 1809, 1494, 1868]. differentiate [1042]. differentiation [656]. différentielle [264, 2104, 152]. difficult [951]. difficulties [195]. diffused [1796]. diffused-gamma [1796]. Diffusion [2146, 1908, 1160, 563, 520]. diffusions [1510, 1514]. digital [699]. diluted [1537]. dimension [2017, 1622, 1594, 2002, 2069]. dimensional [1775, 605, 1920, 405, 926, 902, 1942, 1940, 2053, 1907, 2043, 1943, 1015, 1720, 2010, 2068, 1991, 1874, 1868, 2039, 2008]. Dimensionality [1108]. dimensions [1558, 228]. d’indépendance [872]. d’information [1261]. d’insectes [1325]. Direct [1970, 1868]. direction [2016]. Directional [1058, 1106, 776, 1318, 1770, 1346, 1849, 1807, 433]. directions [439]. Dirichlet [1757, 243, 565, 1161, 1934, 933, 1060, 11, 1015, 901, 29]. Dirichlet-process [243]. disabled [486]. Disciplines [2105, 187]. discontinu [1075]. discontinuous [1075]. discordancy [534]. Discovering [288, 2142]. discovery [2015]. discrepancy [70]. Discrete [1427, 208, 429, 1997, 1085, 1140, 755, 827, 77, 1350, 1995, 1402, 1738, 899, 1077, 160, 841, 959]. Discrete-time [1427, 1995]. discretely [1590]. discriminant [1128, 1045]. discrimination [418, 1397, 1935, 471, 259, 694, 947]. Discussion [865, 866, 1157, 224, 571, 885, 273, 269, 1046, 1326, 271, 1155, 385, 1387, 402, 911, 409, 343, 1119, 912, 500, 1902, 886, 1901, 1226, 344, 572]. Discussions [1065]. disease [1150, 1144, 1153, 1151, 1155, 1889, 1827, 1527, 1903, 1902, 1739, 1154, 1508, 1196, 1901, 1939]. dispersed [1649]. dispersion [792, 918, 1206]. Distance [1807, 1338, 1925, 1197, 833, 453, 928, 1012, 1013, 1182, 1457, 1650, 1953]. Distance-based [1807, 1182]. distance-to-boundary [1338]. distances [820, 453]. distinguishability [23]. distribué [1026]. Distributed [2059, 2137, 1026, 311]. Distribution [653, 1382, 566, 967, 448, 81, 856, 1246, 768, 6, 28, 106, 277, 1747, 444, 840, 23, 19, 1139, 817, 414, 243, 1164, 183, 50, 516, 1867, 176, 54, 577, 983, 547, 589, 1059, 133, 1037, 651, 819, 707, 1410, 845, 230, 1915, 827, 1288, 766, 114, 452, 229, 283, 1663, 38, 1934, 72, 377, 415, 835, 378, 1262, 310, 1393, 82, 182, 242, 1493, 236, 2129, 689, 1, 281, 974, 89, 666, 397, 1060, 1115, 280, 22, 510, 48, 282, 881, 1126, 325, 1168, 426, 706, 101, 172, 2080, 804, 764, 1491, 64, 90, 428, 458, 533, 1114]. distribution [29, 679, 1631, 721, 390, 1247]. Distribution-free [1382, 967, 1867, 48]. Distributions [8, 629, 1270, 396, 1112, 1332, 618, 1597, 665, 1416, 1647, 1085, 1140, 1685, 40, 893, 935, 1395, 755, 725, 1192, 1055, 441, 393, 825, 1722, 590, 27, 296, 641, 513, 668, 1923, 1318, 21, 532, 96, 1316, 567, 766, 1024, 2006, 322, 749, 77, 977, 112, 438, 633, 1350, 1838, 726, 320, 933, 2000, 1538, 1921, 917, 1790, 1472, 716, 1190, 1444, 12, 75, 514, 364, 433, 1121, 160, 841, 130, 304, 1669, 1980, 1550]. disturbances [85]. divergence [1837, 1668, 1379, 1722, 1850, 801, 1433]. divergence-based [1837]. diverging [1729]. diversity [130]. Divide [2060]. dividend [50]. dividendes [50]. divisible [197, 65]. Dixon [148]. DM [464]. DM105.00 [466]. DNA [1035, 947]. Do [538]. d’Oja [737]. domain [1592, 1832, 1612, 1736, 1519]. domains [371, 1260]. Dominance [2110, 205]. dominant [1617]. Dominique [152]. Don [368]. Donald [105]. données [405, 488]. donor [1451]. donors [626]. dont [405]. Door’ [1880]. d’optimisation [2098, 184]. d’ordre [1026, 855, 54, 293]. Dordrecht [134]. d’oscillations [335]. Dose [270, 741, 1465, 1792]. dose-intensity [1465]. Dose-response [270, 741, 1792]. double [59, 46, 1956]. double-threshold [1956]. Doubly [2041, 1981, 1568, 1673, 1419, 1855, 2007, 1503]. down [938]. drift [942]. driven [1793, 817, 1441]. drop [1674, 1249, 1482]. drop-out [1674, 1482]. drop-outs [1249]. Dropping [296]. DT [24]. Dual [240, 2101, 2130, 1806, 59, 307]. dual-classification [59]. dual-record [1806]. Dubuc [184]. Dubuque [290]. due [2064]. Dummy [2008]. d’un [855, 54, 595, 566, 15, 483]. d’une [768, 646, 291]. Duration [1554]. Duxbury [123]. Dwivedi [462]. Dynamic [2063, 1995, 1774, 738, 1671, 71, 1589]. dynamics [1988].

Early [338]. EAS [2071]. Eastern [166, 187, 204, 284]. easy [53]. échantillon [54]. échantillonnale [291]. echelon [688]. echelon-form [688]. Ecology [2134, 265]. econometric [4, 71, 5]. Economic [2040]. economics [217]. Ed [1716, 1724, 1733]. EDF [849, 1662]. EDF-based [1662]. edge [1710]. Edgeworth [711, 733, 690, 1005, 1704]. edit [480]. Edited [219, 104, 222, 165, 168, 307, 289, 368, 462, 205, 200, 369, 134]. Edition [237, 285, 135, 463, 306, 148, 170]. editor [148, 383, 639, 1092, 316, 1466, 361, 470, 42, 1092, 1814, 1202, 1467, 963, 1029, 1384, 1425, 1857, 1141, 1181, 1773, 1968, 1245, 1248, 1292, 1291]. Editor/Lettre [1202, 1245]. Editorial [1656, 1588, 909, 1341, 1425, 1967, 1211, 1587, 1707]. Editors [743, 1342, 2074, 1899, 584]. Edits [2148, 368]. Edmonton [307, 285]. Edward [464, 308]. EEG [1906]. Effect [131, 198, 1590, 1998, 997, 2065, 2007, 2048, 1576, 845, 646, 1488, 1349, 726, 1243, 1655, 1538, 675, 1391, 262, 325, 871, 1823, 1214, 540, 2008, 1629]. effectiveness [1325]. Effects [809, 1062, 609, 613, 781, 59, 1632, 2011, 1198, 1497, 1725, 1361, 1173, 1881, 610, 795, 583, 1370, 1639, 381, 756, 1538, 1754, 1854, 350, 53, 86, 213, 1357, 1491, 1624, 491, 1482, 1030, 1442, 1188]. effet [646]. effets [59]. efficacité [557, 1325]. efficacy [1322]. efficiencies [92, 324]. Efficiency [770, 334, 1635, 705, 557, 346, 588, 558, 447, 918, 1865, 374, 1424]. Efficient [1696, 1923, 628, 1241, 1461, 2068, 931, 788, 83, 1470, 1561, 1238, 399, 1180, 1922, 1824]. eighties [484]. elapsed [1408]. Elastic [1922]. electricity [1818]. electroencephalography [1907, 1853]. élémentaire [335, 293]. elementary [99]. elements [98]. Elephant [946]. Elephant-seal [946]. elicitation [984]. eligibility [595]. ellipsoids [704]. elliptical [1270, 653, 721, 836, 1647, 569, 766]. elliptically [396]. Elsevier [266, 287, 466, 460]. EM-test [1532]. embedding [349]. embeddings [1096]. Empirical [1785, 1830, 1296, 1918, 1185, 1192, 827, 1380, 1492, 972, 347, 1295, 1134, 1621, 1511, 1878, 1415, 1089, 1635, 1001, 1810, 1088, 1164, 1694, 1841, 2026, 413, 2024, 1448, 773, 869, 588, 692, 1659, 1828, 1762, 434, 1530, 2059, 816, 1493, 713, 419, 1620, 1560, 341, 797, 1954, 1033, 709, 1675, 1689, 1102, 1612, 1677, 1714, 1213, 1315, 1171, 1297, 1549, 50]. Empirical-distribution-function [827]. empirique [50]. empiriques [1001]. employee [2054]. employment [471]. employment-discrimination [471]. emulators [1739]. encouru [155]. Encyclopedia [2102, 151]. endemic [2004]. endemic-epidemic [2004]. endogeneity [1524]. endogenous [2008]. energy [5]. Engelman [148]. Engineering [2140, 286, 306]. Engineers [135, 2109]. England [104]. enigmatic [1436]. enlarged [560]. Enrollment [248]. entre [293]. entreprises [488]. entropic [1862]. entropie [264]. entropy [264, 1722]. environment [345, 1842, 675, 1826]. environmental [741, 739, 749, 1131, 272]. environments [370]. environmetrics [2036]. epidemic [2004, 1557, 1988]. epidemiologic [491]. Epidemiology [2152, 519, 1775, 1539]. episode [612]. Epistemic [134]. equal [1984]. equalities [278, 340]. equality [497, 216, 1222, 348, 236, 548, 75, 1641]. equalizer [679]. Equally [321, 228]. equation [87, 78, 2017, 85, 1533, 1971, 1494, 25, 1047]. equations [76, 2007, 56, 1111, 1076, 1610, 1809, 1560, 797, 1215, 1503]. Equator [2161].

Equi [437]. Equi-information [437]. equicorrelation [198, 790]. equity [883, 884]. equivalence [1429, 1505]. Equivariant [449, 591, 373]. Erdos [388, 1175]. ergodic [1087]. ergodicity [1074]. Erik [121]. Errata [542, 623, 543, 1247]. Erratum [1597, 1497, 1444, 1553]. erreur [646]. Error [1008, 1336, 630, 778, 1428, 949, 516, 906, 1920, 1360, 1335, 1090, 1787, 246, 1869, 646, 1871, 1585, 1605, 835, 1829, 356, 1824, 1718, 816, 1243, 594, 1421, 512, 1051, 1539, 774, 128, 1964, 834, 1502, 1689, 161, 804, 1357, 1919, 1753, 1698, 670, 1582, 1700, 1941, 1236, 1089]. error-components [816]. error-contaminated [1941]. error-in-variables [630, 778]. errors [2051, 769, 1284, 1348, 223, 1019, 762, 1267, 1471, 1173, 478, 485, 1262, 1765, 3, 1021, 1464, 1917, 923, 1317, 1377, 1944, 2047, 954]. espaces [98]. Esséen [244, 357]. estimate [1250, 244, 127, 646, 498, 1639, 587, 1562]. Estimated [1948, 935, 1376, 1838, 1147, 1394, 544, 1507]. estimated-score [1507]. estimates [530, 396, 83, 1282, 833, 235, 412, 741, 869, 657, 517, 645, 554, 478, 275, 835, 1592, 439, 735, 719, 494, 687, 1391, 481, 829, 798, 1070, 647, 119, 682, 790, 227, 337]. estimateur [768, 932, 646]. estimateurs [557, 595, 544]. Estimating [479, 1057, 1105, 825, 589, 1725, 637, 1989, 1869, 1605, 576, 1811, 1177, 1233, 1655, 1754, 487, 797, 596, 1482, 360, 1806, 2007, 20, 1064, 1111, 1076, 1610, 701, 1589, 320, 1560, 1540, 634, 1006, 1215, 1503, 806, 1939, 1682]. Estimation [1001, 502, 1250, 1818, 176, 59, 961, 740, 590, 892, 1059, 636, 739, 803, 1889, 1136, 1468, 824, 85, 556, 805, 2101, 703, 72, 1882, 1262, 245, 734, 1524, 342, 1306, 66, 627, 1848, 136, 1388, 297, 1033, 1760, 41, 2132, 1964, 638, 1691, 2050, 1121, 1168, 834, 161, 2005, 1464, 1917, 1941, 1611, 1183, 669, 921, 1022, 1998, 2019, 247, 616, 1265, 1668, 1084, 1451, 767, 769, 1144, 1244, 449, 1284, 1300, 822, 1333, 1050, 1083, 817, 1251, 1197, 1975, 1296, 2079, 1525, 1561, 987, 1238, 1359, 1972, 1633, 680, 1788, 787, 34, 1348, 1454, 156, 550, 1680, 260]. estimation [1504, 927, 773, 1923, 307, 1039, 1072, 651, 729, 1176, 477, 945, 615, 1600, 1228, 1796, 628, 1893, 956, 1851, 1210, 1193, 815, 1007, 131, 395, 1568, 862, 1735, 179, 366, 873, 1800, 1288, 1585, 330, 1916, 4, 1283, 1791, 1498, 1058, 1106, 539, 992, 1087, 1623, 604, 1548, 1370, 1596, 1782, 1630, 1637, 1970, 712, 1701, 1940, 2059, 1204, 1718, 1263, 1937, 858, 816, 1023, 1100, 1218, 1489, 155, 1995, 1766, 1845, 1312, 2000, 1241, 419, 756, 1163, 928, 1012, 1013, 2043, 1566, 25, 1555, 918, 972, 971, 1051, 648, 746, 1745, 1309, 1943, 1512, 785, 1644, 1750, 1734]. estimation [857, 1008, 1086, 1938, 191, 527, 1884, 1872, 8, 376, 1169, 1450, 868, 317, 570, 115, 128, 1604, 1461, 958, 1485, 2002, 1679, 1839, 1689, 160, 1272, 1553, 804, 1253, 1273, 596, 472, 1357, 1432, 1755, 64, 1712, 763, 929, 1196, 1114, 1457, 1736, 1874, 1919, 1650, 1638, 1753, 2031, 1159, 515, 1465, 1171, 1142, 1317, 1952, 1932, 1433, 1519, 1619, 1868, 1956, 1271, 1631, 2008, 1188, 1980, 465, 724, 411, 1001, 504, 892, 902, 204]. estimation-adjusted [1952]. estimator [768, 277, 444, 932, 502, 1187, 1886, 1276, 686, 1003, 794, 1319, 820, 1627, 660, 1664, 1410, 120, 986, 1843, 505, 568, 666, 503, 691, 376, 1964, 706, 1823, 525, 679]. Estimators [302, 300, 1837, 497, 976, 939, 1139, 1925, 930, 479, 978, 511, 557, 263, 1441, 625, 516, 969, 549, 717, 1128, 1397, 1143, 253, 1104, 800, 372, 762, 1409, 595, 399, 1603, 1652, 867, 283, 1663, 175, 1850, 574, 999, 1201, 415, 1824, 642, 594, 988, 1832, 974, 512, 591, 693, 70, 159, 373, 544, 1132, 1812, 1984, 606, 510, 199, 394, 2072, 84, 1502, 934, 915, 993, 389, 406, 455, 575, 180, 17, 931, 802, 1113]. estimés [544]. etiological [162]. étude [482, 1325]. Euler [2002]. Evaluating [1096, 1826, 1482]. Evaluation [1935, 3, 1155, 1621]. evaluations [1549]. event [1808, 1277, 955, 1535, 1337, 2060, 1544, 1994, 1582, 1607, 1797]. events [747, 481]. evidence [1546, 1811]. Exact [1035, 1161, 1099, 22, 1085, 1140, 1429, 443, 457, 792, 648, 397, 671]. Exactly [1186]. example [498, 72, 604, 965, 326, 195]. examples [1473, 1436, 1015]. exchangeable [616, 1658]. exclusive [1960]. Excursion [210]. exemple [633]. existence [39, 737]. existing [2058]. Expanded [237]. expansion [733, 1037, 690, 45, 68, 821, 1612]. expansions [1005, 112, 1704]. expectation [1091, 1999]. expectation-maximization [1999]. expected [9, 600, 1195]. expectile [1979]. expectiles [2072]. experiment [1265]. experimental [305, 501, 772, 206, 2052, 2084, 55, 1629]. Experiments [2114, 2160, 24, 321, 822, 1591, 2029, 2157, 246, 919, 1894, 1148, 1200, 1034, 1132, 579, 1381, 284]. experts [1498, 1895]. explainable [1030]. explained [1488]. explanatory [968]. explicit [253, 503]. exploration [1880]. exploratory [248]. explosion [937]. Exponential [509, 294, 840, 176, 725, 27, 651, 327, 534, 1221, 395, 574, 765, 805, 7, 378, 792, 320, 142, 972, 158, 297, 44, 652, 75, 829, 1126, 160, 533, 1956, 390]. exponentiality [407]. exponents [907]. exposed [1988]. Expositions [240, 203]. exposure [784, 1827, 610, 1939]. exposure-disease [1827, 1939]. exposures [2064]. expressions [263]. Extended [2004, 975, 1094, 1098]. Extending [1612]. Extension [1378, 259, 551]. external [1847, 1985, 1236]. extrêmes [926]. Extrapolation [1139, 1208, 146, 1019]. extrema [1196]. Extreme [1270, 467, 1980, 1143, 1458, 926, 1664, 1410, 749, 1551, 469, 2072, 468]. Extreme-value [467, 1458, 1410, 749, 1551, 468]. extremes [927, 1510, 1160].

F [267, 165, 56, 218]. f17.50 [201]. f3.50 [166]. f4.95 [202]. Factor [215, 1055, 1859, 1572, 99, 1759, 1149, 1629]. factorial [1216, 1242, 1616, 1200, 2062, 177]. factorials [292]. factoring [633]. factorisation [633]. Factors [697, 698, 903, 905]. faible [98]. faibles [1028]. Failure [2128, 60, 238, 1613, 370, 916, 900, 1623, 1996, 889, 2060, 1542, 1252]. failure-time [900, 889]. false [2015]. familial [619, 394]. Families [537, 813, 941, 659, 392, 405, 1221, 395, 574, 765, 1431, 142, 652, 389]. family [1379, 1345, 1705, 547, 138, 415, 972, 638, 1819, 802]. Fast [1633, 1579, 1182]. favourable [950]. feature [1336, 1498, 1926]. features [1885]. fécondité [15]. Federal [255]. Fels [752]. Ferguson [933]. Ferguson-Dirichlet [933]. fertility [15]. fiducial [163, 196]. field [118, 2046, 2002]. fields [552, 902, 1131, 1310]. Fieller [934]. Fienberg [221, 103]. fill [1451]. filling [2053]. filter [1077]. filters [274]. filtrations [560]. Finance [1269, 1857, 1160]. financial [1558, 2079, 483]. financiers [483]. Finding [499, 679]. Findlay [205]. fingerprinting [1905]. Finite [630, 652, 1837, 411, 2027, 1818, 733, 775, 833, 893, 957, 1109, 1532, 1892, 1632, 1673, 175, 1798, 1534, 275, 1262, 701, 850, 586, 1245, 1114, 1743, 778]. Finite-population [630, 701, 778]. finite-sample-size [275]. fire [1056]. firefighters [315]. First [1455, 559, 91, 976, 612, 146, 171, 86, 30, 387]. First-order [1455, 146, 387]. First-passage [559]. first-passage-time [976]. fish [674, 675, 673, 1682]. Fisher [105, 921, 598, 708, 1093, 1024, 1594, 671, 629, 429]. fisheries [965]. fishery [1816]. fit [1793, 891, 1625, 355, 1239, 1202, 997, 1379, 849, 1054, 1846, 1662, 1301, 145, 827, 1579, 324, 998, 378, 1838, 447, 1431, 1861, 689, 1358, 917, 1115, 1617, 154, 881, 1126, 1174, 1096, 1406, 428]. Fitting [875, 1541, 658, 1454, 1880, 1934]. five [468]. fixed [1162, 1632, 358, 360]. Flaherty [201]. flat [1003]. Fleming [1580]. Flexible [1695, 1118, 1934, 1730, 1591]. flexibly [1939]. flows [483]. fluctuation [382]. flurprimidol [809]. flux [483]. follow [1375, 871]. follow-up [1375, 871]. following [894]. followup [1554]. fonctionnel [855]. fonctions [50]. fondé [723]. food [1496]. footed [1392]. force [485]. Forecasting [2143, 1399, 274, 2004, 1321, 431, 459]. forest [1056]. forests [1995]. foreword [221]. form [939, 155, 291, 688, 1494]. formal [1634]. formation [1885]. forme [155, 291]. formed [809]. former [1092]. Forms [2116, 513, 448, 185, 1015, 101, 234]. formula [253, 503]. formulation [1180]. Forthcoming [1259, 1279, 1313, 1327, 1363]. forward [1991]. Foundations [2087, 249, 80, 134, 121, 134]. Fourier [677, 2017]. fractal [2002]. fraction [1565]. fractional [1055, 1242, 1616, 292, 2062]. fractionally [1455]. frailties [1556]. frailty [2049, 1731, 1996]. Frais [1280, 1328, 1290, 1365]. framelet [1972]. framework [1079, 1390]. Francisco [237]. Frane [148]. Fraser [288, 123, 2023]. free [355, 1867, 1382, 967, 48, 151, 1491]. Freedman [124, 520]. French [768, 714, 1022, 830, 1026, 932, 939, 1001, 855, 1300, 264, 1250, 557, 605, 872, 582, 50, 59, 54, 1325, 1075, 723, 1078, 892, 595, 405, 477, 926, 1261, 1103, 646, 98, 977, 566, 251, 737, 633, 1028, 1042, 348, 15, 155, 291, 450, 488, 544, 483, 293, 387, 276]. frequencies [1474, 136]. frequency [356, 1056, 1496]. frequentist [2027, 707, 1093]. frequently [159]. Friedman [702]. fulfillment [241]. full [983, 1618]. Fully [1561, 1286]. Function [2105, 2106, 2161, 1808, 28, 1187, 78, 1276, 188, 613, 1975, 1164, 1397, 922, 127, 588, 524, 1410, 815, 827, 537, 1064, 187, 192, 38, 314, 1087, 1623, 795, 1548, 1970, 1717, 1262, 816, 1976, 713, 2123, 1848, 262, 350, 731, 839, 426, 213, 1432, 1516, 1786, 1114, 1958, 679, 1953]. function-on-scalar [1975, 1976]. Functional [1871, 1944, 1648, 855, 1972, 1162, 552, 1909, 1581, 1670, 1974, 1982, 1460, 1978, 1400, 1873, 1383, 1971, 1979, 2066, 1937, 1189, 1483, 1845, 1984, 1983, 1689, 1096, 1905, 2031, 1826, 1977, 1980, 1968]. Functional-coefficient [1944]. functionals [1867, 933, 536, 1691].

Functions [2126, 217, 2161, 2155, 788, 526, 95, 23, 620, 351, 937, 1525, 50, 1359, 2020, 1409, 56, 692, 707, 859, 1320, 366, 1118, 1951, 766, 1828, 322, 1663, 955, 858, 2025, 1667, 1526, 1555, 1309, 1540, 1047, 158, 1626, 12, 43, 1006, 1517, 600, 2050, 592, 180, 1196, 1405, 1482, 1191, 1631, 522]. fundamental [1722]. Fundy [1816]. Further [331, 230]. fused [2043]. fused-lasso [2043]. Future [1965, 2033].

G [166, 205, 222, 167, 287, 308]. game [36, 91]. games [452, 597]. gamma [475, 870, 1796, 1288, 2025, 829, 804, 1432]. GARCH [1576, 1944, 1956]. gas [5]. Gauss [517, 51]. Gauss-Markov [517, 51]. Gaussian [921, 601, 665, 1559, 1705, 1672, 723, 546, 552, 1796, 1978, 1764, 559, 377, 1782, 656, 689, 1897, 1739, 1750, 280, 2002, 593, 303, 276, 1641, 1999]. Gaussian-based [593]. gaussiennes [276]. gaussiens [723, 110]. GEE [1499]. Geisser [259]. gene [1842]. gene-environment [1842]. General [1256, 1378, 1416, 932, 906, 661, 21, 537, 1082, 1608, 45, 68, 35, 1491, 144, 1405, 679, 1180]. générales [932]. generalised [1169]. généralisées [977, 633]. generalizations [323]. Generalized [953, 775, 818, 1369, 46, 919, 1610, 996, 1245, 1675, 665, 442, 1685, 1911, 1375, 1255, 40, 598, 1567, 1071, 1018, 1063, 432, 2048, 480, 1228, 875, 1145, 175, 1283, 1771, 1934, 977, 576, 633, 1596, 1922, 1347, 2067, 1602, 1580, 177, 11, 12, 1760, 22, 850, 1307, 1404, 1439, 1835, 2050, 834, 1206, 1215, 1479, 1481, 1478, 938, 1679, 1839, 2, 1146, 920, 2068, 1669, 1191, 1199]. generalized-moments [1018, 1063]. generated [454]. generating [858, 713]. generation [1389, 981]. genes [948]. genetic [1150, 1775, 1151, 569, 1155, 1380, 420, 2034, 1152, 1153, 1154]. genetics [1820]. genome [1155, 1910]. genome-wide [1910]. genomic [1346]. genotype [2034]. genotype- [2034]. Gentleman [165]. Geometric [1074, 985, 1233, 1200, 1209]. Géométrie [2104, 152]. geometry [440]. Geophysics [2136, 285]. George [147]. geostatistical [1834, 16]. Geostochastic [125]. Geramita [185]. Gérard [186]. Ghoudi [1458]. Gibbs [960, 649]. Gini [1728, 1714]. Giri [149, 222, 284]. GIS [1323]. given [405, 38, 1220]. glasso [1745]. glasso-based [1745]. glimpse [1175]. Global [1843]. GMANOVA [424]. GMLE [959]. goal [2046]. Godambe [384]. Gómez [1235, 1235]. Goodman [221]. Goodness [1793, 1239, 1202, 353, 1301, 145, 1579, 1838, 1861, 689, 1358, 917, 1115, 1174, 891, 1625, 355, 1379, 849, 1054, 1846, 1662, 827, 324, 1431, 1617, 428]. Goodness-of-fit [1793, 1239, 1202, 1301, 145, 1579, 1838, 1861, 1358, 917, 1115, 1174, 891, 355, 1379, 1846, 1662, 827, 324, 1431, 1617]. Goodness-of-link [353]. Gopal [239]. Gordon [461, 308]. Government [2115, 201]. Grade [308]. Graham [309]. graph [1924, 1557]. graphic [1276]. Graphical [696, 1647, 1634, 16, 723, 1710, 1804, 1782, 1992, 614, 1855, 2071, 1641, 1868, 1999]. graphiques [723]. Graphon [2043]. graphs [1345, 1744, 1709]. Grassmann [266]. Graybill [376]. Graybill-Deal [376]. greater [1424]. Green [615, 537]. Greenwood [461]. Grenander [1003].

Group [2010, 421, 1572, 356, 65, 2031, 1932]. grouped [1906, 423, 955, 1126, 1378]. groupes [488]. groups [488]. Growth [2101, 697, 698, 809, 810, 1535, 726, 1494, 675, 159, 752, 673, 811, 69, 307]. growth-decay [726]. GRP [16]. Grubbs [602]. Guadalupe [1235]. guard [153]. guess [1424]. Guest [1814, 2074, 1857, 1773, 1968, 1899]. Guide [167]. guided [1959, 1475]. Gupta [147, 289]. Guttman [306].

H [170, 219, 220, 265, 288, 221, 147, 201, 187, 135, 200, 151, 1245]. H-function [187]. Håkan [121]. Haar [772]. Hacking [196]. Hadamard [185, 2116]. Hájek [1103]. Halifax [2164, 289]. Halsted [166, 187, 204, 284]. Hamilton [222]. Hammersley [120]. hand [19]. hardes [1042]. Hardy [680]. Hardy-Weinberg [680]. harp [1041, 1040]. Harper [134]. Harrington [1580]. Hartley [1505]. having [1131]. Hazard [1961, 1997, 1996, 1263, 889, 1056, 1994, 1465]. hazards [1304, 1920, 2049, 1652, 1767, 1016, 795, 1362, 1882, 1717, 1940, 1394, 1487, 1461, 1697]. Heading [1416]. Health [1949, 487, 609, 1123, 610, 780, 491]. Heath [205]. heavy [1440, 1980, 1550]. heavy-tailed [1980]. Heidelberg [520]. height [752]. held [2164, 462, 134]. Hellinger [1457, 1650]. helpful [1827]. Hemisphere [522]. herds [1041, 1042, 1040]. heterogeneity [2065, 861, 1170, 2046, 1020, 1030]. heterogeneous [1903, 1900, 2044, 1902, 1901]. Heteroscedastic [1660, 2051, 1137, 1264, 1019, 1787, 1585, 1262, 786, 736, 1491, 1786]. heteroscedasticity [1138, 998, 1295, 1768, 1443, 1693]. heteroscedasticityY [999]. Heuristic [456]. Hidden [920, 1417, 2045, 1177, 1983, 1661]. Hierarchical [658, 1642, 1735, 1817, 616, 1004, 1123, 1316, 1082, 1347, 1496, 1825, 1571]. High [1874, 1558, 1775, 1920, 567, 1942, 1655, 1907, 1943, 1720, 2010, 2068, 1991, 1917, 1868, 2039, 2008]. High-dimensional [1874, 1775, 1920, 1942, 1907, 1943, 1720, 2010, 2068, 1991, 1868, 2039, 2008]. high-order [567]. high-speed [1655]. Higher [915, 842, 788, 1924, 832]. Higher-order [915, 842, 788, 1924, 832]. highly [1148]. Highs [1160]. Hilbert [117, 1383, 45, 68, 1727]. Hilger [188]. histogram [769]. historical [853, 650, 745, 1189, 450, 2031]. histories [494]. historique [450]. history [699, 2032, 1056, 638, 1994]. HIV [1537]. HLA [1154]. Hodges [558]. Hodges-Lehmann [558]. Holden [237]. Holden-Day [237]. Holland [266, 287, 289, 466, 460]. homicide [499]. homicides [379, 380, 382]. Homogeneity [1892, 1109, 1395, 1713, 640, 2018, 996, 1522, 1257, 1286]. homology [948]. honorary [312, 171, 391]. honour [2074, 1531, 1587]. Hooker [134]. hospital [781]. hot [1645]. hot-deck [1645]. hourly [493]. household [486]. Hubert [290]. hulls [536]. human [780]. Hunt [290]. Hunter [306]. Huntington [147]. hybridizing [2052]. hybrids [760]. hyperbolic [1685, 1115]. hypercubic [329]. hypergeometric [56, 12]. hypersphere [146]. hypotheses [1396, 1305, 1520, 684, 2081, 158]. Hypothesis [1798, 1361, 46, 1087, 82, 991, 1, 1955, 1960, 1941, 1768]. hypothesized [145].

I-optimal [2051, 1922]. I.Freedman [265]. i.i.d. [1139]. Ian [218]. ice [944]. Iceberg [358, 359, 360]. identifiability [712, 1614, 1232]. Identifiable [1816]. identification [1344, 688, 1640]. identify [1043]. Identifying [1056]. identiquement [1026]. identities [230]. identity [1810]. ignorable [1631]. II [134, 462, 68, 81]. III [134, 167, 2125, 1149, 200]. illness [783]. Illustration [1011]. image [879, 771]. images [1910, 401, 644]. imaging [1909, 1908, 1820, 1121]. imbalance [2052]. immediate [53]. immigration [1485]. impact [1175, 1391]. Implications [1061, 831, 910, 1195]. importance [981]. improper [1897]. improve [1484, 1323]. Improved [511, 516, 2078, 1681, 539, 1132, 1872, 934, 1085, 1140, 1454, 388, 1036]. Improvement [1231, 517, 919, 1643]. Improving [1548, 829]. impulse [1320]. Imputation [1962, 1302, 1586, 868, 1451, 1830, 1694, 480, 1260, 1673, 1535, 1645, 1337, 1134, 1688]. imputed [1561, 1568, 1645]. in/for [1965]. inactivity [35]. incidence [1438, 1717, 1324, 2050, 2005]. incident [1754]. incidental [245]. Including [954, 878]. inclusion [648, 746]. inclusion-consistent [648, 746]. income [764, 1511]. incomplete [34, 2007, 1069, 956, 992, 1540, 789, 1674, 1481, 65, 1453, 1513, 1688]. incompletely [858, 967]. incorporate [2004, 1847, 1985]. incorporating [1992, 2067, 1855, 1702]. incorporation [418]. incorrect [133]. increasing [1222]. increments [388, 400]. indépendance [293]. indépendantes [276]. independence [2019, 608, 52, 872, 1578, 916, 1172, 532, 140, 1842, 751, 89, 621, 509, 293]. independent [1428, 1594, 276, 1550]. index [1680, 1935, 1664, 1492, 1298, 151, 1945]. indexed [214, 1982]. India [202]. indicators [1489, 1487]. indices [1714]. indirect [1811, 2081, 25]. individual [1739]. individual-level [1739]. individualized [705]. individuals [222]. Inducement [2076]. industrial [5]. industry [283]. Inequalities [475, 569, 2119, 1195, 96, 252, 286]. inequality [1567, 796, 430, 528]. infection [1590]. infectious [1739]. infer [1336]. Inference [649, 2087, 1860, 2091, 2164, 1260, 1421, 134, 1463, 1891, 1405, 1075, 450, 1366, 2027, 1802, 1642, 249, 1785, 2026, 1672, 1226, 813, 941, 2028, 1097, 392, 2078, 432, 730, 870, 1436, 819, 345, 1430, 1460, 818, 1380, 1514, 1586, 1771, 1212, 900, 422, 1506, 1809, 2033, 450, 1000, 2062, 1225, 1595, 2023, 1401, 1938, 1017, 736, 196, 586, 514, 1307, 1068, 2035, 925, 1206, 1720, 1776, 121, 1709, 1295, 1308, 592, 1542, 1582, 1700, 1377, 1598, 1697, 1945, 1590, 1415, 1573, 1089, 1075, 134, 149]. Inferences [1479, 840, 421, 1696, 1299, 1481, 1478, 1675, 1511]. Inferential [570]. inférieures [276]. inferred [494]. infinite [138, 1320, 554, 109, 1015, 1743]. infinite-dimensional [1015]. infinite-variance [1320, 554]. infinitely [197]. Infinitesimal [716]. inflammatory [1150, 1153, 1151, 1154]. inflated [1480, 1170, 2046, 1927]. inflating [1781]. Inflation [1428, 1071]. influence [1563, 1397, 787, 1325, 585, 859, 440, 1403, 763, 1199]. influencing [697, 698]. Information [1707, 1716, 1724, 1733, 1708, 1741, 1749, 1756, 1763, 1772, 1784, 1794, 1803, 1844, 1813, 1822, 1833, 1856, 1864, 1875, 1888, 1898, 1913, 1929, 1946, 1966, 1987, 2003, 2021, 2038, 2057, 2073, 1584, 754, 921, 2019, 174, 208, 1261, 1847, 1985, 2058, 437, 1091, 1024, 1020, 1791, 422, 1076, 1594, 1811, 2067, 701, 2081, 66, 1615, 2013, 621, 1056, 80, 199, 728, 629, 1114, 1213, 980, 429]. informative [1771, 1701, 1674, 1011]. informed [1911]. Ingram [286]. inhibitor [810]. inhomogeneous [1848, 1690]. initial [248]. innovations [1578]. input [1148]. Insect [1322, 1325, 1326]. Insights [1936]. inspection [959]. Institute [168, 1814, 2037]. Instructor [124, 2099]. Instrumental [1851, 2064, 1872, 2008]. instruments [2047]. Insurance [1269]. integer [120, 1963, 1163]. integer-valued [1963, 1163]. integral [208, 446]. integrals [17]. Integrated [879, 1806, 1663, 1455, 594, 666, 1392]. Integrating [2058]. integration [2054, 1606]. intensity [1265, 1420, 1848, 673, 1465, 724]. inter [741, 14]. inter- [741]. inter-birth [14]. interaction [59, 708, 1285, 1116, 1027, 37, 673, 213, 1628]. interactions [1896, 1149, 1826, 1702]. interactive [2011]. Interblock [728]. intercept [1577]. intercept-scale [1577]. interclass [394]. Interdisciplinary [1665]. interest [655]. interesting [874]. interface [165]. interference [2063]. Interferon [662, 663]. Interim [1952, 1583, 982]. intermittent [1554]. intermittently [1994]. internationai [201]. International [2115, 2102, 134, 1269, 151]. interplay [206]. interpolation [784]. interpretation [317]. interpreting [2029]. interrater [964]. Interval [494, 1234, 1309, 119, 227, 1653, 1251, 1613, 869, 14, 955, 1235, 1535, 1882, 1580, 648, 746, 1512, 1750, 899, 64, 959, 1252]. interval-censored [1613, 1235, 1882, 1580, 899, 1252]. interval-censoring [1653]. interval-censorship [959]. interval-grouped [955]. intervals [1426, 1653, 1085, 1140, 761, 1830, 243, 403, 1185, 1501, 261, 585, 223, 1382, 894, 1099, 1789, 973, 26, 860, 1493, 1127, 1897, 1067, 336, 678, 1102, 1253, 90, 1424, 1315]. intra [741]. intra-laboratory [741]. intraclass [547, 974, 394]. Introduction [2022, 2112, 1857, 779, 1899, 2124, 2149, 1323, 2074, 218, 287, 1968, 466, 1814, 1773, 1385]. Introductory [2140, 2097, 170, 306]. intronic [948]. Invalidity [128]. Invariance [454, 2150, 987, 808, 461]. Invariant [366, 1653, 1230, 52, 507, 1663, 1207, 804, 304]. Inverse [77, 665, 1672, 546, 377, 689, 280, 303]. inverse-Gaussian [1672, 546, 377, 303]. inversely [815]. Inverted [29]. investigation [146]. involving [622, 473]. Iowa [290]. irregular [1648, 1973]. Irwin [306]. ISBN [149, 166, 148, 170, 219, 220, 237, 238, 239, 240, 266, 267, 265, 288, 309, 519, 122, 188, 205, 169, 104, 221, 222, 186, 103, 147, 167, 168, 189, 124, 202, 201, 307, 217, 464, 285, 187, 204, 218, 287, 310, 311, 290, 150, 123, 289, 368, 185, 135, 459, 152, 522, 200, 284, 286, 105, 203, 461, 520, 466, 151, 460, 121, 369, 134, 463, 308, 184, 306]. Island [168, 310, 311]. Iso [1135]. Iso-chi-squared [1135]. Isotonic [724, 512, 1932]. ISSN [240, 310, 311]. Issue [1707, 1716, 1724, 1733, 1708, 1741, 1749, 1756, 1763, 1772, 1784, 1794, 1803, 1844, 1813, 1822, 1833, 1856, 1864, 1875, 1888, 1898, 1913, 1929, 1946, 1966, 1987, 2003, 2021, 2038, 2057, 2073, 1899, 1814, 2022, 2074, 1531, 1857, 1773, 1968, 1478, 1587]. issues [490, 1226, 2032, 491]. iterated [563]. iteration [1676]. IV [1149]. ix [188, 189, 522, 284].

J [721, 148, 170, 219, 220, 238, 267, 288, 1202, 1140, 778, 188, 941, 307, 1063, 1210, 897, 745, 1235, 310, 311, 290, 368, 344, 200, 1245, 369, 306, 1013, 1247]. Jack [309, 1531]. Jackknife [736, 1714, 244, 587, 1560, 691, 1954]. Jackknifing [235]. Jacobians [143]. James [263, 253, 503, 1689]. James-Stein [263, 253, 503]. Jan [121]. Jane [165]. Jennifer [185]. Jennrich [148]. Jerome [522]. Jerry [1531]. Jersey [217]. Joffe [222, 150]. Johannes [459]. John [170, 220, 238, 267, 288, 309, 519, 122, 167, 218, 459, 121, 306]. Joint [918, 1437, 393, 1711, 1535, 1926, 1566, 1565, 1853]. Joseph [310]. Journal [1340, 1364, 1947, 898, 1314, 1329]. joy [1699]. Jr [311]. judgment [1615, 2012]. Judith [151]. Jump [2020]. Jump-robust [2020]. jumps [2000]. June [1723]. jurys [1258, 1434, 1477, 1575, 1609, 1656, 1683, 2056].

K1A [200]. Kalbfleisch [238, 1531]. Kale [204]. Kalman [1077]. Kanasewich [285]. kappa [964]. Karol [307]. Kathleen [169]. Keele [104]. Keith [522]. Ken [290]. Ken-dall [290]. Ken-dall/Hunt [290]. Kendall [1642, 1073, 1172, 1601, 1658, 1825]. Kennedy [310, 311]. Kenneth [167]. Kernel [1266, 1718, 768, 788, 530, 932, 930, 1187, 1886, 1441, 1867, 1331, 412, 646, 645, 1283, 1383, 1843, 505, 1979, 735, 594, 647, 1547, 993]. kernel-based [1979, 1547]. kernels [593]. Keyfitz [463]. Khatri [287, 377]. khi [830]. khi-deux [830]. Khoudraji [1458]. Kiado [521]. Kibble [1288]. kidney [524]. Kingston [465]. known [893, 1413]. Kolmogorov [302, 300, 438, 1275, 629, 929, 931]. Kolmogorov-Lévy-type [438]. Kolmogorov-Smirnov [300, 1275]. Kolmogorov-Smirnov-type [302]. Koziol [615, 537]. Koziol-Green [615]. Kr6tki [307]. Krewski [369]. Krieger [147]. Krishnan [104]. Kronecker [541]. Kruskal [221, 368, 151]. Kullback [1722]. Kulperger [344]. Kurtosis [573].

L [148, 238, 222, 134]. laboratory [741]. labour [485]. lack [997, 1400, 998, 154, 1406]. lack-of-fit [997]. lactation [411]. Lagrange [1576]. lags [634]. lake [612, 494, 611]. land [477]. Landsat [332]. Langevin [589]. Laplace [1597, 618, 87, 1934, 890]. Laplacian [818]. Large [1447, 1551, 64, 414, 529, 1599, 1603, 1673, 1579, 1782, 563, 2061, 683, 1156]. Large-sample [1551, 1599, 1579, 2061]. largest [231]. Larry [311]. Laspeyresia [1325]. lasso [1572, 1706, 2043, 1812, 1990]. lassos [1647]. Late [306]. latent [183, 1904, 18, 1999]. latin [53, 966]. Latin-square [966]. Lattice [2121, 2122, 116, 239, 203]. lattices [329]. Law [2093, 414, 529, 211, 388, 683, 166]. Lawless [267, 1531]. Laws [563, 197, 1026, 605, 405, 456, 38]. layout [1027, 882, 246]. layouts [1285]. learning [1965, 2046, 1589, 1779]. Least [330, 108, 1465, 497, 1164, 794, 1777, 820, 432, 439, 1127, 25, 666, 606, 691, 1462]. least-distances [820]. Least-squares [330, 794, 1777, 439, 666, 606, 691]. lecteurs [1715, 1748, 1783, 1821, 1863, 1912, 1986]. Lecture [185]. Ledolter [459]. left [1742, 1493]. left-truncated [1742, 1493]. Lehmann [558]. Leibler [1722]. lemma [1245]. length [1438, 935, 1540, 1755, 525]. length-biased [1438, 1755, 525]. lengths [1127, 336]. lens [2070]. Leo [221, 311]. leptokurtic [1747]. leptokurtic-normal [1747]. Let [2029]. Letac [186]. Letter [1202, 1245]. Lettre [1202, 1245]. level [1770, 1608, 1242, 1373, 292, 1739, 1734, 1646]. levels [1418]. Lévy [1793, 438]. Lévy-driven [1793]. Lexington [205]. Life [2032, 2132, 526, 1277, 322, 7, 1506, 8, 204]. Lifetime [2141, 267, 1235, 982, 1755, 1415]. Lightning [498]. Likelihood [1437, 1781, 190, 1514, 832, 1595, 1401, 1068, 1582, 921, 714, 277, 396, 1366, 444, 767, 1251, 250, 686, 1785, 1830, 1296, 183, 985, 1806, 775, 957, 1109, 1185, 1274, 1694, 1841, 2026, 1192, 549, 922, 1097, 260, 1143, 408, 1962, 1367, 660, 133, 331, 1436, 2048, 729, 1221, 615, 1228, 838, 895, 1603, 956, 1180, 1007, 1098, 1659, 179, 867, 1166, 1649, 1380, 1492, 1369, 1762, 1212, 1235, 1362, 1530, 2059, 82, 1218, 1602, 1493, 1766, 236, 1312, 1371, 719, 1667, 1234, 1620, 1666, 1566, 1555, 1560, 1414, 1617, 797, 1931, 1954, 193, 1765, 317, 570, 958, 925, 1206, 1485, 1245, 1503, 1675]. likelihood [1776, 1195, 1102, 1295, 1612, 1253, 924, 472, 1134, 1308, 1621, 1677, 1714, 425, 1712, 533, 1213, 1315, 1698, 1511, 1878, 1249, 1513, 1297, 1549, 1011, 1191, 1956, 1271, 1415, 1089, 1618, 1113]. Likelihood-based [1514, 1582, 1380, 1765, 1195, 1511, 1249]. likelihood-ratio [396, 775, 957, 1245, 924, 533]. likelihoods [1918, 2082, 1494, 1195]. Limit [263, 661, 554, 294, 113, 844, 1446, 552, 430, 231, 525, 390]. limitations [883]. Limited [284, 1391, 1026]. limités [1026]. Limiting [438, 666]. Limits [2131, 1448, 545, 189, 1840, 1413]. Lindsay [765]. Line [902, 1228, 1914, 1474, 2129, 774, 738, 310]. linéaires [714, 483]. Linear [188, 1162, 1778, 1829, 2106, 687, 857, 483, 714, 847, 497, 1284, 978, 475, 414, 442, 1219, 1375, 1255, 216, 516, 904, 1156, 1613, 1128, 34, 274, 1567, 1071, 1454, 1680, 27, 1632, 796, 1950, 174, 762, 657, 474, 1990, 1072, 2048, 1471, 1301, 1320, 875, 1460, 1488, 1145, 1719, 46, 517, 1577, 38, 1771, 1400, 919, 947, 1596, 1922, 1639, 1970, 1979, 1262, 1298, 1347, 1373, 2067, 2066, 1520, 1937, 1189, 2081, 1845, 2123, 890, 512, 142, 397, 1060, 367, 1009, 1132, 1760, 527, 736, 829, 619, 41, 1307, 1404, 1439, 968, 570, 1964, 1254, 834, 1206, 1479, 1478, 1720, 989, 1679]. linear [1839, 1689, 920, 1134, 2068, 704, 763, 826, 1021, 2031, 1509, 1518, 923, 1171, 1236, 1297, 1377, 1598, 1252, 2039, 1693, 1199, 1980]. linéarité [387]. linearity [299, 301, 387]. Linguistic [2108, 219]. Linguistics [219, 257]. link [659, 1018, 1063, 353, 1081]. linking [1142]. living [486]. Lloyd [105]. Local [1251, 1924, 1454, 1362, 1403, 1273, 389, 1199, 1635, 1255, 844, 1885, 1664, 1850, 1031, 1970, 1931, 907, 710, 1317, 768]. locale [768]. Locally [52, 141, 1824, 1792]. location [355, 444, 1336, 302, 1892, 392, 372, 232, 1990, 399, 642, 320, 928, 1012, 1013, 1604, 1691, 1102, 472, 389, 406, 455, 575, 931, 1378]. location-scale [1892, 472]. locations [468]. locus [1155]. Log [1426, 847, 261, 1722, 1263, 1132, 510, 1195, 458]. log-likelihoods [1195]. log-linear [847, 1132]. log-logistic [510]. log-normal [261]. Log-rank [1426, 1263]. log-zero-Poisson [458]. logarithm [563]. logarithmic [40, 142, 64, 90, 130]. Logistic [1397, 418, 705, 770, 1962, 1044, 1018, 1063, 1118, 1526, 510, 1823, 1939, 29]. logit [876]. lognormal [281, 1550]. logspline [1023]. Loi [54, 768]. Lois [1026, 605, 405, 977]. London [201, 151, 134]. long [1886, 1124, 988, 1984]. long-memory [1984]. long-range [988]. long-term [1124]. Longitudinal [1375, 1659, 1483, 1480, 1230, 1648, 953, 1223, 1499, 1437, 1133, 1554, 1488, 1349, 1881, 851, 897, 1374, 1711, 1548, 1630, 1701, 1476, 1973, 1943, 1122, 1033, 350, 1307, 1674, 1565, 1835, 1215, 1479, 1481, 1478, 1729, 1453, 1249, 1482, 1513, 1700, 1297, 1442]. look [499, 485, 1849]. lorsque [544]. Loss [180, 1196, 906, 1663, 1526, 487, 731, 839, 161, 804]. losses [1554]. Lotteries [255]. lotto [452, 597]. low [1909, 1028, 1874, 1511]. low-rank [1909, 1874]. Lower [806, 78, 906, 1608, 281, 804]. lower-bounded [804]. lower-order [1608]. lows [1160]. Ltd [188, 202, 1038]. lumber [283]. lung [262, 350, 315]. lung-function [262, 350]. lupus [1581]. Luz [1235]. Lyapunov [907]. lying [973].

M [166, 148, 122, 205, 222, 202, 187, 287, 1235, 290, 284, 466, 151, 521, 308]. Machine [1965]. machines [1751]. Mack [598]. Mack-Wolfe [598]. Macmillan [202, 135, 151]. macroeconomic [1686]. Macromolecules [2148, 368]. made [1044]. magnetic [1121]. magnetoencephalography [1853]. Magnus [121]. Major [489]. majorant [1164]. majority [1552]. Majorization [2119, 286]. Making [2110, 205]. Malliavin [152]. Mallows [1834]. MAMSE [1790]. Man [19]. Man-dealt [19]. Management [2135, 1323, 266]. Mann [1037, 1061, 1036]. Manoukian [464]. manova [21, 507, 46]. Mansell [201]. Manual [167, 124, 2099, 306, 167]. many [1185, 1501]. map [1336, 1796]. maple [809, 811]. Mapping [710, 1904, 1484, 1903, 1900, 1902, 1508, 1196, 1901, 489]. Marcel [265, 169, 189, 150, 185, 105]. marges [405]. Marginal [1091, 1249, 618, 1597, 1437, 813, 941, 800, 730, 1086, 364, 1503, 1195, 1542, 1582, 1700]. marginalization [328]. Marginalized [1349]. Marginally [1411]. marginals [665]. marine [358, 360]. mark [1386]. mark-recapture [1386]. marked [879, 1581]. markers [1035, 1263]. market [2079]. markets [1618]. Markov [186, 520, 560, 1668, 257, 2028, 1417, 715, 1408, 2145, 2147, 2162, 517, 2103, 2045, 1177, 333, 342, 1209, 910, 1122, 209, 51, 1983, 920, 1542, 1590, 1611, 1618, 520]. Markov-chain [910]. Marshall [286]. Martingale [323]. martingales [1075]. Mary [1588, 1587]. Mass [205, 103, 123, 368]. massive [2061, 2060]. Masthead [1354, 1716, 1724, 1733, 1708]. match [707]. matched [1239, 750, 756, 1507, 908]. matched-pairs [756, 908]. Matching [2006, 1094, 2083, 2052, 1602, 1423]. materials [60]. Mathai [166, 202, 187]. Mathematical [240, 2113, 222, 168, 2134, 2151, 2092, 310, 311, 2153, 2096, 203, 265, 138, 217, 464, 463, 104]. Mathematics [220, 265, 185, 286, 462, 465, 226]. Mathématiques [184, 186, 152]. matière [582]. Matrices [2116, 259, 143, 1908, 295, 1303, 541, 62, 1641, 185]. Matrix [2126, 2144, 502, 1788, 1962, 56, 18, 766, 1916, 112, 539, 1622, 1594, 82, 2025, 591, 693, 9, 139, 460, 531, 1874, 217]. matrix-convexity [139]. matrix-variate [1962, 2025]. matter [582]. maturities [1682]. maxima [840, 672, 181]. maximal [969]. maximin [1264]. maximization [1999]. maximize [1928]. maximizing [950, 704]. Maximum [396, 767, 729, 1228, 179, 921, 277, 444, 937, 686, 549, 260, 660, 1603, 956, 1007, 867, 310, 1218, 1766, 719, 2129, 1566, 1555, 193, 317, 958, 1149, 472, 1712, 1956, 1113]. maximum-bias [937]. Maximum-likelihood [396, 767, 729, 921, 277, 444, 686, 549, 260, 660, 956, 719, 317, 958, 472]. maximum-relevance [867]. May [462, 134, 37, 94, 8]. McLeish [343]. McMaster [222]. MCMC [1635]. MDT [24]. Mean [1502, 449, 511, 1525, 1185, 1238, 1501, 176, 20, 1725, 372, 1059, 595, 870, 651, 131, 578, 807, 120, 992, 604, 415, 1630, 734, 66, 594, 1832, 1566, 512, 918, 44, 8, 376, 1964, 706, 1295, 144, 929, 1424, 1919, 670, 1415]. mean-covariance [1566]. meander [210]. means [616, 395, 1673, 574, 996, 970, 1502, 1689, 161, 533, 1736]. measure [1184, 805, 478, 292, 1862]. measured [1428, 1994]. Measurement [1539, 949, 1920, 1360, 1871, 1829, 1824, 774, 1528, 834, 1689, 1357, 1464, 1753, 1698, 1582, 1700, 1941, 1317, 1089, 954]. measurement-error [949]. Measurements [63, 2138, 418, 1217, 262, 350, 1565, 1168, 308]. Measures [2117, 433, 964, 831, 1244, 813, 941, 100, 975, 434, 1412, 801, 1195, 429, 221]. Measuring [1347, 130, 1546, 898]. mechanism [726]. Median [1159, 1022, 23, 848, 1788, 416, 737, 1506, 1356]. médiane [737]. medians [126, 1549]. medical [473, 871]. Medicine [2152, 519]. Meeker [311]. mélangeant [932]. member [312, 391]. members [222, 138, 171]. membres [1258, 1434, 1477, 1575, 1609, 1656, 1683, 2056]. memory [1886, 1445, 1984]. meta [1647]. meta-elliptical [1647]. method [714, 1112, 767, 1338, 1972, 1331, 274, 1957, 545, 1847, 1985, 1659, 646, 1649, 1828, 1283, 694, 1530, 1527, 1321, 701, 1440, 1560, 53, 877, 2, 1547, 1780, 1960, 303, 1213, 1698, 1513, 679, 1746]. méthode [714, 646]. méthodes [1022, 251]. methodology [1125, 948]. Methods [2143, 122, 2088, 2141, 2106, 2108, 1022, 696, 814, 267, 1239, 914, 321, 188, 776, 375, 718, 2017, 1962, 346, 1267, 615, 1584, 1468, 1145, 1498, 478, 1535, 251, 1622, 453, 614, 1645, 2123, 2158, 1820, 459, 481, 1337, 1657, 1872, 1404, 667, 115, 764, 398, 1463, 1582, 1870, 219]. Metropolis [1416]. Michel [152]. microarray [1486, 1605]. microstate [1906]. mid [753]. mid- [753]. Miettinen [519]. migration [946]. Mikhail [308]. millennial [1495]. Mills [368]. Minima [181]. minimally [1137, 1011]. minimally-supported [1137]. Minimax [2051, 1779, 1163, 1205, 1169, 804, 575, 1021, 52, 906, 141, 1101, 1663, 539, 415, 835, 1616, 693, 455, 929, 679, 969]. Minimax-variance [575, 929]. Minimisation [2106, 2123, 188]. minimizing [629]. Minimum [1925, 453, 1650, 1197, 833, 1789, 1608, 928, 1012, 1013, 70, 704, 1457]. Minimum-distance [453, 833, 928, 1012, 1013]. misclassification [1962, 1827, 1122, 3, 1232, 1721, 1895]. misclassified [2064, 1890, 1891]. Mises [1395, 755, 983, 1080, 140, 1350, 1393, 881, 455]. mismeasured [1851, 1515]. Missing [2069, 814, 1451, 1802, 1830, 1957, 1569, 729, 1301, 1889, 1652, 1260, 873, 1145, 1687, 1111, 1882, 1306, 1766, 1337, 1865, 1404, 1507, 868, 1487, 1604, 2050, 1134, 1214, 1464, 1878, 1700, 1631]. missing-data [729]. Mississauga [315]. misspecification [845, 1173, 1538, 842, 924]. misspecified [1101, 1205, 1941]. misspecifying [1243]. Mixed [1034, 1567, 546, 1632, 1488, 1881, 851, 897, 1771, 451, 1370, 2067, 1937, 756, 1644, 1760, 1439, 1544, 1835, 925, 1206, 1479, 1478, 84, 1679, 1839, 1357, 1624, 1823, 1214, 1453, 1442, 1199, 1378]. mixed-data [1378]. mixed-effect [1488, 1214]. mixed-effects [1632, 1370, 756, 1624, 1442]. Mixed-scale [1034]. mixing [1139, 88, 845, 1058, 1106, 1236]. mixing-distribution [845]. Mixture [417, 750, 1837, 1757, 511, 1685, 775, 833, 893, 957, 1109, 1395, 1532, 1696, 1348, 590, 1632, 1993, 1351, 845, 1380, 1498, 1764, 1798, 1058, 1106, 1534, 1934, 1908, 712, 2018, 991, 1522, 1640, 1853, 325, 958, 1245, 1895, 1650, 1556, 1746]. mixture-of-experts [1498]. Mixtures [349, 665, 1255, 1892, 859, 179, 1610, 1182, 652, 364, 1070]. MLE [1130]. mnet [1720]. mobile [1022]. mobility [2004]. Modal [1534]. mode [1183, 939, 156, 1829]. Model [1237, 859, 1476, 1234, 1991, 1509, 1795, 60, 1747, 1050, 1304, 1219, 516, 643, 1238, 1274, 1613, 1591, 1672, 59, 1678, 34, 1904, 1417, 1758, 903, 546, 1680, 153, 916, 1948, 1094, 2049, 174, 762, 1581, 1351, 133, 870, 1990, 14, 615, 480, 861, 628, 690, 1880, 1889, 1173, 2058, 1652, 1906, 259, 1048, 772, 537, 1770, 1767, 2046, 1492, 37, 1881, 1789, 517, 1798, 5, 1400, 900, 1711, 1420, 1908, 457, 1235, 795, 1495, 583, 1362, 1370, 1882, 1834, 1639, 1843, 356, 1926, 1971, 1996, 1262, 1506, 2067, 2066, 1520, 1177, 1189, 816, 1524, 1523, 1602, 15, 713, 991, 1684, 1522, 1, 1194]. model [1931, 1009, 1132, 1854, 1122, 1033, 93, 1760, 194, 41, 899, 968, 1853, 774, 1964, 732, 2060, 871, 1461, 925, 35, 1195, 1146, 1096, 842, 924, 472, 1406, 2010, 1712, 763, 1231, 1650, 1698, 1988, 1377, 959, 1697, 1703, 2047, 1415, 1611, 1999, 1188, 1826, 1618, 1378]. model-averaging [2047]. Model-based [1476, 1509, 1747, 1048, 1684, 1964, 1703]. model-fitting [1880]. modèle [59, 15]. modèles [723]. Modeling [1459, 1255, 1123, 1407, 786, 781, 1486, 1118, 1399, 1539]. Modelling [1130, 1336, 1269, 359, 1983, 1825, 1435, 1817, 1230, 318, 1359, 1277, 1335, 1859, 1915, 1557, 1535, 1730, 1483, 1660, 1852, 1565, 2035, 1508, 1392, 1932, 946]. Models [219, 104, 375, 2134, 1857, 2092, 2141, 2108, 2051, 1815, 1808, 1837, 561, 847, 1480, 653, 721, 1117, 267, 265, 836, 1202, 497, 1757, 1559, 83, 1997, 953, 1925, 930, 1333, 677, 630, 778, 442, 249, 770, 1053, 1375, 949, 1255, 1924, 421, 216, 775, 849, 833, 904, 893, 957, 1074, 1109, 1138, 1156, 1437, 1532, 1920, 2020, 1696, 1389, 1226, 2032, 1004, 507, 1567, 723, 1769, 1071, 370, 905, 1137, 1264, 545, 1935, 2004, 1632, 820, 1198, 1355, 1497, 1725, 837, 1993, 1104, 2054, 365, 1671, 1950, 1019, 1334, 417, 588, 474, 1133, 1072, 1361, 2048, 729]. models [353, 1981, 1787, 1301, 524, 1032, 1603, 1237, 1851, 845, 1316, 1170, 443, 2058, 1430, 827, 1101, 875, 1572, 1460, 1488, 1380, 1468, 1016, 1145, 1349, 1082, 1369, 1710, 1719, 1804, 851, 897, 1374, 270, 4, 71, 1498, 1764, 1058, 1106, 1534, 1771, 36, 919, 750, 998, 1536, 1596, 1782, 1805, 1922, 1383, 1834, 712, 1731, 1824, 1896, 1963, 1940, 2018, 1298, 1347, 2045, 981, 1937, 792, 333, 995, 342, 780, 1312, 1861, 756, 1779, 688, 1494, 1394, 1566, 890, 1358, 1352, 142, 666, 1062, 1739, 1617, 1927, 1034, 159, 1047, 158, 108, 687, 1644, 553, 1626, 1385, 1190, 1444, 1402, 1765]. models [1614, 51, 1541, 1017, 509, 527, 691, 736, 1205, 1403, 619, 1872, 1955, 1983, 1169, 1307, 1404, 1439, 876, 364, 1077, 1496, 1174, 1254, 1544, 1835, 2050, 958, 1759, 834, 1206, 1479, 1481, 1478, 213, 1245, 1503, 989, 1776, 758, 1679, 1839, 1689, 1272, 1553, 920, 764, 1134, 1357, 1547, 1562, 1624, 1786, 2068, 1823, 1669, 826, 1021, 2071, 856, 1246, 1247, 1214, 1453, 1641, 2055, 1464, 1721, 1682, 1753, 1917, 2031, 1518, 1405, 1542, 1878, 1249, 1941, 1287, 1816, 1958, 1142, 1236, 1297, 1317, 1598, 1556, 1952, 1944, 1945, 1690, 1792, 1030, 1252, 1442, 1868, 1956, 2047, 2039, 1693, 1199, 1583, 1980, 1378, 649]. moderate [32]. Modern [2153, 226, 464]. modification [496]. Modifications [598, 1605]. Modified [1481, 505, 1602, 751, 706, 1999]. modifier [2007]. Mohanty [239]. molecular [1775]. Moment [96, 79, 1664, 44, 1872, 514, 533]. Moments [601, 618, 1597, 1018, 1063, 77, 556, 843, 45, 68, 1654, 1289]. Monograph [104]. Monographs [105, 265, 169, 189]. monotone [1330, 1201, 1131, 1832, 337]. Monotonic [693]. Monotonicity [57, 1516]. Monte [914, 2028, 232, 1005, 1483, 1209, 910, 1077, 1606]. monthly [382]. Montreal [462, 186, 102, 152, 184]. mortality [1399, 72, 1535]. moth [1322, 1325, 1326, 1324, 1323]. Motion [2146, 520, 456, 359, 509]. motor [908]. motor-vehicle [908]. movements [946]. Moving [1705, 113, 844, 1352, 1086, 207]. moving-average [207]. moyen [15]. moyenne [1022, 595]. MR1128406 [778]. MR1166847 [745]. MR1166850 [721]. MR1309322 [1245]. MR1421864 [897]. MR1438419 [1247]. MR1648490 [1013]. MR1944375 [1202]. MR2016224 [1444]. MSE [502]. multi [1769, 1090, 1615, 1760, 1983, 162, 1542]. multi-agent [1090]. multi-centre [1760]. multi-curve [1769]. multi-observer [1615]. multi-scale [1983]. multi-state [1542]. multi-variate [162]. multicategory [1751]. multiclass [1447]. multidimensional [1640, 728]. multilevel [2054, 1403, 1562]. multimodal [1472]. multinomial [957, 796, 1429, 1222, 182, 61, 1232, 1891]. Multiparameter [160]. multipath [1130, 824]. Multiple [574, 1346, 1337, 664, 1006, 1688, 662, 1546, 1459, 1428, 1785, 663, 1418, 141, 1719, 1971, 1373, 1233, 1744, 359, 2062, 129, 938, 1606, 2068, 1213, 1768, 1870, 2009, 2047, 724]. multiple-response [1459]. Multiple-sclerosis [664]. Multiple-shrinkage [574]. multiple-time-series [359]. multiplication [33]. multiplicative [142]. multiplicity [1744]. multiplier [1576]. multirater [1081]. multisample [799]. multiscale [1191]. multistate [1389, 2032, 1390]. multitype [298]. multivalued [962]. Multivariate [2086, 665, 784, 1909, 2091, 2164, 749, 642, 2044, 1601, 2124, 2149, 67, 1819, 13, 173, 1624, 1183, 396, 914, 1747, 1692, 1001, 1338, 1274, 1613, 1879, 1722, 156, 412, 1069, 27, 820, 1198, 1497, 147, 1752, 57, 2017, 880, 2049, 1318, 692, 532, 96, 791, 1906, 143, 578, 1349, 807, 851, 897, 31, 434, 287, 1551, 1212, 1207, 415, 1536, 1805, 734, 1393, 1108, 1356, 1522, 448, 81, 1943, 1190, 1444, 1852, 433, 1077, 466, 426, 1215, 647, 551, 406, 181, 1453, 1518, 29, 1598, 1619, 1271, 149, 289]. multivariées [1001]. multivariés [279]. multiway [427]. murders [381]. Mutability [947]. mutagenic [740, 741, 739, 742]. mutual [2019]. Myers [135].

N [222, 202, 217, 284, 461, 151, 369]. N2L [165]. Nadaraya [127]. Nadaraya-Watson [127]. naive [1049]. Nancy [2074, 2075]. Nanthi [465]. Narayan [149]. Narayana [203]. Nash [188]. Nathan [463]. national [1593]. natural [1221, 765, 652, 1736]. Naus [310]. near [997, 1454]. near-replicate [997]. nearest [1025, 2043]. nearest-neighbour [2043]. nearly [555, 666]. nécessaire [595]. necessary [595, 802]. need [255, 637, 636]. needs [638]. Neff [310]. Negative [451, 40, 961, 977, 320]. négatives [977]. neighbour [2043]. neighbourhood [332]. neighbourhood-based [332]. neighbourhoods [455, 629, 929, 931]. neighbours [1025]. Nested [1742, 2007, 474, 1585, 1964, 1759, 1503, 2031]. nested-error [1585]. Neter [167]. Netherlands [134]. network [1885]. networks [1745, 1868]. neural [1971, 2080]. Neuroimaging [1899, 1903, 1902, 1901]. next [981, 1880]. Next-Door’ [1880]. Ney [150]. Neyman [1725, 1]. nigripes [1392]. Nikodym [932]. Nishisato [240]. NMR [1640]. No [240, 200, 203, 465, 721, 1202, 1140, 778, 941, 1063, 313, 1210, 2058, 897, 745, 1235, 1013, 1245, 1247, 679]. noise [601, 1224, 1490, 1845, 136]. noise-contaminated [1845]. noisy [1110]. Non [474, 1512, 1679, 23, 1559, 1026, 1925, 1331, 39, 21, 1261, 1320, 1978, 31, 566, 1717, 82, 89, 44, 1353, 1759, 1518, 1482, 1590, 1737, 1631, 81]. non-circular [44]. non-compliance [1482]. non-correlation [1353]. non-existence [39]. non-Gaussian [1559, 1978]. non-ignorable [1631]. non-linear [1320]. Non-nested [474, 1759]. non-normality [1518]. non-normalized [1925]. non-null [21, 82, 89, 81]. Non-parametric [1512, 1679, 23, 31, 1737]. non-proportional [1717]. non-response [1331]. non-stationary [566, 1590]. nonadditivity [183]. nonasthmatic [610]. noncentral [183, 577, 441]. noncentrality [703]. noncompactly [1961]. nonconjugate [649]. noncontiguous [895]. noncorrelation [880]. nondifferential [1827]. nonelliptically [1422]. nonexact [746]. nonexact-confidence-interval [746]. nonexistence [1289]. Nonhomogeneous [837, 1002, 724]. nonidentically [1026]. nonignorable [1404, 1878]. nonignorably [1214]. noninferiority [1371]. noninformative [936]. Nonlinear [1383, 1214, 318, 1563, 1255, 1074, 585, 628, 443, 1370, 1971, 440, 1852, 691, 1179, 599, 1357, 1624, 1878, 1443, 954]. nonnormality [997]. Nonparametric [1203, 669, 1396, 1886, 1095, 776, 2042, 1359, 2065, 1348, 233, 223, 471, 1176, 1545, 1210, 1193, 815, 1883, 862, 1623, 1076, 1630, 1052, 2000, 1051, 1414, 783, 1503, 1406, 1432, 1755, 1638, 1469, 1571, 2019, 1625, 757, 1648, 1084, 1692, 1197, 1107, 1296, 1138, 1529, 1633, 1054, 1769, 1330, 1680, 741, 718, 1923, 1267, 1285, 819, 1974, 1261, 92, 828, 423, 1087, 568, 1926, 981, 1413, 1555, 336, 798, 1884, 1517, 1691, 958, 1281, 1286, 1547, 495, 1917, 1768, 1433, 2034, 1573, 1443]. nonparanormal [1710, 1868]. nonprobability [2026]. nonregular [1216]. Nonresponse [1376, 1938]. Nonrobustness [1020]. nonstandard [1104]. nonstationary [555, 1984, 923, 1236]. norm [1059, 734]. Normal [2093, 2131, 166, 1747, 449, 261, 1486, 1722, 1069, 189, 57, 595, 10, 18, 578, 807, 539, 992, 26, 415, 242, 236, 991, 1413, 974, 376, 731, 325, 839, 970, 101, 1070, 2080, 806, 180, 1424, 931, 1518]. Normalité [939]. normality [620, 769, 939, 250, 1156, 1725, 641, 2017, 372, 331, 1982, 512, 336, 1518, 1220]. normalized [1925]. Normally [2137, 311]. normals [511]. Norman [310]. normed [843]. North [266, 287, 123, 289, 466, 460]. North-Holland [289, 466, 460]. Northern [1949]. Norton [124]. Nostrand [308]. note [76, 1202, 1197, 717, 27, 20, 329, 416, 120, 38, 1235, 735, 97, 9, 139, 1517, 1289, 924, 901, 515, 1597]. Notes [42, 1467, 185, 152, 1448]. noyau [768, 932, 646]. NS [2164]. nugget [1570]. nuisance [355, 1667, 754]. nuisance-parameter-free [355]. Null [81, 1037, 21, 82, 1, 89, 1955, 514, 533]. null-distribution [1037]. Number [1740, 1723, 1732, 257, 775, 893, 961, 1057, 1105, 1673, 1233, 487, 1182, 1877, 1245, 1149, 1629]. numbers [414, 529, 580, 1168]. Numerical [2106, 699, 188, 2123, 2158]. numerics [874]. nutritional [1539].

O [151]. object [1968]. Objective [1360, 1361, 1570, 1626]. objects [657, 1452]. observable [1859]. observables [327]. observation [1554, 1027, 1536, 1701, 1537, 198, 231, 540, 1607, 390]. observational [2065, 1589, 1754]. observations [930, 957, 34, 1069, 296, 534, 131, 1887, 333, 594, 1865, 347, 1890, 774, 1311, 1700, 1611]. observed [1124, 858, 1590]. observer [1615]. Obtaining [83]. occasion [2159]. occupation [825]. occupation-time [825]. occurred [747]. Occurrence [519, 2152]. occurrences [257]. occurring [513]. October [2164, 289]. odds [1194, 802]. Odeh [168, 189, 310, 311, 105]. off [1945]. official [1774]. often [1391]. oil [1618]. Oja [737]. Okanagan [1322, 1326, 1324, 1323]. old [499]. Oldham [522]. Olkin [286]. Oller [1235, 1235]. Olli [519]. Omitted [1133]. omitting [1062]. On-line [1914, 774, 738]. One [762, 1637, 362, 607, 1097, 1361, 1027, 228, 435, 38, 882, 1927, 1435, 683, 1624, 496, 5]. one-day [1435]. One-sided [362, 607, 1097, 435, 1624]. One-stage [1637]. One-step [762]. one-way [1361, 882, 496]. Online [1340, 1314, 1329, 1780, 1364, 2044]. Ontario [240, 222, 165, 368, 200, 203, 134, 465]. operating [1989, 1469]. operation [498]. operator [1778]. operator-based [1778]. operators [328, 143]. opportunities [943, 1014]. Optimal [1416, 1504, 1950, 424, 1951, 1887, 1873, 1845, 1194, 1540, 1217, 2061, 1368, 876, 2051, 1216, 1456, 321, 319, 905, 1137, 1264, 772, 2083, 1058, 1106, 1922, 1616, 1411, 1589, 1186, 1163, 915, 1836, 1945, 1792, 990]. Optimality [1299, 305, 985, 206, 966, 562, 877, 1799, 803]. optimization [1942, 2052]. Optimum [507, 99]. option [1858, 1339, 1861]. orchards [1325, 1324]. Order [165, 991, 788, 79, 526, 1265, 1668, 58, 1026, 855, 1924, 1866, 1226, 54, 640, 1473, 146, 569, 432, 567, 1455, 1608, 1177, 974, 1225, 832, 53, 514, 8, 600, 172, 293, 387, 915, 1311, 842, 683, 1917, 1518]. Order-restricted [991, 1225, 514]. ordered [1398, 1135, 1222, 996, 751, 1615, 1555, 1311, 496]. ordering [452, 252, 917, 256]. orderings [1038, 436, 1736]. orders [557, 605, 111]. ordinal [1692, 353, 524, 1851, 1082, 1081, 1865]. Ordinary [606, 497, 1971]. ordination [1916]. ordre [387]. Ordres [557, 605]. orientations [1368]. origin [115]. Ornstein [1793]. orthant [1039]. Orthogonal [2116, 74, 794, 549, 1320, 45, 68, 996, 1831, 1646, 878, 185]. Orthogonalizing [659]. OS9.1 [200]. Oscillating [2080]. Other [2105, 187, 478, 53]. Ottawa [200]. outcome [1375, 1238, 1569, 1504, 1538, 1697]. outcome-dependent [1569, 1538, 1697]. outcome-related [1375]. outcomes [1911, 1229, 1711, 1643, 1876]. outlier [176, 1049, 2015, 854, 801, 8, 1519]. outliers [153, 773, 94, 1674, 325]. output [1165, 1148]. outs [1249]. over-dispersed [1649]. overdetermined [965]. Overdispersed [780, 846]. overdispersion [561, 1170]. overhaul [837]. overview [450]. Owen [168, 189, 105].

P [104, 2089, 147, 202, 289, 461, 462, 521, 148]. P-series [2089, 148]. P. [633]. paclobutrazol [809]. Pair [1559, 95, 1367, 1523]. Pair-copula [1559, 1523]. pair-specific [95]. paired [861, 846, 473]. pairing [375]. pairs [1152, 1268, 947, 750, 756, 908]. pairwise [1710, 1513]. Pál [1175]. palaeomagnetic [942]. panel [2011, 1649, 1536, 2010, 990, 1619, 2009, 1607]. panel-data [990]. Paperbound [465]. Papers [465, 1259, 1279, 1327, 1363, 343, 1313]. papers/Articles [1259, 1327, 1279, 1363, 1313]. paraître [1259, 1363, 1279, 1313]. paradigm [1778]. paradox [384]. Paradoxes [828]. paraítre [1327]. parallel [431]. Parallelism [1599]. Parameter [1039, 355, 444, 761, 299, 625, 904, 1395, 2078, 1267, 1007, 554, 422, 1010, 1257, 336, 158, 544, 804, 901]. parameter-based [904]. parameters [921, 1022, 355, 549, 34, 590, 803, 1316, 1091, 1570, 999, 973, 703, 72, 1350, 245, 1306, 320, 1667, 928, 1012, 1013, 918, 797, 829, 834, 506, 543, 682, 754, 504]. paramètre [544]. paramètres [1022]. Parametric [1520, 1542, 1202, 23, 1925, 659, 545, 861, 31, 1579, 1358, 1512, 1644, 1174, 1679, 1405, 1433, 1737, 2034]. paramétrique [1261]. parametrization [1441]. Pareto [6, 1335, 211, 327]. park [83]. Parsimonious [1804]. part [1480, 844, 503, 2159, 5, 81]. Partial [1866, 56, 1884, 985, 643, 1777, 569, 388, 422, 2066, 1, 341, 1512, 1646, 1980]. partial-sum [341]. Partially [1242, 1219, 1818, 1124, 1680, 1072, 443, 1370, 1298, 333, 1615, 889, 1236, 1297, 1377, 1598, 2039]. participants [1269]. particular [723]. particuliers [723]. partir [488]. Partitioned [541]. partitioning [1914]. passage [976, 559, 91]. past [2033]. Path [2121, 2122, 239, 116, 203, 1751]. paths [358]. patient [157]. patients [1144, 73]. pattern [401, 1629]. Patterned [531]. patterns [19, 793, 499, 379, 1131, 1110, 16]. Paul [152]. Pay [883, 884]. PBIB [445, 1, 109]. peaks [1454, 469]. peaks-over-threshold [469]. Pearson [891, 2079, 15]. Pearson-type [891]. Pederzoli [166]. Penalized [833, 957, 1319, 2048, 1180, 1048, 1942, 1197, 615, 1076, 1564, 1526, 1776, 1462]. penalties [1048, 1969]. penalty [1596, 1619]. percentile [526, 1333, 545, 322, 356, 337]. perfect [1761]. performance [1475]. period [1533, 786]. periodic [351, 1300, 1883]. periodically [2001]. périodiques [1300]. periods [35]. permutation [1426, 1653, 1576, 251, 1463]. permutations [251, 2103, 186]. personal [764]. personal-income [764]. persons [486]. Perspective [2115, 1842, 1200, 1765, 201]. perspectives [1227, 2125, 200]. Perturbation [794, 1955]. Perturbation-based [1955]. PET [1910]. Peter [150]. pH [494]. phase [1083, 1918, 1651, 1113, 2080]. Philosophy [134]. Phoebastria [1392]. phoques [1042]. photo [1269]. photographic [60]. phylogenetic [2035]. Physical [134]. Pickands [1409]. piecewise [955]. pingos [229]. Pisani [124]. Pitman [558, 805, 576, 915]. pivotals [570]. planar [1574]. Planning [2029, 1528, 1629]. Plans [2131, 189, 528, 738, 1452]. Platek [369]. players [2046]. plural [476]. plus [1269]. Pocket [2095, 105]. point [1837, 1817, 1130, 1265, 1344, 976, 137, 760, 1860, 1879, 1448, 1075, 1581, 1981, 1883, 1568, 1719, 824, 1000, 1848, 1601, 1750, 431, 857, 1401, 352, 1727, 1877, 2068, 1945, 1690]. points [321, 113, 727, 1061, 310, 2129, 1056]. Poisson [665, 879, 40, 1448, 546, 837, 861, 1170, 2046, 1099, 451, 633, 1824, 182, 780, 713, 1848, 1062, 881, 925, 161, 551, 458, 1002, 2047, 724]. poles [862]. policy [1366, 50]. politique [50]. polls [1261]. pollutants [784, 610]. pollution [612, 609, 613, 781, 782, 780, 783]. polygon [973]. polykay [33]. polynômes [1103]. polynomial [321, 1334, 1129, 1317, 1571]. polynomials [1103, 971]. polytomous [1935]. pomme [1325]. pomonelle [1325]. ponctuels [1075]. pondérée [830]. Pooling [328]. pools [1635]. Population [2134, 2101, 265, 411, 2076, 2027, 630, 778, 226, 1818, 1806, 733, 1185, 1501, 307, 345, 1673, 578, 1262, 701, 1412, 850, 1114, 1556]. population-based [1412, 1556]. populations [1473, 10, 807, 31, 1222, 1233, 236, 1257, 1391, 586, 376, 130, 254, 1661]. positional [1336]. Positive [1500, 1600, 961, 1908, 503, 1275]. positive-part [503]. positively [579]. positives [822]. positivity [436]. possible [1268, 437, 1592, 262]. possibly [835, 108, 1941]. Post [1880, 1750, 1776, 1615, 2012]. Post-selection [1750, 1776]. post-stratification [2012]. post-stratified [1615]. postage [465]. posterior [618, 1597, 296, 707, 1346, 1921, 716, 617, 1289, 1220]. posttest [1694]. potency [740, 741, 739, 742]. potential [1392]. Potts [1853]. Potts-mixture [1853]. pour [714, 1022, 1026, 264, 605, 1325, 1075, 977, 1042]. poverty [1489]. Power [2142, 49, 28, 1379, 183, 577, 107, 37, 94, 213, 304, 428, 288]. power-divergence [1379]. power-series [304, 428]. powerful [1472]. powers [576, 9, 11]. pp [149, 166, 148, 170, 219, 220, 237, 238, 239, 240, 266, 267, 265, 288, 309, 519, 122, 188, 205, 169, 104, 221, 222, 186, 103, 147, 165, 167, 168, 189, 124, 202, 201, 307, 217, 464, 285, 187, 204, 218, 287, 102, 310, 311, 290, 150, 123, 289, 368, 185, 135, 459, 152, 522, 200, 284, 286, 105, 203, 461, 520, 466, 151, 460, 462, 121, 369, 134, 463, 308, 184, 306, 465]. practical [1440, 910]. Practice [2148, 1858, 2029, 368]. Practicing [753]. pre [45, 68]. pre-Hilbert [45, 68]. preach [2029]. precipitation [510]. precision [1788, 539]. predict [2054, 1475]. predictable [1552]. Predicting [327]. Prediction [243, 1687, 7, 390, 1801, 855, 630, 778, 1053, 1777, 261, 1935, 1335, 1889, 2058, 1869, 1605, 747, 1131, 1730, 1527, 1717, 2067, 1108, 860, 1686, 1231, 1919, 670, 1171, 990]. predictions [1893]. Predictive [1815, 534, 297, 420, 1546, 296, 446, 1146, 682]. predictor [850]. predictors [823, 1851, 1992, 1855]. Preface [1269, 1478, 1269]. Preface/Préface [1269]. preferences [582, 583, 581, 582]. preliminary [1812]. premier [387]. Prentice [238]. preparation [481]. Prepivoting [1676]. presence [355, 411, 1561, 176, 1261, 131, 1568, 1673, 245, 1524, 136, 1421, 213, 1295, 1491, 1621, 406, 1231, 1482, 1089, 754, 153, 1261]. present [961, 37, 94, 8]. preservation [1278, 323]. Press [166, 148, 219, 239, 240, 103, 307, 217, 285, 187, 204, 123, 284, 286, 203, 369, 521, 149, 151]. Presses [102, 184, 186, 152]. pretest [1694]. prevalence [1438, 1811, 2005]. prevalent [1754]. previous [1342, 1466]. prévision [855]. Price [1858, 1885]. pricing [1858, 1861]. Primer [2118, 169]. Princeton [217]. Principal [1959, 723, 1483, 1933]. principales [723]. Principle [2150, 987, 461]. Principles [519, 2152, 454, 808, 408]. Prior [1795, 984, 1470, 243, 1055, 1004, 1993, 1796, 1316, 422, 2067, 66, 1602, 41, 617]. Prior-based [1795]. prior-data [1993]. priors [608, 1846, 1125, 1993, 1094, 1361, 707, 933, 1423, 936, 970, 1720, 1571]. Privacy [2115, 201]. probabilistic [93]. probabilité [768]. probabilités [2098, 184]. Probabilities [32, 976, 57, 118, 330, 1346, 1386, 1394, 458]. Probability [2112, 1094, 2085, 994, 2118, 2093, 2094, 2120, 1738, 134, 2109, 768, 530, 166, 1244, 1501, 725, 1175, 2024, 454, 950, 211, 1627, 1165, 14, 63, 1800, 1789, 434, 1302, 1376, 1637, 858, 719, 713, 446, 164, 1928, 191, 850, 30, 1423, 521, 704, 360, 169, 218, 202, 123, 135]. Probability-scale [1738]. probit [1851, 1524, 876, 2014]. problem [79, 602, 23, 87, 50, 598, 1238, 1860, 984, 499, 1569, 1025, 1093, 874, 622, 212, 824, 434, 977, 568, 854, 334, 1243, 89, 648, 746, 1928, 709, 564]. problème [50, 977]. Problèmes [2098, 184]. problems [1130, 1456, 58, 1296, 607, 484, 1957, 1504, 1143, 718, 513, 692, 873, 1306, 61, 801, 1540, 162, 272]. procedure [4, 1370, 1310, 1877, 938, 1281, 62, 923]. procedures [58, 1219, 979, 232, 1695, 1027, 1673, 1136, 324, 720, 538, 774, 1700, 1188]. Proceedings [2164, 289, 134, 462, 165]. process [1625, 1817, 601, 1265, 1757, 976, 855, 1810, 299, 301, 243, 113, 565, 661, 1581, 1172, 416, 628, 1430, 1906, 1557, 1161, 556, 1934, 566, 1536, 1537, 1431, 1848, 1739, 1086, 899, 901, 1432, 1542, 1690, 1590, 1607, 724]. processes [1793, 921, 714, 1022, 277, 1344, 318, 879, 1001, 1300, 414, 137, 760, 1088, 363, 413, 852, 2024, 454, 1448, 1075, 260, 837, 1981, 1210, 1193, 345, 1320, 2001, 559, 1971, 400, 450, 1186, 1000, 341, 563, 209, 352, 586, 347, 1485, 495, 1002, 1870, 207, 2163, 150, 465]. processor [1475]. processus [714, 1022, 1001, 855, 1300, 1075, 566, 450]. Product [2137, 556, 311, 1186, 397, 44, 22, 970, 525]. product-limit [525]. production [411]. productivity [898]. Products [256, 107, 33, 11, 2, 541]. professional [2046]. Professors [343]. profile [922, 1212, 1530, 1602, 1650, 1698]. Profiling [1693]. Program [368, 148, 1325, 1322, 1326]. programme [1325]. Programs [2089, 148]. progression [1904]. projected [644]. Projection [1409, 1103, 1250, 1594, 2053]. projects [1814]. proof [79, 107, 942]. proofs [1699]. propensity [1938]. properties [277, 625, 214, 274, 1137, 1367, 660, 926, 867, 283, 435, 1372, 175, 275, 1160, 1010, 974, 512, 12, 280, 850, 510, 126, 323, 1991, 959, 1300, 1642]. property [328]. proportion [822, 1672, 1511]. Proportional [1394, 1304, 1920, 259, 1767, 1016, 795, 1717, 1194, 1461, 1442]. proportional-hazards [795]. proportions [247, 680, 869, 657, 918, 1927, 990]. proposals [1635]. Propriétés [926, 1300]. Prospective [2034]. Prospectus [151]. Protection [773]. Providence [168, 310, 311]. proximal [1837]. proximal-point [1837]. proxy [1859, 1495]. proxy-based [1495]. Pseudo [1274, 2026, 1312, 1689, 1315, 2067, 1667, 1931, 1675, 1171, 1499]. Pseudo-empirical [1689, 1315, 1171]. pseudo-GEE [1499]. Pseudo-likelihood [1274, 1312, 1667, 1931]. pseudo-prior [2067]. Pseudolikelihood [873]. public [1366]. publications [898]. Publishers [461, 151]. Publishing [147, 201, 290, 289, 368, 135, 522, 466, 460, 134]. pulmonary [613, 314]. Pure [265, 185, 465]. purely [1300]. purely-bilinear [1300]. purement [1300]. pursuit [1594]. Purves [124]. Pyke [343].

Q [311]. quadrant [1500, 1600, 1275]. Quadratic [355, 2116, 185, 508, 552, 513, 1221, 453, 448, 1401, 262, 101, 234]. quality [708, 919, 710, 1415]. quality-adjusted [1415]. quality-improvement [919]. quand [1028]. quantal [48]. Quantile [1930, 1958, 2041, 852, 1950, 1681, 1787, 1942, 1934, 805, 505, 587, 1726, 1422, 2061, 586, 798, 2072, 1406, 1755, 1669, 1743, 2009, 1980]. quantiles [1808, 701, 617, 2072, 1631]. quantitative [905]. quantity [708]. quarterly [481]. Quasi [993, 1956, 1098, 1206, 1485]. quasi-likelihood [1098, 1206, 1485]. Quasi-maximum [1956]. Quasi-universal [993]. quasilikelihood [986, 1481]. que [264]. Quebec [462, 186, 102, 152, 184]. Queen [308, 465]. Questions [2127, 486, 290]. queues [1447]. quintic [39]. Quota [241]. Quotient [995].

R [148, 237, 238, 1202, 222, 168, 285, 187, 310, 311, 289, 369]. R. [344]. Radon [932]. railway [1655]. rain [490, 732]. rainbow [699]. rainfall [268, 492]. Ramon [1235]. Randall [167]. Random [727, 535, 2137, 106, 197, 87, 1802, 475, 137, 781, 1830, 1345, 59, 210, 393, 214, 2024, 808, 507, 117, 552, 370, 1632, 1061, 661, 1361, 819, 63, 1664, 690, 1889, 345, 118, 902, 537, 1145, 1349, 298, 2016, 580, 47, 554, 98, 973, 583, 311, 858, 155, 1995, 1766, 933, 1538, 528, 1060, 1931, 256, 1854, 1310, 579, 536, 644, 695, 968, 1487, 774, 871, 101, 213, 1709, 1406, 1030, 1188]. random-effect [871]. random-effects [59, 1854, 1030]. random-walk [774]. randomization [371]. randomization-theory [371]. randomized [1240, 720, 1100, 1, 172, 374, 1876, 1482]. randomly [214, 354, 645, 719]. randomness [71, 631]. rang [872, 387]. range [1570, 988]. rangs [293]. Rank [814, 1802, 1027, 1812, 1426, 1653, 95, 58, 1563, 299, 301, 872, 1909, 1184, 791, 37, 94, 1263, 702, 966, 111, 1615, 367, 44, 347, 172, 374, 387, 763, 1874]. Rank-based [1802, 1027]. rank-invariant [1653]. rank-order [58]. rank-ordered [1615]. Ranked [1398, 1382, 1662, 1761, 1036, 1421, 1615, 2013, 967, 1876]. ranked-set [1382, 1662, 1761, 2013]. ranking [696, 479, 1421, 2013, 488]. ranking-ratio [479]. rankings [1761]. ranks [1256, 293]. Rao [369]. Rapport [1292, 1181, 1248]. rare [1819, 1452]. rate [1426, 530, 937, 1428, 1696, 955, 2015, 159, 1056, 1743, 1638, 1287]. rates [1280, 1328, 976, 197, 781, 872, 643, 1123, 645, 1399, 889, 1927, 1508, 389, 1290, 1365]. rates/Frais [1280, 1328, 1290, 1365]. Rathie [202]. ratio [396, 479, 183, 775, 957, 1109, 1274, 1758, 371, 1367, 133, 1505, 2048, 838, 895, 1369, 1399, 26, 1970, 1263, 1645, 82, 236, 1371, 397, 1617, 115, 1245, 1102, 924, 1712, 533, 1315, 2055, 802, 1549, 1945]. rational [16]. ratios [1097, 1948, 1654, 1191]. Raymond [135]. re [1952]. re-estimation [1952]. reaction [1265]. Reading [368]. realization [724]. reasonable [905]. rebuilding [155]. recall [565]. recapture [1389, 1390, 1386, 1132, 1385, 1391, 1464]. receiver [1469]. recessive [1617]. reciprocity [60]. recognition [401]. Reconstructing [1056, 644]. Reconstruction [291, 1495]. reconstruisant [155]. record [1806, 815]. record-breaking [815]. Recovering [1973]. Recovery [1992, 1390]. recurrent [1808, 955, 1337, 1544, 1582, 1797]. recurrent-event [955]. Recursive [156, 517, 530, 1524]. rédacteur [1181, 1248, 1292]. rédaction [1202, 1245]. rédigées [152]. Reduce [1530, 1127]. reduced [770, 2072].

Reducing [2064, 1667]. reduction [1090, 1622, 1594, 735, 1108, 1413, 1008, 596, 2069, 1261, 1261]. Reed [220]. reexamination [326]. Referee [1683, 1821, 1863, 1912, 2056, 1258, 1343, 1434, 1477, 1575, 1609, 1656, 1715, 1748, 1783]. Referees [1986]. Reference [970]. Refining [403]. reflected [731]. reflecting [76]. reflection [1283]. Reflections [2028, 2030]. Regeneration [715]. Regeneration-based [715]. regime [1861, 1618]. regime-switching [1861, 1618]. regimes [2063]. region [1949, 1154]. regional [1914, 4]. regions [1035, 2042, 1003, 1192, 178, 948, 443, 424, 807, 1621]. registration [1031]. Regression [669, 1569, 1049, 371, 2049, 1915, 1767, 1687, 965, 1303, 2139, 1252, 2051, 891, 1480, 1230, 1648, 1692, 215, 321, 1284, 1563, 978, 817, 299, 1266, 1961, 1428, 770, 1197, 1219, 1975, 1107, 2041, 1930, 1490, 1277, 1788, 717, 1162, 659, 1636, 34, 1769, 546, 905, 1264, 1330, 127, 1957, 1909, 1504, 820, 1752, 585, 1993, 1950, 1019, 1334, 1962, 1044, 1267, 353, 1681, 1974, 1080, 1228, 1664, 1261, 1471, 1301, 524, 628, 1032, 690, 1180, 1545, 230, 1170, 443, 1048, 862, 1136, 1101, 1229, 772, 1118, 1942, 2046, 1082, 1081, 1871, 1585, 1978, 85, 175, 1031, 999, 451, 1111]. regression [583, 835, 998, 1543, 1829, 356, 1979, 1373, 2044, 2066, 1645, 1306, 780, 1976, 1845, 1052, 419, 1779, 890, 1358, 1526, 1051, 336, 1927, 1931, 544, 1726, 440, 1190, 1444, 1422, 191, 1765, 1541, 1017, 850, 736, 2061, 1205, 1884, 154, 1169, 1179, 1593, 195, 1487, 1517, 1964, 732, 1855, 1691, 871, 1994, 2050, 834, 84, 1720, 1959, 678, 1381, 1689, 173, 1134, 1406, 1462, 1516, 1786, 1515, 303, 1669, 1991, 704, 1836, 337, 1743, 1721, 1939, 1159, 1878, 923, 1249, 1465, 1958, 1297, 1377, 1598, 1746, 1944, 1932, 1768, 1870, 1191, 1433, 1737, 2014, 2009, 2034, 1415, 1797, 1936, 2039, 1693, 1826, 1980, 954]. régression [1261, 544, 220]. Regression-spline [669]. regression-type [850]. regressions [705, 1798, 1443]. regressors [1051]. regular [1558, 1608, 495]. regularity [365]. Regularization [1764, 1911, 1788, 1751]. regularly [1841]. regulators [809, 811]. Reid [2074, 2075]. Reidel [134]. reinforced [1751]. Reinhold [308]. reject [1339]. Rejoinder [1158, 225, 410, 386, 887, 1120, 1903, 888, 913]. Rejoinders [1066]. related [1322, 1375, 823, 1722, 569, 873, 252, 1947, 61, 1928, 682]. relatedness [1819]. relating [421]. relation [1438, 558]. Relations [407, 526, 1866]. relationship [782, 1129]. relationships [177, 1939]. Relative [324, 1543, 1299, 979, 1493, 1556]. relaxation [1121]. Release [1322, 1326]. relevance [867]. relevant [893]. Reliability [204, 2132, 874, 498, 281, 972, 158, 553, 1528]. remark [400]. remarks [331]. remedies [1293]. Remerciements [1258, 1748, 1821, 1863, 1912, 2056, 1434, 1477, 1575, 1609, 1656, 1683, 1715, 1783, 1986]. remote [477]. removal [767]. removed [1988]. Rendus [219, 220, 265, 288, 309, 307, 217, 285, 218, 287, 284, 286, 308, 306]. renewal [1210, 1193, 97]. Rényi [388, 429]. Rényi-type [388]. reparametrization [758]. repeated [418, 720, 1217, 1168]. replacement [733, 1627, 1637, 528, 35]. replicate [997, 1627]. replicated [1981, 1242, 966].

Replication [1637, 154, 1646]. Réplique [572]. Reply [866]. Report [1092, 1342, 963, 1029, 1384, 1466, 1141, 1425, 1181, 1291, 698, 1248]. report/rapport [1181, 1248]. reported [747]. reporting [747, 1949]. Reports [1292]. representation [687]. representations [1161, 586]. Reproducing [1979, 1383]. reproduction [1948]. Rerandomization [2083]. Resampling [1677, 586, 1562]. Research [1814, 519, 2152, 134, 1699, 898, 1165, 289, 272]. residential [609]. Residual [442, 526, 416, 322, 71, 1506, 726, 53, 86, 1168, 1755]. residuals [1049, 1032, 1738, 347, 856, 1246, 1247]. résistance [872, 872]. Resolution [384, 1149]. Resolving [671]. resonance [1121]. respect [1736]. respiratory [610, 783, 1590]. Response [1651, 344, 572, 1799, 612, 1333, 1459, 1525, 1162, 1331, 1909, 820, 950, 741, 1962, 1301, 1320, 873, 270, 1302, 1376, 720, 485, 1401, 1541, 1760, 48, 1491, 1792, 1702]. response-adaptive [1760]. response-biased [1541]. Response-dependent [1651]. response-related [873]. responses [1802, 1545, 875, 1419, 1383, 1122, 1865, 1134, 1878]. resting [1906]. resting-state [1906]. restoration [771]. Restricted [916, 1101, 1566, 1226, 1411, 991, 1225, 514, 1463]. restriction [1872]. restrictions [1039, 46, 517, 1934, 1218, 1008]. result [588, 835]. results [632, 260, 651, 457, 860, 295, 341, 910, 234]. Résumés [221, 290, 289, 222]. rétrécisseur [595]. retrospective [684, 2034]. retrospectively [875]. reversals [1225]. reversible [209]. Revesz [521]. Review [149, 166, 148, 170, 219, 220, 237, 238, 239, 240, 266, 267, 265, 288, 309, 1757, 255, 519, 122, 188, 205, 169, 104, 186, 103, 147, 165, 167, 168, 189, 124, 202, 201, 307, 217, 464, 285, 187, 204, 218, 287, 102, 310, 311, 150, 123, 368, 55, 185, 135, 459, 80, 152, 522, 43, 200, 30, 284, 286, 105, 203, 461, 520, 466, 151, 460, 462, 121, 369, 521, 134, 463, 308, 184, 306, 465, 964, 2027, 319, 2032, 1820, 341, 832]. Review/Comptes [219, 220, 265, 288, 309, 307, 217, 285, 218, 287, 284, 286, 308, 306]. Reviews [221, 290, 289, 222]. Reviews/Résumés [221, 290, 289, 222]. Revision [624]. révisions [482]. revisited [1093, 492]. revue [1014]. Reweighting [1652]. Rhode [168, 310, 311]. Richard [288, 2159]. ridge [175, 1812, 195, 1797]. right [1625, 1202, 1828, 1926, 1493, 889, 1755, 1712, 1611]. right-censored [1493, 889, 1755, 1712, 1611]. Risk [2110, 2079, 263, 906, 1678, 253, 417, 1695, 2058, 435, 1543, 334, 155, 503, 1643, 596, 205]. risk-adjusted [1695]. risks [1696, 1935, 916, 1730, 1614, 1461, 806, 358]. risque [155]. Rivest [1458]. RKHS [1648]. RKHS-based [1648]. Road [1321]. Robert [147, 189, 124, 1445, 105]. robin [1116, 87]. Robust [1366, 2030, 616, 930, 1284, 1050, 1975, 1879, 1788, 969, 1128, 927, 1143, 1019, 1576, 945, 1229, 1800, 1701, 1717, 816, 1000, 1526, 1644, 1179, 1307, 1404, 1450, 709, 1604, 989, 764, 406, 455, 826, 1919, 1682, 1703, 1442, 1631, 1188, 1333, 1223, 2041, 2020, 717, 1397, 1636, 676, 1125, 2007, 1335, 1752, 1334, 232, 532, 1449, 1568, 1673, 1101, 437, 1087, 242, 1779, 928, 1012, 1013, 544, 1205, 1503, 1669, 931, 1513, 1746, 1519]. robustes [544]. Robustness [1137, 84, 602, 246, 986, 1102, 1011, 325]. ROC [1192, 1584, 1082, 1527, 1790, 1512]. Roger [124]. role [1885, 883, 858, 884, 1225, 1154, 942]. Ronald [170, 220, 288, 135]. root [697, 698, 1294, 397]. roots [183, 1097, 81, 1006]. Rosenblatt [275]. rotation [1303]. round [1116]. rounding [646]. Routledge [288]. Rs.48.00 [202]. RTI [1038]. rule [31, 679]. Rules [33, 92, 10, 1589]. run [935].

S [1202, 519, 217, 204, 287, 123, 2023, 80, 461, 466, 306]. sacrifice [1034]. Saddlepoint [1653, 1726, 599, 1426, 1037, 1828, 1147, 821]. safe [753]. saisonnier [566]. saisonnière [1028]. salmonella [742]. Sample [1713, 553, 95, 58, 1296, 1156, 1867, 54, 1331, 34, 1599, 260, 20, 819, 1305, 1914, 212, 1460, 766, 175, 434, 1551, 1579, 275, 457, 1036, 512, 863, 1954, 2061, 1872, 514, 282, 199, 1606, 678, 1381, 64, 1712, 1549, 1952, 49]. Sample-size [1713]. sampled [601, 815, 875]. Sampler [2111, 309, 960, 649]. samplers [1606]. samples [1625, 1202, 757, 2027, 1785, 2026, 404, 471, 1869, 1615, 2013, 2012, 1394, 1067, 1750, 1541, 1174, 231, 1231]. Sampling [411, 2087, 283, 2131, 510, 254, 2111, 1438, 1202, 321, 319, 351, 1244, 1083, 1830, 1818, 733, 1398, 1501, 565, 1957, 1569, 20, 189, 1781, 1627, 501, 1382, 1662, 1761, 681, 803, 622, 1771, 1637, 981, 1651, 1704, 1186, 1538, 528, 1421, 1540, 1217, 1931, 191, 850, 325, 115, 1876, 738, 1661, 1142, 1475, 1697, 1703, 1452, 1113, 2165, 291, 309, 121, 369]. San [237]. Sankoff [219, 368]. Särndal [121]. SARS [2005]. SARS-CoV-2 [2005]. Saskatchewan [266]. satellite [1484]. Saxena [187]. SCAD [2041]. SCAD- [2041]. Scalable [1907, 1921]. scalar [1975, 1976]. scale [355, 511, 1892, 1529, 969, 590, 1577, 835, 816, 928, 1012, 1013, 1034, 544, 1310, 829, 1738, 1983, 115, 1691, 804, 472, 901, 1621, 629, 504]. scale-invariant [804]. scale-space [1310]. scaled [414, 1710]. Scaling [2130, 240, 1416]. scallop [1816]. scan [1155]. scatter [642, 406]. scheduled [1686]. scheme [1569, 329, 1475]. Schlee [1078]. Schneider [167]. Science [286, 461, 466, 460, 134, 165, 1014, 864, 1774]. Sciences [1814, 2144, 2037, 943, 222, 1773, 134, 460]. Scientists [2109, 135]. Scituate [123]. sclerosis [662, 663, 664]. Score [1071, 1170, 922, 1676, 44, 1938, 1507]. scores [1256]. scoring [95, 710]. screen [486]. Screening [2131, 822, 626, 1889, 1926, 962, 189]. screening/prediction [1889]. sea [1816, 229]. seal [1045, 946, 1041, 1040]. seals [1042, 1043]. seamless [1596, 1619]. seamless- [1619]. search [229, 810, 382]. Seasonal [484, 381, 555, 274, 1355, 457, 566, 1028, 666, 483]. Seasonality [380]. Seberry [185]. Second [432, 974, 135, 463, 2139, 1518, 220, 146, 86, 2164]. Second-order [432, 974, 1518, 146]. Section [383, 639, 316, 361, 470, 490, 743, 584]. sectional [785, 350, 990]. sectionally [1033]. sections [535]. seedlings [697, 698]. segregation [1345]. Selecta [222]. Selected [310, 311, 2096, 67, 168]. Selecting [428]. Selection [643, 1527, 1646, 847, 1757, 1219, 1975, 1705, 1274, 1920, 823, 723, 903, 1957, 1632, 569, 1267, 1974, 1449, 1136, 1572, 1016, 1710, 1498, 1764, 1798, 1596, 1731, 1940, 1976, 1523, 2040, 1684, 967, 1394, 1750, 1831, 420, 1776, 1146, 1463, 1462, 1729, 993, 1991, 2071, 1700, 1958, 1690, 1619, 1797]. selection-biased [1957]. selections [1332]. selector [1599]. Self [1230, 1525, 1620, 400]. Self-concordance [1620]. Self-consistent [1525]. Self-modelling [1230]. self-similar [400]. Semi [741, 1730, 333, 1644, 1122, 209, 1542, 1611]. semi-competing [1730]. semi-Markov [333, 1122, 209, 1542, 1611]. Semi-nonparametric [741]. semi-parametric [1644]. Séminaire [152, 184, 186]. seminar [289]. Semiparametric [1238, 1430, 1536, 1963, 1996, 1506, 1766, 1657, 1017, 1254, 1381, 1272, 1553, 1712, 1753, 1932, 1945, 1870, 1274, 1672, 1696, 1957, 1649, 1881, 1369, 900, 1908, 1548, 1840, 1639, 1824, 1520, 1524, 1931, 1852, 1835, 2050, 1273, 1755, 1650, 1377, 1113]. semivarying [1711]. sensing [477]. Sensitive [506, 543]. Sensitivity [1055, 617, 2014, 1470, 1667, 716, 1661]. separable [1626]. separation [63]. September [1732]. Sequence [2136, 2148, 822, 285, 368]. sequences [632, 214, 852, 454, 844, 825, 88, 98, 672, 2082, 234, 276, 2055]. Sequential [1591, 565, 981, 1727, 1573, 1240, 950, 651, 838, 1305, 1710, 251, 1894, 1411, 792, 473, 1179, 1991, 826, 1952]. séquentielles [251]. Serge [184]. serial [1355, 1073, 1172, 791]. serially [1377]. Series [2133, 185, 286, 105, 148, 237, 265, 1886, 760, 1255, 555, 40, 499, 221, 117, 2089, 1198, 1355, 1497, 1578, 233, 880, 588, 1670, 1471, 1695, 1883, 1455, 1840, 348, 359, 1352, 785, 481, 431, 1353, 1852, 1984, 798, 1110, 647, 64, 90, 304, 428, 482, 348, 148]. serological [2005]. Services [1683, 1821, 1863, 1912, 2056, 200, 1258, 1343, 1434, 1477, 1575, 1609, 1656, 1715, 1748, 1783, 1986]. Services/Remerciements [1912]. ses [595]. Set [1084, 134, 516, 1398, 796, 1382, 1662, 1761, 1036, 1421, 1615, 2013, 1876, 1936]. set-based [1936]. Sets [1928, 23, 821, 934]. setting [1860, 796]. settings [1538, 1595]. setup [724]. several [157, 1578, 657, 807, 31, 917, 548, 1257, 75, 376, 1407]. sex [1388]. sex-specific [1388]. shadows [695]. Shannon [1722]. shape [1934, 1218, 1008, 1432, 1939]. shaped [256]. share [194]. Sharpe [1970]. shift [300, 1791]. shifts [1373]. Shiryayev [461]. Shizuhiko [240]. short [1886, 741, 739, 1099, 610, 1321]. short-term [741, 739, 610]. Shortest [261, 26]. Shrinkage [1424, 2009, 978, 511, 595, 574, 419]. sib [1152, 1268, 1367]. sided [607, 1097, 435, 362, 1624]. sides [849]. sieve [1649]. sign [1207]. signal [1750, 1310]. signals [1992]. signature [1904]. signé [387]. signed [1563, 301, 1097, 387]. signed-rank [1563]. significance [262]. silver [809, 811]. silver-maple-tree [811]. similar [725, 400]. similarity [1167]. Simon [288]. Simple [1183, 1097, 870, 1373, 848, 1830, 640, 1684, 352, 41, 732]. simplex [727]. simplicial [737]. simpliciale [737]. simplified [4, 1235, 1234]. Simulation [1510, 15, 2029, 1534, 586, 1435, 15]. simultanément [544]. Simultaneous [680, 1632, 1976, 928, 1012, 1173, 395, 1537, 25, 504, 1013]. simultaneously [544]. Single [422, 1546, 1680, 1492, 1298, 49]. single-index [1680, 1492, 1298]. Single-parameter [422]. single-sample [49]. Singular [51, 476, 174, 2025, 108]. Sinha [204]. sinusoids [136]. SIR [1948]. Size [2101, 1112, 1806, 1713, 307, 275, 310, 1028, 2129, 488, 553, 586, 1729, 1952]. sizes [1750, 514]. skew [1722, 1190, 1444, 1669]. skew-normal [1722]. skewed [1332, 1923, 1318]. skewness [831, 472]. Slepian [430]. slippages [129]. Slow [960]. SLSEs [278, 340]. Slutsky [217, 2126]. Small [34, 260, 1585, 1204, 1489, 512, 785, 678, 1142, 1801, 869, 471, 1893, 1735, 1791, 1067, 1644, 1009, 1734, 1450, 1964, 1068, 1168, 1502, 1679, 1839, 1689, 1231, 1919, 1171]. Small-area [785, 869, 1919]. small-circle [1009]. Small-sample [260, 678]. Smirnov [302, 300, 1275]. Smooth [1070, 1224, 1330, 1725, 1519]. Smoothed [1493]. smoother [677]. smoothers [817, 495]. Smoothing [1267, 968, 658, 1095, 1441, 1331, 1237, 1828, 1976, 783, 1077, 1517, 1969, 2014]. Social [864, 318, 673]. Society [168, 312, 171, 310, 311, 391]. Sociological [104]. Sociology [2092, 104]. soit [264]. sojourns [1665]. solution [602, 648, 746, 1751]. Solutions [306, 76, 484, 167]. Some [294, 655, 1333, 943, 226, 699, 711, 393, 1448, 484, 718, 1334, 651, 1227, 532, 10, 175, 275, 457, 333, 1160, 860, 295, 61, 281, 1472, 1009, 910, 1015, 280, 758, 908, 561, 1818, 607, 1226, 2032, 329, 232, 692, 345, 143, 94, 456, 38, 159, 514, 617, 841, 2070]. Somesh [147]. somme [830]. Sondage [102]. sondages [1261]. Sons [170, 220, 238, 267, 288, 309, 519, 122, 218, 459, 121, 306]. sont [405, 1028, 544]. sources [740, 2013, 1870]. Space [1344, 1095, 117, 1914, 192, 851, 897, 1374, 1495, 1148, 2053, 45, 68, 1310, 1727, 1077, 1682, 1816]. space-filling [2053]. Space-time [1344]. spaced [321, 1841, 228]. spaces [625, 98, 1383, 995, 1010]. spacing [944]. spacings [294, 508, 620, 413, 567, 212, 99]. Spanier [522]. Sparse [2079, 2031, 1648, 1752, 1916, 1971, 1992, 1845, 1973, 1933, 1855, 1959, 1958]. Sparsity [1798, 2076]. Spatial [1131, 1834, 1839, 1022, 784, 1345, 1841, 1904, 1360, 1372, 1484, 1570, 1908, 1356, 1739, 1626, 1462, 1917, 1556, 1690]. spatially [714, 1267, 1982]. spatiaux [714, 1022]. Spatio [1508, 1981, 1907, 1402]. Spatio-temporal [1508, 1981, 1907, 1402]. spatiotemporal [1853]. Special [1814, 2074, 490, 1857, 1773, 1968, 1899, 2022, 1531, 1478, 1587]. species [1035, 254]. specific [95, 813, 941, 1081, 1388]. specification [1018, 1063]. specified [354, 1779, 889, 617]. Spectral [274, 2001, 556]. spectroscopy [879]. spectrum [699]. speed [1655]. speeds [469, 467, 468]. sphere [1112, 535]. spheres [618, 1597]. spherical [2016, 1009]. spherically [516, 1059]. Spiring [839]. spline [669, 855, 1266, 643, 1237, 1649, 1969, 2014, 855]. spline-based [1649]. splined [643]. splines [1671, 1261, 1564, 1516]. split [1894]. split-and-conquer [1894]. spraying [1325]. spread [573]. Springer [221, 464, 520, 463]. Springer-Verlag [221, 464, 520, 463]. sprouts [809]. spruce [697, 698]. spurious [131, 198, 540]. square [1668, 830, 641, 966, 751, 397, 53]. squared [577, 1135, 594, 512, 128, 1964, 1502, 161, 804, 856, 1246, 1247, 1919, 670]. squared-error [804]. squares [497, 794, 1777, 432, 330, 439, 1127, 25, 666, 108, 606, 691, 1462, 1465]. Sri [239]. Srivastava [287, 466]. SSC [222]. stabilité [737]. Stability [672, 549, 737]. Stabilizing [1067]. stable [454]. Stacy [1625]. Staffordshire [104]. stage [1219, 1719, 1637, 1704, 2012]. standard [223]. Stanton [222]. star [256]. star-shaped [256]. State [851, 1495, 257, 1095, 1906, 1374, 1537, 1983, 1077, 1682, 1542, 1816, 897]. State-space [851, 1495, 1816]. state-switching [1983]. statements [164, 648, 746]. states [1417]. Stationary [1374, 277, 852, 844, 588, 556, 566, 400, 1401, 1590]. stationnaire [566]. Statist [721, 1202, 1140, 778, 941, 1063, 1210, 897, 745, 1235, 1245, 1013, 1247]. statistic [299, 301, 1037, 140, 1, 924, 533, 1220]. statistica [222]. Statistical [2143, 2086, 1814, 2144, 2027, 1775, 2054, 700, 1269, 1116, 2091, 2150, 2164, 345, 948, 2128, 1771, 2141, 1903, 1900, 2122, 2095, 2084, 1902, 889, 105, 162, 2037, 134, 962, 272, 491, 1901, 1598, 1697, 149, 148, 238, 698, 1546, 943, 490, 613, 1699, 54, 2154, 147, 1094, 365, 346, 1382, 1227, 883, 4, 1586, 1014, 947, 1778, 289, 995, 1820, 295, 1634, 1047, 1640, 562, 203, 461, 460, 2035, 1709, 758, 1905, 592, 249, 312, 926, 1773, 171, 391, 450, 267, 55, 459, 465]. statisticians [884]. Statistics [2113, 169, 168, 189, 2085, 2099, 2100, 898, 2140, 1857, 2102, 310, 311, 2153, 2105, 2120, 2107, 2096, 135, 401, 200, 2124, 2149, 105, 462, 2109, 2097, 2139, 2142, 465, 79, 1256, 248, 1026, 302, 300, 40, 755, 1175, 2024, 407, 221, 165, 1049, 569, 476, 1367, 994, 1165, 567, 951, 357, 143, 187, 112, 438, 1350, 1676, 123, 2094, 631, 1371, 367, 44, 753, 857, 8, 600, 172, 293, 387, 1774, 2, 799, 1311, 599, 683, 942, 1936, 1314, 1329, 338, 168, 170, 220, 222, 167, 202, 464, 287, 466, 151, 462, 521, 306, 288]. Statistics [1340, 1364, 1947, 43, 30]. Statistics. [124]. statistique [54, 1014, 450, 152]. statistiques [1026, 926, 293, 387]. status [2049, 1767, 1543, 1537, 1254, 1461, 1381, 1515]. Stein [502, 263, 421, 2017, 253, 326, 503, 1812, 1689, 596]. Stein-type [1812]. Step [795, 1837, 762, 1430, 1051, 938, 62]. Step-function [795]. step-up-down [938]. step-wise [62]. Stephen [221, 103]. stereograms [1533]. Sterile [1322, 1325, 1326]. steriles [1325]. Stochastic [370, 1671, 2135, 1857, 36, 2110, 921, 76, 318, 78, 1053, 605, 1859, 1981, 956, 1557, 1881, 252, 556, 436, 1809, 1483, 450, 1494, 917, 789, 465, 557, 266, 205]. stochastically [1555]. stochastique [2104, 152]. stochastiques [557, 605, 450]. stock [1682]. storm [1817]. Strassen [113]. strategies [1333, 191]. strategy [597]. stratification [2012, 1703]. stratified [1083, 1830, 733, 20, 1704, 1615]. streaming [2044]. streams [1780]. strength [1546, 1602, 553, 936, 1939]. strength-stress [553]. stress [1602, 553, 936]. stress-strength [1602, 936]. strict [632]. strictement [1300]. Strictly [1330, 1300]. String [2148, 368]. Strong [508, 808, 1320, 17, 1139, 412, 211, 645, 97, 586, 1831, 1236, 802, 521]. Structural [211, 83, 1395, 1129, 2023, 1503, 2009, 954]. structure [1924, 1123, 1787, 1369, 1940, 1992, 1186, 1278, 1855, 35, 841, 1729, 2010, 360].

Structured [164, 1910, 1422]. structures [501, 228, 246, 1630, 726, 448, 358]. Stuart [306]. Student [426, 49]. Studentization [730, 13]. Studentized [244, 357]. studies [258, 1239, 1918, 1694, 2029, 961, 2065, 547, 217, 232, 1521, 1554, 1346, 1543, 1903, 1900, 1842, 1651, 780, 1241, 1754, 1902, 1617, 1640, 1507, 1528, 871, 1994, 491, 1901, 1482]. study [1150, 1322, 248, 612, 697, 1742, 609, 1041, 1151, 708, 1670, 313, 1847, 1985, 37, 1910, 1812, 1040, 1392, 1453, 1816, 1325]. Styan [147]. sub [1661]. sub-populations [1661]. Subclasses [276]. subdistribution [1997, 1717]. Subgroup [2066]. subject [1920, 813, 941, 1827, 1122, 1657, 1464, 1870]. subject-specific [813, 941]. subjected [137]. submarine [229]. Subrahmaniam [169]. subsampling [1139, 1950, 2061]. Subscription [1280, 1290, 1328, 1365]. subset [967]. Subspace [2011, 1660]. sufficiency [754]. sufficient [40, 802, 2069, 264]. suffisante [264]. suites [98, 276]. sum [830, 1161, 1235, 341]. summaries [617]. Summary [221, 290, 289, 1936, 222]. summed [1569]. summer [609]. sums [508, 107, 569, 388, 1992, 1654]. Sun [1202, 1202]. superdiagonal [1300]. superdiagonaux [1300]. superficies [477]. superharmonic [395]. Supérieures [184, 186, 152]. superpopulation [630, 778]. superposition [1002]. Supersaturated [878, 877]. supersmooth [1224, 668]. supervised [2070]. Supply [200]. support [1574, 1751, 489]. supported [1961, 1137]. supremum [843]. Sure [1926]. surface [1401]. suroptimale [1250, 1250]. surprise [1299]. surrogate [1504]. surveillance [1695]. Survey [2087, 2165, 2138, 1736, 319, 486, 1304, 1830, 1561, 2026, 480, 803, 1260, 1568, 478, 485, 1204, 1838, 863, 667, 1593, 748, 638, 121, 369, 1502, 308, 1171, 487]. surveys [1459, 1499, 1331, 1592, 864, 1890, 199, 162, 1675, 1213, 1315]. Survival [1304, 1124, 1426, 669, 914, 1742, 1427, 1276, 1930, 1672, 417, 651, 1430, 537, 1468, 1828, 2006, 423, 1623, 1805, 1639, 1896, 1298, 1995, 719, 1386, 1388, 1034, 1657, 1643, 899, 1565, 1743, 1698, 1941, 1556]. survival/sacrifice [1034]. survivor [1555]. survivors [1124]. Susceptible [1988]. Susceptible-exposed-asymptomatic-symptomatic-active-removed [1988]. switch [1577]. Switching [1769, 36, 1861, 1983, 1618]. Symmetric [1728, 516, 1059, 627, 974, 43]. symmetrical [532, 38, 99]. symmetrically [1311]. symmetry [1203, 1088, 848, 1178, 1634]. symposium [165]. symptomatic [1988]. synthesis [1093, 1811]. Synthetic [1998, 1847, 1985]. System [2101, 2079, 1806, 307, 118, 498, 1655, 25, 936]. Systematic [351, 321]. Systems [2135, 266, 965, 148].

T [147, 203]. table [608, 1269, 751, 93, 483]. tableau [483]. Tables [168, 310, 311, 702, 2095, 2131, 2096, 1135, 100, 105, 189]. tail [976, 1789, 1309, 832]. tailed [1440, 1980]. taille [1028, 488]. tails [1335]. tale [994, 2077]. talk [1043]. Tanur [151]. target [1416]. tau [1073, 1601, 1658]. Taux [872]. Teaching [338]. Team [1814]. technique [614]. Techniques [1636, 1757, 1095, 163, 1009]. télédétection [477]. telephone [580]. telephones [908]. template [1387]. Temporal [379, 613, 499, 1981, 1484, 1131, 1907, 1402, 1508, 1870]. tennis [93]. tensor [1908]. term [1124, 741, 739, 610, 1321]. terminal [809, 1607]. terms [1236]. Test [830, 548, 2019, 1625, 608, 757, 602, 52, 848, 183, 775, 957, 1109, 1138, 1532, 1867, 577, 640, 1049, 1367, 1018, 1063, 1429, 1285, 2048, 1576, 1458, 791, 1027, 94, 1020, 71, 1207, 742, 1843, 702, 1537, 348, 236, 882, 751, 1036, 917, 1617, 671, 1812, 1275, 1640, 1819, 938, 1245, 172, 603, 682, 1286, 1406, 1491, 1960, 1641, 1768, 2039, 830, 167, 496]. Testing [1338, 1574, 216, 1395, 1417, 1473, 1567, 2017, 796, 1505, 1761, 1787, 354, 1982, 204, 1577, 1400, 998, 1222, 2018, 684, 1371, 1522, 1257, 621, 1178, 75, 154, 129, 2132, 1126, 1516, 315, 533, 1737, 1443, 526, 83, 1345, 793, 893, 1892, 2020, 607, 640, 407, 547, 1198, 1355, 1497, 1578, 233, 916, 1361, 1135, 353, 1600, 18, 1827, 1798, 1579, 132, 140, 1346, 7, 378, 792, 82, 713, 991, 1, 1353, 1517, 69, 213, 1941, 1932]. Tests [1088, 880, 1172, 279, 578, 434, 2081, 158, 373, 1426, 1653, 1793, 1203, 891, 1396, 355, 95, 1202, 997, 1296, 872, 598, 849, 1156, 1274, 1713, 1529, 607, 1846, 2065, 507, 1071, 1078, 1725, 916, 1061, 1662, 853, 471, 534, 1500, 141, 532, 1305, 1301, 145, 1170, 828, 827, 94, 1369, 435, 1372, 1551, 132, 650, 745, 251, 1111, 356, 1838, 1263, 447, 1849, 996, 966, 1580, 1358, 89, 1115, 1472, 1726, 1955, 1439, 881, 1174, 989, 374, 962, 1621, 1624, 1905, 1549, 1191, 2034, 1936, 1954]. Textbooks [169, 189, 105, 265]. Theil [84]. their [24, 1525, 2024, 346, 92, 322, 2000, 857, 1311]. Theodorescu [222]. theorem [421, 244, 552, 107, 430, 116, 633, 97, 551]. théorème [633]. Theorems [2153, 294, 844, 1446, 661, 464, 672, 841]. theoretic [1240]. Theoretical [1372, 2152, 910, 519]. Theories [134]. Theory [1216, 2133, 2085, 1269, 874, 123, 2119, 2148, 286, 134, 1156, 2154, 1473, 1061, 371, 212, 1850, 554, 368, 2084, 55, 427, 108, 863, 80, 1812, 196, 254, 134, 237]. there [73, 884]. Thinkers [2156]. Third [1265, 285, 306, 170]. Thivent [1210]. Thomas [170, 220, 288]. Thompson [1588, 1587]. Three [1436, 1608, 34, 1041, 1042, 1537, 1371, 667, 1040]. three-armed [1371]. Three-level [1608]. three-state [1537]. three-variable [34]. threshold [518, 1138, 2048, 1800, 635, 1896, 1052, 469, 634, 1956]. thresholding [1972]. Tightness [529]. tilted [1915]. Time [2133, 760, 2128, 2136, 2148, 899, 1230, 238, 1265, 1344, 976, 1997, 1427, 1886, 1255, 555, 1613, 2020, 1277, 825, 1198, 1355, 1497, 1578, 233, 916, 880, 1408, 588, 1670, 651, 1471, 1883, 1468, 270, 47, 900, 72, 1455, 1535, 1362, 1840, 1996, 1321, 2045, 1809, 348, 359, 1995, 1386, 1352, 889, 785, 1391, 481, 634, 431, 1353, 1852, 209, 1984, 798, 2060, 1121, 1110, 647, 1542, 1252, 1590, 237, 285, 368]. time-dependent [1468, 1386]. Time-discrete [899]. time-invariant [1230]. time-reversible [209]. time-sequential [651]. time-series [588, 785, 481]. time-to-event [1277, 1535, 2060]. time-to-failure [1542]. time-to-response [270]. time-varying [1362, 1809]. times [241, 1053, 1408, 1554, 1623, 1701, 1565, 1743, 959]. Tiwari [1202]. Tolerance [1090, 2131, 189, 178, 231]. Tolerance-based [1090]. too [951]. tool [1594]. tools [1634]. Topics [2165, 655, 369, 462]. topographie [1325]. topography [1325]. topological [98]. topologiques [98]. Toronto [240, 203, 151]. total [411, 1818, 436]. toxicity [270]. trafficking [1321]. train [315]. trajectories [1818]. trajectory [1973]. transdimensional [1209]. Transect [944]. Transect-spacing [944]. transects [902]. transform [849, 346, 37, 765, 1862]. transform-both-sides [849]. transformation [1705, 1613, 659, 1680, 1681, 1449, 894, 1536, 1639, 1731, 1926, 1254, 1786, 1159, 1405, 1252, 1442]. transformation-based [1681]. transformations [1156]. transformed [1088, 1705, 904]. Transforming [1762]. transforms [453, 446]. Transient [363]. transition [78, 1349, 330]. translation [389]. transplanted [524]. transplanted-kidney [524]. transported [609]. travel [1321, 1386]. treatment [2065, 663, 700, 2063, 1655, 1589, 1754, 99, 1491, 1482, 2047, 2008, 1628, 1702]. treatments [2062, 362]. tree [640, 1535, 811]. trees [1427, 809, 1044, 1687]. treillis [714]. trellis [714]. trend [1426, 411, 853, 533]. trends [613, 961, 1670]. trial [73, 157, 700, 1628]. trials [655, 1240, 2048, 1989, 72, 1371, 61, 473, 1760, 2050, 1583]. triangle [1268]. triangulaire [714]. triangular [714, 987, 808]. tricky [854]. trimmed [809, 604, 929, 49]. trimmed-mean [604]. trois [1042]. true [1336]. Truncated [1558, 473, 840, 1742, 40, 625, 96, 324, 1493, 719, 1526, 829]. truncation [1278]. Tryambakeshwar [462]. tumorigenicity [1381]. Tuning [1532]. tuples [452]. twentieth [2156]. twenty [379]. twins [1906]. Two [1480, 1219, 813, 1460, 2053, 2137, 2012, 514, 1549, 1837, 608, 247, 58, 1083, 257, 363, 1296, 605, 216, 1918, 941, 1417, 1330, 73, 547, 880, 1285, 405, 994, 926, 1305, 100, 443, 1027, 902, 1770, 212, 37, 1719, 322, 2001, 434, 26, 1242, 1637, 1896, 1592, 311, 2077, 1651, 348, 1704, 292, 320, 1036, 2043, 1257, 1051, 158, 1353, 191, 1872, 325, 1206, 213, 1149, 1381, 1646, 1714, 1712, 533, 1113, 1022]. two-decision [58]. Two-dimensional [2053, 605, 405, 926, 902, 2043]. two-factor [1149]. two-level [1770, 1242, 292, 1646]. Two-moment [514, 533]. two-parameter [1257, 158]. Two-part [1480]. two-phase [1083, 1918, 1651, 1113]. Two-sample [1549, 58, 1296, 1305, 212, 434, 1036, 1872, 1381, 1712]. Two-stage [1219, 2012, 1719, 1637, 1704]. two-state [257]. two-step [1837, 1051]. two-way [608, 1285, 100, 1027, 37, 1896, 191, 213]. type [891, 302, 300, 113, 727, 1664, 388, 438, 140, 15, 1812, 1275, 850, 1405, 757, 1428, 15]. types [851, 897]. typhoon [1960].

Uhlenbeck [1793]. ultrahigh [1940]. ultrahigh-dimensional [1940]. Ultrastructural [836]. umbrella [1396, 598]. Unaliasing [1474]. unbalanced [1256, 471]. Unbiased [607, 376, 497, 1171]. uncertain [1570]. uncertainty [1798, 1937, 1655, 1388, 2005, 1231]. unconditional [650, 745]. Under-reporting [1949]. underdetermined [968]. underlying [1973]. unequal [1501, 1860, 1627, 1637, 540, 514]. unequal-probability [1501]. unfolding [1112]. unicité [737, 544]. unidirectional [1721]. unified [1332, 1449, 193, 1357, 1287]. Uniform [299, 301, 252, 436, 821, 294, 508, 620, 127, 412, 1037, 645, 98, 917, 17]. uniformes [98]. uniformity [1338, 1574, 407, 2053, 1472]. uniformly [538]. Unifying [2034]. unilatéraux [279]. unimodal [892, 1218, 617]. unimodales [892]. unique [277]. Uniqueness [820, 712, 717, 1599, 737, 544]. Unit [1734, 1331, 1938, 207]. Univariate [1081, 10, 907]. Universal [305, 1268, 993]. universe [1705]. Université [102, 152, 184, 186]. University [148, 219, 220, 240, 266, 265, 104, 222, 165, 307, 217, 2164, 285, 289, 203, 462, 134, 465, 431, 288, 309, 308]. unknown [1574, 511, 848, 516, 1860, 657, 498, 835, 1350, 1262, 970, 1786]. unmatched [1142]. unmet [638]. unobserved [1818]. unreplicated [1200]. unrestricted [2000]. Untruthful [720]. unusual [229]. update [1165]. updating [1780]. Upper [231, 1335]. urn [1952, 1583]. US$14.64 [307]. US$14.95 [284]. Use [486, 713, 481, 83, 1846, 1221, 283, 525]. used [1451, 354, 132, 159]. useful [1778]. uses [143]. Using [1972, 1484, 1263, 1127, 634, 1709, 1876, 608, 95, 914, 449, 1095, 649, 1128, 1397, 983, 1935, 741, 2054, 1505, 1576, 739, 1796, 1659, 498, 1302, 1376, 1771, 1212, 1346, 955, 1910, 1527, 1811, 701, 2064, 816, 587, 2081, 1218, 15, 2062, 801, 971, 1790, 1897, 789, 1056, 41, 199, 1606, 1502, 1273, 1096, 1463, 1516, 1874, 1171, 1142, 2014, 1936, 2008, 1042, 1424]. Utilisation [1042]. utility [603]. utilizing [66].

V [222, 168, 185, 203]. vaccine [2050]. valeurs [926]. validation [1117, 1470, 1827, 163, 1991, 1753]. Validity [1865, 1268, 1235, 1832, 1234, 821]. Valley [1323]. value [1143, 1458, 926, 1664, 1410, 749, 1551, 682, 467, 468]. valued [1963, 1163]. values [1426, 1451, 1185, 1501, 742, 9, 868, 600]. Variability [1488, 740, 268, 1484, 492]. Variable [1920, 1974, 1572, 1596, 1731, 1940, 1462, 1700, 1690, 1619, 1797, 1757, 449, 1975, 823, 34, 723, 412, 1449, 1851, 1016, 1916, 735, 1976, 2040, 1684, 1872, 647, 1463, 1491, 1958]. variable-kernel [735, 647]. Variables [2137, 197, 475, 630, 778, 1428, 1375, 643, 987, 1398, 214, 808, 723, 1330, 1133, 63, 145, 1173, 622, 2016, 1263, 2064, 397, 256, 579, 968, 101, 1780, 1688, 2008, 1322, 311]. Variance [1451, 1244, 1083, 1114, 479, 511, 686, 969, 1004, 20, 1627, 945, 1221, 1261, 1316, 1568, 1320, 554, 576, 1548, 1637, 587, 1660, 44, 850, 527, 691, 376, 1439, 868, 282, 1068, 213, 119, 806, 1562, 455, 575, 929, 180, 227, 496, 1030]. variances [216, 516, 1074, 1505, 26, 2077, 970, 1286, 540]. variants [1819]. variate [1962, 2025, 162]. variates [22, 173]. Variation [752, 2108, 219, 552, 741, 1418, 742, 1028, 991, 1877]. variational [1928]. variations [908]. various [922, 1836]. Varying [1805, 2049, 1787, 1492, 1362, 1843, 1610, 1809, 1835, 1465, 1317]. Varying-association [1805]. varying-coefficient [1492, 1843, 1465, 1317]. varying-coefficients [1835]. vecteur [595]. vecteurs [293]. vector [511, 1198, 1497, 1351, 595, 807, 992, 1060, 2071, 1751]. vectors [18, 293]. vehicle [908]. vergers [1325]. verification [553, 1621]. Verlag [221, 464, 520, 463]. versus [19, 969]. VI [310]. via [2019, 2042, 1296, 1788, 1993, 1382, 1880, 1572, 430, 1585, 1687, 1399, 1370, 1483, 656, 2043, 1660, 1309, 1517, 2002, 1462, 1191, 1571, 1619, 2034]. victimization [1592]. Victoria [220]. Victoria. [219]. view [1332]. vignettes [1407]. VII [311, 204, 310, 465]. viii [168, 289, 200, 461]. vine [1523]. vines [1558, 1804]. visualization [1534]. vital [481]. vitamin [1590]. vocalisations [1042]. vocalization [1041, 1040]. vocalizations [1042]. vol [169, 221, 186, 189, 185, 152, 184]. volatilities [1860]. Volatility [1686, 2020, 1859]. Volterra [345]. Volume [1280, 1290, 1328, 1365, 1740, 1723, 1732, 168, 310, 311, 286, 105, 151, 134, 265, 222]. voting [975]. vraisemblance [714].

W [148, 147, 124, 310, 311, 286, 105, 134]. waiting [1053]. Wald [791]. Wald-Wolfowitz [791]. walk [210, 2024, 298, 47, 554, 774]. walks [393, 858]. Walpole [135]. Warkentyne [219]. Warps [2148, 368]. Was [1552]. Washington [522]. Wasserman [167]. water [710]. Waterloo [165]. Watson [127]. wavelet [1112, 1282, 1163]. way [608, 1285, 1361, 100, 1027, 37, 1896, 882, 191, 213, 496]. Weak [852, 645, 586, 1446, 98]. weakly [808, 2055]. Weibull [114, 283, 1602]. weighing [564]. weight [33]. Weighted [2070, 413, 1921, 1422, 1203, 830, 2041, 1788, 1143, 1752, 661, 1221, 838, 1572, 867, 1166, 1024, 1579, 1111, 843, 1676, 1127, 1790, 606, 161, 1308, 1960, 455, 1661]. Weighting [1593, 1376]. weights [1019, 1414, 1169, 1481, 1502, 1171]. Weinberg [680]. well [538]. Wesley [368]. Western [134]. whether [354]. Which [437, 118]. white [601]. white-noise [601]. Whitmore [205, 167]. Whitney [1036, 1061, 1037]. Who [1044]. whose [405, 364]. wide [1910]. Wiener [113]. Wilcoxon [1037, 1036]. wild [2042]. Wiley [166, 170, 220, 238, 267, 288, 309, 519, 122, 187, 204, 218, 459, 284, 121, 306]. Wilks [306]. will [1043]. William [220, 221, 167, 311, 151]. Williams [309]. wind [469, 467, 468]. Winfried [266]. winning [597]. Winsorized [49]. wise [62]. Wishart [183, 230, 2025]. within [2079]. without [733, 565, 296, 1627, 1637, 154, 1702]. WLSEs [278, 340]. Wolfe [598]. Wolfowitz [791]. Wonnacott [170, 220, 288]. work [1446]. worlds [994]. Worldwide [1165]. worth [417]. Wretman [121].

X [220, 306, 166, 265, 221, 147, 150, 105, 520, 466, 274, 167, 307, 152, 105, 134]. X-11-ARIMA [274]. xi [238, 239, 267, 122, 123, 121, 134, 306]. xii [237, 266, 217, 187, 102, 135, 203, 520, 308]. xiii [148, 240, 165, 285, 460]. xiv [205, 520]. xix [220, 167]. xv [309, 124, 202, 459, 369]. xvi [149, 464, 287]. xvii [219]. xviii [288, 307]. xx [286]. xxi [519, 151, 463]. xxii [170].

Yatchew [1078]. years [379]. Yogendra [462]. York [166, 170, 219, 220, 238, 239, 266, 267, 265, 288, 309, 519, 122, 169, 221, 147, 189, 124, 464, 187, 204, 218, 287, 150, 185, 135, 459, 284, 286, 105, 461, 520, 466, 151, 460, 121, 369, 521, 463, 308, 306]. Youden [1945]. Young [111].

Z [105, 151]. Zero [2046, 1480, 1185, 1501, 1071, 1170, 1927, 683, 458, 1550]. zero-heavy [1550]. Zero-inflated [2046, 1480, 1170]. zero-one [683]. zeroor-one [1927]. zeros [1673, 862].

References

Ogawa:1973:AND

[1] Junjiro Ogawa and Sadao Ikeda. On the asymptotic null distribution of the F -statistic for testing a partial null hypothesis in a randomized PBIB design with m associate classes under the Neyman model. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2): 1–23, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Tracy:1973:CMP

[2] D. S. Tracy and B. C. Gupta. A combinatorial method for products of generalized h-statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):25–33, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Srivastava:1973:EME

[3] M. S. Srivastava. Evaluation of misclassification errors. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):35–50, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Karasek:1973:SSE

[4] Mirek Karasek. A simplified statistical estimation procedure for regional econometric models. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):51–67, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Khazzoom:1973:EMD

[5] J. Daniel Khazzoom. An econometric model of the demand for energy in Canada. Part One: The industrial demand for gas. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):69–107, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Ahsanullah:1973:CPD

[6] M. Ahsanullah and A. B. M. Lutful Kabir. A characterization of the Pareto distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):109–112, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Lingappaiah:1973:PEL

[7] G. S. Lingappaiah. Prediction in exponential life testing. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):113–117, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Sinha:1973:DOS

[8] S. K. Sinha. Distributions of order statistics and estimation of mean life when an outlier may be present. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):119–121, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Rahman:1973:NEV

[9] M. Rahman and M. Ahsanullah. A note on the expected values of powers of a matrix. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):123–125, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic). See erratum [42].

Gupta:1973:SNC

[10] A. K. Gupta and Z. Govindarajulu. Some new classification rules for c univariate normal populations. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):139–157, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Rogers:1973:PPG

[11] G. S. Rogers and D. L. Young. On the products of powers of generalized Dirichlet components with an application. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):159–169, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Saxena:1973:CPB

[12] R. K. Saxena and P. L. Sethi. Certain properties of bivariate distributions associated with generalized hypergeometric functions. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):171–180, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Tan:1973:MSA

[13] W. Y. Tan. Multivariate Studentization and its applications. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):181–199, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

George:1973:PMI

[14] A. George. A probability model for inter-birth interval. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):201–209, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Mitra:1973:SFA

[15] S. Mitra and A. Romaniuk. Simulation de la fécondité au moyen d’un modèle basé sur la courbe du type I de Pearson. (French) [Simulation of fertility using a model based on the Pearson curve of Type I]. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):211–229, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Yam:1973:GRP

[16] J. Yam. Graphical rational patterns (GRP) application to Canadian geostatistical data. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):231–246, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Winter:1973:SUC

[17] B. B. Winter. Strong uniform consistency of integrals of density estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):247–253, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Gupta:1973:TLV

[18] Rameshwar D. Gupta. On testing latent vectors of a normal covariance matrix. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):255–260, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Berger:1973:DHP

[19] Paul D. Berger. On the distribution of hand patterns in bridge: Mandealt versus computer-dealt. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):261–266, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

DuMouchel:1973:NEV

[20] W. H. DuMouchel, Z. Govindarajulu, and E. Rothman. A note on estimating the variance of the sample mean in stratified sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 1(1–2):267–274, ???? 1973. ISSN 0319-5724 (print), 1708-945X (electronic).

Fujikoshi:1974:ANN

[21] Yasunori Fujikoshi. On the asymptotic non-null distributions of the LR criterion in a general manova. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):1–12, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Shah:1974:EDP

[22] M. C. Shah and P. N. Rathie. Exact distribution of product of generalized F -variates. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):13–24, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Basu:1974:DSD

[23] A. K. Basu. On distinguishability of sets of distribution functions and non-parametric median problem. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):25–44, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Aggarwal:1974:DMD

[24] K. R. Aggarwal. DT and MDT designs and their applications as breeding experiments. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):61–73, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Park:1974:ILS

[25] Soo-Bin Bin Park. On indirect least squares estimation of a simultaneous equation system. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):75–82, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Levy:1974:SCI

[26] K. J. Levy and S. C. Narula. Shortest confidence intervals for the ratio of two normal variances. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):83–87, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Doss:1974:NML

[27] D. C. Doss. A note on multivariate linear exponential distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1– 2):89–94, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Ahsanullah:1974:CPF

[28] M. Ahsanullah and A. B. M. Lutful Kabir. A characterization of the power function distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):95–98, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Yassaee:1974:IDD

[29] Hedayat Yassaee. Inverted Dirichlet distribution and multivariate logistic distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):99–105, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Shklov:1974:BRB

[30] N. Shklov. Book review: A first course in probability and statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1– 2):127–128, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Kanazawa:1974:NPC

[31] Mitsuyo Kanazawa. A non-parametric classification rule for several multivariate populations. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):145–156, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Ghosh:1974:PMD

[32] Malay Ghosh. Probabilities of moderate deviations under m-dependence. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1– 2):157–168, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Nagambal:1974:RMP

[33] P. N. Nagambal and Derrick S. Tracy. Rules for multiplication of any polykay by polykay products of weight 5. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):169–180, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Dagenais:1974:SSE

[34] Marcel G. Dagenais. Small sample estimation of regression parameters in the three-variable linear model, with incomplete observations. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):181–195, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Tilquin:1974:BRM

[35] C. Tilquin and R. Cléroux. The block replacement model with inactivity periods and general cost structure. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):197–213, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Laughland:1974:SGM

[36] A. R. Laughland and K. P. K. Nair. Stochastic game models of brand switching. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):215–225, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Iman:1974:PSR

[37] Ronald L. Iman. A power study of a rank transform for the two-way classification model when interaction may be present. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):227–239, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Khatri:1974:NCS

[38] C. G. Khatri. A note on characterization of some laws by the symmetrical distribution of one linear function given another. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):241–247, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Duthie:1974:ENE

[39] Audrey I. Duthie. On the existence and non-existence of quintic designs. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1– 2):249–259, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Charalambides:1974:DSS

[40] Ch. A. Charalambides. The distributions of sufficient statistics of truncated generalized logarithmic series, Poisson and negative binomial distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):261–275, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Singh:1974:ESL

[41] Jagbir Singh. Estimation in a simple linear model using prior beliefs. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1– 2):277–283, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1974:ENE

[42] Anonymous. Editor’s notes. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):284, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic). See [9].

Shah:1974:BRB

[43] N. D. Shah. Book review: Symmetric functions in statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 2(1–2):285–286, ???? 1974. ISSN 0319-5724 (print), 1708-945X (electronic).

Rodden:1975:MVR

[44] B. E. Rodden and J. P. Landolt. The mean and variance of rank and exponential score product moment statistics in the non-circular case. Canadian Journal of Statistics = Revue canadienne de statistique, 3(1): 1–11, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Mathai:1975:BMT

[45] A. M. Mathai. Bounds for moments through a general orthogonal expansion in a pre-Hilbert space — I. Canadian Journal of Statistics = Revue canadienne de statistique, 3(1):13–34, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Kabe:1975:GMD

[46] D. G. Kabe. Generalized MANOVA double linear hypothesis with double linear restrictions. Canadian Journal of Statistics = Revue canadienne de statistique, 3(1):35–44, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Kapadia:1975:CRW

[47] Asha Seth Kapadia. The correlated random walk in continuous time with barriers. Canadian Journal of Statistics = Revue canadienne de statistique, 3(1):45–56, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Shuster:1975:DFA

[48] Jonathan J. Shuster and Mark C. K. Yang. A distribution-free approach to quantal response assays. Canadian Journal of Statistics = Revue canadienne de statistique, 3(1):57–70, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Yuen:1975:PCS

[49] Karen K. Yuen. Power comparisons of single-sample Winsorized T, trimmed T and Student’s T. Canadian Journal of Statistics = Revue canadienne de statistique, 3(1):71–80, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Chateau:1975:CEA

[50] Jean-Pierre D. Chateau. Contribution empirique au problème d’agrégation de fonctions autoregressives: Le cas de la politique de distribution de dividendes. (French) [Empirical contribution to the problem of aggregation of autoregressive functions: The case of dividend distribution policy]. Canadian Journal of Statistics = Revue canadienne de statistique, 3(1):81–104, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Scobey:1975:SGM

[51] P. Scobey. Singular Gauss-Markov models. Canadian Journal of Statistics = Revue canadienne de statistique, 3(1):105–110, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Basu:1975:LBI

[52] Sujit Kumar Basu and Bimal Kumar Sinha. Locally best invariant and locally minimax test of independence. Canadian Journal of Statistics = Revue canadienne de statistique, 3(1):111–118, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Sharma:1975:EMC

[53] V. K. Sharma. An easy method of constructing latin square designs balanced for the immediate residual and other order effects. Canadian Journal of Statistics = Revue canadienne de statistique, 3(1):119–124, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Colin:1975:LCS

[54] Bernard Colin. Loi conditionnelle de la statistique d’ordre 1 d’un échantillon: Application. (French) [Conditional distribution of statistical order 1 of a sample: Application]. Canadian Journal of Statistics = Revue canadienne de statistique, 3(1):125–132, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Ogawa:1975:BRB

[55] J. Ogawa. Book review: Statistical theory of the analysis of experimental designs. Canadian Journal of Statistics = Revue canadienne de statistique, 3(1):133–152, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Fujikoshi:1975:PDE

[56] Yasunori Fujikoshi. Partial differential equations for hypergeometric functions _{3}F_{2} of matrix argument. Canadian Journal of Statistics = Revue canadienne de statistique, 3(2):153–163, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Eaton:1975:MCM

[57] Morris L. Eaton and Michael D. Perlman. Monotonicity of certain multivariate normal probabilities. Canadian Journal of Statistics = Revue canadienne de statistique, 3(2):165–173, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Baskerville:1975:BTD

[58] J. C. Baskerville and D. L. Solomon. Bayes two-decision procedures for two-sample problems and rank-order data. Canadian Journal of Statistics = Revue canadienne de statistique, 3(2):187–202, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Clement:1975:EBE

[59] Bernard Clement. Estimation bayesienne des effets dans le modèle à effets aléatoires de classification double avec interaction. (French) [Bayesian estimation of the effects in the random-effects model with dualclassification interaction]. Canadian Journal of Statistics = Revue canadienne de statistique, 3(2):203–221, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Anderson:1975:MRF

[60] William J. Anderson. On a model for reciprocity failure in photographic materials. Canadian Journal of Statistics = Revue canadienne de statistique, 3(2):223–234, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Neuts:1975:SCP

[61] Marcel F. Neuts and Charles C. Carson. Some computational problems related to multinomial trials. Canadian Journal of Statistics = Revue canadienne de statistique, 3(2):235–248, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Wadycki:1975:CCM

[62] Walter J. Wadycki. Clustering of correlation matrices: a step-wise procedure. Canadian Journal of Statistics = Revue canadienne de statistique, 3(2):249–266, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Glick:1975:MSA

[63] Ned Glick. Measurements of separation among probability densities or random variables. Canadian Journal of Statistics = Revue canadienne de statistique, 3(2):267–276, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Wani:1975:LSI

[64] J. K. Wani and H. P. Lo. Large sample interval estimation for the logarithmic series distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 3(2):277–284, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Trivedi:1975:CGD

[65] H. T. Trivedi and V. K. Sharma. On the construction of group divisible incomplete block designs. Canadian Journal of Statistics = Revue canadienne de statistique, 3(2):285–288, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Mehta:1975:EMU

[66] J. S. Mehta and K. Raghunandanan. Estimation of the mean utilizing prior information. Canadian Journal of Statistics = Revue canadienne de statistique, 3(2):289–298, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Subrahmaniam:1975:MAS

[67] Kocherlakota Subrahmaniam and Kathleen Subrahmaniam. Multivariate analysis a selected and abstracted bibliography. 1957-1972. Canadian Journal of Statistics = Revue canadienne de statistique, 3(2):299–300, ???? 1975. ISSN 0319-5724 (print), 1708-945X (electronic).

Mathai:1976:BMT

[68] A. M. Mathai. Bounds for moments through a general orthogonal expansion in a pre-Hilbert space. II. Canadian Journal of Statistics = Revue canadienne de statistique, 4(1):1–12, November 1976. ISSN 0319-5724 (print), 1708-945X (electronic).

Tan:1976:TCB

[69] W. Y. Tan. On testing correlation between growth curves. Canadian Journal of Statistics = Revue canadienne de statistique, 4(1):13–32, November 1976. ISSN 0319-5724 (print), 1708-945X (electronic).

Ponnapalli:1976:DMD

[70] R. Ponnapalli. Deficiencies of minimum discrepancy estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 4(1):33–50, November 1976. ISSN 0319-5724 (print), 1708-945X (electronic).

Kenward:1976:NTR

[71] Lloyd R. Kenward. A new test for residual randomness in a class of dynamic autocorrelated econometric models. Canadian Journal of Statistics = Revue canadienne de statistique, 4(1):51–64, November 1976. ISSN 0319-5724 (print), 1708-945X (electronic).

Lampkin:1976:EDP

[72] H. Lampkin and J. Ogawa. Estimation of distribution parameters in time mortality trials: an example of time mortality analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 4(1):65–93, November 1976. ISSN 0319-5724 (print), 1708-945X (electronic).

Donner:1976:DCT

[73] Allan Donner. The design of a clinical trial when there are two categories of patients. Canadian Journal of Statistics = Revue canadienne de statistique, 4(1):95–108, November 1976. ISSN 0319-5724 (print), 1708-945X (electronic).

Chakravarty:1976:CBO

[74] Rupak Chakravarty and A. Dey. On the construction of balanced and orthogonal arrays. Canadian Journal of Statistics = Revue canadienne de statistique, 4(1):109–117, November 1976. ISSN 0319-5724 (print), 1708-945X (electronic).

Shah:1976:TES

[75] S. M. Shah and V. R. Rathod. Testing equality of several exponential distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 4(1):119–126, November 1976. ISSN 0319-5724 (print), 1708-945X (electronic).

Anderson:1976:NRB

[76] William J. Anderson. A note on reflecting boundaries for solutions of stochastic differential equations. Canadian Journal of Statistics = Revue canadienne de statistique, 4(1):127–131, November 1976. ISSN 0319-5724 (print), 1708-945X (electronic).

Kabe:1976:IMD

[77] D. G. Kabe. Inverse moments of discrete distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 4(1):133–141, November 1976. ISSN 0319-5724 (print), 1708-945X (electronic).

Boyarsky:1976:LBT

[78] Abraham Boyarsky. A lower bound for the transition density function of a stochastic differential equation. Canadian Journal of Statistics = Revue canadienne de statistique, 4(1):143–149, November 1976. ISSN 0319-5724 (print), 1708-945X (electronic).

Ali:1976:APO

[79] Mir M. Ali. An alternative proof of order statistics moment problem. Canadian Journal of Statistics = Revue canadienne de statistique, 4 (1):151–153, November 1976. ISSN 0319-5724 (print), 1708-945X (electronic).

Rohrer:1976:BRB

[80] Hermann Rohrer. Book review: Basic concepts in information theory and S axiomatic foundations and application. Canadian Journal of Statistics = Revue canadienne de statistique, 4(1):155–156, November 1976. ISSN 0319-5724 (print), 1708-945X (electronic).

Pillai:1977:DCR

[81] K. C. S. Pillai. Distribution of characteristic roots in multivariate analysis. Part II. Non-null distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 5(1):1–62, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Mathai:1977:NND

[82] A. M. Mathai and W. Y. Tan. The non-null distribution of the likelihood ratio criterion for testing the hypothesis that the covariance matrix is diagonal. Canadian Journal of Statistics = Revue canadienne de statistique, 5(1):63–74, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Beaman:1977:OEE

[83] J. Beaman, H. K. Cheung, and J. Knetsch. Obtaining efficient estimates of park use and testing for the structural adequacy of models. Canadian Journal of Statistics = Revue canadienne de statistique, 5(1):75–92, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Swamy:1977:RTM

[84] P. A. V. B. Swamy and J. S. Mehta. Robustness of Theil’s mixed regression estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 5(1):93–109, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Kadiyala:1977:ERE

[85] K. R. Kadiyala and K. S. R. Murthy. Estimation of regression equation with Cauchy disturbances. Canadian Journal of Statistics = Revue canadienne de statistique, 5(1):111–120, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Sharma:1977:CDC

[86] V. K. Sharma. Change-over designs with complete balance for first and second residual effects. Canadian Journal of Statistics = Revue canadienne de statistique, 5(1):121–132, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Becus:1977:RRP

[87] Georges A. Becus. The random Robin problem for Laplace’s equation. Canadian Journal of Statistics = Revue canadienne de statistique, 5(1): 133–139, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Deo:1977:DAM

[88] Chandrakant M. Deo. Delayed averages of mixing sequences. Canadian Journal of Statistics = Revue canadienne de statistique, 5(1):141–145, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Pederzoli:1977:NND

[89] G. Pederzoli. The non-null distribution for the problem of tests of independence. Canadian Journal of Statistics = Revue canadienne de statistique, 5(1):147–151, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Wani:1977:CCI

[90] J. K. Wani and H. P. Lo. Comparing confidence intervals for the logarithmic series distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):153–158, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Nair:1977:FPG

[91] K. P. K. Nair, A. R. Laughland, and S. Srinivasan. First passage game. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2): 159–166, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Govindarajulu:1977:CNC

[92] Z. Govindarajulu and A. K. Gupta. Certain nonparametric classification rules and their asymptotic efficiencies. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):167–178, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Schulman:1977:PMT

[93] R. S. Schulman and M. A. Hamdan. A probabilistic model for table tennis. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):179–186, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Iman:1977:PTS

[94] Ronald L. Iman and W. J. Conover. On the power of the t-test and some rank tests when outliers may be present. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):187–193, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Archambault:1977:SRT

[95] W. A. T. Archambault, Jr., G. A. Mack, and D. A. Wolfe. k-sample rank tests using pair-specific scoring functions. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):195–207, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Gupta:1977:MIT

[96] A. K. Gupta and D. S. Tracy. Moment inequalities for truncated multivariate distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):209–212, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Mohan:1977:NSR

[97] N. R. Mohan. A note on a strong renewal theorem. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):213–218, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Kohler:1977:CFE

[98] Boris Kohler. Convergence faible sur des espaces topologiques uniformes et suites aléatoires d’éléments aléatoires. (French) [Weak convergence in topological spaces of uniform and random sequences of random elements]. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2): 219–225, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Sia:1977:OSE

[99] L. L. Sia. Optimum spacings of elementary treatment contrasts in symmetrical 2-factor P A block designs. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):227–234, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Hamdan:1977:CTM

[100] M. A. Hamdan. Comparison of two measures of association in two-way contingency tables. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):235–240, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Tan:1977:DQF

[101] W. Y. Tan. On the distribution of quadratic forms in normal random variables. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):241–250, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Lepage:1977:BRB

[102] Yves Lepage. Book review: Sondage: Compilation et Corrélation. Par Adolf Diegel. Les Presses de l’Université du Québec, Montréal, Québec, 1977. xii + 162 pp., $9.95. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):263, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Dunnett:1977:BRB

[103] Charles W. Dunnett. Book review: The Analysis of Cross-Classified Categorical Data. Stephen E. Fienberg. The MIT Press, Cambridge, Mass., 1977. 151 pp., US $10.95. ISBN 0-262-06063-9. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):263–264, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Clarkson:1977:BRB

[104] Douglas B. Clarkson. Book review: Mathematical Models of Sociology. Sociological Review Monograph 24. Edited by P. Krishnan. University of Keele, Keele, Staffordshire, England, 1977. 229 pp., $9.00. ISBN 0-904425-03-7. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):264–265, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Stanley:1977:BRB

[105] Richard M. Stanley. Book review: Pocket Book of Statistical Tables. Statistics: Textbooks and Monographs Series, Volume 22. Compiled by Robert E. Odeh, Donald B. Owen, Z. W. Birnbaum, and Lloyd Fisher. Marcel Dekker, Inc., New York, 1977. x + 166 pp., US $8.75. ISBN 0-8247-6515-X. Canadian Journal of Statistics = Revue canadienne de statistique, 5(2):265–266, ???? 1977. ISSN 0319-5724 (print), 1708-945X (electronic).

Alagar:1978:DRD

[106] V. S. Alagar. The distribution of random determinants. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):1–9, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Dwyer:1978:CPT

[107] P. S. Dwyer, N. N. Mikhail, and D. S. Tracy. A concise proof of a theorem on products of power sums. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):11–17, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Rao:1978:LST

[108] C. Radhakrishna Rao. Least squares theory for possibly singular models. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1): 19–23, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Saha:1978:ICP

[109] G. M. Saha and A. D. Das. An infinite class of PBIB designs. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):25–32, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Deo:1978:CAG

[110] C. M. Deo and S. F. Wong. Champs aléatoires gaussiens. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):33–40, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Narayana:1978:YCR

[111] T. V. Narayana, I. Richard Savage, and K. M. Lal Saxena. Young chains and rank orders. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):41–47, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Konishi:1978:AED

[112] Sadanori Konishi. Asymptotic expansions for the distributions of statistics based on a correlation matrix. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):49–56, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Chan:1978:STL

[113] A. H. C. Chan, M. Csörgő, and P. Révész. Strassen type limit points for moving averages of a Wiener process. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):57–75, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Janardan:1978:ACW

[114] K. G. Janardan and D. J. Schaeffer. Another characterization of the Weibull distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):77–78, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Srivenkataramana:1978:COS

[115] T. Srivenkataramana. Change of origin and scale in ratio and difference methods of estimation in sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):79–86, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Kern:1978:BTL

[116] Malcolm Kern and Stanley Walter. Ballot theorem and lattice path crossings. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):87–90, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Deo:1978:RSH

[117] Chandrakant M. Deo. On a random series in a Hilbert space. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):91–94, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Hamilton:1978:CSC

[118] M. Hamilton and W. J. Anderson. A consistent system of conditional probabilities which is not compatible with any random field. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):95–101, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Venables:1978:IEV

[119] W. Venables and A. T. James. Interval estimates for variance components. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):103–111, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Khan:1978:NAH

[120] Rasul A. Khan. A note on the admissibility of Hammersley’s estimator of an integer mean. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):113–119, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Thompson:1978:BRB

[121] Mary E. Thompson. Book review: Foundations of inference in survey sampling. by Claes-Magnus Cassel, Carl-Erik Särndal, Jan Håkan Wretman, John Wiley & Sons, Inc., New York, 1977. xi + 192 pp., $21.85. ISBN 0-471-02563-1. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):133–134, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Broekhoven:1978:BRB

[122] Louis H. Broekhoven. Book review: Computational Methods for Data Analysis, by John M. Chambers. John Wiley & Sons, Inc., New York, 1977. xi + 268 pp., $18.40. ISBN 0-471-02772-3. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):134–135, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Mackay:1978:BRB

[123] R. J. Mackay. Book review: Probability and statistics: Theory and applications. by D. A. S. Fraser. Duxbury Press, North Scituate, Mass., 1976. xi + 623 pp., US $18.95. ISBN 0-87872-099-5. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):135–137, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Endrenyl:1978:BRB

[124] Laszlo Endrenyl. Book review: Statistics. by David Freedman, Robert Pisani, Roger Purves. W. W. Norton & Co., Inc., New York, 1978. xv + 506 + A83 pp., US $13.95. ISBN 0-393-09076-0. [Instructor’s Manual, 135 pp. ISBN 0-393-09041-8.]. Canadian Journal of Statistics = Revue canadienne de statistique, 6(1):137, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Dawson:1978:GC

[125] D. A. Dawson. Geostochastic calculus. Canadian Journal of Statistics = Revue canadienne de statistique, 6(2):143–168, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Tomkins:1978:CPC

[126] R. J. Tomkins. Convergence properties of conditional medians. Canadian Journal of Statistics = Revue canadienne de statistique, 6(2):169–177, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Devroye:1978:UCN

[127] Luc P. Devroye. The uniform convergence of the Nadaraya-Watson regression function estimate. Canadian Journal of Statistics = Revue canadienne de statistique, 6(2):179–191, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Steele:1978:IAS

[128] J. Michael Steele. Invalidity of average squared error criterion in density estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 6(2):193–200, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Singh:1978:TMS

[129] Ashok K. Singh. Testing multiple slippages. Canadian Journal of Statistics = Revue canadienne de statistique, 6(2):201–218, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Wani:1978:MDB

[130] Jagannath K. Wani. Measuring diversity in biological populations with logarithmic abundance distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 6(2):219–228, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Guttman:1978:ECE

[131] Irwin Guttman and G. C. Tiao. Effect of correlation on the estimation of a mean in the presence of spurious observations. Canadian Journal of Statistics = Revue canadienne de statistique, 6(2):229–247, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Koul:1978:CTT

[132] Hira L. Koul. A class of tests for testing “new is better than used”. Canadian Journal of Statistics = Revue canadienne de statistique, 6(2): 249–271, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Foutz:1978:ADL

[133] Robert V. Foutz and R. C. Srivastava. The asymptotic distribution of the likelihood ratio when the model is incorrect. Canadian Journal of Statistics = Revue canadienne de statistique, 6(2):273–279, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Tracy:1978:BRB

[134] Derrick S. Tracy. Book review: Foundations of Probability Theory, Statistical Inference, and Statistical Theories of Science. Proceedings of an International Research Colloquium held at the University of Western Ontario, London, Canada, 10–13 May 1973. Volume I: Foundations and Philosophy of Epistemic Applications of Probability Theory, xi + 308 pp.; Volume II: Foundations and Philosophy of Statistical Inference, xi + 455 pp.; Volume III: Foundations and Philosophy of Statistical Theories in the Physical Sciences, xi + 308 pp. edited by W. L. Harper, C. A. Hooker, D. Reidel Publishing Company, Dordrecht, The Netherlands, 1976. Set: US $105.00. ISBN 90-277-0614-X. Canadian Journal of Statistics = Revue canadienne de statistique, 6(2):281–283, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Reilly:1978:BRB

[135] Park M. Reilly. Book review: Probability and Statistics for Engineers and Scientists by Ronald E. Walpole, Raymond H. Myers. Second Edition. Macmillan Publishing Co., Inc., New York, 1978. xii + 580 pp. US $16.95. ISBN 0-02-424110-5. Canadian Journal of Statistics = Revue canadienne de statistique, 6(2):283–284, ???? 1978. ISSN 0319-5724 (print), 1708-945X (electronic).

Ong:1979:EFS

[136] B. G. Ong and L. L. Campbell. Estimation of frequencies of sinusoids in the presence of noise. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):11–19, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Brillinger:1979:APP

[137] David R. Brillinger. Analyzing point processes subjected to random deletions. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):21–27, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

DuBien:1979:MCM

[138] Janice L. DuBien and William D. Warde. A mathematical comparison of the members of an infinite family of agglomerative clustering algorithms. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1): 29–38, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Rizvi:1979:NMC

[139] M. Haseeb Rizvi and R. W. Shorrock. A note on matrix-convexity. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):39–41, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Koziol:1979:CMT

[140] James A. Koziol and Amanda F. Nemec. On a Cramér–von Mises type statistic for testing bivariate independence. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):43–52, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Giri:1979:LMT

[141] N. Giri. Locally minimax tests for multiple correlations. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):53–60, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Patil:1979:AMD

[142] G. P. Patil and M. V. Ratnaparkhi. On additive and multiplicative damage models and the characterizations of linear and logarithmic exponential families. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):61–64, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Henderson:1979:OMS

[143] Harold V. Henderson and S. R. Searle. Vec and vech operators for matrices, with some uses in Jacobians and multivariate statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):65–81, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Weerahandi:1979:CGM

[144] Samaradasa Weerahandi and James V. Zidek. A characterization of the general mean. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):83–90, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Gupta:1979:GFT

[145] R. D. Gupta and D. G. Kabe. Goodness-of-fit tests of hypothesized canonical variables. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):91–96, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Draper:1979:IFO

[146] Norman R. Draper and Agnes M. Herzberg. An investigation of firstorder and second-order designs for extrapolation outside a hypersphere. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1): 97–101, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Dunnett:1979:BRBa

[147] Charles W. Dunnett. Book review: A bibliography of multivariate statistical analysis, T. W. Anderson, Somesh Das Gupta, George P. H. Styan, Robert E. Krieger Publishing Co., Huntington, New York, 1977. x + 642 pp. US $28.50. ISBN 0-88275-477-7. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):111, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Anderson:1979:BRB

[148] Gary D. Anderson. Book review: BMDP-77, biomedical computer programs, P-series. by W. J. Dixon, series editor, and M. B. Brown editor 1977 edition. Systems, program and statistical development: L. Engelman, J. W. Frane and R. I. Jennrich. University of California Press, Berkeley, 1977. xiii + 880 pp. C $16.20, US $10.50. ISBN 0-520-03569-0. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1): 111–113, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Ahmad:1979:BRB

[149] Ibrahim Ahmad and M. S. Srivastava. Book review: Multivariate statistical inference, by Narayan C. Giri academic press, 1977. xvi + 319 pp. C $34.20, US $25.00. ISBN 0-12-285650-3. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):113–115, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Macdonald:1979:BRB

[150] D. M. Macdonald. Book review: Branching processes, Anatole Joffe, Peter Ney. Marcel Dekker Inc., New York, 1978. x + 322 pp. US $34.50. ISBN 0-8247-6800-0. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):115–116, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Styan:1979:BRB

[151] George P. H. Styan. Book review: International encyclopedia of statistics, William H. Kruskal, Judith M. Tanur. The free press, New York, and Collier Macmillan, London and Toronto, 1978. Volume 1, A–N, and volume 2, O–Z and index, xxi + 1350 pp. $100.00. ISBN 0-02-917960-2. [16 pp. Prospectus available free from the publishers.]. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):116–117, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Santalo:1979:BRB

[152] Luis A. Santaló. Book review: Géométrie différentielle stochastique, par Paul Malliavin Notes rédigées par Danièle Dehen et Dominique Michel. Séminaire de mathématiques supérieures, Département de mathématiques et de statistique, Université de Montréal, vol. 64. Les presses de l’université de Montréal, Montréal, Québec, 1978. 144 pp. $8.00. ISBN 0-8405-0425-X. Canadian Journal of Statistics = Revue canadienne de statistique, 7(1):117–118, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Draper:1979:DGA

[153] Norman R. Draper and Agnes M. Herzberg. Designs to guard against outliers in the présence or absence of model bias. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):127–135 (1980), ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Shillington:1979:TLF

[154] E. Richard Shillington. Testing lack of fit in regression without replication. Canadian Journal of Statistics = Revue canadienne de statistique, 7 (2):137–146, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Moore:1979:EDR

[155] Marc Moore and Maurice Legault. Sur l’estimation du risque encouru en reconstruisant une forme aléatoire. (French) [On the estimation of risk in rebuilding a random form]. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):147–158, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Devroye:1979:REM

[156] Luc Devroye. Recursive estimation of the mode of a multivariate density. Canadian Journal of Statistics = Revue canadienne de statistique, 7 (2):159–167 (1980), ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Donner:1979:DCT

[157] Allan Donner. The design of a clinical trial with several patient categories. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):169–176, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Ranganathan:1979:THR

[158] J. Ranganathan and B. K. Kale. Tests of hypotheses for reliability functions in two-parameter exponential models. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):177–184 (1980), ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Ramasubban:1979:CAR

[159] T. A. Ramasubban. A comparison of alternative rate estimators in some frequently used growth models. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):185–192, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Tsui:1979:MED

[160] Kam-Wah Wah Tsui. Multiparameter estimation of discrete exponential distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):193–200 (1980), ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Tsui:1979:EPM

[161] Kam-Wah Wah Tsui. Estimation of Poisson means under weighted squared error loss. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):201–204 (1980), ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Sterling:1979:SPM

[162] Theodor D. Sterling. Statistical problems in multi-variate (etiological) surveys. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):205–215, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Plante:1979:VFT

[163] André Plante. On the validation of fiducial techniques. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):217–226 (1980), ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Plante:1979:SPS

[164] André Plante. Structured probability statements. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):227–232 (1980), ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Dunnett:1979:BRBb

[165] Charles W. Dunnett. Book review: Proceedings of the computer science and statistics: 12th annual symposium on the interface. Edited by Jane F. Gentleman. University of Waterloo, Waterloo, Ontario, 1979. xiii + 500 pp. C $25.00, US $22.50. (Order from Jane F. Gentleman. Department of Statistics, University of Waterloo, Waterloo, Ontario N2L 3G1.). Canadian Journal of Statistics = Revue canadienne de statistique, 7(2): 233, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Ali:1979:BRB

[166] Mir Maswood Ali. Book review: Characterizations of the normal probability law, by A. M. Mathai and G. Pederzoli. Wiley Eastern, New Delhi, and Halsted Press, New York, 1977. x + 149 pp, US $8.50, f3.50. ISBN 0-470-99322-7. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):233–234, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Dunnett:1979:BRBc

[167] Charles W. Dunnett. Book review: Applied statistics. by John Neter, William Wasserman and G. A. Whitmore. Allyn and Bacon, Inc., Boston, 1978. xix + 743 pp., $23.50. ISBN 0-205-05982-1. Solutions Manual and Guide, iii + 143 pp. ISBN 0-205-06038-2. Test Manual by Kenneth Schneider, Randall Byers and Daniel Brick, [4] + 130 pp. ISBN 0-205-05983-X. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):234–235, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Dunnett:1979:BRBd

[168] Charles W. Dunnett. Book review: Selected Tables in Mathematical Statistics, Volume V. Edited by D. B. Owen and R. E. Odeh for The Institute of Mathematical Statistics. American Mathematical Society, Providence, Rhode Island, 1977. viii + 263 pp. US $16.80. ISBN 0-8218-1905-4.

Canadian Journal of Statistics = Revue canadienne de statistique, 7(2): 235, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Chorneyko:1979:BRB

[169] I. Chorneyko. Book review: A primer in probability. By Kathleen Subrahmaniam. Statistics: Textbooks and monographs, vol. 28. Marcel Dekker, Inc., New York, 1979. 344 pp. US $12.50. ISBN 0-8247-6836-1. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):236–237, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1979:BRB

[170] Anonymous. Book review: Introductory statistics. by Thomas H. Wonnacott and Ronald J. Wonnacott. third edition. John Wiley & Sons, Inc., New York, 1977. xxii + 650 pp. US $18.95. ISBN 0-471-95982-0. Canadian Journal of Statistics = Revue canadienne de statistique, 7(2):237, ???? 1979. ISSN 0319-5724 (print), 1708-945X (electronic).

Maag:1980:FHM

[171] Urs R. Maag. The first honorary members of the Statistical Society of Canada. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):6, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Tardif:1980:ADC

[172] Serge Tardif. On the asymptotic distribution of a class of aligned rank order test statistics in randomized block designs. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):7–25, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

VanDerMerwe:1980:MRA

[173] A. Van Der Merwe and J. V. Zidek. Multivariate regression analysis and canonical variates. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):27–39, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Feuerverger:1980:CIS

[174] A. Feuerverger and D. A. S. Fraser. Categorical information and the singular linear model. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):41–45, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Kadiyala:1980:SFS

[175] Krishna Kadiyala. Some finite sample properties of generalized ridge regression estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):47–58, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Chikkagoudar:1980:EME

[176] M. S. Chikkagoudar and S. H. Kunchur. Estimation of the mean of an exponential distribution in the presence of an outlier. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):59–63, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Raktoe:1980:AGD

[177] B. L. Raktoe, H. Pesotan, and W. T. Federer. On aliasing and generalized defining relationships of factorial arrangements. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):65–77, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Evans:1980:CTR

[178] M. Evans. γ-confidence tolerance regions. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):79–85, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Hill:1980:MLE

[179] David L. Hill, Roy Saunders, and Purushottam W. Laud. Maximum likelihood estimation for mixtures. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):87–93, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Wild:1980:LFA

[180] C. J. Wild. Loss functions and admissibility of normal variance estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):95–101, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Wong:1980:MMM

[181] Chi Song Wong and Kai Sang Wong. Minima and maxima in multivariate analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):103–113, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

McDonald:1980:PAM

[182] David R. McDonald. On the Poisson approximation to the multinomial distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):115–118, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Carter:1980:ADL

[183] E. M. Carter and M. S. Srivastava. Asymptotic distribution of the latent roots of the noncentral Wishart distribution and the power of the likelihood ratio test for nonadditivity. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):119–134, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Wets:1980:BRB

[184] Roger Wets. Book review: Problèmes d’optimisation en calcul des probabilités, par Serge Dubuc. Séminaire de Mathématiques Supérieures, vol.

62. Les Presses de l’Université de Montréal, Québec, 1978. 137 pp. $8.00. ISBN 0-8405-0413-6. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):141, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Raktoe:1980:BRB

[185] B. Leo Raktoe. Book review: Orthogonal designs: Quadratic forms and Hadamard matrices. by Anthony V. Geramita and Jennifer Seberry. Lecture Notes in Pure and Applied Mathematics Series, vol. 45. Marcel Dekker, Inc. New York, 1979. 480 pp. US $35.00. ISBN 0-8247-6774-8. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1): 141–143, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Dubuc:1980:BRB

[186] Critique Par Serge Dubuc. Book review: Chaînes de Markov sur les permutations, par Gérard Letac. séminaire de mathématiques supérieures, vol. 63. les presses de l’université de Montréal, Montréal, Québec, 1978. 117 pp. $8.00. ISBN 0-8405-0412-8. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):143, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Ismail:1980:BRB

[187] Mourad E. H. Ismail. Book review: The H-function with applications in statistics and other disciplines. by A. M. Mathai and R. K. Saxena. Wiley Eastern, New Delhi, and Halsted Press, New York, 1978. xii + 192 pp. US $9.95. ISBN 0-470-26380-6. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):143–145, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Broekhoven:1980:BRB

[188] L. H. Broekhoven. Book review: Compact numerical methods for computers: Linear algebra and function minimisation. by J. C. Nash. Adam Hilger Ltd., Bristol, 1979. ix + 227 pp. £12.50. ISBN 0-85274-330-0. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1): 145–146, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Dunnett:1980:BRB

[189] Charles W. Dunnett. Book review: Tables for normal tolerance limits, sampling plans, and screening. by Robert E. Odeh and D. B. Owen. Statistics: Textbooks and monographs, vol. 32. Marcel Dekker, Inc., New York, 1980. ix + 316 pp. US $49.75. ISBN 0-8247-6944-9. Canadian Journal of Statistics = Revue canadienne de statistique, 8(1):146–148, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Hinkley:1980:L

[190] David V. Hinkley. Likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):151–163 (1981), ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Sarndal:1980:TWC

[191] Carl-Erik Särndal. A two-way classification of regression estimation strategies in probability sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):165–177 (1981), ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Ivanoff:1980:FS

[192] B. Gail Ivanoff. The function space D([0, ∞)^{q}, E). Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):179–191 (1981), ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Scholz:1980:TUD

[193] F.-W. W. Scholz. Towards a unified definition of maximum likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2): 193–203 (1981), ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Sherif:1980:BSM

[194] F. Sherif and Mary E. Thompson. A brand share model for Canadian cigarettes. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):205–215, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Smith:1980:ERR

[195] Gary Smith. An example of ridge regression difficulties. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):217–225, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Sharma:1980:HFT

[196] Shyam S. Sharma. On Hacking’s fiducial theory of inference. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):227–233 (1981), ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Basu:1980:RCI

[197] A. K. Basu. On rates of convergence to infinitely divisible laws for dependent random variables. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):235–247 (1981), ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Srivastava:1980:EED

[198] M. S. Srivastava. Effect of equicorrelation in detecting a spurious observation. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):249–251 (1981), ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Srivastava:1980:CEU

[199] Surendra K. Srivastava. A class of estimators using auxiliary information in sample surveys. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):253–254 (1981), ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Sharma:1980:BRB

[200] Shashi N. Sharma. Book review: Perspectives Canada III, edited by H. J. Adler and D. Brusegard. Statistics Canada, Ottawa, Ontario, 1980. viii + 312 pp. C $9.95. ISBN 0-660-90202-8. [Catalogue No. 11-511E, Supply and Services Canada, Ottawa, Ontario K1A OS9.1]. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):255–256, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Fellegi:1980:BRB

[201] Ivan P. Fellegi. Book review: Privacy and government data banks: an internationai perspective. By David H. Flaherty. Mansell Publishing, London, 1979. 352 pp. f17.50, US. $37.00. ISBN 0-7201-0930-2. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):256–258, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Enns:1980:BRB

[202] Ernest G. Enns. Book review: Probability and statistics. by A. M. Mathai and P. N. Rathie. The Macmillan Company of India Ltd., Bombay, 1977. xv + 406 pp. f4.95, US $13.85, C$20.85, Rs.48.00. ISBN 0-333-19056-4, ISBN 0-333-90115-0. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):258–259, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Stanton:1980:BRB

[203] Ralph G. Stanton. Book review: Lattice path combinatorics with statistical applications. by T. V. Narayana. Mathematical Expositions No. 23, University of Toronto Press, Toronto, Ontario, 1979. xii + 106 pp. C $15.00. ISBN 0-8020-5405-6. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):259, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Joe:1980:BRB

[204] Harry Joe. Book review: Life Testing and Reliability Estimation. By S. K. Sinha and 6. K. Kale. Wiley Eastern, New Delhi, and Halsted Press, New York, 1980. vii + 196 pp. US $21.20, $4.50. ISBN 0-470-26911-1. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2): 259–260, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Callen:1980:BRB

[205] Jeffrey L. Callen. Book review: Stochastic dominance: an approach to decision-making under risk. edited by G. A. Whitmore and M. C. Findlay. Lexington Books (D. C. Heath and Co.), Lexington, Mass., 1978. xiv + 398 pp. C $22.00. ISBN 0-669-01310-2. Canadian Journal of Statistics = Revue canadienne de statistique, 8(2):260–261, ???? 1980. ISSN 0319-5724 (print), 1708-945X (electronic).

Kiefer:1981:IOC

[206] Jack C. Kiefer. The interplay of optimality and combinatorics in experimental design. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1):1–10, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Zwiers:1981:AMA

[207] F. W. Zwiers. Autoregressive moving-average convolution processes on the unit circle. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1):11–25, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Ghurye:1981:DAI

[208] S. G. Ghurye and B. R. Johnson. Discrete approximations to the information integral. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1):27–37, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Schaufele:1981:CTR

[209] Ronald A. Schaufele. A class of time-reversible semi-Markov processes. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1): 39–56, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Csaki:1981:EMR

[210] E. Csáki and S. G. Mohanty. Excursion and meander in random walk. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1): 57–70, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Dyer:1981:SPB

[211] Danny Dyer. Structural probability bounds for the strong Pareto law. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1): 71–77, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Holst:1981:AST

[212] Lars Holst and J. S. Rao. Asymptotic spacings theory with applications to the two-sample problem. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1):79–89, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Tan:1981:PFA

[213] W. Y. Tan. The power function and an approximation for testing variance components in the presence of interaction in two-way random effects models. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1):91–99, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Csorgo:1981:APR

[214] M. Csörgő and Z. Rychlik. Asymptotic properties of randomly indexed sequences of random variables. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1):101–107, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Basilevsky:1981:FAR

[215] Alexander Basilevsky. Factor analysis regression. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1):109–117, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Chaubey:1981:TEV

[216] Y. P. Chaubey. Testing the equality of the variances of two linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1): 119–127, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Ferguson:1981:BRC

[217] D. G. Ferguson. Book review/comptes rendus critiques: Demand Functions and the Slutsky Matrix. By S. N. Afriat. Princeton studies in mathematical economics. Princeton University Press, Princeton, New Jersey, 1980. xii + 269 pp. US $16.50. ISBN 0-691-04222-5. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1):129–130, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Kocherlakota:1981:BRC

[218] Kathleen Kocherlakota. Book review/comptes rendus critiques: An introduction to applied probability. By Ian F. Blake. John Wiley & Sons, Inc., New York, 1979. [16] + 528 pp. US $17.10. ISBN 0-471-03210-7. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1): 130–131, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1981:BRCa

[219] Anonymous. Book review/comptes rendus critiques: Linguistic variation: Models and methods. Edited by David Sankoff. Academic Press, New York, 1978. xvii + 296 pp. US $22.50. ISBN 0-12-618850-5. Review by H. J, Warkentyne, Dept. of Linguistics, University of Victoria. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1): 131–133, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1981:BRCb

[220] Anonymous. Book review/comptes rendus critiques: Regression: a second course in statistics. By Thomas H. Wonnacott and Ronald J. Wonnacott. John Wiley & Sons, Inc., New York, 1981. xix + 556 pp. US $26.95, £12.10. ISBN 0-471-95974-X. Review by William J. Reed, Dept. of Mathematics, University of Victoria. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1):133–135, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Davidson:1981:SRRa

[221] Roger R. Davidson. Summary–reviews/résumés bibliographiques: Measures of association for cross classifications. By Leo A. Goodman and William H. Kruskal. (With a foreword by Stephen E. Fienberg.) Springer series in statistics, vol. 1. Springer-Verlag, New York, 1979. x + 146 pp. US $12.00. ISBN 0-387-90443-3. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1):135–136, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Davidson:1981:SRRb

[222] Roger R. Davidson. Summary–Reviews/Résumés bibliographiques: Selecta statistica canadiana, volume V: Mathematical statistics. Edited by M. Behara, C. Blyth, L. L. Campbell, M. Csörgő;, D. A. Dawson, N. Giri, A. Joffe, R. G. Stanton, and R. Theodorescu. Dept. of mathematical sciences, McMaster University, Hamilton, Ontario, 1979. 150 pp. C $50.00 (C $40.00 for individuals, C $20.00 for SSC members). ISBN 0-920788-04-1. Canadian Journal of Statistics = Revue canadienne de statistique, 9(1):136, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Efron:1981:NSE

[223] Bradley Efron. Nonparametric standard errors and confidence intervals. Canadian Journal of Statistics = Revue canadienne de statistique, 9(2): 139–158, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic). See discussion [224] and reply [225].

Barnard:1981:D

[224] George A. Barnard. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 9(2):158–170, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic). See [223].

Efron:1981:R

[225] Bradley Efron. Rejoinder. Canadian Journal of Statistics = Revue canadienne de statistique, 9(2):170–172, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic). See [223].

Brillinger:1981:SAM

[226] David R. Brillinger. Some aspects of modern population mathematics. Canadian Journal of Statistics = Revue canadienne de statistique, 9(2): 173–194, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Wild:1981:IEV

[227] C. J. Wild. Interval estimates for variance components. Canadian Journal of Statistics = Revue canadienne de statistique, 9(2):195–201, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Herzberg:1981:CES

[228] Agnes M. Herzberg and S. Huda. A comparison of equally spaced designs with different correlation structures in one and more dimensions. Canadian Journal of Statistics = Revue canadienne de statistique, 9(2): 203–208, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Johnson:1981:SSP

[229] B. R. Johnson. The search for submarine pingos in the Beaufort Sea: an unusual application of the binomial distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 9(2):209–214, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Haff:1981:FIW

[230] L. R. Haff. Further identities for the Wishart distribution with applications in regression. Canadian Journal of Statistics = Revue canadienne de statistique, 9(2):215–224, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Takada:1981:UTL

[231] Yoshikazu Takada. Upper tolerance limit for the largest observation from censored samples. Canadian Journal of Statistics = Revue canadienne de statistique, 9(2):225–228, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Forst:1981:MCS

[232] Frederick R. Forst and Mir Masoom Ali. Monte Carlo studies of some adaptive robust procedures for location. Canadian Journal of Statistics = Revue canadienne de statistique, 9(2):229–235, ???? 1981. ISSN 0319-5724 (print), 1708-945X (electronic).

Dufour:1982:NTT

[233] Jean-Marie Dufour, Yves Lepage, and Hanna Zeidan. Nonparametric testing for time series: a bibliography. Canadian Journal of Statistics = Revue canadienne de statistique, 10(1):1–38, March 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Wilmesmeier:1982:CRS

[234] James M. Wilmesmeier and F. T. Wright. Convergence results for sequences of quadratic forms. Canadian Journal of Statistics = Revue canadienne de statistique, 10(1):39–47, March 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Cheng:1982:JE

[235] Kuang-Fu Fu Cheng. Jackknifing L-estimates. Canadian Journal of Statistics = Revue canadienne de statistique, 10(1):49–58, March 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Muirhead:1982:DLR

[236] Robb J. Muirhead. On the distribution of the likelihood ratio test of equality of normal populations. Canadian Journal of Statistics = Revue canadienne de statistique, 10(1):59–62, March 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1982:BRBa

[237] Anonymous. Book review: Time series: Data analysis and theory. Expanded Edition. By David R. Brillinger. Holden-Day, Inc., San Francisco, 1981. xii + 540 pp. US $28.95. ISBN 0-8162-1150-7. Canadian Journal of Statistics = Revue canadienne de statistique, 10(1):63, March 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1982:BRBb

[238] Anonymous. Book review: The statistical analysis of failure time data. By J. D. Kalbfleisch and R. L. Prentice. John Wiley & Sons, Inc., New York, 1980. xi + 321 pp. US $31.50, C $40.35. ISBN 0-471-05519-0. Canadian Journal of Statistics = Revue canadienne de statistique, 10 (1):64–66, March 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1982:BRBc

[239] Anonymous. Book review: Lattice path counting and applications. By Sri Gopal Mohanty. Academic Press, Inc., New York, 1979. xi + 185 pp. US $24.00. ISBN 0-12-504050-4. Canadian Journal of Statistics = Revue canadienne de statistique, 10(1):66–67, March 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1982:BRBd

[240] Anonymous. Book review: Analysis of categorical data: Dual scaling and its applications. By Shizuhiko Nishisato. Mathematical Expositions No. 24, University of Toronto Press, Toronto, Ontario, 1980. xiii + 276 pp. C$25.00. ISBN 0-8020-5489-7, ISSN 0076-5333. Canadian Journal of Statistics = Revue canadienne de statistique, 10(1):67–69, March 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Anderson:1982:QFT

[241] Keaven Anderson, Milton Sobel, and V. R. R. Uppuluri. Quota fulfillment times. Canadian Journal of Statistics = Revue canadienne de statistique, 10(2):73–88, June 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

McLeish:1982:RAN

[242] D. L. McLeish. A robust alternative to the normal distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 10(2):89–102, June 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Campbell:1982:PID

[243] Gregory Campbell and Myles Hollander. Prediction intervals with a Dirichlet-process prior distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 10(2):103–111, June 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Cheng:1982:BET

[244] Kuang-Fu Fu Cheng. On a Berry–Esséen theorem for a Studentized jackknife L-estimate. Canadian Journal of Statistics = Revue canadienne de statistique, 10(2):113–119, June 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Mak:1982:EPI

[245] Tak K. Mak. Estimation in the presence of incidental parameters. Canadian Journal of Statistics = Revue canadienne de statistique, 10(2):121– 132, June 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Herzberg:1982:DEC

[246] Agnes M. Herzberg. The design of experiments for correlated error structures: Layout and robustness. Canadian Journal of Statistics = Revue canadienne de statistique, 10(2):133–138, June 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Alvo:1982:BED

[247] Mayer Alvo and Paul Cabilio. Bayesian estimation of the difference between two proportions. Canadian Journal of Statistics = Revue canadienne de statistique, 10(2):139–145, June 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Baggs:1982:EIC

[248] Ivan Baggs. Enrollment in initial courses in statistics: an exploratory study. Canadian Journal of Statistics = Revue canadienne de statistique, 10(2):147–153, June 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Brenner:1982:FSM

[249] D. Brenner and D. A. S. Fraser. On the foundations: Statistical models and inference. Canadian Journal of Statistics = Revue canadienne de statistique, 10(3):155–161, September 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Brenner:1982:ANL

[250] D. Brenner, D. A. S. Fraser, and Philip McDunnough. On asymptotic normality of likelihood and conditional analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 10(3):163–172, September 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Lefebvre:1982:AMS

[251] Mario Lefebvre. Une application des méthodes séquentielles aux tests de permutations. (French) [An application of sequential methods to permutation tests]. Canadian Journal of Statistics = Revue canadienne de statistique, 10(3):173–180, September 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Keilson:1982:USO

[252] Julian Keilson and Ushio Sumita. Uniform stochastic ordering and related inequalities. Canadian Journal of Statistics = Revue canadienne de statistique, 10(3):181–198, September 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Egerton:1982:EFR

[253] M. F. Egerton and P. J. Laycock. An explicit formula for the risk of James-Stein estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 10(3):199–205, September 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Wani:1982:STS

[254] J. K. Wani and G. A. Watterson. Sampling theory for species abundances in certain biological populations. Canadian Journal of Statistics = Revue canadienne de statistique, 10(3):207–211, September 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Bellhouse:1982:NFL

[255] David R. Bellhouse. The need for a Federal Lotteries Review Board. Canadian Journal of Statistics = Revue canadienne de statistique, 10 (3):213–217, September 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Rivest:1982:PRV

[256] Louis-Paul Rivest. Products of random variables and star-shaped ordering. Canadian Journal of Statistics = Revue canadienne de statistique, 10 (3):219–223, September 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Brainerd:1982:NOT

[257] Barron Brainerd and Sun Man Chang. Number of occurrences in twostate Markov chains, with an application in linguistics. Canadian Journal of Statistics = Revue canadienne de statistique, 10(3):225–231, September 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1982:CSD

[258] Anonymous. Case studies in data analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 10(4):237–259, ???? 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Hawkins:1982:EGD

[259] Douglas M. Hawkins and E. Liefde Raath. An extension of Geisser’s discrimination model to proportional covariance matrices. Canadian Journal of Statistics = Revue canadienne de statistique, 10(4):261–270, ???? 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Dion:1982:SSR

[260] Jean-Pierre Dion, Gilbert Labelle, and Alain Latour. Small-sample results for maximum-likelihood estimation in branching processes. Canadian Journal of Statistics = Revue canadienne de statistique, 10(4):271– 276, ???? 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Dahiya:1982:SCP

[261] Ram C. Dahiya and Irwin Guttman. Shortest confidence and prediction intervals for the log-normal. Canadian Journal of Statistics = Revue canadienne de statistique, 10(4):277–291, ???? 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Schulzer:1982:PSQ

[262] Michael Schulzer, Moira Chan-Yeung, and Felisa Tan. On the possible significance of the quadratic effect of age on lung-function measurements. Canadian Journal of Statistics = Revue canadienne de statistique, 10(4): 293–303, ???? 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Casella:1982:LER

[263] George Casella and Jiunn Tzon Hwang. Limit expressions for the risk of James-Stein estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 10(4):305–309, ???? 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Blais:1982:CSP

[264] Michel Blais. Condition suffisante pour que l’entropie différentielle soit < ∞. (French) [Sufficient condition for differential entropy < ∞]. Canadian Journal of Statistics = Revue canadienne de statistique, 10(4):311–313, ???? 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1982:BRC

[265] Anonymous. Book review/comptes rendus critiques: Deterministic mathematical models in population ecology. By H. I.Freedman, The University of Alberta. Pure and Applied Mathematics: a series of monographs and textbooks, volume 57. Marcel Dekker, Inc., New York, 1980. x + 254 pp. US $29.75. ISBN 0-8247-6653-9. Canadian Journal of Statistics = Revue canadienne de statistique, 10(4):315, ???? 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1982:BRBe

[266] Anonymous. Book review: Stochastic systems for management. by Winfried K. Grassmann, University of Saskatchewan. Elsevier North Holland, Inc., New York, 1981. xii + 358 pp. US $27.95. ISBN 0-444-00449-1. Canadian Journal of Statistics = Revue canadienne de statistique, 10 (4):315–316, ???? 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1982:BRBf

[267] Anonymous. Book review: Statistical models and methods for lifetime data. By J. F. Lawless. John Wiley & Sons, New York, 1982. xi + 580 pp. C $50.95, US $36.95. ISBN 0471-08544-8. Canadian Journal of Statistics = Revue canadienne de statistique, 10(4):316–317, ???? 1982. ISSN 0319-5724 (print), 1708-945X (electronic).

Eynon:1983:VRA

[268] Barrett P. Eynon and Paul Switzer. The variability of rainfall acidity. Canadian Journal of Statistics = Revue canadienne de statistique, 11(1): 11–23, March 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Dawson:1983:D

[269] Donald A. Dawson. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 11(1):23–24, March 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Kalbfleisch:1983:DRM

[270] John D. Kalbfleisch, Daniel R. Krewski, and John Van Ryzin. Doseresponse models for time-to-response toxicity data. Canadian Journal of Statistics = Revue canadienne de statistique, 11(1):25–46, March 1983. ISSN 0319-5724 (print), 1708-945X (electronic). With discussion by V. T. Farewell and J. F. Lawless.

Farewell:1983:D

[271] V. T. Farewell. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 11(1):46–49, March 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Ware:1983:SPE

[272] James H. Ware and Thomas A. Louis. Statistical problems in environmental research. Canadian Journal of Statistics = Revue canadienne de statistique, 11(1):51–68, March 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Bray:1983:D

[273] David F. Bray. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 11(1):68–70, March 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Dagum:1983:SPC

[274] Estela Bee Dagum. Spectral properties of the concurrent and forecasting seasonal linear filters of the X-11-ARIMA method. Canadian Journal of Statistics = Revue canadienne de statistique, 11(1):73–90, March 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Kraft:1983:SFS

[275] Charles H. Kraft, Yves Lepage, and Constance Van Eeden. Some finitesample-size properties of Rosenblatt density estimates. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):95–104, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Wong:1983:CIS

[276] H. S. F. Wong. Classes inférieures de suites gaussiennes asymptotiquement indépendantes. (French) [Subclasses of Gaussian asymptoticallyindependent sequences]. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):105–108, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Allen:1983:APM

[277] O. Brian Allen. Asymptotic properties of the maximum-likelihood estimator for a class of birth-and-death processes admitting a unique stationary distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):109–118, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Baksalary:1983:EBB

[278] J. K. Baksalary and R. Kala. On equalities between BLUEs, WLSEs, and SLSEs. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):119–123, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Hudon:1983:TUM

[279] Georges Hudon. Tests unilatéraux multivariés. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):125–129, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Seshadri:1983:IGD

[280] V. Seshadri. The inverse Gaussian distribution: Some properties and characterizations. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):131–136, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Padgett:1983:SBL

[281] W. J. Padgett and M. P. Johnson. Some Bayesian lower bounds on reliability in the lognormal distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):137–147, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Solomon:1983:ADS

[282] Herbert Solomon and Michael A. Stephens. An approximation to the distribution of the sample variance. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):149–154, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Johnson:1983:SPE

[283] Richard A. Johnson and James H. Haskell. Sampling properties of estimators of a Weibull distribution of use in the lumber industry. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):155–169, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Shklov:1983:BRC

[284] Norman Shklov. Book review/comptes rendus critiques: Design and analysis of experiments. By M. N. Das and N. C. Giri. Wiley Eastern Limited, New Delhi, 1979, and Halsted Press, New York, 1981. ix + 295 pp. US$14.95, $4.50. ISBN 0-85226-158-6, 0-470-26861-1. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):171, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Horita:1983:BRC

[285] R. E. Horita. Book review/comptes rendus critiques: Time sequence analysis in geophysics. Third Edition. By E. R. Kanasewich. The University of Alberta Press, Edmonton, Alberta, 1981. xiii + 480 pp. C$30.00. ISBN 0-88864-074-9. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):171–172, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Solomon:1983:BRC

[286] Daniel L. Solomon. Book review/comptes rendus critiques: Inequalities: Theory of majorization and its applications. By Albert W. Marshall and Ingram Olkin. Volume 143 in the Mathematics in Science and Engineering Series. Academic Press, Inc., New York, 1979. xx + 569 pp. ISBN 0-12-473750-1. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):172–173, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Kocherlakota:1983:BRC

[287] S. Kocherlakota. Book review/comptes rendus critiques: An introduction to multivariate statistics. By M. S. Srivastava and C. G. Khatri. Elsevier North Holland, Inc., New York, 1979. xvi + 350 pp. US $19.50. ISBN 0-444-00302-9. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):173–175, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1983:BRCa

[288] Anonymous. Book review/comptes rendus critiques: Statistics: Discovering its power. By Ronald J. Wonnacott and Thomas H. Wonnacott. John Wiley and Sons, Inc., New York, 1982. xviii + 378 pp. US $22.95. ISBN 0-471-01412-5. Review by Richard D. Routledge, Simon Fraser University. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):175–177, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Mathai:1983:SRR

[289] A. M. Mathai. Summary–reviews/résumés bibliographiques: Multivariate statistical analysis. Proceedings of the research seminar at Dalhousie University, Halifax, October 5–7, 1979. Edited by R. P. Gupta, North-Holland Publishing Company, Amsterdam, 1980. viii + 289 pp. Df1.85.00, US $41.50. ISBN 0-444-86019-3. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):177, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

MacKay:1983:SRR

[290] R. Jock MacKay. Summary–reviews/résumés bibliographiques: Biostatistics: 1064 Questions. By J. J. Hubert and E. M. Carter. Kendall/Hunt Publishing Company, Dubuque, Iowa, 1980. 155 pp. C$6.95, US $5.70. ISBN 0-8403-2287-9. Canadian Journal of Statistics = Revue canadienne de statistique, 11(2):178, June 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Moore:1983:REF

[291] Marc Moore and Normand Laniel. Reconstruction échantillonnale d’une forme convexe. (French) [Sampling reconstruction of a convex form]. Canadian Journal of Statistics = Revue canadienne de statistique, 11 (3):181–197, September 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Monette:1983:MAA

[292] Georges Monette. A measure of aliasing and applications to two-level fractional factorials. Canadian Journal of Statistics = Revue canadienne de statistique, 11(3):199–206, September 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Tardif:1983:DEI

[293] Serge Tardif. Une démonstration élémentaire de l’indépendance entre les vecteurs des rangs et des statistiques d’ordre. (French) [A basic demonstration of independence between the vectors of ranks and order statistics]. Canadian Journal of Statistics = Revue canadienne de statistique, 11(3):207–209, September 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Aly:1983:SLT

[294] Emad-Eldin A. A. Aly. Some limit theorems for uniform and exponential spacings. Canadian Journal of Statistics = Revue canadienne de statistique, 11(3):211–219, September 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Neudecker:1983:SRC

[295] H. Neudecker and Tom Wansbeek. Some results on commutation matrices, with statistical applications. Canadian Journal of Statistics = Revue canadienne de statistique, 11(3):221–231, September 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Draper:1983:DOA

[296] Norman R. Draper and Irwin Guttman. Dropping observations without affecting posterior and predictive distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 11(3):233–237, September 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Rinco:1983:EPA

[297] Stefan Rinco. Estimation of Pry_{p} > max(y_{1}, …, y_{p−1}): Predictive approach in exponential case. Canadian Journal of Statistics = Revue canadienne de statistique, 11(3):239–244, September 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Ivanoff:1983:MBR

[298] B. Gail Ivanoff. The multitype branching random walk. I. Canadian Journal of Statistics = Revue canadienne de statistique, 11(3):245–257, September 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Boulanger:1983:UALa

[299] Alain Boulanger. Uniform asymptotic linearity in a regression parameter of a process based on a rank statistic. Canadian Journal of Statistics = Revue canadienne de statistique, 11(4):265–269, December 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Boulanger:1983:ESB

[300] Alain Boulanger. Estimators of shift based on statistics of the Kolmogorov-Smirnov type. Canadian Journal of Statistics = Revue canadienne de statistique, 11(4):271–284, December 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Boulanger:1983:UALb

[301] Alain Boulanger and Constance Van Eeden. Uniform asymptotic linearity of a process based on a signed rank statistic. Canadian Journal of Statistics = Revue canadienne de statistique, 11(4):285–292, December 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Boulanger:1983:ELB

[302] Alain Boulanger and Constance Van Eeden. Estimators of location based on Kolmogorov-Smirnov-type statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 11(4):293–303, December 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Whitmore:1983:RMC

[303] G. A. Whitmore. A regression method for censored inverse-Gaussian data. Canadian Journal of Statistics = Revue canadienne de statistique, 11(4):305–315, December 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Wani:1983:CIP

[304] Jagannath K. Wani and Hing Po Lo. A characterization of invariant power-series abundance distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 11(4):317–323, December 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Bondar:1983:UOE

[305] James V. Bondar. Universal optimality of experimental designs: Definitions and a criterion. Canadian Journal of Statistics = Revue canadienne de statistique, 11(4):325–331, December 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Young:1983:BRC

[306] J. Clij Young. Book review/comptes rendus critiques: Introductory engineering statistics. Third Edition. By Irwin Guttman, The Late S. S. Wilks, and J. Stuart Hunter. John Wiley and Sons, Inc., New York, 1982. xi + 580 pp. US $33.95. ISBN 0-471-07859-X. [Solutions Manual available.]. Canadian Journal of Statistics = Revue canadienne de statistique, 11(4):333, December 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Fellegi:1983:BRC

[307] Ivan P. Fellegi. Book review/comptes rendus critiques: Developments in dual system estimation of population size and growth. Edited by Karol J. Kr6tki. The University of Alberta Press, Edmonton, Alberta, 1978. xviii + 260 pp. C$17.50, US$14.64. ISBN 0-88864-017-X. Canadian Journal of Statistics = Revue canadienne de statistique, 11(4):334–339, December 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Watts:1983:BRC

[308] Donald G. Watts. Book review/comptes rendus critiques: Analysis and adjustment of survey measurements. By Edward M. Mikhail and Gordon Grade. Van Nostrand Reinhold Company, New York, 1981. xii + 340 pp. US $28.50. ISBN 0-442-25369-9. Review by Donald G. Watts, Queen’s University. Canadian Journal of Statistics = Revue canadienne de statistique, 11(4):338–339, December 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1983:BRCb

[309] Anonymous. Book review/comptes rendus critiques: A sampler on sampling. By Bill Williams. John Wiley and Sons, Inc., New York, 1978. xv + 254 pp. C$37.50, US $26.95. ISBN 0-471-03036-8. Review by Jack E. Graham, Carleton University. Canadian Journal of Statistics = Revue canadienne de statistique, 11(4):339–341, December 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

MacKay:1983:BRBa

[310] R. Jock MacKay. Book review: The distribution of the size of the maximum cluster of points on a line. By Norman D. Neff and Joseph I. Naus. Selected Tables in Mathematical Statistics, Volume VI (W. J. Kennedy and R. E. Odeh, coeditors). American Mathematical Society, Providence, Rhode Island, 1980. vii + 207 pp. US $12.80. ISBN 0-8218-1906-2, ISSN 0094-8837. Canadian Journal of Statistics = Revue canadienne de statistique, 11(4):342, December 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

MacKay:1983:BRBb

[311] R. Jock MacKay. Book review: The product of two normally distributed random variables. By William Q. Meeker, Jr., Larry W. Cornwell, and Leo A. Aroian. Selected Tables in Mathematical Statistics, Volume VII (W. J. Kennedy and R. E. Odeh, coeditors). American Mathematical Society, Providence, Rhode Island, 1981. vii + 256 pp. US $12.40. ISBN 0-8218-1907-0, ISSN 0094-8837. Canadian Journal of Statistics = Revue canadienne de statistique, 11(4):342–343, December 1983. ISSN 0319-5724 (print), 1708-945X (electronic).

Fellegi:1984:NHM

[312] Ivan P. Fellegi. New honorary member of the Statistical Society of Canada. Canadian Journal of Statistics = Revue canadienne de statistique, 12(1):1–6, March 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Gentleman:1984:CSD

[313] Jane F. Gentleman and G. A. Whitmore. Case study in data analysis, no. 2. Canadian Journal of Statistics = Revue canadienne de statistique, 12(1):7–10, March 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Kusiak:1984:APF

[314] Robert Kusiak and Jaan Roos. Analysis of pulmonary function data. Canadian Journal of Statistics = Revue canadienne de statistique, 12 (1):10–15, March 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Worsley:1984:TLD

[315] Keith J. Worsley. Testing for lung damage to firefighters at the Mississauga train derailment. Canadian Journal of Statistics = Revue canadienne de statistique, 12(1):16–21, March 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Gentleman:1984:SAE

[316] Jane F. Gentleman. Section associate editors’ comments. Canadian Journal of Statistics = Revue canadienne de statistique, 12(1):21–25, March 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Sprott:1984:IML

[317] David A. Sprott and Román Viveros-Aguilera. The interpretation of maximum-likelihood estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 12(1):27–38, March 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Bartholomew:1984:RDN

[318] D. J. Bartholomew. Recent developments in nonlinear stochastic modelling of social processes. Canadian Journal of Statistics = Revue canadienne de statistique, 12(1):39–52, March 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Bellhouse:1984:ROD

[319] David R. Bellhouse. A review of optimal designs in survey sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 12 (1):53–65, March 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Mukhopadhyay:1984:EDL

[320] Nitis Mukhopadhyay and Hosny I. Hamdy. On estimating the difference of location parameters of two negative exponential distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 12 (1):67–76, March 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Bellhouse:1984:ESD

[321] David R. Bellhouse and Agnes M. Herzberg. Equally spaced design points in polynomial regression: a comparison of systematic sampling methods with the optimal design of experiments. Canadian Journal of Statistics = Revue canadienne de statistique, 12(2):77–90, June 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Joe:1984:CTL

[322] Harry Joe and Frank Proschan. Comparison of two life distributions on the basis of their percentile residual life functions. Canadian Journal of Statistics = Revue canadienne de statistique, 12(2):91–97, June 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Tomkins:1984:MGP

[323] R. James Tomkins. Martingale generalizations and preservation of martingale properties. Canadian Journal of Statistics = Revue canadienne de statistique, 12(2):99–106, June 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Koziol:1984:REG

[324] James A. Koziol and A. John Petkau. Relative efficiencies of goodnessof-fit procedures with truncated data. Canadian Journal of Statistics = Revue canadienne de statistique, 12(2):107–117, June 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Srivastava:1984:DCC

[325] M. S. Srivastava and G. C. Lee. On the distribution of the correlation coefficient when sampling from a mixture of two bivariate normal densities: Robustness and the effect of outliers. Canadian Journal of Statistics = Revue canadienne de statistique, 12(2):119–133, June 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Plante:1984:RSA

[326] André Plante. A reexamination of Stein’s antifiducial example. Canadian Journal of Statistics = Revue canadienne de statistique, 12(2):135–141, June 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Geisser:1984:PPE

[327] Seymour Geisser. Predicting Pareto and exponential observables. Canadian Journal of Statistics = Revue canadienne de statistique, 12(2):143– 152, June 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Genest:1984:POM

[328] Christian Genest. Pooling operators with the marginalization property. Canadian Journal of Statistics = Revue canadienne de statistique, 12(2): 153–163, June 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Duthie:1984:NCS

[329] Audrey I. Duthie. A note on the construction of some lattices for the hypercubic association scheme. Canadian Journal of Statistics = Revue canadienne de statistique, 12(2):165–167, June 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Kalbfleisch:1984:LSE

[330] J. D. Kalbfleisch and J. F. Lawless. Least-squares estimation of transition probabilities from aggregate data. Canadian Journal of Statistics = Revue canadienne de statistique, 12(3):169–182, September 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Fraser:1984:FRA

[331] D. A. S. Fraser and P. McDunnough. Further remarks on asymptotic normality of likelihood and conditional analyses. Canadian Journal of Statistics = Revue canadienne de statistique, 12(3):183–190, September 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Owen:1984:NBC

[332] Art Owen. A neighbourhood-based classifier for Landsat data. Canadian Journal of Statistics = Revue canadienne de statistique, 12(3):191–200, September 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Matthews:1984:SOS

[333] David E. Matthews. Some observations on semi-Markov models for partially censored data. Canadian Journal of Statistics = Revue canadienne de statistique, 12(3):201–205, September 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Matthews:1984:ECA

[334] David E. Matthews. Efficiency considerations in the analysis of a competing-risk problem. Canadian Journal of Statistics = Revue canadienne de statistique, 12(3):207–210, September 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Kohler:1984:CEO

[335] Boris A. B. Kohler. Un cas élémentaire d’oscillations. Canadian Journal of Statistics = Revue canadienne de statistique, 12(3):211–216, September 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Puri:1984:ANL

[336] Madan L. Puri and Tiee-Jian Wu. Asymptotic normality of the lengths of a class of nonparametric confidence intervals for a regression parameter. Canadian Journal of Statistics = Revue canadienne de statistique, 12 (3):217–228, September 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Wright:1984:ABM

[337] F. T. Wright. The asymptotic behavior of monotone percentile regression estimates. Canadian Journal of Statistics = Revue canadienne de statistique, 12(3):229–236, September 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Watts:1984:TSC

[338] Donald G. Watts. Teaching Statistics in Canada: The Early Days. Canadian Journal of Statistics = Revue canadienne de statistique, 12(3):237– 239, September 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1984:C

[339] Anonymous. Corrigendum. Canadian Journal of Statistics = Revue canadienne de statistique, 12(3):240, September 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Baksalary:1984:CEB

[340] J. K. Baksalary and R. Kala. Corrigendum: “On equalities between BLUEs, WLSEs, and SLSEs”. Canadian Journal of Statistics = Revue canadienne de statistique, 12(3):240, 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Pyke:1984:ARE

[341] Ronald Pyke. Asymptotic results for empirical and partial-sum processes: a review. Canadian Journal of Statistics = Revue canadienne de statistique, 12(4):241–264, 283–287, December 1984. ISSN 0319-5724 (print), 1708-945X (electronic). With discussion by R. J. Kulperger and a reply by the author.

McLeish:1984:EAM

[342] D. L. McLeish. Estimation for aggregate models: the aggregate Markov chain. Canadian Journal of Statistics = Revue canadienne de statistique, 12(4):265–282, 283–285, December 1984. ISSN 0319-5724 (print), 1708-945X (electronic). With discussion by R. J. Kulperger.

Kulperger:1984:DPP

[343] R. J. Kulperger. Discussion of the papers of Professors Pyke and McLeish. Canadian Journal of Statistics = Revue canadienne de statistique, 12(4):283–285, December 1984. ISSN 0319-5724 (print), 1708-945X (electronic). See [242, 341, 344].

Pyke:1984:RDR

[344] Ronald Pyke. Response to the discussion of R. J. Kulperger. Canadian Journal of Statistics = Revue canadienne de statistique, 12(4):286–287, December 1984. ISSN 0319-5724 (print), 1708-945X (electronic). See [343].

Guttorp:1984:SIS

[345] Peter Guttorp and Reg Kulperger. Statistical inference for some Volterra population processes in a random environment. Canadian Journal of Statistics = Revue canadienne de statistique, 12(4):289–302, December 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Feuerverger:1984:STM

[346] Andrey Feuerverger and Philip McDunnough. On statistical transform methods and their efficiency. Canadian Journal of Statistics = Revue canadienne de statistique, 12(4):303–317, December 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Shorack:1984:ERP

[347] Galen R. Shorack. Empirical and rank processes of observations and residuals. Canadian Journal of Statistics = Revue canadienne de statistique, 12(4):319–332, December 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Melard:1984:TEA

[348] Guy Mélard and Roch Roy. Sur un test d’égalité des autocovariances de deux séries chronologiques. (French) [On an equality test of autocovariances of two time series]. Canadian Journal of Statistics = Revue canadienne de statistique, 12(4):333–342, December 1984. ISSN 0319-5724 (print), 1708-945X (electronic).

Evans:1985:MEA

[349] M. Evans, D. A. S. Fraser, and G. Monette. Mixtures, embedding, and ancillarity. Canadian Journal of Statistics = Revue canadienne de statistique, 13(1):1–6, March 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Schulzer:1985:ACS

[350] Michael Schulzer, Donald A. Enarson, and Moira Chan-Yeung. Analyzing cross-sectional and longitudinal lung-function measurements: the effects of age. Canadian Journal of Statistics = Revue canadienne de statistique, 13(1):7–15, March 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Bellhouse:1985:SSP

[351] D. R. Bellhouse. Systematic sampling of periodic functions. Canadian Journal of Statistics = Revue canadienne de statistique, 13(1):17–28, March 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Schiopu-Kratina:1985:CCS

[352] Ioana Schiopu-Kratina. Convergence of compensators of simple point processes. Canadian Journal of Statistics = Revue canadienne de statistique, 13(1):29–35, March 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Genter:1985:GLT

[353] Frederic C. Genter and Vernon T. Farewell. Goodness-of-link testing in ordinal regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 13(1):37–44, March 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Hollander:1985:TWN

[354] Myles Hollander, Dong Ho Park, and Frank Proschan. Testing whether new is better than used of a specified age, with randomly censored data. Canadian Journal of Statistics = Revue canadienne de statistique, 13(1): 45–52, March 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Aly:1985:QNP

[355] Emad-Eldin A. A. Aly and Miklós Csörgő. Quadratic nuisanceparameter-free goodness-of-fit tests in the presence of location and scale parameters. Canadian Journal of Statistics = Revue canadienne de statistique, 13(1):53–70, March 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Liski:1985:PRM

[356] Erkki P. Liski and Simo Puntanen. A percentile regression model with an application to error frequency in group conversation tests. Canadian Journal of Statistics = Revue canadienne de statistique, 13(1):71– 78, March 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Helmers:1985:BEB

[357] R. Helmers. The Berry–Esséen bound for Studentized U-statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 13(1):79– 82, March 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Whitmore:1985:IPC

[358] G. A. Whitmore and Jane F. Gentleman. Iceberg paths and collision risks for fixed marine structures. Canadian Journal of Statistics = Revue canadienne de statistique, 13(2):83–87, June 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Moore:1985:MIM

[359] Marc Moore. Modelling iceberg motion: a multiple-time-series approach. Canadian Journal of Statistics = Revue canadienne de statistique, 13(2): 88–94, June 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Zwiers:1985:EPC

[360] F. W. Zwiers. Estimating the probability of collision between an iceberg and a fixed marine structure. Canadian Journal of Statistics = Revue canadienne de statistique, 13(2):94–105, June 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Whitmore:1985:SAE

[361] G. À. Whitmore. Section associate editors’ comments. Canadian Journal of Statistics = Revue canadienne de statistique, 13(2):106–108, June 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Robertson:1985:OSC

[362] Tim Robertson and F. T. Wright. One-sided comparisons for treatments with a control. Canadian Journal of Statistics = Revue canadienne de statistique, 13(2):109–122, June 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Campbell:1985:TAB

[363] L. Lorne Campbell. Transient analysis of birth-death processes with two boundaries. Canadian Journal of Statistics = Revue canadienne de statistique, 13(2):123–130, June 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Small:1985:DMW

[364] Christopher G. Small. Decomposition of models whose marginal distributions are mixtures. Canadian Journal of Statistics = Revue canadienne de statistique, 13(2):131–136, June 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Evans:1985:RSM

[365] M. Evans, D. A. S. Fraser, and G. Monette. On regularity for statistical models. Canadian Journal of Statistics = Revue canadienne de statistique, 13(2):137–144, June 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Hooper:1985:IEF

[366] Peter M. Hooper. Invariant estimation of functions. Canadian Journal of Statistics = Revue canadienne de statistique, 13(2):145–154, June 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Reid:1985:CLR

[367] N. Reid. Curvature and linear rank statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 13(2):155–165, June 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Mullin:1985:BRB

[368] R. Mullin. Book review: Time warps, string edits, and macromolecules: the theory and practice of sequence comparison. Edited by D. Sankoff and J. B. Kruskal. Addison-Wesley Publishing Company, Inc., Advanced Book Program, Reading, Mass., Don Mills, Ontario, 1983. 300 pp. US $31.95. ISBN 0-201-07809-0. Canadian Journal of Statistics = Revue canadienne de statistique, 13(2):167–168, June 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Thompson:1985:BRB

[369] E. Thompson. Book review: Current topics in survey sampling. edited by D. Krewski, R. Platek, and J. N. K. Rao. Academic Press, New York, 1981. xv + 509 pp. US $35.00. ISBN 0-12-426280-5. Canadian Journal of Statistics = Revue canadienne de statistique, 13(2):168–169, June 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Desmond:1985:SMF

[370] Anthony Desmond. Stochastic models of failure in random environments. Canadian Journal of Statistics = Revue canadienne de statistique, 13 (3):171–183, September 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Elvers:1985:RAR

[371] Eva Elvers, Carl-Erik Erik Särndal, Jan H. Wretman, and Göran Örnberg. Regression analysis and ratio analysis for domains: a randomization-theory approach. Canadian Journal of Statistics = Revue canadienne de statistique, 13(3):185–199, September 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Field:1985:ANM

[372] Christopher Field. Approach to normality of mean and M-estimators of location. Canadian Journal of Statistics = Revue canadienne de statistique, 13(3):201–210, September 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Rivest:1985:TBE

[373] Louis-Paul Rivest. Tests based on equivariant estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 13(3):211–216, September 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Tardif:1985:AEA

[374] Serge Tardif. On the asymptotic efficiency of aligned-rank tests in randomized block designs. Canadian Journal of Statistics = Revue canadienne de statistique, 13(3):217–232, September 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

DeJong:1985:MMP

[375] Piet De Jong and Malcolm Greig. Models and methods for pairing data. Canadian Journal of Statistics = Revue canadienne de statistique, 13 (3):233–241, September 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Sinha:1985:UEV

[376] Bimal Kumar Sinha. Unbiased estimation of the variance of the Graybill-Deal estimator of the common mean of several normal populations. Canadian Journal of Statistics = Revue canadienne de statistique, 13(3):243– 247, September 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Letac:1985:KCI

[377] Gérard Letac and Vanamamalai Seshadri. On Khatri’s characterization of the inverse-Gaussian distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 13(3):249–252, September 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Lockhart:1985:ADC

[378] R. A. Lockhart. The asymptotic distribution of the correlation coefficient in testing fit to the exponential distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 13(3):253–256, September 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Gentleman:1985:TPT

[379] Jane F. Gentleman and G. A. Whitmore. Temporal patterns in twenty years of Canadian homicides. Canadian Journal of Statistics = Revue canadienne de statistique, 13(4):262–266, December 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

McKie:1985:SCH

[380] Craig McKie. Seasonality in Canadian homicides, 1961–1980. Canadian Journal of Statistics = Revue canadienne de statistique, 13(4):266–268, December 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

McLeod:1985:SEC

[381] A. Ian McLeod, Ian B. MacNeill, and Jahnabimala D. Bhattacharyya. Seasonal effects in Canadian murders. Canadian Journal of Statistics = Revue canadienne de statistique, 13(4):269–275, December 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Nakamura:1985:SMF

[382] Alice Nakamura and Masao Nakamura. A search for monthly fluctuation in Canadian homicides: 1965–1980. Canadian Journal of Statistics = Revue canadienne de statistique, 13(4):275–280, December 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1985:SEC

[383] Anonymous. Section editors’ comments. Canadian Journal of Statistics = Revue canadienne de statistique, 13(4):281–291, December 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Genest:1985:RGP

[384] Christian Genest and Mark J. Schervish. Resolution of Godambe’s paradox. Canadian Journal of Statistics = Revue canadienne de statistique, 13(4):293–297, December 1985. ISSN 0319-5724 (print), 1708-945X (electronic). See discussion [385] and rejoinder [386].

Fraser:1985:D

[385] Donald A. S. Fraser. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 13(4):298–300, December 1985. ISSN 0319-5724 (print), 1708-945X (electronic). See [384, 386].

Genest:1985:R

[386] Christain Genest and Mark J. Schervish. Rejoinder. Canadian Journal of Statistics = Revue canadienne de statistique, 13(4):301, December 1985. ISSN 0319-5724 (print), 1708-945X (electronic). See [385].

Tardif:1985:LAD

[387] Serge Tardif. Sur la linéarité asymptotique du premier ordre de statistiques de rang signé. (French) [On the first-order asymptotic linearity of signed rank statistics]. Canadian Journal of Statistics = Revue canadienne de statistique, 13(4):303–310, December 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Huse:1985:IER

[388] Vera R. Huse and Josef Steinebach. On an improved Erdős–Rényi-type law for increments of partial sums. Canadian Journal of Statistics = Revue canadienne de statistique, 13(4):311–315, December 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Weiss:1985:LRC

[389] Günter M. T. Weiss. Local rates of convergence for consistent estimators of location of translation families. Canadian Journal of Statistics = Revue canadienne de statistique, 13(4):317–324, December 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

Takada:1985:PLO

[390] Yoshikazu Takada. Prediction Limit for Observation from the Exponential Distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 13(4):325–330, December 1985. ISSN 0319-5724 (print), 1708-945X (electronic).

MacNeill:1986:NHMa

[391] Ian B. MacNeill. New honorary member of the Statistical Society of Canada. Canadian Journal of Statistics = Revue canadienne de statistique, 14(1):1–3, March 1986. ISSN 0319-5724 (print), 1708-945X (electronic). With a French translation.

Diciccio:1986:ACI

[392] Thomas J. Diciccio. Approximate conditional inference for location families. Canadian Journal of Statistics = Revue canadienne de statistique, 14(1):5–18, March 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Csaki:1986:SJD

[393] E. Csáki and S. G. Mohanty. Some joint distributions for conditional random walks. Canadian Journal of Statistics = Revue canadienne de statistique, 14(1):19–28, March 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Srivastava:1986:CEI

[394] M. S. Srivastava and R. S. Katapa. Comparison of estimators of interclass and intraclass correlations from familial data. Canadian Journal of Statistics = Revue canadienne de statistique, 14(1):29–42, March 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Haff:1986:SCS

[395] L. R. Haff and R. W. Johnson. The superharmonic condition for simultaneous estimation of means in exponential families. Canadian Journal of Statistics = Revue canadienne de statistique, 14(1):43–54, March 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Anderson:1986:MLE

[396] T. W. Anderson, Kai Tai Fang, and Huang Hsu. Maximum-likelihood estimates and likelihood-ratio criteria for multivariate elliptically contoured distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 14(1):55–59, March 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Provost:1986:EDR

[397] Serge B. Provost. The exact distribution of the ratio of a linear combination of chi-square variables over the root of a product of chi-square variables. Canadian Journal of Statistics = Revue canadienne de statistique, 14(1):61–67, March 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:1986:CMA

[398] Y. H. Wang. Coupling methods in approximations. Canadian Journal of Statistics = Revue canadienne de statistique, 14(1):69–74, March 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Froda:1986:CEE

[399] Sorana Froda. Comparison of efficient L- and R-estimators of location. Canadian Journal of Statistics = Revue canadienne de statistique, 14(1): 75–79, March 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Maejima:1986:RSS

[400] Makoto Maejima. A remark on self-similar processes with stationary increments. Canadian Journal of Statistics = Revue canadienne de statistique, 14(1):81–82, March 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Ripley:1986:SIP

[401] Brian D. Ripley. Statistics, images, and pattern recognition. Canadian Journal of Statistics = Revue canadienne de statistique, 14(2):83–102, June 1986. ISSN 0319-5724 (print), 1708-945X (electronic). See discussion [402].

Hinkley:1986:D

[402] David V. Hinkley, Carl N. Morris, Marc Moore, Art Owen, and R. Chellappa. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 14(2):102–111, June 1986. ISSN 0319-5724 (print), 1708-945X (electronic). See [401].

Casella:1986:RBC

[403] George Casella. Refining binomial confidence intervals. Canadian Journal of Statistics = Revue canadienne de statistique, 14(2):113–129, June 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Csorgo:1986:CBC

[404] Sándor Csörgő and Lajos Horváth. Confidence bands from censored samples. Canadian Journal of Statistics = Revue canadienne de statistique, 14(2):131–144, June 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Genest:1986:CAF

[405] Christian Genest and R. Jock MacKay. Copules archimédiennes et families de lois bidimensionnelles dont les marges sont données. (French) [Archimedean copulas and families of two-dimensional laws whose bounds are given]. Canadian Journal of Statistics = Revue canadienne de statistique, 14(2):145–159, June 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Wiens:1986:REM

[406] D. P. Wiens and Z. Zheng. Robust M-estimators of multivariate location and scatter in the presence of asymmetry. Canadian Journal of Statistics = Revue canadienne de statistique, 14(2):161–176, June 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Currie:1986:RBS

[407] Iain D. Currie and Michael A. Stephens. Relations between statistics for testing exponentiality and uniformity. Canadian Journal of Statistics = Revue canadienne de statistique, 14(2):177–180, June 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Evans:1986:PAL

[408] Michael J. Evans, Donald A. S. Fraser, and Georges Monette. On principles and arguments to likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 14(3):181–194, September 1986. ISSN 0319-5724 (print), 1708-945X (electronic). See discussion [409] and rejoinder [410].

Kalbfleisch:1986:D

[409] J. D. Kalbfleisch. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 14(3):194–199, September 1986. ISSN 0319-5724 (print), 1708-945X (electronic). See [408].

Evans:1986:R

[410] Michael J. Evans, Donald A. S. Fraser, and Georges Monette. Rejoinder. Canadian Journal of Statistics = Revue canadienne de statistique, 14(3): 199, September 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Bartlett:1986:SFP

[411] Roy F. Bartlett. Sampling a finite population in the presence of trend and correlation: Estimation of total 305-day lactation production in cattle. Canadian Journal of Statistics = Revue canadienne de statistique, 14 (3):201–210, September 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Devroye:1986:SUC

[412] Luc Devroye and Clark S. Penrod. The strong uniform convergence of multivariate variable kernel estimates. Canadian Journal of Statistics = Revue canadienne de statistique, 14(3):211–219, September 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Csorgo:1986:WES

[413] Miklós Csörgő and Lajos Horváth. Weighted empirical spacings processes. Canadian Journal of Statistics = Revue canadienne de statistique, 14(3):221–232, September 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Bose:1986:LLN

[414] A. Bose. A law of large numbers for the scaled age distribution of linear birth-and-death processes. Canadian Journal of Statistics = Revue canadienne de statistique, 14(3):233–244, September 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:1986:FAM

[415] Tze Fen Li and Dinesh S. Bhoj. A family of admissible minimax estimators of the mean of a multivariate, normal distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 14(3):245–250, September 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Ghosh:1986:NRM

[416] J. K. Ghosh and C. K. Mustafi. A note on the residual median process. Canadian Journal of Statistics = Revue canadienne de statistique, 14 (3):251–255, September 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Farewell:1986:MMS

[417] Vernon T. Farewell. Mixture models in survival analysis: Are they worth the risk? Canadian Journal of Statistics = Revue canadienne de statistique, 14(3):257–262, September 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Andrews:1986:BIR

[418] D. F. Andrews, R. Brant, and M. E. Percy. Bayesian incorporation of repeated measurements in logistic discrimination. Canadian Journal of Statistics = Revue canadienne de statistique, 14(3):263–266, September 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Nebebe:1986:BEB

[419] Fassil Nebebe and T. W. F. Stroud. Bayes and empirical Bayes shrinkage estimation of regression coefficients. Canadian Journal of Statistics = Revue canadienne de statistique, 14(4):267–280, December 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Tallis:1986:PDS

[420] G. M. Tallis. Predictive densities and selection with genetic applications. Canadian Journal of Statistics = Revue canadienne de statistique, 14(4):281–288, December 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Chang:1986:TSR

[421] Ted Chang and Cesareo Villegas. On a theorem of Stein relating Bayesian and classical inferences in group models. Canadian Journal of Statistics = Revue canadienne de statistique, 14(4):289–296, December 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Lambert:1986:SPI

[422] Diane Lambert and George T. Duncan. Single-parameter inference based on partial prior information. Canadian Journal of Statistics = Revue canadienne de statistique, 14(4):297–305, December 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Johnson:1986:BNS

[423] Wesley Johnson and Ronald Christensen. Bayesian nonparametric survival analysis for grouped data. Canadian Journal of Statistics = Revue canadienne de statistique, 14(4):307–314, December 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Hooper:1986:OCR

[424] Peter M. Hooper and Wai Kwok Yau. Optimal confidence regions in GMANOVA. Canadian Journal of Statistics = Revue canadienne de statistique, 14(4):315–322, December 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Wasserman:1986:CCL

[425] Larry A. Wasserman. Censoring, conditionality, and likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 14(4):323– 327, December 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Sutradhar:1986:CFM

[426] Brajendra C. Sutradhar. On the characteristic function of multivariate Student t-distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 14(4):329–337, December 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Pukelsheim:1986:ATM

[427] Friedrich Pukelsheim. Approximate theory of multiway block designs. Canadian Journal of Statistics = Revue canadienne de statistique, 14 (4):339–346, December 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Wani:1986:SPS

[428] Jagannath K. Wani and Hing-Po Po Lo. Selecting a power-series distribution for goodness of fit. Canadian Journal of Statistics = Revue canadienne de statistique, 14(4):347–353, December 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Zografos:1986:DAC

[429] K. Zografos, K. Ferentinos, and T. Papaioannou. Discrete approximations to the Csiszár, Rényi, and Fisher measures of information. Canadian Journal of Statistics = Revue canadienne de statistique, 14(4):355– 366, December 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Huffer:1986:SIC

[430] Fred W. Huffer. Slepian’s inequality via the central limit theorem. Canadian Journal of Statistics = Revue canadienne de statistique, 14(4):367– 370, December 1986. ISSN 0319-5724 (print), 1708-945X (electronic).

Rubin:1987:BBP

[431] D. B. Rubin and T. W. F. Stroud. Bayesian break-point forecasting in parallel time series, with application to university admissions. Canadian Journal of Statistics = Revue canadienne de statistique, 15(1):1–19, March 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Fraser:1987:SOI

[432] D. A. S. Fraser and H. Massam. Second-order inference for generalized least squares. Canadian Journal of Statistics = Revue canadienne de statistique, 15(1):21–30, March 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Small:1987:MCM

[433] Christopher G. Small. Measures of centrality for multivariate and directional distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 15(1):31–39, March 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Kim:1987:TMT

[434] Kang-Kyun Kim and Robert V. Foutz. Tests for the multivariate twosample problem based on empirical probability measures. Canadian Journal of Statistics = Revue canadienne de statistique, 15(1):41–51, March 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Johnson:1987:APB

[435] Bruce R. Johnson and Donald R. Truax. Asymptotic properties of Bayes risk for one-sided tests. Canadian Journal of Statistics = Revue canadienne de statistique, 15(1):53–61, March 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Lynch:1987:USO

[436] James Lynch, Gillian Mimmack, and Frank Proschan. Uniform stochastic orderings and total positivity. Canadian Journal of Statistics = Revue canadienne de statistique, 15(1):63–69, March 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Herzberg:1987:EIR

[437] Agnes M. Herzberg, P. Prescott, and M. Akhtar. Equi-information robust designs: Which designs are possible? Canadian Journal of Statistics = Revue canadienne de statistique, 15(1):71–76, March 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Kozek:1987:LDK

[438] Andrzej Kozek. Limiting distributions of Kolmogorov-Lévy-type statistics under the alternative. Canadian Journal of Statistics = Revue canadienne de statistique, 15(1):77–85, March 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Massam:1987:CDL

[439] Hélène Massam. Consistent directions for least-squares estimates. Canadian Journal of Statistics = Revue canadienne de statistique, 15(1):87– 90, March 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Ross:1987:GCD

[440] William H. Ross. The geometry of case deletion and the assessment of influence in nonlinear regression. Canadian Journal of Statistics = Revue canadienne de statistique, 15(2):91–103, June 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Cox:1987:AND

[441] David R. Cox and Nancy Reid. Approximations to noncentral distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 15(2):105–114, June 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Brant:1987:RCG

[442] Rollin Brant. Residual components in generalized linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 15(2):115– 126, June 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Hamilton:1987:CTE

[443] David Hamilton. A comparison of two exact confidence regions for partially nonlinear regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 15(2):127–135, June 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Bai:1987:MLE

[444] Z. D. Bai and J. C. Fu. On the maximum-likelihood estimator for the location parameter of a Cauchy distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 15(2):137–146, June 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Baksalary:1987:CPD

[445] Jerzy K. Baksalary and Zenon Tabis. Connectedness of PBIB designs. Canadian Journal of Statistics = Revue canadienne de statistique, 15(2): 147–150, June 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

OReilly:1987:CPP

[446] Federico J. O’Reilly and Cesareo Villegas. Conditional and predictive probability integral transforms. Canadian Journal of Statistics = Revue canadienne de statistique, 15(2):151–158, June 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

McLaren:1987:AEC

[447] C. G. McLaren and Richard Lockhart. On the asymptotic efficiency of certain correlation tests of fit. Canadian Journal of Statistics = Revue canadienne de statistique, 15(2):159–167, June 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Pavur:1987:DMQ

[448] Robert J. Pavur. Distribution of multivariate quadratic forms under certain covariance structures. Canadian Journal of Statistics = Revue canadienne de statistique, 15(2):169–176, June 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Berry:1987:EEN

[449] J. Calvin Berry. Equivariant estimation of a normal mean using a normal concomitant variable for covariance adjustment. Canadian Journal of Statistics = Revue canadienne de statistique, 15(2):177–183, June 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Moore:1987:ISP

[450] Marc Moore. Inférence statistique dans les processus stochastiques: Aperçu historique. (French) [Statistical inference in stochastic processes: historical overview]. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3):185–207, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Lawless:1987:NBM

[451] Jerald F. Lawless. Negative binomial and mixed Poisson regression. Canadian Journal of Statistics = Revue canadienne de statistique, 15 (3):209–225, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Joe:1987:ODD

[452] Harry Joe. An ordering of dependence for distribution of k-tuples, with applications to lotto games. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3):227–238, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Luong:1987:MDM

[453] A. Luong and M. E. Thompson. Minimum-distance methods based on quadratic distances for transforms. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3):239–251, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Dabrowski:1987:IPP

[454] André Robert Dabrowski. Invariance principles in probability for stable processes generated by a class of dependent sequences. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3):253–267, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Wiens:1987:RWC

[455] Douglas P. Wiens. Robust weighted Cramér-von Mises estimators of location, with minimax variance in ϵ-contamination neighbourhoods. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3):269– 278, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Imhof:1987:HAS

[456] J.-P. P. Imhof. Heuristic approach to some laws for Brownian motion. Canadian Journal of Statistics = Revue canadienne de statistique, 15 (3):279–282, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Latour:1987:SER

[457] Alain Latour and Roch Roy. Some exact results on the sample autocovariances of a seasonal ARIMA model. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3):283–291, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Willmot:1987:PLZ

[458] Gordon E. Willmot. On the probabilities of the log-zero-Poisson distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3):293–297, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Reinsel:1987:BRB

[459] Gregory C. Reinsel. Book review: Statistical methods for forecasting. By Bovas Abraham and Johannes Ledolter. John Wiley and Sons, Inc., New York, 1983. xv + 445 pp. US $38.95. ISBN 0-471-86764-0. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3):299–300, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Styan:1987:BRBa

[460] George P. H. Styan. Book review: Applied matrix algebra in the statistical sciences. By Alexander Basilevsky. Elsevier Science Publishing Co.

(North-Holland), Inc., New York, 1983. xiii + 389 pp. ISBN 0-444-00756-3. Canadian Journal of Statistics = Revue canadienne de statistique, 15 (3):300–302, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Stoyanov:1987:BRBa

[461] Jordan M. Stoyanov. Book review: Contiguity and the statistical invariance principle. By P. E. Greenwood and A. N. Shiryayev. Gordon and Breach Science Publishers S. A., New York, 1985. viii + 236 pp. US $39.00. ISBN 2-88124-013-5. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3):302–304, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Styan:1987:BRBb

[462] George P. H. Styan. Book review: Topics in applied statistics. Edited by Yogendra P. Chaubey and Tryambakeshwar D. Dwivedi. Proceedings of Statistics ’81 Canada Conference held at Concordia University, Montreal, Canada, April 29–May 1, 1981. Dept. of Mathematics, Concordia University, Montreal, Quebec. ii + 808 + [7] pp. 1984, C. $120.00. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3):302–304, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Turner:1987:BRB

[463] James E. Turner. Book review: Applied mathematical demography. Second Edition. By Nathan Keyfitz. Springer-Verlag, New York, 1985. xxi + 441 pp. ISBN 0-387-96155-0. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3):304–305, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Genest:1987:BRB

[464] Christian Genest. Book review: Modern concepts and theorems of mathematical statistics. By Edward B. Manoukian. Springer-Verlag, New York, 1986. xvi + 156 pp. US $29.80, DM 58.00, ISBN 0-387-96186-0. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3):305–307, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Zielinski:1987:BRB

[465] Jan. M. Zielinski. Book review: Statistical estimation for stochastic processes. By K. Nanthi. Queen’s Papers in Pure and Applied Mathematics, No. 62, Department of Mathematics and Statistics, Queen’s University, Kingston, Ontario, 1983. vii + 269 pp. C. $13.95 + postage. Paperbound. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3): 307, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Stroud:1987:BRB

[466] Thomas W. F. Stroud. Book review: An introduction to applied multivariate statistics. By M. S. Srivastava and E. M. Carter. Elsevier Science Publishing Co. (North-Holland) Inc., New York, 1983. 394 + x pp. DM105.00, US $34.50. ISBN 0-444-00621-4. Canadian Journal of Statistics = Revue canadienne de statistique, 15(3):308–309, September 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Whitmore:1987:EVA

[467] G. A. Whitmore and Jane F. Gentleman. Extreme-value analysis of Canadian wind speeds. Canadian Journal of Statistics = Revue canadienne de statistique, 15(4):312–316, December 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Zwiers:1987:EVA

[468] F. W. Zwiers. An extreme-value analysis of wind speeds at five Canadian locations. Canadian Journal of Statistics = Revue canadienne de statistique, 15(4):317–327, December 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Ross:1987:PTA

[469] W. H. Ross. A peaks-over-threshold analysis of extreme wind speeds. Canadian Journal of Statistics = Revue canadienne de statistique, 15 (4):328–335, December 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Whitmore:1987:SEC

[470] G. A. Whitmore. Section editors’ comments. Canadian Journal of Statistics = Revue canadienne de statistique, 15(4):335–337, December 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Gastwirth:1987:NTS

[471] Joseph L. Gastwirth and Jane-Ling Wang. Nonparametric tests in small unbalanced samples: Application in employment-discrimination cases. Canadian Journal of Statistics = Revue canadienne de statistique, 15 (4):339–348, December 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Viveros:1987:ASM

[472] Román Viveros and David A. Sprott. Allowance for skewness in maximum-likelihood estimation with application to the location-scale model. Canadian Journal of Statistics = Revue canadienne de statistique, 15(4):349–361, December 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Petkau:1987:TSM

[473] A. John Petkau. Truncated sequential medical trials involving paired data. Canadian Journal of Statistics = Revue canadienne de statistique, 15(4):363–374, December 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Fraser:1987:NNL

[474] D. A. S. Fraser and R. J. Gebotys. Non-nested linear models: a conditional confidence approach. Canadian Journal of Statistics = Revue canadienne de statistique, 15(4):375–386, December 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Bock:1987:ILC

[475] M. E. Bock, P. Diaconis, F. W. Huffer, and M. D. Perlman. Inequalities for linear combinations of gamma random variables. Canadian Journal of Statistics = Revue canadienne de statistique, 15(4):387–395, December 1987. ISSN 0319-5724 (print), 1708-945X (electronic).

Fellegi:1988:SSP

[476] I. P. Fellegi and M. B. Wilk. Is statistics singular or plural? Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1):1–8, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Germain:1988:LTE

[477] Marie-France Germain and Claude Julien. L’application de la télédétection à l’estimation des superficies agricoles. (French) [The application of remote sensing in the estimation of agricultural land]. Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1):9–23, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Kovar:1988:BOM

[478] J. G. Kovar, J. N. K. Rao, and C.-F. J. Wu. Bootstrap and other methods to measure errors in survey estimates. Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1):25–45, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Binder:1988:EVR

[479] D. A. Binder and A. Théberge. Estimating the variance of rankingratio estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1):47–55, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Giles:1988:MGE

[480] Philip Giles. A model for generalized edit and imputation of survey data. Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1):57–73, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Roberts:1988:UTS

[481] Georgia R. Roberts. Use of time-series methods in the preparation of quarterly estimates of vital events. Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1):75–82, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Lefrancois:1988:ASC

[482] Bernard Lefrançois. Application des séries chronologiques à l’étude des révisions. Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1):83–96, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Taillon:1988:DTA

[483] Jacques Taillon. Dessaisonalisation d’un tableau avec contraintes linéaires: Application aux flux financiers canadiens. (French) [Seasonal adjustment of a table with linear constraints: application to Canadian financial flows]. Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1):97–107, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Dagum:1988:SAE

[484] Estela Bee Dagum, Guy Huot, and Marietta Morry. Seasonal adjustment in the eighties: Some problems and solutions. Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1):109–126, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Lemaitre:1988:LRE

[485] Georges Lemaître. A look at response errors in the labour force survey. Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1):127–141, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Binder:1988:UQA

[486] David A. Binder and Jean-Pierre Morin. Use of questions on activities of daily living to screen for disabled persons in a household survey. Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1): 143–156, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Paton:1988:EAN

[487] David G. Paton. Estimating the average number of loss days for acute conditions from the Canada Health Survey. Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1):157–162, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Proulx:1988:CEP

[488] Joanne Proulx and Ioana Şchiopu-Kratina. Classement des entreprises par groupes de taille à partir de données administratives. (French) [Ranking of companies by size groups from administrative data]. Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1):163–169, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Yan:1988:MBC

[489] Joel Yan, Kaveri Chinnappa, Rennie Molnar, and Pamela Tallon. Major benefits from computer-assisted collection mapping to support the Census of Canada. Canadian Journal of Statistics = Revue canadienne de statistique, 16(S1):171–184, August 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Burnett:1988:SSS

[490] R. Burnett, D. Krewski, and D. L. McLeish. Special section on statistical issues in acid rain. Canadian Journal of Statistics = Revue canadienne de statistique, 16(1):1–4, March 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Ware:1988:SIE

[491] James H. Ware and Daniel O. Stram. Statistical issues in epidemiologic studies of the health effects of ambient acid aerosols. Canadian Journal of Statistics = Revue canadienne de statistique, 16(1):5–13, March 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Le:1988:VRA

[492] D. Nhu Le and A. John Petkau. The variability of rainfall acidity revisited. Canadian Journal of Statistics = Revue canadienne de statistique, 16(1):15–38, March 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Switzer:1988:AHD

[493] Paul Switzer. An analysis of hourly deposition data. Canadian Journal of Statistics = Revue canadienne de statistique, 16(1):39–50, March 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Oehlert:1988:IED

[494] Gary W. Oehlert. Interval estimates for diatom-inferred lake pH histories. Canadian Journal of Statistics = Revue canadienne de statistique, 16(1): 51–60, March 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Weerahandi:1988:BNS

[495] S. Weerahandi and J. V. Zidek. Bayesian nonparametric smoothers for regular processes. Canadian Journal of Statistics = Revue canadienne de statistique, 16(1):61–74, March 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Wright:1988:OWA

[496] F. T. Wright. The one-way analysis of variance with ordered alternatives: a modification of Bartholomew’s Ē^{2} Test. Canadian Journal of Statistics = Revue canadienne de statistique, 16(1):75–85, March 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Baksalary:1988:CEB

[497] Jerzy K. Baksalary. Criteria for the equality between ordinary least squares and best linear unbiased estimators under certain linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 16(1): 97–102, March 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Johnson:1988:LDB

[498] Bruce R. Johnson. Lightning detection in British Columbia: an example of using system operation at unknown reliability to estimate component reliability. Canadian Journal of Statistics = Revue canadienne de statistique, 16(1):105–115, March 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Dagum:1988:NLO

[499] Estela Bee Dagum, Guy Huot, and Marietta Morry. A new look at an old problem: Finding temporal patterns in homicide series — the Canadian case. Canadian Journal of Statistics = Revue canadienne de statistique, 16(2):117–131, June 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

McLeod:1988:D

[500] A. I. McLeod, I. B. MacNeill, Alice Nakamura, and Masao Nakamura. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 16(2):131–134, June 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Fienberg:1988:IOC

[501] Stephen E. Fienberg and Judith M. Tanur. From the inside out and the outside in: Combining experimental and sampling structures. Canadian Journal of Statistics = Revue canadienne de statistique, 16(2):135–151, June 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Bilodeau:1988:EMM

[502] M. Bilodeau and M. S. Srivastava. Estimation of the MSE matrix of the Stein estimator. Canadian Journal of Statistics = Revue canadienne de statistique, 16(2):153–159, June 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Robert:1988:EFR

[503] Christian Robert. An explicit formula for the risk of the positive-part James-Stein estimator. Canadian Journal of Statistics = Revue canadienne de statistique, 16(2):161–168, June 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Bilodeau:1988:SES

[504] Martin Bilodeau. On the Simultaneous Estimation of Scale Parameters. Canadian Journal of Statistics = Revue canadienne de statistique, 16(2): 169–174, June 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Lio:1988:MKQ

[505] Y. L. Lio and W. J. Padgett. A modified kernel quantile estimator under censoring. Canadian Journal of Statistics = Revue canadienne de statistique, 16(2):175–183, June 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Tibshirani:1988:SP

[506] Robert Tibshirani and Larry A. Wasserman. Sensitive parameters. Canadian Journal of Statistics = Revue canadienne de statistique, 16(2):185– 192, June 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Das:1988:OIT

[507] Rita Das and Bimal K. Sinha. Optimum invariant tests for random MANOVA models. Canadian Journal of Statistics = Revue canadienne de statistique, 16(2):193–200, June 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Aly:1988:SAQ

[508] Emad-Eldin A. A. Aly. Strong approximations of quadratic sums of uniform spacings. Canadian Journal of Statistics = Revue canadienne de statistique, 16(2):201–207, June 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Seshadri:1988:EMB

[509] V. Seshadri. Exponential models, Brownian motion, and independence. Canadian Journal of Statistics = Revue canadienne de statistique, 16 (3):209–221, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Shoukri:1988:SPE

[510] M. M. Shoukri, I. U. H. Mian, and D. S. Tracy. Sampling properties of estimators of the log-logistic distribution with application to Canadian precipitation data. Canadian Journal of Statistics = Revue canadienne de statistique, 16(3):223–236, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Bravo:1988:ISE

[511] Gina Bravo and Brenda MacGibbon. Improved shrinkage estimators for the mean vector of a scale mixture of normals with unknown variance. Canadian Journal of Statistics = Revue canadienne de statistique, 16 (3):237–245, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Parsons:1988:SSP

[512] Van L. Parsons and Thu M. Hoang. Small sample properties of isotonic estimators — bias and mean squared error of linear combinations under normality. Canadian Journal of Statistics = Revue canadienne de statistique, 16(3):247–258, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Erasmus:1988:ADQ

[513] C. M. Erasmus and F. Lombard. Asymptotic distributions of quadratic forms occurring in changepoint problems. Canadian Journal of Statistics = Revue canadienne de statistique, 16(3):259–268, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Singh:1988:TMA

[514] Bahadur Singh and F. T. Wright. Two-moment approximations to some null distributions in order-restricted inference: Unequal sample sizes. Canadian Journal of Statistics = Revue canadienne de statistique, 16 (3):269–282, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Yatracos:1988:NCE

[515] Yannis G. Yatracos. A note on L_{1} consistent estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 16(3):283–292, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:1988:ISE

[516] Jeesen Chen and Jiunn T. Hwang. Improved set estimators for the coefficients of a linear model when the error distribution is spherically symmetric with unknown variances. Canadian Journal of Statistics = Revue canadienne de statistique, 16(3):293–299, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Kala:1988:RIE

[517] Radosław Kala and Krzysztof Kłaczyński. Recursive improvement of estimates in a Gauss-Markov model with linear restrictions. Canadian Journal of Statistics = Revue canadienne de statistique, 16(3):301–305, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Bennett:1988:CCD

[518] G. W. Bennett. Consultant’s corner: Determination of anaerobic threshold. Canadian Journal of Statistics = Revue canadienne de statistique, 16 (3):307–316, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Bertrand:1988:BRB

[519] Philip V. Bertrand. Book review: Theoretical epidemiology: Principles of Occurrence Research in Medicine. by Olli S. Miettinen. John Wiley & Sons, New York, 1985. xxi + 359 pp. ISBN 0-471-86769-1. Canadian Journal of Statistics = Revue canadienne de statistique, 16(3):317–318, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Stoyanov:1988:BRB

[520] Jordan M. Stoyanov. Book review: Markov chains. by David Freedman. Springer-Verlag, New York, Heidelberg, Approximating Countable Markov Chains, by David Freedman. Springer-Verlag, Brownian Motion and Diffusion, by David Freedman. Springer-Verlag, New York, Berlin, 1983. xiv + 382 pp. ISBN 0-387-9008-0. New York, Heidelberg, Berlin, 1983. x + 140 pp. ISBN 0-387-90804-8. Heidelberg, Berlin, 1983. xii + 231 pp. ISBN 0-387-90805-6. Canadian Journal of Statistics = Revue canadienne de statistique, 16(3):318–320, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Tomkins:1988:BRB

[521] R. J. Tomkins. Book review: Strong approximations in probability and statistics. by M. Csörgő and P. Revesz. Akademia Kiado, Budapest — Academic Press, New York, 1981. Canadian Journal of Statistics = Revue canadienne de statistique, 16(3):320–321, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Seshadri:1988:BRB

[522] V. Seshadri. Book review: An atlas of functions. by Jerome Spanier and Keith B. Oldham. Hemisphere Publishing Corporation, Washington, D. C., 1987. ix + 700 pp. ISBN 0-89116-573-8. Canadian Journal of Statistics = Revue canadienne de statistique, 16(3):321, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1988:A

[523] Anonymous. Addendum to [????]. Canadian Journal of Statistics = Revue canadienne de statistique, 16(3):323, September 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Greenwood:1988:CRM

[524] C. Greenwood and V. Farewell. A comparison of regression models for ordinal data in an analysis of transplanted-kidney function. Canadian Journal of Statistics = Revue canadienne de statistique, 16(4):325–335, December 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Winter:1988:PLE

[525] B. B. Winter and A. Földes. A product-limit estimator for use with length-biased data. Canadian Journal of Statistics = Revue canadienne de statistique, 16(4):337–355, December 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Aly:1988:CTO

[526] Emad-Eldin A. A. Aly. Comparing and testing order relations between percentile residual life functions. Canadian Journal of Statistics = Revue canadienne de statistique, 16(4):357–369, December 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Shao:1988:BVB

[527] Jun Shao. Bootstrap variance and bias estimation in linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 16(4):371– 382, December 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

OBrien:1988:IRR

[528] George L. O’Brien and Augustine Wong. An inequality for random replacement sampling plans. Canadian Journal of Statistics = Revue canadienne de statistique, 16(4):383–391, December 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Daffer:1988:TLL

[529] Pete Daffer and Ioana Schiopu-Kratina. L^{1}-tightness and the law of large numbers in D(IR). Canadian Journal of Statistics = Revue canadienne de statistique, 16(4):393–397, December 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Abdul-Al:1988:RCR

[530] K. I. Abdul-Al. On the rate of convergence of recursive kernel estimates of probability densities. Canadian Journal of Statistics = Revue canadienne de statistique, 16(4):399–409, December 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Tracy:1988:PMD

[531] D. S. Tracy and K. G. Jinadasa. Patterned matrix derivatives. Canadian Journal of Statistics = Revue canadienne de statistique, 16(4):411–418, December 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Giri:1988:SRT

[532] Narayan C. Giri. Some robust tests of independence in symmetrical multivariate distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 16(4):419–428, December 1988. ISSN 0319-5724 (print), 1708-945X (electronic).

Wright:1989:TTE

[533] F. T. Wright and James M. Guffey. Testing for a trend in exponential means: a two-moment approximation to the null distribution of the likelihood-ratio statistic. Canadian Journal of Statistics = Revue canadienne de statistique, 17(1):9–18, March 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Geisser:1989:PDT

[534] Seymour Geisser. Predictive discordancy tests for exponential observations. Canadian Journal of Statistics = Revue canadienne de statistique, 17(1):19–26, March 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Coleman:1989:RSS

[535] Rodney Coleman. Random sections of a sphere. Canadian Journal of Statistics = Revue canadienne de statistique, 17(1):27–39, March 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Small:1989:FRC

[536] Christopher G. Small. On functionals of random convex hulls. Canadian Journal of Statistics = Revue canadienne de statistique, 17(1):41– 43, March 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Hollander:1989:FCB

[537] Myles Hollander and Edsel Peña. Families of confidence bands for the survival function under the general random censorship model and the Koziol–Green model. Canadian Journal of Statistics = Revue canadienne de statistique, 17(1):59–74, March 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Romano:1989:DBC

[538] Joseph P. Romano. Do bootstrap confidence procedures behave well uniformly in P? Canadian Journal of Statistics = Revue canadienne de statistique, 17(1):75–80, March 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Krishnamoorthy:1989:IME

[539] K. Krishnamoorthy and A. K. Gupta. Improved minimax estimation of a normal precision matrix. Canadian Journal of Statistics = Revue canadienne de statistique, 17(1):91–102, March 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Young:1989:ECU

[540] Dean M. Young, Robert Pavur, and Virgil R. Marco. On the effect of correlation and unequal variances in detecting a spurious observation. Canadian Journal of Statistics = Revue canadienne de statistique, 17(1): 103–105, March 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Tracy:1989:PKP

[541] Derrick S. Tracy and Kankanam G. Jinadasa. Partitioned Kronecker products of matrices and applications. Canadian Journal of Statistics = Revue canadienne de statistique, 17(1):107–120, March 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1989:E

[542] Anonymous. Errata. Canadian Journal of Statistics = Revue canadienne de statistique, 17(1):121, March 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Tibshirani:1989:ESP

[543] Robert Tibshirani and Larry A. Wasserman. Errata: “Sensitive parameters”. Canadian Journal of Statistics = Revue canadienne de statistique, 17(1):121, 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Rivest:1989:UER

[544] Louis-Paul Rivest. De l’unicité des estimateurs robustes en régression lorsque le paramètre d’échelle et le paramètre de la régression sont estimés simultanément. (French) [On the uniqueness of robust regression estimators when the scale parameter and the regression parameter are estimated simultaneously]. Canadian Journal of Statistics = Revue canadienne de statistique, 17(2):141–153, June 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Diciccio:1989:APM

[545] Thomas J. Diciccio and Joseph P. Romano. The automatic percentile method: Accurate confidence limits in parametric models. Canadian Journal of Statistics = Revue canadienne de statistique, 17(2):155–169, June 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Dean:1989:MPI

[546] C. Dean, J. F. Lawless, and G. E. Willmot. A mixed Poisson–inverse-Gaussian regression model. Canadian Journal of Statistics = Revue canadienne de statistique, 17(2):171–181, June 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Donner:1989:TAD

[547] Allan Donner, George A. Wells, and Michael Eliasziw. On two approximations to the F -distribution: Application to testing for intraclass correlation in family studies. Canadian Journal of Statistics = Revue canadienne de statistique, 17(2):209–215, June 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Paul:1989:TES

[548] S. R. Paul. Test for the equality of several correlation coefficients. Canadian Journal of Statistics = Revue canadienne de statistique, 17(2):217– 227, June 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Cox:1989:SML

[549] D. R. Cox and N. Reid. On the stability of maximum-likelihood estimators of orthogonal parameters. Canadian Journal of Statistics = Revue canadienne de statistique, 17(2):229–233, June 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Devroye:1989:CDD

[550] Luc Devroye. Consistent deconvolution in density estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 17(2):235–239, June 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:1989:MEP

[551] Y. H. Wang. A multivariate extension of Poisson’s theorem. Canadian Journal of Statistics = Revue canadienne de statistique, 17(2):241–245, June 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Deo:1989:FCL

[552] C. M. Deo. A functional central limit theorem for the quadratic variation of a class of Gaussian random fields. Canadian Journal of Statistics = Revue canadienne de statistique, 17(2):247–251, June 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Reiser:1989:SSC

[553] B. Reiser and I. Guttman. Sample size choice for reliability verification in strength-stress models. Canadian Journal of Statistics = Revue canadienne de statistique, 17(3):253–259, September 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Knight:1989:LTA

[554] Keith Knight. Limit theory for autoregressive-parameter estimates in an infinite-variance random walk. Canadian Journal of Statistics = Revue canadienne de statistique, 17(3):261–278, September 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Chan:1989:NNS

[555] Ngai Hang Chan. On the nearly nonstationary seasonal time series. Canadian Journal of Statistics = Revue canadienne de statistique, 17 (3):279–284, September 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Kim:1989:EPM

[556] Peter T. Kim. Estimation of product moments of a stationary stochastic process with application to estimation of cumulants and cumulant spectral densities. Canadian Journal of Statistics = Revue canadienne de statistique, 17(3):285–299, September 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Caperaa:1989:OSE

[557] Philippe Capéraà. Ordres stochastiques et efficacité asymptotique des L-estimateurs. (French) [Stochastic orders and asymptotic efficiency of L-estimators]. Canadian Journal of Statistics = Revue canadienne de statistique, 17(3):301–310, September 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Kourouklis:1989:RBH

[558] Stavros Kourouklis. On the relation between Hodges-Lehmann efficiency and Pitman efficiency. Canadian Journal of Statistics = Revue canadienne de statistique, 17(3):311–318, September 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Lefebvre:1989:FPD

[559] Mario Lefebvre. First-passage densities of controlled Gaussian processes. Canadian Journal of Statistics = Revue canadienne de statistique, 17 (3):319–327, September 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Al-Hussaini:1989:MBE

[560] Ata N. Al-Hussaini and Robert J. Elliott. Markov bridges and enlarged filtrations. Canadian Journal of Statistics = Revue canadienne de statistique, 17(3):329–332, September 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Albert:1989:BAS

[561] James H. Albert and Patricia A. Pepple. A Bayesian approach to some overdispersion models. Canadian Journal of Statistics = Revue canadienne de statistique, 17(3):333–344, September 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Shah:1989:COC

[562] K. R. Shah and B. K. Sinha. On the choice of optimality criteria in comparing statistical designs. Canadian Journal of Statistics = Revue canadienne de statistique, 17(3):345–348, September 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Remillard:1989:LIL

[563] Bruno Rémillard and Donald A. Dawson. Laws of the iterated logarithm and large deviations for a class of diffusion processes. Canadian Journal of Statistics = Revue canadienne de statistique, 17(4):349–376, December 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Yu:1989:PCW

[564] Kai Fun Yu. On a problem of counting by weighing. Canadian Journal of Statistics = Revue canadienne de statistique, 17(4):377–389, December 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Christensen:1989:SSR

[565] Ronald Christensen. Sequential sampling without recall from a Dirichlet process. Canadian Journal of Statistics = Revue canadienne de statistique, 17(4):391–397, December 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Latour:1989:DAA

[566] Alain Latour and Roch Roy. Distribution asymptotique des autocorrélations d’un processus saisonnier non stationnaire. (French) [Asymptotic distribution of the autocorrelations of a seasonal non-stationary process]. Canadian Journal of Statistics = Revue canadienne de statistique, 17(4):399–417, December 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Guttorp:1989:ADH

[567] P. Guttorp and R. A. Lockhart. On the asymptotic distributions of high-order spacings statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 17(4):419–426, December 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:1989:CND

[568] Ming Chung Liu and Robert L. Taylor. A consistent nonparametric density estimator for the deconvolution problem. Canadian Journal of Statistics = Revue canadienne de statistique, 17(4):427–438, December 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Fang:1989:IPS

[569] Kai-Tai Tai Fang and Jia-Juang Juang Liang. Inequalities for the partial sums of elliptical order statistics related to genetic selection. Canadian Journal of Statistics = Revue canadienne de statistique, 17(4):439–446, December 1989. ISSN 0319-5724 (print), 1708-945X (electronic).

Sprott:1990:IEL

[570] D. A. Sprott. Inferential estimation, likelihood, and linear pivotals. Canadian Journal of Statistics = Revue canadienne de statistique, 18 (1):1–10, March 1990. ISSN 0319-5724 (print), 1708-945X (electronic). See discussion [571] and response [572].

Barnard:1990:DSC

[571] G. A. Barnard, O. E. Barndorff-Nielsen, and D. A. S. Fraser. Discussion /commentaries. Canadian Journal of Statistics = Revue canadienne de statistique, 18(1):10–15, March 1990. ISSN 0319-5724 (print), 1708-945X (electronic). See [570].

Sprott:1990:RDS

[572] D. A. Sprott. Response to the discussion /réplique aux commentaires. Canadian Journal of Statistics = Revue canadienne de statistique, 18 (1):15, March 1990. ISSN 0319-5724 (print), 1708-945X (electronic). See [570].

Balanda:1990:KS

[573] Kevin P. Balanda and H. L. MacGillivray. Kurtosis and spread. Canadian Journal of Statistics = Revue canadienne de statistique, 18(1):17– 30, March 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Ki:1990:MSE

[574] Fanny Ki and Kam-Wah Tsui. Multiple-shrinkage estimators of means in exponential families. Canadian Journal of Statistics = Revue canadienne de statistique, 18(1):31–46, March 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Wiens:1990:MVE

[575] Douglas P. Wiens. Minimax-variance L- and R-estimators of location. Canadian Journal of Statistics = Revue canadienne de statistique, 18(1): 47–57, March 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Kubokawa:1990:EPG

[576] Tatsuya Kubokawa. Estimating powers of the generalized variance under the Pitman closeness criterion. Canadian Journal of Statistics = Revue canadienne de statistique, 18(1):59–62, March 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Conaway:1990:PCC

[577] Mark Conaway, Carolyn Pillers, Tim Robertson, and James Sconing. The power of the circular cone test: a noncentral chi-bar-squared distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 18(1): 63–70, March 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Hudon:1990:TMM

[578] Georges Hudon. Tests for the mean of a multivariate normal population. Canadian Journal of Statistics = Revue canadienne de statistique, 18(1): 71–77, March 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Shaked:1990:CEC

[579] Moshe Shaked and Y. L. Tong. Comparison of experiments for a class of positively dependent random variables. Canadian Journal of Statistics = Revue canadienne de statistique, 18(1):79–86, March 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Johnston:1990:CCC

[580] Bruce Johnston. Consultant’s corners: The case of the random telephone numbers. Canadian Journal of Statistics = Revue canadienne de statistique, 18(1):87–92, March 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Whitmore:1990:BCC

[581] G. A. Whitmore and Jane F. Gentleman. Beer chemistry and Canadians’ beer preferences. Canadian Journal of Statistics = Revue canadienne de statistique, 18(2):94–98, June 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Carmichael:1990:APC

[582] Jean-Pierre Carmichael, Gaéatan Daigle, and Louis-Paul Rivest. Analyse des préférences des Canadiens en matière de bière. (French) [Analysis of preferences of Canadians in the matter of beer]. Canadian Journal of Statistics = Revue canadienne de statistique, 18(2):99–108, June 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:1990:RMR

[583] Bing Li and A. John Petkau. A regression model with random effects for beer chemistry and Canadians’ beer preferences. Canadian Journal of Statistics = Revue canadienne de statistique, 18(2):108–121, June 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Whitmore:1990:SEC

[584] G. A. Whitmore. Section Editors’ comments. Canadian Journal of Statistics = Revue canadienne de statistique, 18(2):121–125, June 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Dzieciolowski:1990:ACI

[585] Krzysztof Dzieciolowski and William H. Ross. Assessing case influence on confidence intervals in nonlinear regression. Canadian Journal of Statistics = Revue canadienne de statistique, 18(2):127–139, June 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Shi:1990:WSR

[586] Xiquan Shi, C. F. J. Wu, and Jiahua Chen. Weak and strong representations for quantile processes from finite populations with application to simulation size in resampling inference. Canadian Journal of Statistics = Revue canadienne de statistique, 18(2):141–148, June 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Martin:1990:UJE

[587] Michael A. Martin. On using the jackknife to estimate quantile variance. Canadian Journal of Statistics = Revue canadienne de statistique, 18(2): 149–153, June 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Feuerverger:1990:ERE

[588] Andrey Feuerverger. An efficiency result for the empirical characteristic function in stationary time-series models. Canadian Journal of Statistics = Revue canadienne de statistique, 18(2):155–161, June 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Ducharme:1990:ECL

[589] Gilles R. Ducharme and Philip Milasevic. Estimating the concentration of the Langevin distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 18(2):163–169, June 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Dey:1990:ESP

[590] Dipak K. Dey. Estimation of scale parameters in mixture distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 18(2): 171–178, June 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Perron:1990:EEC

[591] François Perron. Equivariant estimators of the covariance matrix. Canadian Journal of Statistics = Revue canadienne de statistique, 18(2):179– 182, June 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Wasserman:1990:BFS

[592] Larry A. Wasserman. Belief functions and statistical inference. Canadian Journal of Statistics = Revue canadienne de statistique, 18(3):183–196, September 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Wand:1990:GBK

[593] Matthew P. Wand and William R. Schucany. Gaussian-based kernels. Canadian Journal of Statistics = Revue canadienne de statistique, 18 (3):197–204, September 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Meloche:1990:ABM

[594] J. Meloche. Asymptotic behaviour of the mean integrated squared error of kernel density estimators for dependent observations. Canadian Journal of Statistics = Revue canadienne de statistique, 18(3):205–211, September 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Fraisse:1990:CNA

[595] Anne-Marie Fraisse, Jean-Pierre Raoult, Christian Robert, and Madeleine Roy. Une condition nécessaire d’admissibilité et ses conséquences sur les estimateurs à rétrécisseur de la moyenne d’un vecteur normal. (French) [A necessary condition for eligibility and its consequences for the estimators of the shrinkage of the mean of a normal vector]. Canadian Journal of Statistics = Revue canadienne de statistique, 18(3):213–220, September 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Venter:1990:ERR

[596] J. H. Venter and S. J. Steel. Estimating risk reduction in Stein estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 18 (3):221–232, September 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Joe:1990:WSL

[597] Harry Joe. A winning strategy for lotto games? Canadian Journal of Statistics = Revue canadienne de statistique, 18(3):233–244, September 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:1990:MMW

[598] Y. I. Chen and D. A. Wolfe. Modifications of the Mack-Wolfe umbrella tests for a generalized Behrens-Fisher problem. Canadian Journal of Statistics = Revue canadienne de statistique, 18(3):245–253, September 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:1990:SAN

[599] Suojin Wang. Saddlepoint approximations for nonlinear statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 18(3):255– 264, September 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Stephens:1990:ACF

[600] M. A. Stephens. Asymptotic calculations of functions of expected values and covariances of order statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 18(3):265–270, September 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Anderson:1990:MSA

[601] Oliver D. Anderson. Moments of the sampled autocovariances and autocorrelations for a Gaussian white-noise process. Canadian Journal of Statistics = Revue canadienne de statistique, 18(3):271–284, September 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Baksalary:1990:CSP

[602] Jerzy K. Baksalary and Simo Puntanen. A complete solution to the problem of robustness of Grubbs’s test. Canadian Journal of Statistics = Revue canadienne de statistique, 18(3):285–287, September 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Viana:1990:TDU

[603] M. Viana and V. Farewell. A test for diagnostic utility. Canadian Journal of Statistics = Revue canadienne de statistique, 18(4):289–295, December 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Leger:1990:BAE

[604] Christian Léger and Joseph P. Romano. Bootstrap adaptive estimation: the trimmed-mean example. Canadian Journal of Statistics = Revue canadienne de statistique, 18(4):297–314, December 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Caperaa:1990:CDO

[605] Philippe Capéraà and Christian Genest. Concepts de dépendance et ordres stochastiques pour des lois bidimensionnelles. (French) [Dependence concepts and stochastic orders for two-dimensional laws]. Canadian Journal of Statistics = Revue canadienne de statistique, 18(4):315–326, December 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Shao:1990:OWL

[606] Jun Shao. Ordinary and weighted least-squares estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 18(4):327–336, December 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Cohen:1990:UTS

[607] Arthur Cohen and Harold B. Sackrowitz. Unbiased tests for some onesided testing problems. Canadian Journal of Statistics = Revue canadienne de statistique, 18(4):337–346, December 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Albert:1990:BTT

[608] James H. Albert. A Bayesian test for a two-way contingency table using independence priors. Canadian Journal of Statistics = Revue canadienne de statistique, 18(4):347–363, December 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Burnett:1990:CCCa

[609] Richard Burnett, Mark Raizenne, and Daniel Krewski. Consultant’s corner/le coin du consultant: Acute health effects of transported air pollution: a study of children attending a residential summer camp. Canadian Journal of Statistics = Revue canadienne de statistique, 18(4):367–373, December 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Kan:1990:CCC

[610] Lisa Kan, Barry Wiggs, Mark Irwin, and Irene Yee. Consultant’s corner/ le coin du consultant: Acute respiratory health effects in asthmatic and nonasthmatic children associated with short-term exposure to air pollutants. Canadian Journal of Statistics = Revue canadienne de statistique, 18(4):373–377, December 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Brasher:1990:CCC

[611] Penny Brasher, Sheilah Hogg-Johnson, and Sharon-Lise Normand. Consultant’s corner/le coin du consultant: Analysis of the Lake Couchiching data. Canadian Journal of Statistics = Revue canadienne de statistique, 18(4):377–382, December 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Baskerville:1990:CCC

[612] Jon C. Baskerville and Larry Stitt. Consultant’s corner/le coin du consultant: An analysis of associations between camper characteristics and response to the first pollution episode in the lake couchiching study. Canadian Journal of Statistics = Revue canadienne de statistique, 18 (4):383–388, December 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Burnett:1990:CCCb

[613] Richard Burnett, Daniel Krewski, and Jennifer Shedden. Consultant’s corner/le coin du consultant: Adjusting for temporal trends in the statistical analysis of the effects of air pollution on pulmonary function in children. Canadian Journal of Statistics = Revue canadienne de statistique, 18(4):389–395, December 1990. ISSN 0319-5724 (print), 1708-945X (electronic).

Manchester:1991:TCG

[614] Lise Manchester. A technique for comparing graphical methods. Canadian Journal of Statistics = Revue canadienne de statistique, 19(1):1–22, March 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Gijbels:1991:DEU

[615] I. Gijbels and V. K. Klonias. Density estimation under the Koziol-Green model of censoring by penalized likelihood methods. Canadian Journal of Statistics = Revue canadienne de statistique, 19(1):23–38, March 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Angers:1991:RHB

[616] Jean-François Angers and James O. Berger. Robust hierarchical Bayes estimation of exchangeable means. Canadian Journal of Statistics = Revue canadienne de statistique, 19(1):39–56, March 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Sivaganesan:1991:SSP

[617] S. Sivaganesan. Sensitivity of some posterior summaries when the prior is unimodal with specified quantiles. Canadian Journal of Statistics = Revue canadienne de statistique, 19(1):57–65, March 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Bagchi:1991:LAP

[618] Parthasarathy Bagchi and Joseph B. Kadane. Laplace approximations to posterior moments and marginal distributions on circles, spheres, and cylinders. Canadian Journal of Statistics = Revue canadienne de statistique, 19(1):67–77, March 1991. ISSN 0319-5724 (print), 1708-945X (electronic). See erratum [1597].

Shoukri:1991:CLM

[619] M. M. Shoukri, I. U. H. Mian, and D. S. Tracy. Correlated linear models for the analysis of familial correlations. Canadian Journal of Statistics = Revue canadienne de statistique, 19(1):79–91, March 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Beirlant:1991:ANF

[620] J. Beirlant, P. Janssen, and N. Veraverbeke. On the asymptotic normality of functions of uniform spacings. Canadian Journal of Statistics = Revue canadienne de statistique, 19(1):93–101, March 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Perron:1991:TIA

[621] François Perron. Testing independence with additional information. Canadian Journal of Statistics = Revue canadienne de statistique, 19(1): 103–108, March 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Hamilton:1991:CCC

[622] D. C. Hamilton and M. L. Lesperance. Consultant’s corner: a consulting problem involving bivariate acceptance sampling by variables. Canadian Journal of Statistics = Revue canadienne de statistique, 19(1):109–117, March 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1991:E

[623] Anonymous. Errata. Canadian Journal of Statistics = Revue canadienne de statistique, 19(1):119, March 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1991:AR

[624] Anonymous. Author’s revision. Canadian Journal of Statistics = Revue canadienne de statistique, 19(1):119–120, March 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Charras:1991:BAP

[625] Alec Charras and Constance Van Eeden. Bayes and admissibility properties of estimators in truncated parameter spaces. Canadian Journal of Statistics = Revue canadienne de statistique, 19(2):121–134, June 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Gastwirth:1991:BAS

[626] Joseph L. Gastwirth, Wesley O. Johnson, and Dana M. Reneau. Bayesian analysis of screening data: Application to AIDS in blood donors. Canadian Journal of Statistics = Revue canadienne de statistique, 19(2):135– 150, June 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Meloche:1991:ESD

[627] J. Meloche. Estimation of a symmetric density. Canadian Journal of Statistics = Revue canadienne de statistique, 19(2):151–164, June 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Greenwood:1991:EEN

[628] P. E. Greenwood and W. Wefelmeyer. Efficient estimation in a nonlinear counting-process regression model. Canadian Journal of Statistics = Revue canadienne de statistique, 19(2):165–178, June 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Wiens:1991:DMF

[629] Douglas P. Wiens and K. H. Eden Wu. Distributions minimizing Fisher information for scale in Kolmogorov neighbourhoods. Canadian Journal of Statistics = Revue canadienne de statistique, 19(2):179–190, June 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Bolfarine:1991:FPP

[630] Heleno Bolfarine. Finite-population prediction under error-in-variables superpopulation models. Canadian Journal of Statistics = Revue canadienne de statistique, 19(2):191–207, June 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

McDonald:1991:ARS

[631] D. McDonald. On the asymptotics of randomness statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 19(2):209–217, June 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Bouzar:1991:CRS

[632] Nadjib Bouzar. Convergence results for strict C-sequences. Canadian Journal of Statistics = Revue canadienne de statistique, 19(2):219–227, June 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Letac:1991:CEA

[633] Gérard Letac. Contre-exemple au théorème de P. C. Consul sur la factorisation des distributions de Poisson généralisées. (French) [Counterexample to the theorem of P. C. Consul on factoring generalized Poisson distributions]. Canadian Journal of Statistics = Revue canadienne de statistique, 19(2):229–232, June 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Routledge:1991:CCU

[634] R. D. Routledge. Consultant’s corner: Using time lags in estimating anaerobic threshold. Canadian Journal of Statistics = Revue canadienne de statistique, 19(2):233–236, June 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Jones:1991:CCD

[635] M. C. Jones and M. S. Handcock. Consultant’s corner: Determination of anaerobic threshold: What anaerobic threshold? Canadian Journal of Statistics = Revue canadienne de statistique, 19(2):236–239, June 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Gentleman:1991:ENC

[636] Jane F. Gentleman and G. A. Whitmore. Estimation of the need for child care in Canada. Canadian Journal of Statistics = Revue canadienne de statistique, 19(3):242–249, September 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Gee:1991:ENC

[637] Ellen M. Gee and James G. Mcdaniel. Estimating the need for child care: the Canadian case. Canadian Journal of Statistics = Revue canadienne de statistique, 19(3):249–265, September 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Struthers:1991:EMU

[638] C. A. Struthers. Estimation of met and unmet child-care needs in Canada based on the family history survey. Canadian Journal of Statistics = Revue canadienne de statistique, 19(3):265–279, September 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1991:SEC

[639] Anonymous. Section editors’ comments. Canadian Journal of Statistics = Revue canadienne de statistique, 19(3):279–282, September 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Conaway:1991:CCT

[640] Mark Conaway, Carolyn Pillers, Tim Robertson, and Jim Sconing. A circular-cone test for testing homogeneity against a simple tree order. Canadian Journal of Statistics = Revue canadienne de statistique, 19 (3):283–296, September 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Dykstra:1991:ANC

[641] Richard Dykstra. Asymptotic normality for chi-bar-square distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 19 (3):297–306, September 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Lopuhaa:1991:MEL

[642] Hendrik P. Lopuhaä. Multivariate τ-estimators for location and scatter. Canadian Journal of Statistics = Revue canadienne de statistique, 19 (3):307–321, September 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:1991:SSV

[643] Hung Chen and Keh-Wei Chen. Selection of the splined variables and convergence rates in a partial spline model. Canadian Journal of Statistics = Revue canadienne de statistique, 19(3):323–339, September 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Small:1991:RCB

[644] Christopher G. Small. Reconstructing convex bodies from random projected images. Canadian Journal of Statistics = Revue canadienne de statistique, 19(4):341–347, December 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Karunamuni:1991:WSU

[645] R. J. Karunamuni and Song Yang. Weak and strong uniform consistency rates of kernel density estimates for randomly censored data. Canadian Journal of Statistics = Revue canadienne de statistique, 19(4):349–359, December 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Houde:1991:LEA

[646] Louis Houde and Yves Lepage. L’effet de l’erreur d’arrondi sur l’estimateur d’une densité par la méthode du noyau. (French) [The effect of rounding error on the density estimate by a kernel method]. Canadian Journal of Statistics = Revue canadienne de statistique, 19(4):361–369, December 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Tran:1991:MVK

[647] Lanh Tat Tran. On multivariate variable-kernel density estimates for time series. Canadian Journal of Statistics = Revue canadienne de statistique, 19(4):371–387, December 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Plante:1991:ICS

[648] André Plante. An inclusion-consistent solution to the problem of absurd confidence statements: 1. Consistent exact confidence-interval estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 19(4):389–397, December 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Carlin:1991:INB

[649] Bradley P. Carlin and Nicholas G. Polson. Inference for nonconjugate Bayesian Models using the Gibbs sampler. Canadian Journal of Statistics = Revue canadienne de statistique, 19(4):399–405, December 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Krewski:1991:CUT

[650] Daniel Krewski, Robert T. Smythe, Karen Y. Fung, and Richard Burnett. Conditional and unconditional tests with historical controls. Canadian Journal of Statistics = Revue canadienne de statistique, 19(4):407–423, December 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Gardiner:1991:SAD

[651] J. C. Gardiner and V. Susarla. Some asymptotic distribution results in time-sequential estimation of the mean exponential survival time. Canadian Journal of Statistics = Revue canadienne de statistique, 19(4):425– 436, December 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Seshadri:1991:FMN

[652] V. Seshadri. Finite mixtures of natural exponential families. Canadian Journal of Statistics = Revue canadienne de statistique, 19(4):437–445, December 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Ali:1991:DCC

[653] Mir M. Ali and Anwarul H. Joarder. Distribution of the correlation coefficient for the class of bivariate elliptical models. Canadian Journal of Statistics = Revue canadienne de statistique, 19(4):447–452, December 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1991:AC

[654] Anonymous. Author’s corrections. Canadian Journal of Statistics = Revue canadienne de statistique, 19(4):453, December 1991. ISSN 0319-5724 (print), 1708-945X (electronic).

Armitage:1992:STC

[655] Peter Armitage. Some topics of current interest in clinical trials. Canadian Journal of Statistics = Revue canadienne de statistique, 20(1):1–8, March 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Masry:1992:GDD

[656] Elias Masry and John A. Rice. Gaussian deconvolution via differentiation. Canadian Journal of Statistics = Revue canadienne de statistique, 20(1):9–21, March 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Fortier:1992:BLC

[657] Jean-J. Fortier. Best linear corrector of classification estimates of proportions of objects in several unknown classes. Canadian Journal of Statistics = Revue canadienne de statistique, 20(1):23–33, March 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Angers:1992:HBC

[658] Jean-François Angers and Mohan Delampady. Hierarchical Bayesian curve fitting and smoothing. Canadian Journal of Statistics = Revue canadienne de statistique, 20(1):35–49, March 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Czado:1992:OPL

[659] Claudia Czado and Thomas J. Santner. Orthogonalizing parametric link transformation families in binary regression analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 20(1):51–61, March 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Ferguson:1992:APC

[660] H. Ferguson. Asymptotic properties of a conditional maximum-likelihood estimator. Canadian Journal of Statistics = Revue canadienne de statistique, 20(1):63–75, March 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Einmahl:1992:LTG

[661] John H. J. Einmahl and Alex J. Koning. Limit theorems for a general weighted process under random censoring. Canadian Journal of Statistics = Revue canadienne de statistique, 20(1):77–89, March 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Berkowitz:1992:CCC

[662] Jonathan Berkowitz. Consultant’s corner/le coin du consultant: Interferon in multiple sclerosis. Canadian Journal of Statistics = Revue canadienne de statistique, 20(1):91–96, March 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Dean:1992:ITM

[663] C. Dean and M. L. Lesperance. Interferon as a treatment for multiple sclerosis? Canadian Journal of Statistics = Revue canadienne de statistique, 20(1):96–102, March 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Schwarz:1992:MSD

[664] Carl J. Schwarz, Jeff A. Sloan, Linda R. Neden, and John F. Brewster. Multiple-sclerosis dataset. Canadian Journal of Statistics = Revue canadienne de statistique, 20(1):102–108, March 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Barndorff-Nielsen:1992:MDG

[665] O. E. Barndorff-Nielsen, P. Blæsild, and V. Seshadri. Multivariate distributions with generalized inverse Gaussian marginals, and associated Poisson mixtures. Canadian Journal of Statistics = Revue canadienne de statistique, 20(2):109–120, June 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Perron:1992:LDL

[666] Pierre Perron. The limiting distribution of the least-squares estimator in nearly integrated seasonal models. Canadian Journal of Statistics = Revue canadienne de statistique, 20(2):121–134, June 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Sitter:1992:CTB

[667] R. R. Sitter. Comparing three bootstrap methods for survey data. Canadian Journal of Statistics = Revue canadienne de statistique, 20(2):135– 154, June 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Fan:1992:DSD

[668] Jianqing Fan. Deconvolution with supersmooth distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 20(2):155–169, June 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Abrahamowicz:1992:NDE

[669] Michał Abrahamowicz, Antonio Ciampi, and James O. Ramsay. Nonparametric density estimation for censored survival data: Regressionspline approach. Canadian Journal of Statistics = Revue canadienne de statistique, 20(2):171–185, June 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Yatracos:1992:PMS

[670] Yannis G. Yatracos. On prediction and mean squared error. Canadian Journal of Statistics = Revue canadienne de statistique, 20(2):187–200, June 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Routledge:1992:RCF

[671] R. D. Routledge. Resolving the conflict over Fisher’s exact test. Canadian Journal of Statistics = Revue canadienne de statistique, 20(2):201–209, June 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:1992:STM

[672] Michael Z. F. Li and R. J. Tomkins. Stability theorems for maxima of dependent sequences. Canadian Journal of Statistics = Revue canadienne de statistique, 20(2):211–219, June 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Ruzzante:1992:CCC

[673] Daniel E. Ruzzante and Roger W. Doyle. Consultant’s corner/le coin du consultant: Behaviour, growth, and intensity of social interaction among fish. Canadian Journal of Statistics = Revue canadienne de statistique, 20(2):221–227, June 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Hamilton:1992:AFB

[674] David C. Hamilton. Analysis of fish behaviour data. Canadian Journal of Statistics = Revue canadienne de statistique, 20(2):228–233, June 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Raboud:1992:EEG

[675] Janet Raboud and Melania Pintilie. The effect of environment on the growth of fish. Canadian Journal of Statistics = Revue canadienne de statistique, 20(2):233–239, June 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Daigle:1992:RB

[676] Gaétan Daigle and Louis-Paul Rivest. A robust biplot. Canadian Journal of Statistics = Revue canadienne de statistique, 20(3):241–255, September 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Bilodeau:1992:FSA

[677] Martin Bilodeau. Fourier smoother and additive models. Canadian Journal of Statistics = Revue canadienne de statistique, 20(3):257–269, September 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Tingley:1992:SSI

[678] M. A. Tingley. Small-sample intervals for regression. Canadian Journal of Statistics = Revue canadienne de statistique, 20(3):271–280, September 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Yu:1992:GMF

[679] Qi Qing Yu. A general method of finding a minimax estimator of a distribution function when no equalizer rule is available. Canadian Journal of Statistics = Revue canadienne de statistique, 20(3):281–290, September 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Chow:1992:SEH

[680] Mo Suk Chow and Duncan K. H. Fong. Simultaneous estimation of the Hardy-Weinberg proportions. Canadian Journal of Statistics = Revue canadienne de statistique, 20(3):291–296, September 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Geisser:1992:CS

[681] Seymour Geisser. On the curtailment of sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 20(3):297–309, September 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Viana:1992:BEP

[682] M. A. G. Viana and V. Ramakrishnan. Bayesian estimates of predictive value and related parameters of a diagnostic test. Canadian Journal of Statistics = Revue canadienne de statistique, 20(3):311–321, September 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:1992:ZOL

[683] H. Wang and R. J. Tomkins. A zero-one law for large order statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 20 (3):323–334, September 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Maymin:1992:TRH

[684] Zakhar Maymin and Sam Gutmann. Testing retrospective hypotheses. Canadian Journal of Statistics = Revue canadienne de statistique, 20 (3):335–345, September 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Shah:1992:BDA

[685] K. R. Shah and Ashish Das. Binary designs are not always the best. Canadian Journal of Statistics = Revue canadienne de statistique, 20 (3):347–351, September 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Brown:1992:AML

[686] Lawrence D. Brown, Mo Suk Chow, and Duncan K. H. Fong. On the admissibility of the maximum-likelihood estimator of the binomial variance. Canadian Journal of Statistics = Revue canadienne de statistique, 20(4):353–358, December 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Rao:1992:LRE

[687] C. Radhakrishna Rao and L. C. Zhao. Linear representation of Mestimates in linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 20(4):359–368, December 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Nsiri:1992:IAE

[688] Saïd Nsiri and Roch Roy. On the identification of ARMA echelon-form models. Canadian Journal of Statistics = Revue canadienne de statistique, 20(4):369–386, December 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

OReilly:1992:GFI

[689] Federico J. O’Reilly and Raúl Rueda. Goodness of fit for the inverse Gaussian distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 20(4):387–397, December 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Gu:1992:EEB

[690] Minggao Gu. On the Edgeworth expansion and bootstrap approximation for the Cox regression model under random censorship. Canadian Journal of Statistics = Revue canadienne de statistique, 20(4):399–414, December 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Shao:1992:CLS

[691] Jun Shao. Consistency of least-squares estimator and its jackknife variance estimator in nonlinear models. Canadian Journal of Statistics = Revue canadienne de statistique, 20(4):415–428, December 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Ghosh:1992:AEC

[692] Sucharita Ghosh and Frits H. Ruymgaart. Applications of empirical characteristic functions in some multivariate problems. Canadian Journal of Statistics = Revue canadienne de statistique, 20(4):429–440, December 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Perron:1992:MME

[693] François Perron. Monotonic minimax estimators of a 2 × 2 covariance matrix. Canadian Journal of Statistics = Revue canadienne de statistique, 20(4):441–449, December 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Lavine:1992:BMC

[694] Michael Lavine and Mike West. A Bayesian method for classification and discrimination. Canadian Journal of Statistics = Revue canadienne de statistique, 20(4):451–461, December 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Small:1992:CCR

[695] Christopher G. Small. A counterexample to a conjecture on random shadows. Canadian Journal of Statistics = Revue canadienne de statistique, 20(4):463–468, December 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Alvo:1992:GMR

[696] Mayer Alvo and Kadir Ertas. Graphical methods for ranking data. Canadian Journal of Statistics = Revue canadienne de statistique, 20(4):469– 482, December 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Bergerud:1992:FIR

[697] Wendy Bergerud. Factors influencing root growth capacity of spruce seedlings: Case-study description. Canadian Journal of Statistics = Revue canadienne de statistique, 20(4):483–487, December 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Bernier-Cardou:1992:FIR

[698] Michéle Bernier-Cardou and Christian Genest. Factors influencing root growth capacity of spruce seedlings: Report on a statistical analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 20 (4):488–501, December 1992. ISSN 0319-5724 (print), 1708-945X (electronic).

Brillinger:1993:DRS

[699] David R. Brillinger. The digital rainbow: Some history and applications of numerical spectrum analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 21(1):1–19, March 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Farewell:1993:SCA

[700] V. T. Farewell, R. Viveros, and D. A. Sprott. Statistical consequences of an adaptive treatment allocation in a clinical trial. Canadian Journal of Statistics = Revue canadienne de statistique, 21(1):21–27, March 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Mak:1993:NME

[701] Tak K. Mak and Anthony Kuk. A new method for estimating finitepopulation quantiles using auxiliary information. Canadian Journal of Statistics = Revue canadienne de statistique, 21(1):29–38, March 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Martin:1993:TFR

[702] Louise Martin, Raymond Leblanc, and Nguyen Ky Toan. Tables for the Friedman rank test. Canadian Journal of Statistics = Revue canadienne de statistique, 21(1):39–43, March 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Kubokawa:1993:ENP

[703] T. Kubokawa, C. P. Robert, and A. K. Md. E. Saleh. Estimation of noncentrality parameters. Canadian Journal of Statistics = Revue canadienne de statistique, 21(1):45–57, March 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Wiens:1993:DAL

[704] Douglas P. Wiens. Designs for approximately linear regression: Maximizing the minimum coverage probability of confidence ellipsoids. Canadian Journal of Statistics = Revue canadienne de statistique, 21(1):59– 70, March 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Bull:1993:CEI

[705] Shelley B. Bull and Allan Donner. A characterization of the efficiency of individualized logistic regressions. Canadian Journal of Statistics = Revue canadienne de statistique, 21(1):71–78, March 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Swanepoel:1993:MBE

[706] J. W. H. Swanepoel and C. F. De Beer. A modified bootstrap estimator for the mean of an asymmetric distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 21(1):79–87, March 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Ghosh:1993:PMP

[707] J. K. Ghosh and Rahul Mukerjee. On priors that match posterior and frequentist distribution functions. Canadian Journal of Statistics = Revue canadienne de statistique, 21(1):89–96, March 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Fick:1993:DIQ

[708] Gordon Hilton Fick. Detecting the interaction of quantity and quality: a study of Fisher’s analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 21(1):97–106, March 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Sivaganesan:1993:RBA

[709] S. Sivaganesan and James Berger. Robust Bayesian analysis of the binomial empirical Bayes problem. Canadian Journal of Statistics = Revue canadienne de statistique, 21(1):107–119, March 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Solow:1993:MWQ

[710] Andrew R. Solow and Arthur G. Gaines. Mapping water quality by local scoring. Canadian Journal of Statistics = Revue canadienne de statistique, 21(2):123–130, June 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Cheah:1993:SAE

[711] P. K. Cheah, D. A. S. Fraser, and N. Reid. Some alternatives to Edgeworth. Canadian Journal of Statistics = Revue canadienne de statistique, 21(2):131–138, June 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Lindsay:1993:UEI

[712] Bruce G. Lindsay and Kathryn Roeder. Uniqueness of estimation and identifiability in mixture models. Canadian Journal of Statistics = Revue canadienne de statistique, 21(2):139–147, June 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Nakamura:1993:UEP

[713] Miguel Nakamura and Víctor Pérez-Abreu. Use of an empirical probability generating function for testing a Poisson model. Canadian Journal of Statistics = Revue canadienne de statistique, 21(2):149–156, June 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Adjengue:1993:MVP

[714] Luc D. Adjengue and Marc Moore. La méthode de vraisemblance pour les processus linéaires spatiaux définis sur un treillis triangulaire. (French) [The likelihood method for linear processes spatially defined on a triangular trellis]. Canadian Journal of Statistics = Revue canadienne de statistique, 21(2):157–179, June 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Datta:1993:RBB

[715] Somnath Datta and William P. McCormick. Regeneration-based bootstrap for Markov chains. Canadian Journal of Statistics = Revue canadienne de statistique, 21(2):181–193, June 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Ruggeri:1993:ISP

[716] Fabrizio Ruggeri and Larry Wasserman. Infinitesimal sensitivity of posterior distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 21(2):195–203, June 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Crisp:1993:NUE

[717] Adam Crisp and Jim Burridge. A note on the uniqueness of M-estimators in robust regression. Canadian Journal of Statistics = Revue canadienne de statistique, 21(2):205–208, June 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Eastwood:1993:SNM

[718] Vera R. Eastwood. Some nonparametric methods for changepoint problems. Canadian Journal of Statistics = Revue canadienne de statistique, 21(2):209–222, June 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Murdoch:1993:PDM

[719] D. J. Murdoch and A. J. Benjamin. The probability of degenerate maximum-likelihood estimates with randomly truncated survival data. Canadian Journal of Statistics = Revue canadienne de statistique, 21 (2):223–231, June 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Krishnamoorthy:1993:UAR

[720] K. Krishnamoorthy and D. Raghavarao. Untruthful answering in repeated randomized response procedures. Canadian Journal of Statistics = Revue canadienne de statistique, 21(2):233–236, June 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Ali:1993:AAD

[721] M. M. Ali and A. H. Joarder. Author’s addendum: “Distribution of the correlation coefficient for the class of bivariate elliptical models” [Canad. J. Statist. 19 (1991), no. 4, 447–452; MR1166850 (93e:62141)]. Canadian Journal of Statistics = Revue canadienne de statistique, 21(2):237, 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1993:AAa

[722] Anonymous. Authors’ addendum. Canadian Journal of Statistics = Revue canadienne de statistique, 21(2):237, June 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

DeFalguerolles:1993:CCV

[723] Antoine De Falguerolles and Saïd Jmel. Un critère de choix de variables en analyse en composantes principales fondé sur des modèles graphiques gaussiens particuliers. (French) [A criterion for variable selection in the analysis of principal components based on particular graphical Gaussian models]. Canadian Journal of Statistics = Revue canadienne de statistique, 21(3):239–256, September 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Zielinski:1993:IEI

[724] Jan M. Zielinski, David B. Wolfson, Lakshmi Nilakantan, and Christian Confavreux. Isotonic estimation of the intensity of a nonhomogeneous Poisson process: the multiple realization setup. Canadian Journal of Statistics = Revue canadienne de statistique, 21(3):257–268, September 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Chukova:1993:CPD

[725] S. Chukova, B. Dimitrov, and Z. Khalil. A characterization of probability distributions similar to the exponential. Canadian Journal of Statistics = Revue canadienne de statistique, 21(3):269–276, September 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Mathai:1993:REG

[726] A. M. Mathai. The residual effect of a growth-decay mechanism and the distributions of covariance structures. Canadian Journal of Statistics = Revue canadienne de statistique, 21(3):277–283, September 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Chu:1993:RCS

[727] David P. T. Chu. Random r-content of an r-simplex from beta-type-2 random points. Canadian Journal of Statistics = Revue canadienne de statistique, 21(3):285–293, September 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Stewart:1993:IIM

[728] Frances P. Stewart and Ralph A. Bradley. Interblock information in multidimensional block designs. Canadian Journal of Statistics = Revue canadienne de statistique, 21(3):295–302, September 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Gelfand:1993:MLE

[729] Alan E. Gelfand and Bradley P. Carlin. Maximum-likelihood estimation for constrained- or missing-data models. Canadian Journal of Statistics = Revue canadienne de statistique, 21(3):303–311, September 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Fraser:1993:ASM

[730] D. A. S. Fraser and A. C. M. Wong. Approximate Studentization with marginal and conditional inference. Canadian Journal of Statistics = Revue canadienne de statistique, 21(3):313–320, September 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Spiring:1993:RNL

[731] Fred A. Spiring. The reflected normal loss function. Canadian Journal of Statistics = Revue canadienne de statistique, 21(3):321–330, September 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Stein:1993:ASR

[732] Michael L. Stein, Xiaotong Shen, and Patricia E. Styer. Applications of a simple regression model to acid rain data. Canadian Journal of Statistics = Revue canadienne de statistique, 21(3):331–346, September 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:1993:EEB

[733] Jia Hua Chen and R. R. Sitter. Edgeworth expansion and the bootstrap for stratified sampling without replacement from a finite population. Canadian Journal of Statistics = Revue canadienne de statistique, 21(4):347–357, December 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Marchand:1993:EMM

[734] Eric Marchand. Estimation of a multivariate mean with constraints on the norm. Canadian Journal of Statistics = Revue canadienne de statistique, 21(4):359–366, December 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

McKay:1993:NBR

[735] Ian McKay. A note on bias reduction in variable-kernel density estimates. Canadian Journal of Statistics = Revue canadienne de statistique, 21 (4):367–375, December 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Shao:1993:JIH

[736] Jun Shao and J. N. K. Rao. Jackknife inference for heteroscedastic linear regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 21(4):377–395, December 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Leon:1993:MSO

[737] Carlos Antonio León and Jean-Claude Massé. La médiane simpliciale d’Oja: Existence, unicité et stabilité. (French) [The simplicial median of Oja. Existence, uniqueness, and stability]. Canadian Journal of Statistics = Revue canadienne de statistique, 21(4):397–408, December 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:1993:DSP

[738] Yan Hong Wu and M. S. Srivastava. Dynamic sampling plans in online control charts. Canadian Journal of Statistics = Revue canadienne de statistique, 21(4):409–419, December 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Gentleman:1993:EMP

[739] Jane F. Gentleman and G. A. Whitmore. Estimation of the mutagenic potency of environmental chemicals using short-term bioassay. Canadian Journal of Statistics = Revue canadienne de statistique, 21(4):422–428, December 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Darlington:1993:EMP

[740] Gerarda A. Darlington. Estimation of mutagenic potency and determination of sources of variability for Ames bioassay data. Canadian Journal of Statistics = Revue canadienne de statistique, 21(4):428–436, December 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Eastwood:1993:SNE

[741] Brian J. Eastwood. Semi-nonparametric estimates of dose-response mutagenic potency of environmental chemicals and an assessment of interand intra-laboratory variation using a short-term bioassay. Canadian Journal of Statistics = Revue canadienne de statistique, 21(4):436–448, December 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Leroux:1993:CVM

[742] B. G. Leroux and D. Krewski. Components of variation in mutagenic potency values based on the Ames salmonella test. Canadian Journal of Statistics = Revue canadienne de statistique, 21(4):448–459, December 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1993:SEC

[743] Anonymous. Section Editors’ comments. Canadian Journal of Statistics = Revue canadienne de statistique, 21(4):459–465, December 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1993:AAb

[744] Anonymous. Authors’ addendum. Canadian Journal of Statistics = Revue canadienne de statistique, 21(4):466, December 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Krewski:1993:AAC

[745] Daniel Krewski, Robert T. Smythe, Karen Y. Fung, and Richard Burnett. Author’s addendum: “Conditional and unconditional tests with historical controls” [Canad. J. Statist. 19 (1991), no. 4, 407–423; MR1166847 (93g:62034)]. Canadian Journal of Statistics = Revue canadienne de statistique, 21(4):466, 1993. ISSN 0319-5724 (print), 1708-945X (electronic).

Plante:1994:ICS

[746] André Plante. An inclusion-consistent solution to the problem of absurd confidence statements: 2. Consistent nonexact-confidence-interval estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 22(1):1–13, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Lawless:1994:ARD

[747] J. F. Lawless. Adjustments for reporting delays and the prediction of occurred but not reported events. Canadian Journal of Statistics = Revue canadienne de statistique, 22(1):15–31, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Stroud:1994:BAB

[748] T. W. F. Stroud. Bayesian analysis of binary survey data. Canadian Journal of Statistics = Revue canadienne de statistique, 22(1):33–45, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Joe:1994:MEV

[749] Harry Joe. Multivariate extreme-value distributions with applications to environmental data. Canadian Journal of Statistics = Revue canadienne de statistique, 22(1):47–64, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Lesperance:1994:MMM

[750] Mary L. Lesperance and John D. Kalbfleisch. Mixture models for matched pairs. Canadian Journal of Statistics = Revue canadienne de statistique, 22(1):65–74, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Oluyede:1994:MCS

[751] Broderick O. Oluyede. A modified chi-square test of independence against a class of ordered alternatives in an r × c contingency table. Canadian Journal of Statistics = Revue canadienne de statistique, 22 (1):75–87, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Ramsay:1994:VHA

[752] J. O. Ramsay, N. Altman, and R. Darrell Bock. Variation in height acceleration in the Fels growth data. Canadian Journal of Statistics = Revue canadienne de statistique, 22(1):89–102, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Routledge:1994:PSS

[753] R. D. Routledge. Practicing safe statistics with the mid-p. Canadian Journal of Statistics = Revue canadienne de statistique, 22(1):103–110, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhu:1994:IAS

[754] Yiliang Zhu and Nancy Reid. Information, ancillarity, and sufficiency in the presence of nuisance parameters. Canadian Journal of Statistics = Revue canadienne de statistique, 22(1):111–123, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Choulakian:1994:CMS

[755] V. Choulakian, R. A. Lockhart, and M. A. Stephens. Cramér-von Mises statistics for discrete distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 22(1):125–137, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Neuhaus:1994:CCE

[756] J. M. Neuhaus, J. D. Kalbfleisch, and W. W. Hauck. Conditions for consistent estimation in mixed-effects models for binary matched-pairs data. Canadian Journal of Statistics = Revue canadienne de statistique, 22(1): 139–148, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Atkinson:1994:NTT

[757] G. F. Atkinson and K. Mount. A nonparametric test for Type I censored data: r samples. Canadian Journal of Statistics = Revue canadienne de statistique, 22(1):149–162, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Tibshirani:1994:SAR

[758] Robert Tibshirani and Larry Wasserman. Some aspects of the reparametrization of statistical models. Canadian Journal of Statistics = Revue canadienne de statistique, 22(1):163–173, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1994:AAa

[759] Anonymous. Authors’ addendums. Canadian Journal of Statistics = Revue canadienne de statistique, 22(1):174, March 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Brillinger:1994:TSP

[760] David R. Brillinger. Time series, point processes, and hybrids. Canadian Journal of Statistics = Revue canadienne de statistique, 22(2):177–206, June 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Bohning:1994:BAC

[761] Dankmar Böhning. Better approximate confidence intervals for a binomial parameter. Canadian Journal of Statistics = Revue canadienne de statistique, 22(2):207–218, June 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Field:1994:OSE

[762] Christopher A. Field and Douglas P. Wiens. One-step M-estimators in the linear model, with dependent errors. Canadian Journal of Statistics = Revue canadienne de statistique, 22(2):219–231, June 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Wiens:1994:BIR

[763] Douglas Wiens and Julie Zhou. Bounded-influence rank estimation in the linear model. Canadian Journal of Statistics = Revue canadienne de statistique, 22(2):233–245, June 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Victoria-Feser:1994:RMP

[764] Maria-Pia Victoria-Feser and Elvezio Ronchetti. Robust methods for personal-income distribution models. Canadian Journal of Statistics = Revue canadienne de statistique, 22(2):247–258, June 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Kokonendji:1994:LTN

[765] C. C. Kokonendji and V. Seshadri. The Lindsay transform of natural exponential families. Canadian Journal of Statistics = Revue canadienne de statistique, 22(2):259–272, June 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Iwashita:1994:ADF

[766] Toshiya Iwashita and Minoru Siotani. Asymptotic distributions of functions of a sample covariance matrix under the elliptical distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 22 (2):273–283, June 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Bedrick:1994:MLE

[767] Edward J. Bedrick. Maximum-likelihood estimation for the removal method. Canadian Journal of Statistics = Revue canadienne de statistique, 22(2):285–293, June 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Abdous:1994:CLE

[768] Belkacem Abdous and Alain Berlinèt. Correction locale de l’estimateur à noyau de la densité d’une loi de probabilité. (French) [Local correction of the kernel estimator of density of a probability distribution]. Canadian Journal of Statistics = Revue canadienne de statistique, 22(2):295–308, June 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Beirlant:1994:ANE

[769] Jan Beirlant, László Györfi, and Gábor Lugosi. On the asymptotic normality of the L_{1}- and L_{2}-errors in histogram density estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 22(3):309– 318, September 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Bull:1994:ERL

[770] Shelley B. Bull, Celia M. T. Greenwood, and Allan Donner. Efficiency of reduced logistic regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 22(3):319–334, September 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Meloche:1994:BIR

[771] J. Meloche and R. H. Zamar. Binary-image restoration. Canadian Journal of Statistics = Revue canadienne de statistique, 22(3):335–355, September 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Herzberg:1994:OED

[772] Agnes M. Herzberg and William N. Traves. An optimal experimental design for the Haar regression model. Canadian Journal of Statistics = Revue canadienne de statistique, 22(3):357–364, September 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Farrell:1994:PAO

[773] Patrick J. Farrell, Brenda MacGibbon, and Thomas J. Tomberlin. Protection against outliers in empirical Bayes estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 22(3):365–376, September 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Srivastava:1994:LCP

[774] M. S. Srivastava and Yanhong Wu. On-line control procedures under the random-walk model with measurement error and attribute observations. Canadian Journal of Statistics = Revue canadienne de statistique, 22 (3):377–386, September 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:1994:GLR

[775] Jia Hua Chen. Generalized likelihood-ratio test of the number of components in finite mixture models. Canadian Journal of Statistics = Revue canadienne de statistique, 22(3):387–399, September 1994. ISSN 0319-5724 (print), 1708-945X (electronic). See comment [1245].

Brunner:1994:NBM

[776] Lawrence J. Brunner and Albert Y. Lo. Nonparametric Bayes methods for directional data. Canadian Journal of Statistics = Revue canadienne de statistique, 22(3):401–412, September 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1994:AAb

[777] Anonymous. Authors’ addendum. Canadian Journal of Statistics = Revue canadienne de statistique, 22(3):413, September 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Bolfarine:1994:AFP

[778] H. Bolfarine. Addendum: “Finite-population prediction under errorin-variables superpopulation models” [Canad. J. Statist. 19 (1991), no. 2, 191–207; MR1128406 (92k:62018)]. Canadian Journal of Statistics = Revue canadienne de statistique, 22(3):413, 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Krewski:1994:I

[779] D. Krewski, R. T. Burnett, and L. K. Chan. Introduction. Canadian Journal of Statistics = Revue canadienne de statistique, 22(4):415–419, December 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

McNeney:1994:OPR

[780] Brad McNeney and John Petkau. Overdispersed Poisson regression models for studies of air pollution and human health. Canadian Journal of Statistics = Revue canadienne de statistique, 22(4):421–440, December 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Burnett:1994:APE

[781] Richard Burnett and Daniel Krewski. Air pollution effects on hospital admission rates: a random effects modeling approach. Canadian Journal of Statistics = Revue canadienne de statistique, 22(4):441–458, December 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Corey:1994:ARB

[782] Paul N. Corey, Wen-Yi Lou, and Irv Broder. Analysis of the relationship between air pollution and asthma. Canadian Journal of Statistics = Revue canadienne de statistique, 22(4):459–470, December 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Schwartz:1994:NSA

[783] Joel Schwartz. Nonparametric smoothing in the analysis of air pollution and respiratory illness. Canadian Journal of Statistics = Revue canadienne de statistique, 22(4):471–487, December 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Brown:1994:MSI

[784] Philip J. Brown, Nhu D. Le, and James V. Zidek. Multivariate spatial interpolation and exposure to air pollutants. Canadian Journal of Statistics = Revue canadienne de statistique, 22(4):489–509, December 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Rao:1994:SAE

[785] J. N. K. Rao and Mingyu Yu. Small-area estimation by combining timeseries and cross-sectional data. Canadian Journal of Statistics = Revue canadienne de statistique, 22(4):511–528, December 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

MacNeill:1994:MHA

[786] I. B. MacNeill, Y. Mao, and L. Xie. Modeling heteroscedastic age-period-cohort cancer data. Canadian Journal of Statistics = Revue canadienne de statistique, 22(4):529–539, December 1994. ISSN 0319-5724 (print), 1708-945X (electronic).

Cuevas:1995:EIC

[787] Antonio Cuevas and Juan Romo. On the estimation of the influence curve. Canadian Journal of Statistics = Revue canadienne de statistique, 23(1):1–9, March 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Abdous:1995:CEC

[788] Belkacem Abdous. Computationally efficient classes of higher-order kernel functions. Canadian Journal of Statistics = Revue canadienne de statistique, 23(1):21–27, March 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Rai:1995:AID

[789] S. N. Rai and D. E. Matthews. The analysis of incomplete data using stochastic covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 23(1):29–42, March 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Viana:1995:CBE

[790] Marios A. G. Viana. Combined Bayesian estimates for the equicorrelation coefficient. Canadian Journal of Statistics = Revue canadienne de statistique, 23(1):43–53, March 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Hallin:1995:MWW

[791] Marc Hallin and Madan L. Puri. A multivariate Wald-Wolfowitz rank test against serial dependence. Canadian Journal of Statistics = Revue canadienne de statistique, 23(1):55–65, March 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Massam:1995:EDC

[792] Hélène Massam. An exact decomposition for a class of exponential dispersion models: Application to sequential testing. Canadian Journal of Statistics = Revue canadienne de statistique, 23(1):67–84, March 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Chan:1995:BCC

[793] Lai K. Chan and Guo-Ying Li. Bivariate control charts for testing x̄ data patterns. Canadian Journal of Statistics = Revue canadienne de statistique, 23(1):85–99, March 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Chapman:1995:PAO

[794] G. R. Chapman. Perturbation analysis for an orthogonal least-squares estimator. Canadian Journal of Statistics = Revue canadienne de statistique, 23(2):101–108, June 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

LeBlanc:1995:SFC

[795] Michael LeBlanc and John Crowley. Step-function covariate effects in the proportional-hazards model. Canadian Journal of Statistics = Revue canadienne de statistique, 23(2):109–129, June 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

ElBarmi:1995:TAS

[796] Hammou El Barmi and Richard Dykstra. Testing for and against a set of linear inequality constraints in a multinomial setting. Canadian Journal of Statistics = Revue canadienne de statistique, 23(2):131–143, June 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Qin:1995:EEE

[797] Jing Qin and Jerry Lawless. Estimating equations, empirical likelihood and constraints on parameters. Canadian Journal of Statistics = Revue canadienne de statistique, 23(2):145–159, June 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Shi:1995:ABN

[798] Pei De Shi. Asymptotic behaviour of nonparametric conditional quantile estimates for time series. Canadian Journal of Statistics = Revue canadienne de statistique, 23(2):161–169, June 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Vaillancourt:1995:AMS

[799] Jean Vaillancourt. Asymptotics for multisample statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 23(2):171–184, June 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Fan:1995:AMD

[800] Rocky Y. K. Fan and Christopher A. Field. Approximations for marginal densities of M-estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 23(2):185–197, June 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Peng:1995:BAO

[801] Fengchun Peng and Dipak K. Dey. Bayesian analysis of outlier problems using divergence measures. Canadian Journal of Statistics = Revue canadienne de statistique, 23(2):199–213, June 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Yu:1995:NSC

[802] Kai Fun Yu. A necessary and sufficient condition for the strong consistency of a family of estimators of the common odds ratio. Canadian Journal of Statistics = Revue canadienne de statistique, 23(2):215–225, June 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Godambe:1995:EPS

[803] V. P. Godambe. Estimation of parameters in survey sampling: Optimality. Canadian Journal of Statistics = Revue canadienne de statistique, 23 (3):227–243, September 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

VanEeden:1995:MEL

[804] Constance Van Eeden. Minimax estimation of a lower-bounded scale parameter of a gamma distribution for scale-invariant squared-error loss. Canadian Journal of Statistics = Revue canadienne de statistique, 23 (3):245–256, September 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Kourouklis:1995:EEQ

[805] S. Kourouklis. Estimation of an exponential quantile under Pitman’s measure of closeness. Canadian Journal of Statistics = Revue canadienne de statistique, 23(3):257–268, September 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Vidakovic:1995:LBB

[806] Brani Vidakovic and Anirban Dasgupta. Lower bounds on Bayes risks for estimating a normal variance: With applications. Canadian Journal of Statistics = Revue canadienne de statistique, 23(3):269–282, September 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Jordan:1995:CRC

[807] Scott M. Jordan and K. Krishnamoorthy. Confidence regions for the common mean vector of several multivariate normal populations. Canadian Journal of Statistics = Revue canadienne de statistique, 23(3):283–297, September 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Dabrowski:1995:SIP

[808] André Robert Dabrowski and Abdelhak Zoglat. Strong invariance principles for triangular arrays of weakly dependent random variables. Canadian Journal of Statistics = Revue canadienne de statistique, 23(3):299– 310, September 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Camacho:1995:ERP

[809] Fernando Camacho and Geoffrey Arron. Effects of the regulators paclobutrazol and flurprimidol on the growth of terminal sprouts formed on trimmed silver maple trees. Canadian Journal of Statistics = Revue canadienne de statistique, 23(3):312–322, September 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Lee:1995:SBG

[810] Hyun Suk Lee and Bob Philips. In search of the “best” growth inhibitor. Canadian Journal of Statistics = Revue canadienne de statistique, 23 (3):322–325, September 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Sloan:1995:DSM

[811] Jef F. A. Sloan, Carl J. Schwarz, and Linda R. Neden. Dataset on silvermaple-tree growth regulators. Canadian Journal of Statistics = Revue canadienne de statistique, 23(3):325–329, September 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Schwarz:1995:CA

[812] C. Schwarz and N. Reid. Comments on the analyses. Canadian Journal of Statistics = Revue canadienne de statistique, 23(3):329–332, September 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Cook:1995:CIS

[813] Richard J. Cook and Vern T. Farewell. Conditional inference for subjectspecific and marginal agreement: Two families of agreement measures. Canadian Journal of Statistics = Revue canadienne de statistique, 23 (4):333–344, December 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Alvo:1995:RCM

[814] Mayer Alvo and Paul Cabilio. Rank correlation methods for missing data. Canadian Journal of Statistics = Revue canadienne de statistique, 23(4):345–358, December 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Gulati:1995:NFE

[815] Sneh Gulati and W. J. Padgett. Nonparametric function estimation from inversely sampled record-breaking data. Canadian Journal of Statistics = Revue canadienne de statistique, 23(4):359–368, December 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Markatou:1995:RSE

[816] Marianthi Markatou and Joel L. Horowitz. Robust scale estimation in the error-components model using the empirical characteristic function. Canadian Journal of Statistics = Revue canadienne de statistique, 23 (4):369–381, December 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Boente:1995:ADD

[817] Graciela Boente and Ricardo Fraiman. Asymptotic distribution of datadriven smoothers in density and regression estimation under dependence. Canadian Journal of Statistics = Revue canadienne de statistique, 23 (4):383–397, December 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Hsu:1995:GLA

[818] John S. J. Hsu. Generalized Laplacian approximations in Bayesian inference. Canadian Journal of Statistics = Revue canadienne de statistique, 23(4):399–410, December 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Gelfand:1995:NBI

[819] Alan E. Gelfand and Saurabh Mukhopadhyay. On nonparametric Bayesian inference for the distribution of a random sample. Canadian Journal of Statistics = Revue canadienne de statistique, 23(4):411–420, December 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Ducharme:1995:ULD

[820] Gilles R. Ducharme. Uniqueness of the least-distances estimator in regression models with multivariate response. Canadian Journal of Statistics = Revue canadienne de statistique, 23(4):421–424, December 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Routledge:1995:UVS

[821] Rick Routledge and Min Tsao. Uniform validity of saddlepoint expansion on compact sets. Canadian Journal of Statistics = Revue canadienne de statistique, 23(4):425–431, December 1995. ISSN 0319-5724 (print), 1708-945X (electronic).

Bickis:1996:EPP

[822] Mikelis Bickis, Susana Bleuer, and Daniel Krewski. On the estimation of the proportion of positives in a sequence of screening experiments. Canadian Journal of Statistics = Revue canadienne de statistique, 24 (1):1–15, March 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Chipman:1996:BVS

[823] Hugh Chipman. Bayesian variable selection with related predictors. Canadian Journal of Statistics = Revue canadienne de statistique, 24 (1):17–36, March 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Joseph:1996:EMC

[824] Lawrence Joseph, Alain C. Vandal, and David B. Wolfson. Estimation in the multipath change point problem for correlated data. Canadian Journal of Statistics = Revue canadienne de statistique, 24(1):37–53, March 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Dabrowski:1996:ECO

[825] André Robert Dabrowski and Herold Dehling. Estimating conditional occupation-time distributions for dependent sequences. Canadian Journal of Statistics = Revue canadienne de statistique, 24(1):55–65, March 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Wiens:1996:RSD

[826] Douglas P. Wiens. Robust sequential designs for approximately linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 24(1):67–79, March 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Henze:1996:EDF

[827] Norbert Henze. Empirical-distribution-function goodness-of-fit tests for discrete models. Canadian Journal of Statistics = Revue canadienne de statistique, 24(1):81–93, March 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Haunsperger:1996:PNT

[828] Deanna B. Haunsperger. Paradoxes in nonparametric tests. Canadian Journal of Statistics = Revue canadienne de statistique, 24(1):95–104, March 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Shao:1996:ITL

[829] Peter Yi-Shi Shao and William E. Strawderman. Improving on truncated linear estimates of exponential and gamma scale parameters. Canadian Journal of Statistics = Revue canadienne de statistique, 24(1):105–114, March 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Bayomog:1996:TDC

[830] Samuel Bayomog, Xavier Guyon, Cécile Hardouin, and Jianfeng Yao. Test de différence de contrastes et somme pondérée de khi-deux. (French) [Test of the contrast difference and weighted chi-square sum]. Canadian Journal of Statistics = Revue canadienne de statistique, 24(1):115–130, March 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Benjamini:1996:CMS

[831] Yoav Benjamini and Abba M. Krieger. Concepts and measures for skewness with data-analytic implications. Canadian Journal of Statistics = Revue canadienne de statistique, 24(1):131–140, March 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Reid:1996:LHO

[832] N. Reid. Likelihood and higher-order approximations to tail areas: a review and annotated bibliography. Canadian Journal of Statistics = Revue canadienne de statistique, 24(2):141–166, June 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:1996:PMD

[833] Jiahua Chen and J. D. Kalbfleisch. Penalized minimum-distance estimates in finite mixture models. Canadian Journal of Statistics = Revue canadienne de statistique, 24(2):167–175, June 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Sutradhar:1996:ERP

[834] B. C. Sutradhar and J. N. K. Rao. Estimation of regression parameters in generalized linear models for cluster correlated data with measurement error. Canadian Journal of Statistics = Revue canadienne de statistique, 24(2):177–192, June 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:1996:ERW

[835] Bing Li and Ruben H. Zamar. M-estimates of regression when the scale is unknown and the error distribution is possibly asymmetric: a minimax result. Canadian Journal of Statistics = Revue canadienne de statistique, 24(2):193–206, June 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Arellano-Valle:1996:UEM

[836] R. B. Arellano-Valle, H. Bolfarine, and F. Vilca-Labra. Ultrastructural elliptical models. Canadian Journal of Statistics = Revue canadienne de statistique, 24(2):207–216, June 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Dykstra:1996:NPP

[837] R. L. Dykstra, Hammou El Barmi, James M. Guffey, and F. T. Wright. Nonhomogeneous Poisson processes as overhaul models. Canadian Journal of Statistics = Revue canadienne de statistique, 24(2):217–228, June 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Gombay:1996:WSL

[838] Edit Gombay. The weighted sequential likelihood ratio. Canadian Journal of Statistics = Revue canadienne de statistique, 24(2):229–239, June 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Sun:1996:SNL

[839] Feng-Bin Sun, Jean-Yves Laramée, and John S. Ramberg. On Spiring’s normal loss function. Canadian Journal of Statistics = Revue canadienne de statistique, 24(2):241–249, June 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Bain:1996:CMI

[840] Lee J. Bain and Gaoxiong Gan. Conditional maxima and inferences for the truncated exponential distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 24(2):251–256, June 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:1996:CTS

[841] Y. Wang, A. K. Gupta, and T. T. Nguyen. Characterization theorems for some discrete distributions based on conditional structure. Canadian Journal of Statistics = Revue canadienne de statistique, 24(2):257–262, June 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Viraswami:1996:HOA

[842] K. Viraswami and N. Reid. Higher-order asymptotics under model misspecification. Canadian Journal of Statistics = Revue canadienne de statistique, 24(2):263–278, June 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:1996:MSN

[843] Deli Li. On moments of the supremum of normed weighted averages. Canadian Journal of Statistics = Revue canadienne de statistique, 24 (3):279–292, September 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Dabrowski:1996:ABP

[844] André Robert Dabrowski and David McDonald. An application of the Bernoulli part to local limit theorems for moving averages on stationary sequences. Canadian Journal of Statistics = Revue canadienne de statistique, 24(3):293–305, September 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Gustafson:1996:EMD

[845] Paul Gustafson. The effect of mixing-distribution misspecification in conjugate mixture models. Canadian Journal of Statistics = Revue canadienne de statistique, 24(3):307–318, September 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Lee:1996:AOP

[846] Hyun Suk Lee. Analysis of overdispersed paired count data. Canadian Journal of Statistics = Revue canadienne de statistique, 24(3):319–326, September 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Albert:1996:BSL

[847] James H. Albert. Bayesian selection of log-linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 24(3):327–347, September 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Cabilio:1996:STS

[848] Paul Cabilio and Joe Masaro. A simple test of symmetry about an unknown median. Canadian Journal of Statistics = Revue canadienne de statistique, 24(3):349–361, September 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:1996:ETG

[849] Gemai Chen. EDF tests of goodness of fit for transform-both-sides models. Canadian Journal of Statistics = Revue canadienne de statistique, 24 (3):363–372, September 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Shah:1996:APG

[850] D. N. Shah and P. A. Patel. Asymptotic properties of a generalized regression-type predictor of a finite population variance in probability sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 24(3):373–384, September 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Jorgensen:1996:SSM

[851] Bent Jørgensen, Søren Lundbye-Christensen, Peter Xue-Kun Song, and Li Sun. State-space models for multivariate longitudinal data of mixed types. Canadian Journal of Statistics = Revue canadienne de statistique, 24(3):385–402, September 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Csorgo:1996:WAQ

[852] Miklós Csörgő and Hao Yu. Weak approximations for quantile processes of stationary sequences. Canadian Journal of Statistics = Revue canadienne de statistique, 24(4):403–430, December 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Fung:1996:CTT

[853] Karen Y. Fung, D. Krewski, and R. T. Smythe. A comparison of tests for trend with historical controls in carcinogen bioassay. Canadian Journal of Statistics = Revue canadienne de statistique, 24(4):431–454, December 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Manchester:1996:WCO

[854] Lise Manchester and Wade Blanchard. When is a curve an outlier? An account of a tricky problem. Canadian Journal of Statistics = Revue canadienne de statistique, 24(4):455–466, December 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Besse:1996:ASP

[855] Philippe C. Besse and Hervé Cardot. Approximation spline de la prévision d’un processus fonctionnel autorégressif d’ordre 1. (French) [Spline approximation of the prediction of a process of autoregressive functional order 1]. Canadian Journal of Statistics = Revue canadienne de statistique, 24(4):467–487, December 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Wong:1996:DCC

[856] H. Wong and W. K. Li. Distribution of the cross-correlations of squared residuals in ARIMA models. Canadian Journal of Statistics = Revue canadienne de statistique, 24(4):489–502, December 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Rukhin:1996:LSC

[857] A. L. Rukhin. Linear statistics in change-point estimation and their asymptotic behaviour. Canadian Journal of Statistics = Revue canadienne de statistique, 24(4):503–515, December 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Maki:1996:RPG

[858] Ellen Maki and Philip McDunnough. The role of probability generating functions for estimation in incompletely observed random walks. Canadian Journal of Statistics = Revue canadienne de statistique, 24(4):517– 534, December 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Gustafson:1996:MIF

[859] Paul Gustafson. Model influence functions based on mixtures. Canadian Journal of Statistics = Revue canadienne de statistique, 24(4):535–548, December 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Mojirsheibani:1996:SRB

[860] Majid Mojirsheibani and Robert Tibshirani. Some results on bootstrap prediction intervals. Canadian Journal of Statistics = Revue canadienne de statistique, 24(4):549–568, December 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Goutis:1996:PMH

[861] Constantinos Goutis and Rex F. Galbraith. A parametric model for heterogeneity in paired Poisson counts. Canadian Journal of Statistics = Revue canadienne de statistique, 24(4):569–581, December 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Hengartner:1996:NRE

[862] Nicolas W. Hengartner and Oliver B. Linton. Nonparametric regression estimation at design poles and zeros. Canadian Journal of Statistics = Revue canadienne de statistique, 24(4):583–591, December 1996. ISSN 0319-5724 (print), 1708-945X (electronic).

Rao:1997:DSS

[863] J. N. K. Rao. Developments in sample survey theory: an appraisal. Canadian Journal of Statistics = Revue canadienne de statistique, 25 (1):1–21, March 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Smith:1997:SSS

[864] T. M. Fred Smith. Social surveys and social science. Canadian Journal of Statistics = Revue canadienne de statistique, 25(1):23–38, March 1997. ISSN 0319-5724 (print), 1708-945X (electronic). See discussion [865].

Anonymous:1997:D

[865] Anonymous. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 25(1):38–44, March 1997. ISSN 0319-5724 (print), 1708-945X (electronic). See [864].

Anonymous:1997:RD

[866] Anonymous. Reply to the discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 25(1):44, March 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Hu:1997:APM

[867] F. Hu. The asymptotic properties of the maximum-relevance weighted likelihood estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 25(1):45–59, March 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Sitter:1997:IMV

[868] R. R. Sitter and J. N. K. Rao. Imputation for missing values and corresponding variance estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 25(1):61–73, March 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Farrell:1997:BAE

[869] Patrick J. Farrell, Brenda Macgibbon, and Thomas J. Tomberlin. Bootstrap adjustments for empirical Bayes interval estimates of small-area proportions. Canadian Journal of Statistics = Revue canadienne de statistique, 25(1):75–89, March 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Fraser:1997:SAI

[870] D. A. S. Fraser, N. Reid, and A. Wong. Simple and accurate inference for the mean of the gamma model. Canadian Journal of Statistics = Revue canadienne de statistique, 25(1):91–99, March 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Sun:1997:RER

[871] Jianguo Sun and David E. Matthews. A random-effect regression model for medical follow-up studies. Canadian Journal of Statistics = Revue canadienne de statistique, 25(1):101–111, March 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Caperaa:1997:TRT

[872] Philippe Capéraà and Ana Isabel Garralda Guillem. Taux de résistance des tests de rang d’indépendance. (French) [Resistance rates of independence rank tests]. Canadian Journal of Statistics = Revue canadienne de statistique, 25(1):113–124, March 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Hu:1997:PEC

[873] X. Joan Hu and Jerald F. Lawless. Pseudolikelihood estimation in a class of problems with response-related missing covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 25(2):125–142, June 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Guttman:1997:BAR

[874] Irwin Guttman and George D. Papandonatos. A Bayesian approach to a reliability problem: Theory, analysis and interesting numerics. Canadian Journal of Statistics = Revue canadienne de statistique, 25(2):143–158, June 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Hines:1997:FGL

[875] R. J. O’Hara Hines. Fitting generalized linear models to retrospectively sampled clusters with categorical responses. Canadian Journal of Statistics = Revue canadienne de statistique, 25(2):159–174, June 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Sitter:1997:ODL

[876] R. R. Sitter and I. Fainaru. Optimal designs for the logit and probit models for binary data. Canadian Journal of Statistics = Revue canadienne de statistique, 25(2):175–190, June 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Tang:1997:MCS

[877] Boxin Tang and C. F. J. Wu. A method for constructing supersaturated designs and its Es^{2} optimality. Canadian Journal of Statistics = Revue canadienne de statistique, 25(2):191–201, June 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Yamada:1997:SDI

[878] Shu Yamada and Dennis K. J. Lin. Supersaturated design including an orthogonal base. Canadian Journal of Statistics = Revue canadienne de statistique, 25(2):203–213, June 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Benn:1997:IMP

[879] A. G. Benn and R. J. Kulperger. Integrated marked Poisson processes with application to image correlation spectroscopy. Canadian Journal of Statistics = Revue canadienne de statistique, 25(2):215–231, June 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

ElHimdi:1997:TNT

[880] Khalid El Himdi and Roch Roy. Tests for noncorrelation of two multivariate ARMA time series. Canadian Journal of Statistics = Revue canadienne de statistique, 25(2):233–256, June 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Spinelli:1997:CMT

[881] John J. Spinelli and Michael A. Stephens. Cramér-von Mises tests of fit for the Poisson distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 25(2):257–268, June 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

OGorman:1997:ATO

[882] Thomas W. O’Gorman. An adaptive test for the one-way layout. Canadian Journal of Statistics = Revue canadienne de statistique, 25(2):269– 279, June 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Gray:1997:PER

[883] M. Gray. Pay equity: the role and limitations of statistical analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 25 (3):281–292, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Monette:1997:TRS

[884] Georges Monette. Is there a role for statisticians in pay equity? Canadian Journal of Statistics = Revue canadienne de statistique, 25(3):293–310, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Bellhouse:1997:D

[885] D. R. Bellhouse. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 25(3):311–314, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Spiring:1997:D

[886] Fred A. Spiring. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 25(3):315–319, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Gray:1997:R

[887] M. Gray. Rejoinder. Canadian Journal of Statistics = Revue canadienne de statistique, 25(3):321–322, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Monette:1997:R

[888] G. Monette. Rejoinder. Canadian Journal of Statistics = Revue canadienne de statistique, 25(3):323, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Qin:1997:SAR

[889] Jing Qin and Jianguo Sun. Statistical analysis of right-censored failuretime data with partially specified hazard rates. Canadian Journal of Statistics = Revue canadienne de statistique, 25(3):325–336, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Papandonatos:1997:LAC

[890] George D. Papandonatos and Seymour Geisser. Laplace approximations for censored linear regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 25(3):337–358, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Akritas:1997:PTG

[891] M. G. Akritas and A. F. Torbeyns. Pearson-type goodness-of-fit tests for regression. Canadian Journal of Statistics = Revue canadienne de statistique, 25(3):359–374, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Fougeres:1997:EDU

[892] Anne-Laure Fougères. Estimation de densités unimodales. (French) [Estimation of unimodal densities]. Canadian Journal of Statistics = Revue canadienne de statistique, 25(3):375–387, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:1997:TNC

[893] Jiahua Chen and Ping Cheng. On testing the number of components in finite mixture models with known relevant component distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 25 (3):389–400, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Hooper:1997:CIF

[894] Peter M. Hooper and Zhenlin Yang. Confidence intervals following Box– Cox transformation. Canadian Journal of Statistics = Revue canadienne de statistique, 25(3):401–416, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Gombay:1997:LRU

[895] Edit Gombay. The likelihood ratio under noncontiguous alternatives. Canadian Journal of Statistics = Revue canadienne de statistique, 25 (3):417–423, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1997:AA

[896] Anonymous. Authors’ addendum. Canadian Journal of Statistics = Revue canadienne de statistique, 25(3):425, September 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Jorgensen:1997:ASS

[897] Bent Jørgensen, Søren Lundbye-Christensen, Peter Xue-Kun Song, and Li Sun. Addendum: “State space models for multivariate longitudinal data of mixed types” [Canad. J. Statist. 24 (1996), no. 3, 385–402; MR1421864 (97f:62144)]. Canadian Journal of Statistics = Revue canadienne de statistique, 25(3):425, 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Genest:1997:SSM

[898] Christian Genest. Statistics on statistics: Measuring research productivity by journal publications between 1985 and 1995. Canadian Journal of Statistics = Revue canadienne de statistique, 25(4):427–443, December 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Sinha:1997:TDB

[899] Debajyoti Sinha. Time-discrete beta-process model for interval-censored survival data. Canadian Journal of Statistics = Revue canadienne de statistique, 25(4):445–456, December 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Kuo:1997:BSI

[900] Lynn Kuo and Bani Mallick. Bayesian semiparametric inference for the accelerated failure-time model. Canadian Journal of Statistics = Revue canadienne de statistique, 25(4):457–472, December 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Walker:1997:NSP

[901] Stephen G. Walker and Bani K. Mallick. A note on the scale parameter of the Dirichlet process. Canadian Journal of Statistics = Revue canadienne de statistique, 25(4):473–479, December 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Heckman:1997:LTT

[902] Nancy Heckman and John Rice. Line transects of two-dimensional random fields: Estimation and design. Canadian Journal of Statistics = Revue canadienne de statistique, 25(4):481–501, December 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

DeSantis:1997:ABF

[903] Fulvio De Santis and Fulvio Spezzaferri. Alternative Bayes factors for model selection. Canadian Journal of Statistics = Revue canadienne de statistique, 25(4):503–515, December 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:1997:BCT

[904] Gemai Chen and Richard A. Lockhart. Box–Cox transformed linear models: a parameter-based asymptotic approach. Canadian Journal of Statistics = Revue canadienne de statistique, 25(4):517–529, December 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Dette:1997:ODR

[905] Holger Dette. E-optimal designs for regression models with quantitative factors — a reasonable choice? Canadian Journal of Statistics = Revue canadienne de statistique, 25(4):531–543, December 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:1997:GLB

[906] Jeesen Chen. A general lower bound of minimax risk for absolute-error loss. Canadian Journal of Statistics = Revue canadienne de statistique, 25(4):545–558, December 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Seifu:1997:ABU

[907] Yodit Seifu and N. Reid. Applications of bivariate and univariate local Lyapunov exponents. Canadian Journal of Statistics = Revue canadienne de statistique, 25(4):559–579, December 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Tibshirani:1997:CTM

[908] Robert Tibshirani and Donald A. Redelmeier. Cellular telephones and motor-vehicle collisions: Some variations on matched-pairs analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 25 (4):581–591, December 1997. ISSN 0319-5724 (print), 1708-945X (electronic).

Genest:1998:E

[909] Christian Genest. Editorial. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):1–4, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Roberts:1998:MCM

[910] Gareth O. Roberts and Jeffrey S. Rosenthal. Markov-chain Monte Carlo: Some practical implications of theoretical results. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):5–20, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic). See discussion [911, 912] and rejoinder [913].

Ishwaran:1998:D

[911] Hemant Ishwaran. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):20–27, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic). See [910].

Madras:1998:D

[912] Neal Madras. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):27–29, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic). See [910].

Roberts:1998:R

[913] Gareth O. Roberts and Jeffery S. Rosenthal. Rejoinder. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):29–31, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Aslanidou:1998:BAM

[914] Helen Aslanidou, Dipak K. Dey, and Debajyoti Sinha. Bayesian analysis of multivariate survival data using Monte Carlo methods. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):33– 48, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Ventura:1998:HOA

[915] Laura Ventura. Higher-order approximations for Pitman estimators and for optimal compromise estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):49–55, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Dykstra:1998:RTT

[916] Richard Dykstra, Subhash Kochar, and Tim Robertson. Restricted tests for testing independence of time to failure and cause of failure in a competing-risks model. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):57–68, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Park:1998:GFT

[917] Chul Gyu Park, Chu-In Charles Lee, and Tim Robertson. Goodnessof-fit test for uniform stochastic ordering among several distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 26 (1):69–81, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Paul:1998:JEM

[918] S. R. Paul and A. S. Islam. Joint estimation of the mean and dispersion parameters in the analysis of proportions: a comparison of efficiency and bias. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):83–94, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Lee:1998:GLM

[919] Y. Lee and J. A. Nelder. Generalized linear models for the analysis of quality-improvement experiments. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):95–105, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Turner:1998:HMC

[920] T. Rolf Turner, Murray A. Cameron, and Peter J. Thomson. Hidden Markov chains in generalized linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):107–125, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Abt:1998:FIM

[921] Markus Abt and William J. Welch. Fisher information and maximumlikelihood estimation of covariance parameters in Gaussian stochastic processes. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):127–137, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Davison:1998:SFC

[922] A. C. Davison and J. E. Stafford. The score function and a comparison of various adjustments of the profile likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):139–148, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Yeh:1998:BPL

[923] Arthur B. Yeh. A bootstrap procedure in linear regression with nonstationary errors. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):149–160, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Viraswami:1998:NLR

[924] K. Viraswami and N. Reid. A note on the likelihood-ratio statistic under model misspecification. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):161–168, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Sutradhar:1998:ALI

[925] Brajendra C. Sutradhar and Zhende Qu. On approximate likelihood inference in a Poisson mixed model. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):169–186, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Ghoudi:1998:PSC

[926] Kilani Ghoudi, Abdelhaq Khoudraji, and Louis-Paul Rivest. Propriétés statistiques des copules de valeurs extrêmes bidimensionnelles. (French) [Statistical properties of two-dimensional extreme value copulas]. Canadian Journal of Statistics = Revue canadienne de statistique, 26(1):187– 197, March 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Dupuis:1998:REE

[927] D. J. Dupuis and C. A. Field. Robust estimation of extremes. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2):199–215, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Ozturk:1998:SRE

[928] Ömer Öztürk and Thomas P. Hettmansperger. Simultaneous robust estimation of location and scale parameters: a minimum-distance approach. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2): 217–229, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic). See addendum [1012].

Wiens:1998:TMM

[929] Douglas P. Wiens, Eden K. H. Wu, and Julie Zhou. On the trimmed mean and minimax-variance L-estimation in Kolmogorov neighbourhoods. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2):231–238, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Bianco:1998:RKE

[930] Ana Bianco and Graciela Boente. Robust kernel estimators for additive models with dependent observations. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2):239–255, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:1998:EBR

[931] Eden K. H. Wu and Julie Zhou. Efficient bias-robust M-estimators of location in Kolmogorov normal neighbourhoods. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2):257–266, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Bensaid:1998:CEN

[932] Nadia Bensaïd and Jean-Pierre Fabre. Convergence de l’estimateur à noyau de derivées de Radon–Nikodym générales dans le cas mélangeant. (French) [Convergence of the kernel estimator of general Radon–Nikodym derivatives in the general case]. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2):267–282, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Muliere:1998:ADR

[933] Pietro Muliere and Luca Tardella. Approximating distributions of random functionals of Ferguson-Dirichlet priors. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2):283–297, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Tsao:1998:ICE

[934] C. Andy Tsao and J. T. Gene Hwang. Improved confidence estimators for Fieller’s confidence sets. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2):299–310, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:1998:RLD

[935] Gemai Chen. The run length distributions of the R, s and s^{2} control charts when σ is estimated. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2):311–322, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Sun:1998:BAS

[936] Dongchu Sun, Malay Ghosh, and Asit P. Basu. Bayesian analysis for a stress-strength system under noninformative priors. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2):323–332, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Berrendero:1998:ERM

[937] José R. Berrendero, Sonia Mazzi, Juan Romo, and Ruben H. Zamar. On the explosion rate of maximum-bias functions. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2):333–351, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Tamhane:1998:GSM

[938] Ajit C. Tamhane, Wei Liu, and Charles W. Dunnett. A generalized step-up-down multiple test procedure. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2):353–363, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Berlinet:1998:NAE

[939] Alain Berlinet, Ali Gannoun, and Eric Matzner-Løber. Normalité asymptotique d’estimateurs convergents du mode conditionnel. (French) [Asymptotic normality of convergent estimators in the conditional form]. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2): 365–380, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1998:AAa

[940] Anonymous. Authors’ addendum. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2):381, June 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Cook:1998:AAC

[941] R. J. Cook and V. T. Farewell. Authors’ addendum: “Conditional inference for subject-specific and marginal agreement: two families of agreement measures” [Canad. J. Statist. 23 (1995), no. 4, 333–344]. Canadian Journal of Statistics = Revue canadienne de statistique, 26(2):381, 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Watson:1998:RSP

[942] Geoffrey Watson. On the role of statistics in the palaeomagnetic proof of continental drift. Canadian Journal of Statistics = Revue canadienne de statistique, 26(3):383–392, September 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Billard:1998:SSO

[943] L. Billard. Some statistical opportunities in the agricultural sciences. Canadian Journal of Statistics = Revue canadienne de statistique, 26 (3):393–404, September 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Cressie:1998:TSD

[944] Noel Cressie. Transect-spacing design of ice cores on the Antarctic continent. Canadian Journal of Statistics = Revue canadienne de statistique, 26(3):405–418, September 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Gervini:1998:REV

[945] Daniel Gervini and Víctor J. Yohai. Robust estimation of variance components. Canadian Journal of Statistics = Revue canadienne de statistique, 26(3):419–430, September 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Brillinger:1998:ESM

[946] David R. Brillinger and Brent S. Stewart. Elephant-seal movements: Modelling migration. Canadian Journal of Statistics = Revue canadienne de statistique, 26(3):431–443, September 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Leong:1998:MDB

[947] Phaik Mooi Leong and Stephan Morgenthaler. Mutability of DNA base pairs: a statistical approach based on linear discrimination. Canadian Journal of Statistics = Revue canadienne de statistique, 26(3):445–454, September 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Hall:1998:SMA

[948] Deborah L. Hall, Karen Kafadar, and Alvin M. Malkinson. Statistical methodology for assessing homology of intronic regions of genes. Canadian Journal of Statistics = Revue canadienne de statistique, 26(3):455– 465, September 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Carroll:1998:NCM

[949] Raymond J. Carroll, Laurence S. Freedman, Victor Kipnis, and Li Li. A new class of measurement-error models, with applications to dietary data. Canadian Journal of Statistics = Revue canadienne de statistique, 26(3):467–477, September 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Durham:1998:SDM

[950] S. D. Durham, N. Flournoy, and W. Li. A sequential design for maximizing the probability of a favourable response. Canadian Journal of Statistics = Revue canadienne de statistique, 26(3):479–495, September 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Hampel:1998:STD

[951] Frank Hampel and Eth Zurich. Is statistics too difficult? Canadian Journal of Statistics = Revue canadienne de statistique, 26(3):497–513, September 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:1998:AAb

[952] Anonymous. Authors’ addendums. Canadian Journal of Statistics = Revue canadienne de statistique, 26(3):515, September 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Berhane:1998:GAM

[953] Kiros Berhane and Robert J. Tibshirani. Generalized additive models for longitudinal data. Canadian Journal of Statistics = Revue canadienne de statistique, 26(4):517–535, December 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Zidek:1998:ISM

[954] J. V. Zidek, N. D. Le, H. Wong, and R. T. Burnett. Including structural measurement errors in the nonlinear regression analysis of clustered data. Canadian Journal of Statistics = Revue canadienne de statistique, 26(4):537–548, December 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Lawless:1998:AIG

[955] J. F. Lawless and M. Zhan. Analysis of interval-grouped recurrent-event data using piecewise constant rate functions. Canadian Journal of Statistics = Revue canadienne de statistique, 26(4):549–565, December 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Gu:1998:SAA

[956] Ming Gao Gu and Shaolin Li. A stochastic approximation algorithm for maximum-likelihood estimation with incomplete data. Canadian Journal of Statistics = Revue canadienne de statistique, 26(4):567–582, December 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:1998:PLR

[957] Jiahua Chen. Penalized likelihood-ratio test for finite mixture models with multinomial observations. Canadian Journal of Statistics = Revue canadienne de statistique, 26(4):583–599, December 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Susko:1998:CNM

[958] E. Susko, J. D. Kalbfleisch, and J. Chen. Constrained nonparametric maximum-likelihood estimation for mixture models. Canadian Journal of Statistics = Revue canadienne de statistique, 26(4):601–617, December 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Yu:1998:APG

[959] Qiqing Yu, Anton Schick, Linxiong Li, and George Y. C. Wong. Asymptotic properties of the GMLE in the case 1 interval-censorship model with discrete inspection times. Canadian Journal of Statistics = Revue canadienne de statistique, 26(4):619–627, December 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Belisle:1998:SCG

[960] Claude Bélisle. Slow convergence of the Gibbs sampler. Canadian Journal of Statistics = Revue canadienne de statistique, 26(4):629–641, December 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Crump:1998:ENS

[961] Kenny S. Crump and Daniel Krewski. Estimation of the number of studies with positive trends when studies with negative trends are present. Canadian Journal of Statistics = Revue canadienne de statistique, 26 (4):643–655, December 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Viana:1998:SAM

[962] Marios A. G. Viana, André Rogatko, and Timothy R. Rebbeck. Statistical assessment of multivalued screening tests. Canadian Journal of Statistics = Revue canadienne de statistique, 26(4):657–668, December 1998. ISSN 0319-5724 (print), 1708-945X (electronic).

Genest:1999:ER

[963] Christian Genest. Editor’s report. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):1–2, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Banerjee:1999:BKR

[964] Mousumi Banerjee, Michelle Capozzoli, Laura McSweeney, and Debajyoti Sinha. Beyond kappa: a review of interrater agreement measures. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1): 3–23, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Manchester:1999:ROS

[965] L. Manchester, C. A. Field, and A. McDougall. Regression for overdetermined systems: a fisheries example. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):25–39, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Monga:1999:AOC

[966] E. Monga and S. Tardif. Asymptotic optimality of a class of rank tests for replicated Latin-square designs. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):41–52, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Pan:1999:DFS

[967] Guohua Pan and Winson Taam. Distribution-free subset selection for incompletely ranked data. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):53–62, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Sneddon:1999:SUL

[968] Gary Sneddon. Smoothing in an underdetermined linear model with random explanatory variables. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):63–79, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Collins:1999:RES

[969] John R. Collins. Robust M-estimators of scale: Minimax bias versus maximal variance. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):81–96, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Sun:1999:RPP

[970] Dongchu Sun and Keying Ye. Reference priors for a product of normal means when variances are unknown. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):97–103, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Petrone:1999:BDE

[971] Sonia Petrone. Bayesian density estimation using Bernstein polynomials. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1): 105–126, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Pensky:1999:EBE

[972] M. Pensky and R. S. Singh. Empirical Bayes estimation of reliability characteristics for an exponential family. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):127–136, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Kolassa:1999:CIP

[973] John E. Kolassa. Confidence intervals for parameters lying in a random polygon. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):149–161, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Pal:1999:SOP

[974] Nabendu Pal and Wooi K. Lim. Second-order properties of intraclass correlation estimators for a symmetric normal distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):163–172, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Heard:1999:EVM

[975] Andrew Heard and Tim Swartz. Extended voting measures. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):173–182, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Baron:1999:CRC

[976] Michael Baron. Convergence rates of change-point estimators and tail probabilities of the first-passage-time process. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):183–197, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Kokonendji:1999:PAP

[977] Célestin C. Kokonendji. Le problème d’Anscombe pour les lois binomiales négatives généralisées. (French) [The Anscombe problem of generalized negative binomial distributions]. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):199–205, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Blaker:1999:CSE

[978] Helge Blaker. A class of shrinkage estimators in linear regression. Canadian Journal of Statistics = Revue canadienne de statistique, 27(1):207– 220, March 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Feuerverger:1999:RAC

[979] Andrey Feuerverger, John Robinson, and Augustine Wong. On the relative accuracy of certain bootstrap procedures. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):225–236, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:1999:BAI

[980] Biao Zhang. Bootstrapping with auxiliary information. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):237–249, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

MacEachern:1999:SIS

[981] Steven N. MacEachern, Merlise Clyde, and Jun S. Liu. Sequential importance sampling for nonparametric Bayes models: the next generation. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2): 251–267, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Papandonatos:1999:BIA

[982] George D. Papandonatos and Seymour Geisser. Bayesian interim analysis of lifetime data. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):269–289, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Damien:1999:FBA

[983] Paul Damien and Stephen Walker. A full Bayesian analysis of circular data using the von Mises distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):291–298, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Cooley:1999:PEC

[984] Craig A. Cooley and Steven N. Maceachern. Prior elicitation in the classification problem. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):299–313, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Chan:1999:GOC

[985] Schultz Chan and Malay Ghosh. A geometric optimality of Cox’s partial likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):315–320, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Lee:1999:RQE

[986] Y. Lee and John A. Nelder. The robustness of the quasilikelihood estimator. Canadian Journal of Statistics = Revue canadienne de statistique, 27 (2):321–327, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:1999:IPT

[987] Hui Chen and Joseph P. Romano. An invariance principle for triangular arrays of dependent variables with application to autocovariance estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):329–343, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Mukherjee:1999:ABC

[988] Kanchan Mukherjee. The asymptotic behaviour of a class of L-estimators under long-range dependence. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):345–360, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Taplin:1999:RTL

[989] Ross H. Taplin. Robust F -tests for linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):361–371, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:1999:OPC

[990] Ping Zhang. The optimal prediction of cross-sectional proportions in categorical panel-data analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):373–382, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Nettleton:1999:ORH

[991] Dan Nettleton. Order-restricted hypothesis testing in a variation of the normal mixture model. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):383–394, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Krishnamoorthy:1999:CEN

[992] K. Krishnamoorthy and Maruthy K. Pannala. Confidence estimation of a normal mean vector with incomplete data. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):395–407, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Wegkamp:1999:QUB

[993] Marten H. Wegkamp. Quasi-universal bandwidth selection for kernel density estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):409–420, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Genest:1999:PST

[994] Christian Genest. Probability and statistics: a tale of two worlds? Canadian Journal of Statistics = Revue canadienne de statistique, 27(2):421– 444, June 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

McCullagh:1999:QSS

[995] Peter McCullagh. Quotient spaces and statistical models. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):447–456, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Mcdermott:1999:GOC

[996] Michael P. Mcdermott. Generalized orthogonal contrast tests for homogeneity of ordered means. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):457–470, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Bedrick:1999:ENN

[997] Edward J. Bedrick and Ronald Christensen. The effect of nonnormality on near-replicate lack-of-fit tests. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):471–484, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:1999:TLF

[998] Chin-Shang Li. Testing lack of fit of regression models under heteroscedasticity. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):485–496, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Knight:1999:AER

[999] Keith Knight. Asymptotics for L_{1}-estimators of regression parameters under heteroscedasticityY. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):497–507, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Ng:1999:RIB

[1000] Edmund T. M. Ng and Richard J. Cook. Robust inference for bivariate point processes. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):509–524, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Berlinet:1999:EDC

[1001] Alain Berlinet and Christian Francq. Estimation du comportement asymptotique des autocovariances et autocorrélations empiriques de processus multivariées. (French) [Estimation of the asymptotic behavior of autocovariances and empirical autocorrelations of multivariate processes]. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):525–546, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Yang:1999:BCS

[1002] Tae Young Yang and Lynn Kuo. Bayesian computation for the superposition of nonhomogeneous Poisson processes. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):547–556, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Carolan:1999:ABG

[1003] Chris Carolan and Richard Dykstra. Asymptotic behavior of the Grenander estimator at density flat regions. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):557–566, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Daniels:1999:PVH

[1004] Michael J. Daniels. A prior for the variance in hierarchical models. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):567– 578, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Hall:1999:MCA

[1005] Peter Hall, Michael A. Martin, and Shan Sun. Monte Carlo approximation to Edgeworth expansions. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):579–584, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Small:1999:MRE

[1006] Christopher G. Small and Zejiang Yang. Multiple roots of estimating functions. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):585–598, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Gupta:1999:MLE

[1007] A. K. Gupta, T. T. Nguyen, and Yining Wang. On maximum likelihood estimation of the binomial parameter n. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):599–606, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Rychlik:1999:ERD

[1008] Tomasz Rychlik. Error reduction in density estimation under shape restrictions. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):607–622, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Rivest:1999:SLM

[1009] Louis-Paul Rivest. Some linear model techniques for analyzing smallcircle spherical data. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):623–638, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Nettleton:1999:CPE

[1010] Dan Nettleton. Convergence properties of the EM algorithm in constrained parameter spaces. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):639–648, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Yuan:1999:MIL

[1011] Ao Yuan and Bertrand S. Clarke. A minimally informative likelihood for decision analysis: Illustration and robustness. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):649–665, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Ozturk:1999:AAS

[1012] Ömer Öztürk and Thomas P. Hettmansperger. Authors’ addendum to “Simultaneous robust estimation of location and scale parameters: a minimum-distance approach”. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):667, September 1999. ISSN 0319-5724 (print), 1708-945X (electronic). See [928].

Ozturk:1999:ASR

[1013] Ömer Öztürk and Thomas P. Hettmansperger. Addendum to: “Simultaneous robust estimation of location and scale parameters: a minimum-distance approach” [Canad. J. Statist. 26 (1998), no. 2, 217– 229; MR1648490 (99f:62046)]. Canadian Journal of Statistics = Revue canadienne de statistique, 27(3):667, 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Lawless:1999:RCS

[1014] Jerald F. Lawless. La revue canadiennede statistique statistical science: Concepts, opportunities and challenges. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):671–682, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Schmuland:1999:DFS

[1015] Byron Schmuland. Dirichlet forms: Some infinite-dimensional examples. Canadian Journal of Statistics = Revue canadienne de statistique, 27 (4):683–700, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Ibrahim:1999:BVS

[1016] Joseph G. Ibrahim, Ming-Hui Chen, and Steven N. MacEachern. Bayesian variable selection for proportional hazards models. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):701–717, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Seifu:1999:SBI

[1017] Yodit Seifu, Thomas A. Severini, and Martin A. Tanner. Semiparametric Bayesian inference for regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):719–734, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Fokianos:1999:GMS

[1018] Konstantinos Fokianos, Amy Peng, and Jing Qin. A generalized-moments specification test for the logistic link. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):735–750, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic). See addendum [1063].

Fang:1999:RED

[1019] Zhide Fang and Douglas P. Wiens. Robust extrapolation designs and weights for biased regression models with heteroscedastic errors. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):751– 770, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Jones:1999:NIT

[1020] Michael P. Jones. Nonrobustness of the information test in detecting heterogeneity. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):771–779, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Wiens:1999:MDA

[1021] Douglas P. Wiens and Julie Zhou. Minimax designs for approximately linear models with AR(1) errors. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):781–794, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Adjengue:1999:DME

[1022] Luc D. Adjengue and Marc Moore. Deux méthodes d’estimation pour les paramètres de processus moyenne mobile spatiaux. (French) [Two estimation methods for parameters of spatial median processes]. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):795–818, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Masse:1999:CLD

[1023] Benoît R. Mâsse and Young K. Truong. Conditional logspline density estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):819–832, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Iyengar:1999:FIW

[1024] Satish Iyengar, Paul Kvam, and Harshinder Singh. Fisher information in weighted distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):833–841, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Enns:1999:CPD

[1025] E. G. Enns, P. F. Ehlers, and T. Misi. A cluster problem as defined by nearest neighbours. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):843–851, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Belbachir:1999:LLP

[1026] Mohammadine Belbachir. Lois limités pour les statistiques d’ordre dans le cas non identiquement distribué. (French) [Limited laws for order statistics in the nonidentically distributed case]. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):853–862, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Hartlaub:1999:RBT

[1027] Bradley A. Hartlaub, Angela M. Dean, and Douglas A. Wolfe. Rankbased test procedures for interaction in the two-way layout with one observation per cell. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):863–874, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Marrero:1999:LVS

[1028] Osvaldo Marrero. L’analyse de la variation saisonnière quand l’amplitude et la taille sont faibles. (French) [The analysis of seasonal variation when the amplitude and the size are low]. Canadian Journal of Statistics = Revue canadienne de statistique, 27(4):875–882, December 1999. ISSN 0319-5724 (print), 1708-945X (electronic).

Genest:2000:ER

[1029] Christian Genest. Editor’s report. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):1–2, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2000:DEH

[1030] Fan Zhang and Robert E. Weiss. Diagnosing explainable heterogeneity of variance in random-effects models. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):3–18, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Kneip:2000:CRL

[1031] A. Kneip, X. Li, K. B. MacGibbon, and J. O. Ramsay. Curve registration by local regression. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):19–29, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Grenier:2000:BRM

[1032] Michèle Grenier and Christian Léger. Bootstrapping regression models with BLUS residuals. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):31–43, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Sashegyi:2000:EEB

[1033] Andreas I. Sashegyi, K. Stephen Brown, and Patrick J. Farrell. Estimation in an empirical Bayes model for longitudinal and cross-sectionally clustered binary data. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):45–63, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Rai:2000:MSM

[1034] S. N. Rai, D. E. Matthews, and D. R. Krewski. Mixed-scale models for survival/sacrifice experiments. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):65–80, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Almudevar:2000:ECR

[1035] Anthony Almudevar. Exact confidence regions for species assignment based on DNA markers. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):81–95, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Ozturk:2000:IRS

[1036] Ömer Öztürk and Douglas A. Wolfe. An improved ranked set two-sample Mann–Whitney–Wilcoxon test. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):123–135, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Froda:2000:USE

[1037] Sorana Froda and Constance Van Eeden. A uniform saddlepoint expansion for the null-distribution of the Wilcoxon–Mann–Whitney statistic. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1): 137–149, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Averous:2000:LRD

[1038] Jean Averous and Jean-Luc Dortet-Bernadet. LTD and RTI dependence orderings. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):151–157, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Fernandez:2000:PEU

[1039] Miguel A. Fernández, Cristina Rueda, and Bonifacio Salvador. Parameter estimation under orthant restrictions. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):171–181, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Terhune:2000:CSD

[1040] Jack Terhune. Case study in data analysis: Vocalization differences among three harp seal herds. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):183–185, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Esterby:2000:CSD

[1041] Sylvia R. Esterby et al. Case study in data analysis: vocalization differences among three harp seal herds. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):183–219, 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Mathieu:2000:AUV

[1042] Patrice Mathieu, Anne-Marie Robert, and Urs. R. Maag. Analyse 1: Utilisation des vocalisations pour différencier trois hardes de phoques. (French) [Analysis 1: Using vocalizations to differentiate three herds of seals]. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):185–189, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Chan:2000:AST

[1043] Keith Chan, Jaroslaw Harezlak, Robert St-Aubin, and Marc Vallée. Analysis 2: Seals, talk to us and we will identify you. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):189–197, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Farruggia:2000:AWM

[1044] Joseph Farruggia, Peter D. M. MacDonald, and Román Viveros-Aguilera. Analysis 3: Who made the call? Classification based on logistic regression and trees. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):197–205, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Allard:2000:ADA

[1045] Jacques Allard, Vartan Choulakian, René LeBlanc, Stephanie J. Mac-Neill, and Smail Mahdi. Analysis 4: Discriminant analysis of seal data. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1): 205–212, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Esterby:2000:DA

[1046] Sylvia R. Esterby. Discussion of the analyses. Canadian Journal of Statistics = Revue canadienne de statistique, 28(1):213–219, March 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Ramsay:2000:DEM

[1047] James O. Ramsay. Differential equation models for statistical functions. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2): 225–240, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Heckman:2000:PRM

[1048] Nancy E. Heckman and James O. Ramsay. Penalized regression with model-based penalties. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):241–258, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Dupuis:2000:RRT

[1049] Debbie J. Dupuis and David C. Hamilton. Regression residuals and test statistics: Assessing naive outlier deletion. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):259–275, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Bilodeau:2000:REM

[1050] Martin Bilodeau and Pierre Duchesne. Robust estimation of the SUR model. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):277–288, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Pinkse:2000:NTS

[1051] Joris Pinkse. Nonparametric two-step regression estimation when regressors and error are dependent. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):289–300, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Muller:2000:NRT

[1052] Ursula U. Müller. Nonparametric regression for threshold data. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):301–310, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Butler:2000:BPW

[1053] Ronald W. Butler and Aparna V. Huzurbazar. Bayesian prediction of waiting times in stochastic models. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):311–325, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Conigliani:2000:BAG

[1054] Caterina Conigliani, J. Iván Castro, and Anthony O’Hagan. Bayesian assessment of goodness of fit against nonparametric alternatives. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):327–342, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Conigliani:2000:SFB

[1055] Caterina Conigliani and Anthony O’Hagan. Sensitivity of the fractional Bayes factor to prior distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):343–352, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Reed:2000:RHF

[1056] William J. Reed. Reconstructing the history of forest fire frequency: Identifying hazard rate change points using the Bayes information criterion. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):353–365, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Cuevas:2000:ENC

[1057] Antonio Cuevas, Manuel Febrero, and Ricardo Fraiman. Estimating the number of clusters. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):367–382, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic). See corrigenda [1105].

Kim:2000:DMM

[1058] Peter T. Kim and Ja-Yong Koo. Directional mixture models and optimal estimation of the mixing density. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):383–398, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic). See addendum [1106].

Fourdrinier:2000:EMS

[1059] Dominique Fourdrinier and Idir Ouassou. Estimation of the mean of a spherically symmetric distribution with constraints on the norm. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):399– 415, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Provost:2000:DLC

[1060] Serge B. Provost and Young-Ho Cheong. On the distribution of linear combinations of the components of a Dirichlet random vector. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):417–425, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Edwardes:2000:IRC

[1061] Michael D. Deb. Edwardes. Implications of random cut-points theory for the Mann–Whitney and binomial tests. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):427–438, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Petersen:2000:EOC

[1062] Martin R. Petersen and James A. Deddens. Effects of omitting a covariate in Poisson models when the data are balanced. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):439–445, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Fokianos:2000:AAG

[1063] Konstantinos Fokianos, Amy Peng, and Jing Qin. Authors’ addendum to: “A generalized-moments specification test for the logistic link” [Canad. J. Statist. 27 (1999), no. 4, 735–750; 1767144]. Canadian Journal of Statistics = Revue canadienne de statistique, 28(2):446, June 2000. ISSN 0319-5724 (print), 1708-945X (electronic). See [1018].

Hu:2000:EFB

[1064] Feifang Hu and John D. Kalbfleisch. The estimating function bootstrap. Canadian Journal of Statistics = Revue canadienne de statistique, 28 (3):449–481, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic). With discussion and rejoinder by the authors.

Anonymous:2000:D

[1065] Anonymous. Discussions. Canadian Journal of Statistics = Revue canadienne de statistique, 28(3):482–495, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2000:R

[1066] Anonymous. Rejoinders. Canadian Journal of Statistics = Revue canadienne de statistique, 28(3):496–499, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Polansky:2000:SBC

[1067] Alan M. Polansky. Stabilizing bootstrap-t confidence intervals for small samples. Canadian Journal of Statistics = Revue canadienne de statistique, 28(3):501–516, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Stern:2000:LIS

[1068] Steven E. Stern and A. H. Welsh. Likelihood inference for small variance components. Canadian Journal of Statistics = Revue canadienne de statistique, 28(3):517–532, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Dominici:2000:CAM

[1069] Francesca Dominici, Giovanni Parmigiani, and Merlise Clyde. Conjugate analysis of multivariate normal data with incomplete observations. Canadian Journal of Statistics = Revue canadienne de statistique, 28 (3):533–550, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Tong:2000:SEN

[1070] Barbara Tong and Kert Viele. Smooth estimates of normal mixtures. Canadian Journal of Statistics = Revue canadienne de statistique, 28 (3):551–561, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Deng:2000:STZ

[1071] Dianliang Deng and Sudhir R. Paul. Score tests for zero inflation in generalized linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 28(3):563–570, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Gao:2000:AEP

[1072] Jiti Gao and Thomas Yee. Adaptive estimation in partially linear autoregressive models. Canadian Journal of Statistics = Revue canadienne de statistique, 28(3):571–586, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Ferguson:2000:KTS

[1073] Thomas S. Ferguson, Christian Genest, and Marc Hallin. Kendall’s tau for serial dependence. Canadian Journal of Statistics = Revue canadienne de statistique, 28(3):587–604, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2000:GEN

[1074] Min Chen and Gemai Chen. Geometric ergodicity of nonlinear autoregressive models with changing conditional variances. Canadian Journal of Statistics = Revue canadienne de statistique, 28(3):605–613, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Dauxois:2000:IPM

[1075] Jean-Yves Dauxois. Inférence par les martingales pour des processus ponctuels à compensateur discontinu. (French) [Inference by martingales for point processes with discontinuous compensator]. Canadian Journal of Statistics = Revue canadienne de statistique, 28(3):615–620, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2000:NEE

[1076] Bing Li. Nonparametric estimating equations based on a penalized information criterion. Canadian Journal of Statistics = Revue canadienne de statistique, 28(3):621–639, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Song:2000:MCK

[1077] Peter Xue-Kun Song. Monte Carlo Kalman filter and smoothing for multivariate discrete state space models. Canadian Journal of Statistics = Revue canadienne de statistique, 28(3):641–652, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Diack:2000:CTS

[1078] Cheikh A. T. Diack. Sur la convergence des tests de Schlee et de Yatchew. (French) [On the convergence tests of Schlee and Yatchew]. Canadian Journal of Statistics = Revue canadienne de statistique, 28(3):653–668, September 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

DasGupta:2000:CBF

[1079] Anirban DasGupta, George Casella, Mohan Delampady, Christian Genest, Herman Rubin, and William E. Strawderman. Correlation in a Bayesian framework. Canadian Journal of Statistics = Revue canadienne de statistique, 28(4):675–687, December 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Ghoudi:2000:CMR

[1080] Kilani Ghoudi and David Mcdonald. Cramér-von Mises regression. Canadian Journal of Statistics = Revue canadienne de statistique, 28(4):689– 714, December 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Ishwaran:2000:UMO

[1081] Hemant Ishwaran. Univariate and multirater ordinal cumulative link regression with covariate specific cutpoints. Canadian Journal of Statistics = Revue canadienne de statistique, 28(4):715–730, December 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Ishwaran:2000:GCH

[1082] Hemant Ishwaran and Constantine A. Gatsonis. A general class of hierarchical ordinal regression models with applications to correlated ROC analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 28(4):731–750, December 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Binder:2000:VET

[1083] David A. Binder, Colin Babyak, Marie Brodeur, Michel Hidiroglou, and Wisner Jocelyn. Variance estimation for two-phase stratified sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 28 (4):751–764, December 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Baillo:2000:SEN

[1084] Amparo Baíllo, Antonio Cuevas, and Ana Justel. Set estimation and nonparametric detection. Canadian Journal of Statistics = Revue canadienne de statistique, 28(4):765–782, December 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Blaker:2000:CCI

[1085] Helge Blaker. Confidence curves and improved exact confidence intervals for discrete distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 28(4):783–798, December 2000. ISSN 0319-5724 (print), 1708-945X (electronic). See corrigenda [1140].

Saavedra:2000:EMD

[1086] Angeles Saavedra and Ricardo Cao. On the estimation of the marginal density of a moving average process. Canadian Journal of Statistics = Revue canadienne de statistique, 28(4):799–815, December 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Laib:2000:RNE

[1087] Naämane Laïb and Elias Ould-S ïd. A robust nonparametric estimation of the autoregression function under an ergodic hypothesis. Canadian Journal of Statistics = Revue canadienne de statistique, 28(4):817–828, December 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Cabana:2000:TSB

[1088] Alejandra Cabaña and Enrique M. Cabaña. Tests of symmetry based on transformed empirical processes. Canadian Journal of Statistics = Revue canadienne de statistique, 28(4):829–839, December 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhong:2000:ELI

[1089] Bob Zhong, Jiahua Chen, and J. N. K. Rao. Empirical likelihood inference in the presence of measurement error. Canadian Journal of Statistics = Revue canadienne de statistique, 28(4):841–852, December 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Frey:2000:TBE

[1090] Michael R. Frey and Carolin M. Frey. Tolerance-based error reduction in multi-agent bioassays. Canadian Journal of Statistics = Revue canadienne de statistique, 28(4):853–874, December 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Hudson:2000:MIE

[1091] Suzanne Hudson and Paul W. Vos. Marginal information for expectation parameters. Canadian Journal of Statistics = Revue canadienne de statistique, 28(4):875–886, December 2000. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2001:ERF

[1092] Anonymous. Editor: Report from the former editor. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1):3–4, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Ghosh:2001:BFP

[1093] Malay Ghosh and Yeong-Hwa Kim. The Behrens-Fisher problem revisited: a Bayes-frequentist synthesis. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1):5–17, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Eno:2001:PMP

[1094] Daniel R. Eno and Keying Ye. Probability matching priors for an extended statistical calibration model. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1):19–35, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Brown:2001:NSU

[1095] Patrick E. Brown and Piet De Jong. Nonparametric smoothing using state space techniques. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1):37–50, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Viele:2001:EFF

[1096] Kert Viele. Evaluating fit in functional data analysis using model embeddings. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1):51–66, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

DiCiccio:2001:SAO

[1097] Thomas J. DiCiccio, Michael A. Martin, and Steven E. Stern. Simple and accurate one-sided inference from signed roots of likelihood ratios. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1): 67–76, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Hall:2001:AEQ

[1098] Daniel B. Hall. On the application of extended quasi-likelihood to the clustered data case. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1):77–97, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Kabaila:2001:ESP

[1099] Paul Kabaila and John Byrne. Exact short Poisson confidence intervals. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1): 99–106, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Melfi:2001:ARD

[1100] Vincent F. Melfi, Connie Page, and Margarida Geraldes. An adaptive randomized design with application to estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1):107–116, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Heo:2001:RMR

[1101] Giseon Heo, Byron Schmuland, and Douglas P. Wiens. Restricted minimax robust designs for misspecified regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1):117–128, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Tsao:2001:REL

[1102] Min Tsao and Julie Zhou. On the robustness of empirical likelihood ratio confidence intervals for location. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1):129–140, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Hallin:2001:PHP

[1103] Marc Hallin, Amal Mellouk, and Khalid Rifi. Projection de Hájek et polynômes de Bernstein. (French) [Hájek projection and Bernstein polynomials]. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1):141–154, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

ElBantli:2001:ABE

[1104] Faouzi El Bantli and Marc Hallin. Asymptotic behaviour of M-estimators in AR(p) models under nonstandard conditions. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1):155–168, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Cuevas:2001:CEN

[1105] Antonio Cuevas, Manuel Febrero, and Ricardo Fraiman. Corrigenda: “Estimating the number of clusters”. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1):170, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic). See [1057].

Kim:2001:ADM

[1106] Peter T. Kim and Ja-Yong Koo. Addendum: “Directional mixture models and optimal estimation of the mixing density”. Canadian Journal of Statistics = Revue canadienne de statistique, 29(1):170, March 2001. ISSN 0319-5724 (print), 1708-945X (electronic). See [1058].

Canette:2001:BNR

[1107] Isabel Cañette. Blind nonparametric regression. Canadian Journal of Statistics = Revue canadienne de statistique, 29(2):173–189, June 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Merola:2001:DRA

[1108] Giovanni M. Merola and Bovas Abraham. Dimensionality reduction approach to multivariate prediction. Canadian Journal of Statistics = Revue canadienne de statistique, 29(2):191–200, June 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2001:LRT

[1109] Hanfeng Chen and Jiahua Chen. The likelihood ratio test for homogeneity in finite mixture models. Canadian Journal of Statistics = Revue canadienne de statistique, 29(2):201–215, June 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Tingley:2001:DPN

[1110] M. A. Tingley and L. McLean. Detection of patterns in noisy time series. Canadian Journal of Statistics = Revue canadienne de statistique, 29(2): 217–237, June 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Lei:2001:DTB

[1111] S.-Y. Claire Lei and Suojin Wang. Diagnostic tests for bias of estimating equations in weighted regression with missing covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 29(2):239–250, June 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Antoniadis:2001:WMU

[1112] Anestis Antoniadis, Jianqing Fan, and Irène Gijbels. A wavelet method for unfolding sphere size distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 29(2):251–268, June 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

vanderVaart:2001:CSM

[1113] Aad van der Vaart and Jon A. Wellner. Consistency of semiparametric maximum likelihood estimators for two-phase sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 29(2):269–288, June 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:2001:VEF

[1114] Changbao Wu and Randy R. Sitter. Variance estimation for the finite population distribution function with complete auxiliary information. Canadian Journal of Statistics = Revue canadienne de statistique, 29 (2):289–307, June 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Puig:2001:GFT

[1115] Pedro Puig and Michael A. Stephens. Goodness-of-fit tests for the hyperbolic distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 29(2):309–320, June 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Gill:2001:SAR

[1116] Paramjit S. Gill and Tim B. Swartz. Statistical analyses for round robin interaction data. Canadian Journal of Statistics = Revue canadienne de statistique, 29(2):321–331, June 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Alqallaf:2001:CVB

[1117] Fatemah Alqallaf and Paul Gustafson. On cross-validation of Bayesian models. Canadian Journal of Statistics = Revue canadienne de statistique, 29(2):333–340, June 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Hooper:2001:FRM

[1118] Peter M. Hooper. Flexible regression modeling with adaptive logistic basis functions. Canadian Journal of Statistics = Revue canadienne de statistique, 29(3):343–364, September 2001. ISSN 0319-5724 (print), 1708-945X (electronic). See discussion [1119] and rejoinder [1120].

Lindstrom:2001:D

[1119] Mary J. Lindstrom, James O. Ramsay, Nancy E. Heckman, Hugh A. Chipman, and Hong Gu. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 29(3):365–374, September 2001. ISSN 0319-5724 (print), 1708-945X (electronic). See [1118].

Hooper:2001:R

[1120] Peter M. Hooper. Rejoinder. Canadian Journal of Statistics = Revue canadienne de statistique, 29(3):374–378, September 2001. ISSN 0319-5724 (print), 1708-945X (electronic). See [1118].

Susko:2001:ERT

[1121] Edward Susko, Michael J. Bronskill, Simon J. Graham, and Robert J. Tibshirani. Estimation of relaxation time distributions in magnetic resonance imaging. Canadian Journal of Statistics = Revue canadienne de statistique, 29(3):379–394, September 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Rosychuk:2001:SMM

[1122] Rhonda J. Rosychuk and Mary E. Thompson. A semi-Markov model for binary longitudinal responses subject to misclassification. Canadian Journal of Statistics = Revue canadienne de statistique, 29(3):395–404, September 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Dean:2001:MRH

[1123] Charmaine B. Dean and Ying Cai MacNab. Modeling of rates over a hierarchical health administrative structure. Canadian Journal of Statistics = Revue canadienne de statistique, 29(3):405–419, September 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Cho:2001:SAL

[1124] Meehyung Cho, Nathaniel Schenker, Jeremy M. G. Taylor, and Dongliang Zhuang. Survival analysis with long-term survivors and partially observed covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 29(3):421–436, September 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Delampady:2001:NAD

[1125] Mohan Delampady, Anirban DasGupta, George Casella, Herman Rubin, and William E. Strawderman. A new approach to default priors and robust Bayes methodology. Canadian Journal of Statistics = Revue canadienne de statistique, 29(3):437–450, September 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Spinelli:2001:TFG

[1126] John J. Spinelli. Testing fit for the grouped exponential distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 29 (3):451–458, September 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

OGorman:2001:UAW

[1127] Thomas W. O’Gorman. Using adaptive weighted least squares to reduce the lengths of confidence intervals. Canadian Journal of Statistics = Revue canadienne de statistique, 29(3):459–471, September 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Croux:2001:RLD

[1128] Christophe Croux and Catherine Dehon. Robust linear discriminant analysis using S-estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 29(3):473–493, September 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Huang:2001:PSR

[1129] Y. H. Steve Huang and Longcheen Huwang. On the polynomial structural relationship. Canadian Journal of Statistics = Revue canadienne de statistique, 29(3):495–512, September 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Asgharian:2001:CMC

[1130] Masoud Asgharian and David B. Wolfson. Covariates in multipath change-point problems: Modelling and consistency of the MLE. Canadian Journal of Statistics = Revue canadienne de statistique, 29(4):515– 528, December 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Le:2001:SPT

[1131] Nhu D. Le, Li Sun, and James V. Zidek. Spatial prediction and temporal backcasting for environmental fields having monotone data patterns. Canadian Journal of Statistics = Revue canadienne de statistique, 29 (4):529–554, December 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Rivest:2001:ILL

[1132] Louis-Paul Rivest and Tina Lévesque. Improved log-linear model estimators of abundance in capture-recapture experiments. Canadian Journal of Statistics = Revue canadienne de statistique, 29(4):555–572, December 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Frees:2001:OVL

[1133] Edward W. Frees. Omitted variables in longitudinal data models. Canadian Journal of Statistics = Revue canadienne de statistique, 29(4):573– 595, December 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2001:ELL

[1134] Qihua Wang and J. N. K. Rao. Empirical likelihood for linear regression models under imputation for missing responses. Canadian Journal of Statistics = Revue canadienne de statistique, 29(4):597–608, December 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Gautam:2001:ICS

[1135] Shiva Gautam, Allan Sampson, and Harshinder Singh. Iso-chi-squared testing of 2 × k ordered tables. Canadian Journal of Statistics = Revue canadienne de statistique, 29(4):609–619, December 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Hengartner:2001:ESP

[1136] Nicolas Hengartner and Marten Wegkamp. Estimation and selection procedures in regression: an L_{1} approach. Canadian Journal of Statistics = Revue canadienne de statistique, 29(4):621–632, December 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Dette:2001:RPM

[1137] Holger Dette, Dale Song, and Weng Kee Wong. Robustness properties of minimally-supported Bayesian D-optimal designs for heteroscedastic models. Canadian Journal of Statistics = Revue canadienne de statistique, 29(4):633–647, December 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2001:NTC

[1138] Min Chen and Gemai Chen. A nonparametric test of conditional autoregressive heteroscedasticity for threshold autoregressive models. Canadian Journal of Statistics = Revue canadienne de statistique, 29(4):649–666, December 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Bertail:2001:ESD

[1139] Patrice Bertail and Dimitris N. Politis. Extrapolation of subsampling distribution estimators: the i.i.d. and strong mixing cases. Canadian Journal of Statistics = Revue canadienne de statistique, 29(4):667–680, December 2001. ISSN 0319-5724 (print), 1708-945X (electronic).

Blaker:2001:CCC

[1140] Helge Blaker. Corrigenda: “Confidence curves and improved exact confidence intervals for discrete distributions” [Canad. J. Statist. 28 (2000), no. 4, 783–798; 1 821 434]. Canadian Journal of Statistics = Revue canadienne de statistique, 29(4):681, December 2001. ISSN 0319-5724 (print), 1708-945X (electronic). See [1085].

Lockhart:2002:ER

[1141] Richard A. Lockhart. Editor’s report. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1):1–2, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

You:2002:SAE

[1142] Yong You and J. N. K. Rao. Small area estimation using unmatched sampling and linking models. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1):3–15, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Dupuis:2002:RWL

[1143] Debbie J. Dupuis and Stephan Morgenthaler. Robust weighted likelihood estimators with an application to bivariate extreme value problems. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1): 17–36, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Belisle:2002:BEC

[1144] Patrick Bélisle, Lawrence Joseph, David B. Wolfson, and Xiaojie Zhou. Bayesian estimation of cognitive decline in patients with Alzheimer’s disease. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1):37–54, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Ibrahim:2002:BMG

[1145] Joseph G. Ibrahim, Ming-Hui Chen, and Stuart R. Lipsitz. Bayesian methods for generalized linear models with covariates missing at random. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1): 55–78, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Trottini:2002:GPC

[1146] Mario Trottini and Fulvio Spezzaferri. A generalized predictive criterion for model selection. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1):79–96, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Ohman-Strickland:2002:AES

[1147] Pamela Ohman-Strickland and George Casella. Approximate and estimated saddlepoint approximations. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1):97–108, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Lim:2002:DAC

[1148] Yong B. Lim, Jerome Sacks, W. J. Studden, and William J. Welch. Design and analysis of computer experiments when the output is highly correlated over the input space. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1):109–126, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Tang:2002:BMN

[1149] Boxin Tang, Fengshi Ma, Debra Ingram, and Hong Wang. Bounds on the maximum number of clear two-factor interactions for 2^{m−p} designs of resolution III and IV. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1):127–136, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2002:CSD

[1150] Anonymous. Case study in data analysis: the genetic analysis of inflammatory bowel disease. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1):137–145, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Esterby:2002:CSD

[1151] Sylvia R. Esterby et al. Case study in data analysis: the genetic analysis of inflammatory bowel disease. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1):137–174, 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2002:AGD

[1152] Jiahua Chen, John D. Kalbfleisch, and Sandra Romero-Hidalgo. Analysis 2: Genetic data analysis of affected sib pairs. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1):145–152, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Darlington:2002:AGA

[1153] Gerarda A. Darlington and Andrew D. Paterson. Analysis 3: Genetic analysis of chromosome 6 in inflammatory bowel disease. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1):152–158, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Roslin:2002:AGA

[1154] Nicole M. Roslin, J. C. Loredo-Osti, Celia M. T. Greenwood, and Kenneth Morgan. Analysis 4: Genetic analysis of the role of the HLA region in inflammatory bowel disease. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1):158–167, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Field:2002:DEC

[1155] Christopher A. Field and Bruce Smith. Discussion of the evaluation of a candidate genetic locus in a genome scan of complex disease. Canadian Journal of Statistics = Revue canadienne de statistique, 30(1):167–169, March 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2002:BCT

[1156] Gemai Chen, Richard A. Lockhart, and Michael A. Stephens. Box– Cox transformations in linear models: Large sample theory and tests of normality. Canadian Journal of Statistics = Revue canadienne de statistique, 30(2):177–209, June 2002. ISSN 0319-5724 (print), 1708-945X (electronic). See discussion [1157] and rejoinder [1158].

Anonymous:2002:D

[1157] Anonymous. Discussion. Canadian Journal of Statistics = Revue canadienne de statistique, 30(2):210–225, June 2002. ISSN 0319-5724 (print), 1708-945X (electronic). See [1156].

Chen:2002:R

[1158] Gemai Chen, Richard A. Lockhart, and Michael A. Stephens. Rejoinder. Canadian Journal of Statistics = Revue canadienne de statistique, 30(2): 226–234, June 2002. ISSN 0319-5724 (print), 1708-945X (electronic). See [1156].

Yang:2002:MET

[1159] Zhenlin Yang. Median estimation through a regression transformation. Canadian Journal of Statistics = Revue canadienne de statistique, 30(2): 235–242, June 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

McLeish:2002:HLS

[1160] Donald L. McLeish. Highs and lows: Some properties of the extremes of a diffusion and applications in finance. Canadian Journal of Statistics = Revue canadienne de statistique, 30(2):243–267, June 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Ishwaran:2002:EAS

[1161] Hemant Ishwaran and Mahmoud Zarepour. Exact and approximate sum representations for the Dirichlet process. Canadian Journal of Statistics = Revue canadienne de statistique, 30(2):269–283, June 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Cuevas:2002:LFR

[1162] Antonio Cuevas, Manuel Febrero, and Ricardo Fraiman. Linear functional regression: the case of fixed design and functional response. Canadian Journal of Statistics = Revue canadienne de statistique, 30(2):285– 300, June 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Oyet:2002:MOI

[1163] Alwell J. Oyet. Minimax A- and D-optimal integer-valued wavelet designs for estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 30(2):301–316, June 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Carolan:2002:LCM

[1164] Christopher A. Carolan. The least concave majorant of the empirical distribution function. Canadian Journal of Statistics = Revue canadienne de statistique, 30(2):317–328, June 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Genest:2002:WRO

[1165] Christian Genest and Mireille Guay. Worldwide research output in probability and statistics: an update. Canadian Journal of Statistics = Revue canadienne de statistique, 30(2):329–342, June 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Hu:2002:WL

[1166] Feifang Hu and James V. Zidek. The weighted likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 30(3):347–371, September 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Maldonado:2002:SAC

[1167] Yolanda Muñoz Maldonado, Joan G. Staniswalis, Louis N. Irwin, and Donna Byers. A similarity analysis of curves. Canadian Journal of Statistics = Revue canadienne de statistique, 30(3):373–381, September 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Susko:2002:ERD

[1168] Edward Susko and Robert Nadon. Estimation of a residual distribution with small numbers of repeated measurements. Canadian Journal of Statistics = Revue canadienne de statistique, 30(3):383–400, September 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Sinha:2002:MWG

[1169] Sanjoy Sinha and Douglas P. Wiens. Minimax weights for generalised Mestimation in biased regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 30(3):401–414, September 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Hall:2002:STH

[1170] Daniel B. Hall and Kenneth S. Berenhaut. Score tests for heterogeneity and overdispersion in zero-inflated Poisson and binomial regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 30 (3):415–430, September 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

You:2002:PEB

[1171] Yong You and J. N. K. Rao. A pseudo-empirical best linear unbiased prediction approach to small area estimation using survey weights. Canadian Journal of Statistics = Revue canadienne de statistique, 30(3):431–439, September 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Genest:2002:TSI

[1172] Christian Genest, Jean-François Quessy, and Bruno Rémillard. Tests of serial independence based on Kendall’s process. Canadian Journal of Statistics = Revue canadienne de statistique, 30(3):441–461, September 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Gustafson:2002:SEM

[1173] Paul Gustafson. On the simultaneous effects of model misspecification and errors in variables. Canadian Journal of Statistics = Revue canadienne de statistique, 30(3):463–474, September 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Sun:2002:GFT

[1174] Yanqing Sun, Sufang Cui, and Ram C. Tiwari. Goodness-of-fit tests for parametric models based on biased samples. Canadian Journal of Statistics = Revue canadienne de statistique, 30(3):475–490, September 2002. ISSN 0319-5724 (print), 1708-945X (electronic). See letter [1202].

Csorgo:2002:GIP

[1175] Miklós Csörgő. A glimpse of the impact of Pál Erdős on probability and statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 30(4):493–556, December 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Gentleman:2002:NEB

[1176] Robert Gentleman and Alain C. Vandal. Nonparametric estimation of the bivariate CDF for arbitrarily censored data. Canadian Journal of Statistics = Revue canadienne de statistique, 30(4):557–571, December 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

MacKay:2002:EOH

[1177] Rachel J. MacKay. Estimating the order of a hidden Markov model. Canadian Journal of Statistics = Revue canadienne de statistique, 30 (4):573–589, December 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Pewsey:2002:TCS

[1178] Arthur Pewsey. Testing circular symmetry. Canadian Journal of Statistics = Revue canadienne de statistique, 30(4):591–600, December 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Sinha:2002:RSD

[1179] Sanjoy Sinha and Douglas P. Wiens. Robust sequential designs for nonlinear regression. Canadian Journal of Statistics = Revue canadienne de statistique, 30(4):601–618, December 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Gu:2002:PLR

[1180] Chong Gu and Young-Ju Kim. Penalized likelihood regression: General formulation and efficient approximation. Canadian Journal of Statistics = Revue canadienne de statistique, 30(4):619–628, December 2002. ISSN 0319-5724 (print), 1708-945X (electronic).

Lockhart:2003:ERR

[1181] Richard A. Lockhart. Editor’s report/rapport du rédacteur en chef. Canadian Journal of Statistics = Revue canadienne de statistique, 31 (1):1–2, March 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Sahu:2003:FDB

[1182] Sujit K. Sahu and Russell C. H. Cheng. A fast distance-based approach for determining the number of components in mixtures. Canadian Journal of Statistics = Revue canadienne de statistique, 31(1):3–22, March 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Abraham:2003:SEM

[1183] Christophe Abraham, Gérard Biau, and Benoît Cadre. Simple estimation of the mode of a multivariate density. Canadian Journal of Statistics = Revue canadienne de statistique, 31(1):23–34, March 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Genest:2003:BMR

[1184] Christian Genest and Jean-François Plante. On Blest’s measure of rank correlation. Canadian Journal of Statistics = Revue canadienne de statistique, 31(1):35–52, March 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2003:ELC

[1185] Jiahua Chen, Shun-Yi Chen, and J. N. K. Rao. Empirical likelihood confidence intervals for the mean of a population containing many zero values. Canadian Journal of Statistics = Revue canadienne de statistique, 31(1): 53–68, March 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Mukherjee:2003:EOS

[1186] Bhramar Mukherjee. Exactly optimal sampling designs for processes with a product covariance structure. Canadian Journal of Statistics = Revue canadienne de statistique, 31(1):69–87, March 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Bouezmarni:2003:CBK

[1187] Taoufik Bouezmarni and Jean-Marie Rolin. Consistency of the beta kernel density function estimator. Canadian Journal of Statistics = Revue canadienne de statistique, 31(1):89–98, March 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhou:2003:RED

[1188] Julie Zhou and Hongtu Zhu. Robust estimation and design procedures for the random effects model. Canadian Journal of Statistics = Revue canadienne de statistique, 31(1):99–110, March 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Malfait:2003:HFL

[1189] Nicole Malfait and James O. Ramsay. The historical functional linear model. Canadian Journal of Statistics = Revue canadienne de statistique, 31(2):115–128, June 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Sahu:2003:NCM

[1190] Sujit K. Sahu, Dipak K. Dey, and Márcia D. Branco. A new class of multivariate skew distributions with applications to Bayesian regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 31(2):129–150, June 2003. ISSN 0319-5724 (print), 1708-945X (electronic). See erratum [1444].

Zhang:2003:ATR

[1191] Chunming M. Zhang. Adaptive tests of regression functions via multiscale generalized likelihood ratios. Canadian Journal of Statistics = Revue canadienne de statistique, 31(2):151–171, June 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Claeskens:2003:ELC

[1192] Gerda Claeskens, Bing-Yi Jing, Liang Peng, and Wang Zhou. Empirical likelihood confidence regions for comparison distributions and ROC curves. Canadian Journal of Statistics = Revue canadienne de statistique, 31(2):173–190, June 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Guedon:2003:NER

[1193] Yann Guédon and Christiane Cocozza-Thivent. Nonparametric estimation of renewal processes from count data. Canadian Journal of Statistics = Revue canadienne de statistique, 31(2):191–223, June 2003. ISSN 0319-5724 (print), 1708-945X (electronic). See correction [1210].

Perevozskaya:2003:ODP

[1194] Inna Perevozskaya, William F. Rosenberger, and Linda M. Haines. Optimal design for the proportional odds model. Canadian Journal of Statistics = Revue canadienne de statistique, 31(2):225–235, June 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Trevisani:2003:IBE

[1195] Matilde Trevisani and Alan E. Gelfand. Inequalities between expected marginal log-likelihoods, with implications for likelihood-based model complexity and comparison measures. Canadian Journal of Statistics = Revue canadienne de statistique, 31(3):239–250, September 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Wright:2003:LFE

[1196] Deanne L. Wright, Hal S. Stern, and Noel Cressie. Loss functions for estimation of extrema with an application to disease mapping. Canadian Journal of Statistics = Revue canadienne de statistique, 31(3):251–266, September 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Bunea:2003:NPM

[1197] Florentina Bunea and Marten H. Wegkamp. A note on penalized minimum distance estimation in nonparametric regression. Canadian Journal of Statistics = Revue canadienne de statistique, 31(3):267–274, September 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Duchesne:2003:TMA

[1198] Pierre Duchesne and Simon Lalancette. On testing for multivariate ARCH effects in vector time series models. Canadian Journal of Statistics = Revue canadienne de statistique, 31(3):275–292, September 2003. ISSN 0319-5724 (print), 1708-945X (electronic). See erratum [1497].

Zhu:2003:LIG

[1199] Hong-Tu Zhu and Sik-Yum Lee. Local influence for generalized linear mixed models. Canadian Journal of Statistics = Revue canadienne de statistique, 31(3):293–309, September 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Miller:2003:AUF

[1200] Arden E. Miller. The analysis of unreplicated factorial experiments from a geometric perspective. Canadian Journal of Statistics = Revue canadienne de statistique, 31(3):311–327, September 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Kolkiewicz:2003:BME

[1201] Adam W. Kolkiewicz. Best monotone M-estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 31(3):329–347, September 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Asgharian:2003:LEL

[1202] Masoud Asgharian. Letter to the Editor/lettre à la rédaction: Biased sampling with right censoring: a note on Sun, Cui & Tiwari (2002) [“Goodness-of-fit tests for parametric models based on biased samples” [Canad. J. Statist. 30 (2002), no. 3, 475–490; MR1944375] by Y. Sun, S. Cui and R. C. Tiwari]. Canadian Journal of Statistics = Revue canadienne de statistique, 31(3):349–350, September 2003. ISSN 0319-5724 (print), 1708-945X (electronic). See [1174].

Abdous:2003:NWS

[1203] Belkacem Abdous, Kilani Ghoudi, and Bruno Rémillard. Nonparametric weighted symmetry tests. Canadian Journal of Statistics = Revue canadienne de statistique, 31(4):357–381, December 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Lohr:2003:SAE

[1204] Sharon L. Lohr and N. G. N. Prasad. Small area estimation with auxiliary survey data. Canadian Journal of Statistics = Revue canadienne de statistique, 31(4):383–396, December 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Shi:2003:MRD

[1205] Peilin Shi, Jane J. Ye, and Julie Zhou. Minimax robust designs for misspecified regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 31(4):397–414, December 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Sutradhar:2003:QLI

[1206] Brajendra C. Sutradhar and R. Prabhakar Rao. On quasi-likelihood inference in generalized linear mixed models with two components of dispersion. Canadian Journal of Statistics = Revue canadienne de statistique, 31(4):415–435, December 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Larocque:2003:AIM

[1207] Denis Larocque. An affine-invariant multivariate sign test for cluster correlated data. Canadian Journal of Statistics = Revue canadienne de statistique, 31(4):437–455, December 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Fang:2003:EDC

[1208] Zhide Fang. Extrapolation designs with constraints. Canadian Journal of Statistics = Revue canadienne de statistique, 31(4):457–468, December 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Petris:2003:GAT

[1209] Giovanni Petris and Luca Tardella. A geometric approach to transdimensional Markov chain Monte Carlo. Canadian Journal of Statistics = Revue canadienne de statistique, 31(4):469–482, December 2003. ISSN 0319-5724 (print), 1708-945X (electronic).

Guedon:2003:CNE

[1210] Yann Guédon and Christiane Cocozza-Thivent. Correction: “Nonparametric estimation of renewal processes from count data” [Canad. J. Statist. 31 (2003), no. 2, 191–223; Cno. 2016227] by Y. Guédon and C. Cocozza-Thivent. Canadian Journal of Statistics = Revue canadienne de statistique, 31(4):483, December 2003. ISSN 0319-5724 (print), 1708-945X (electronic). See [1193].

Wines:2004:E

[1211] Douglas P. Wines. Editorial. Canadian Journal of Statistics = Revue canadienne de statistique, 32(1):1–2, March 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Kolassa:2004:AMC

[1212] John E. Kolassa. Approximate multivariate conditional inference using the adjusted profile likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 32(1):5–14, March 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:2004:CIM

[1213] Changbao Wu. Combining information from multiple surveys through the empirical likelihood method. Canadian Journal of Statistics = Revue canadienne de statistique, 32(1):15–26, March 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:2004:NME

[1214] Lang Wu. Nonlinear mixed-effect models with nonignorably missing covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 32(1):27–37, March 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Sutradhar:2004:AML

[1215] Brajendra C. Sutradhar and Patrick J. Farrell. Analyzing multivariate longitudinal binary data: a generalized estimating equations approach. Canadian Journal of Statistics = Revue canadienne de statistique, 32(1): 39–55, March 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Ai:2004:TOB

[1216] Mingyao Ai and Runchu Zhang. Theory of optimal blocking of nonregular factorial designs. Canadian Journal of Statistics = Revue canadienne de statistique, 32(1):57–72, March 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Quintana:2004:OSR

[1217] Fernando A. Quintana and Peter Müller. Optimal sampling for repeated binary measurements. Canadian Journal of Statistics = Revue canadienne de statistique, 32(1):73–84, March 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Meyer:2004:CML

[1218] Mary C. Meyer and Michael Woodroofe. Consistent maximum likelihood estimation of a unimodal density using shape restrictions. Canadian Journal of Statistics = Revue canadienne de statistique, 32(1):85–100, March 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Bunea:2004:TSM

[1219] Florentina Bunea and Marten H. Wegkamp. Two-stage model selection procedures in partially linear regression. Canadian Journal of Statistics = Revue canadienne de statistique, 32(2):105–118, June 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Yuan:2004:ANP

[1220] Ao Yuan and Bertrand Clarke. Asymptotic normality of the posterior given a statistic. Canadian Journal of Statistics = Revue canadienne de statistique, 32(2):119–137, June 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Ghosh:2004:UWL

[1221] Malay Ghosh, James V. Zidek, Tapabrata Maiti, and Rick White. The use of the weighted likelihood in the natural exponential families with quadratic variance. Canadian Journal of Statistics = Revue canadienne de statistique, 32(2):139–157, June 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2004:TEM

[1222] Xinsheng Liu and Jinde Wang. Testing the equality of multinomial populations ordered by increasing convexity under the alternative. Canadian Journal of Statistics = Revue canadienne de statistique, 32(2):159–168, June 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Cantoni:2004:RAL

[1223] Eva Cantoni. A robust approach to longitudinal data analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 32(2):169–180, June 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Butucea:2004:DSD

[1224] Cristina Butucea. Deconvolution of supersmooth densities with smooth noise. Canadian Journal of Statistics = Revue canadienne de statistique, 32(2):181–192, June 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Perlman:2004:RRO

[1225] Michael D. Perlman and Sanjay Chaudhuri. The role of reversals in orderrestricted inference. Canadian Journal of Statistics = Revue canadienne de statistique, 32(2):193–198, June 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Cohen:2004:DSI

[1226] Arthur Cohen and Harold B. Sackrowitz. A discussion of some inference issues in order restricted models. Canadian Journal of Statistics = Revue canadienne de statistique, 32(2):199–205, June 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Gentleman:2004:SPS

[1227] Robert Gentleman. Some perspectives on statistical computing. Canadian Journal of Statistics = Revue canadienne de statistique, 32(3):209– 226, September 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Gill:2004:MLE

[1228] Ryan Gill. Maximum likelihood estimation in generalized broken-line regression. Canadian Journal of Statistics = Revue canadienne de statistique, 32(3):227–238, September 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Heritier:2004:RBR

[1229] Stéphane Heritier and Elvezio Ronchetti. Robust binary regression with continuous outcomes. Canadian Journal of Statistics = Revue canadienne de statistique, 32(3):239–249, September 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Altman:2004:SMR

[1230] Naomi S. Altman and Julio C. Villarreal. Self-modelling regression for longitudinal data with time-invariant covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 32(3):251–268, September 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Wong:2004:IBP

[1231] Hubert Wong and Bertrand Clarke. Improvement over Bayes prediction in small samples in the presence of model uncertainty. Canadian Journal of Statistics = Revue canadienne de statistique, 32(3):269–283, September 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Swartz:2004:BIM

[1232] Tim B. Swartz, Yoel Haitovsky, Albert Vexler, and Tae Y. Yang. Bayesian identifiability and misclassification in multinomial data. Canadian Journal of Statistics = Revue canadienne de statistique, 32(3):285– 302, September 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Mao:2004:ENC

[1233] Chang Xuan Mao and Bruce G. Lindsay. Estimating the number of classes in multiple populations: a geometric analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 32(3):303–314, September 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Oller:2004:ICM

[1234] Ramon Oller, Guadalupe Gómez, and M. Luz Calle. Interval censoring: Model characterizations for the validity of the simplified likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 32 (3):315–326, September 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Lawless:2004:NIC

[1235] Jerald F. Lawless. A note on interval-censored lifetime data and the constant-sum condition of Oller, Gómez & Calle (2004) [“Interval censoring: model characterizations for the validity of the simplified likelihood” [Canad. J. Statist. 32 (2004), no. 3, 315–326; Cno. 2101759] by Ramon Oller, Guadalupe Gómez, and M. Luz Calle]. Canadian Journal of Statistics = Revue canadienne de statistique, 32(3):327–331, September 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

You:2004:BEB

[1236] Jinhong You and Gemai Chen. Block external bootstrap in partially linear models with nonstationary strong mixing error terms. Canadian Journal of Statistics = Revue canadienne de statistique, 32(4):335–346, December 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Gu:2004:MDS

[1237] Chong Gu. Model diagnostics for smoothing spline ANOVA models. Canadian Journal of Statistics = Revue canadienne de statistique, 32 (4):347–358, December 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2004:SEE

[1238] Jinbo Chen and Norman E. Breslow. Semiparametric efficient estimation for the auxiliary outcome problem with the conditional mean model. Canadian Journal of Statistics = Revue canadienne de statistique, 32 (4):359–372, December 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Arbogast:2004:GFM

[1239] Patrick G. Arbogast and Danyu Y. Lin. Goodness-of-fit methods for matched case-control studies. Canadian Journal of Statistics = Revue canadienne de statistique, 32(4):373–386, December 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Christen:2004:BRC

[1240] J. Andrés Christen, Peter Müller, Kyle Wathen, and Judith Wolf. Bayesian randomized clinical trials: a decision-theoretic sequential design. Canadian Journal of Statistics = Revue canadienne de statistique, 32(4):387–402, December 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Nan:2004:EEC

[1241] Bin Nan. Efficient estimation for case-cohort studies. Canadian Journal of Statistics = Revue canadienne de statistique, 32(4):403–419, December 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Liao:2004:PRT

[1242] Chen-Tuo Liao and Feng-Shun Chai. Partially replicated two-level fractional factorial designs. Canadian Journal of Statistics = Revue canadienne de statistique, 32(4):421–438, December 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Meister:2004:EME

[1243] Alexander Meister. On the effect of misspecifying the error density in a deconvolution problem. Canadian Journal of Statistics = Revue canadienne de statistique, 32(4):439–449, December 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Berger:2004:VEM

[1244] Yves G. Berger. Variance estimation for measures of change in probability sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 32(4):451–467, December 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Tan:2004:LEL

[1245] Xianming Tan and Runchu Zhang. Letter to the Editor/lettre à la rédaction: Generalized likelihood-ratio test of the number of components in finite mixture models (Chen 1994, lemma 1): correction [Canad. J. Statist. 22 (1994), no. 3, 387–399; MR1309322] by J. H. Chen. Canadian Journal of Statistics = Revue canadienne de statistique, 32(4):469, December 2004. ISSN 0319-5724 (print), 1708-945X (electronic). See [775].

Wong:2004:DCC

[1246] Heung Wong and W. K. Li. Distribution of the cross-correlations of squared residuals in ARIMA models. Canadian Journal of Statistics = Revue canadienne de statistique, 32(4):470, December 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Wong:2004:EDC

[1247] Heung Wong and W. K. Li. Errata: “Distribution of the crosscorrelations of squared residuals in ARIMA models” [Canad. J. Statist. 24 (1996), no. 4, 489–502; mr1438419]. Canadian Journal of Statistics = Revue canadienne de statistique, 32(4):470, 2004. ISSN 0319-5724 (print), 1708-945X (electronic).

Wiens:2005:ERR

[1248] Douglas P. Wiens. Editor’s report/Rapport du rédacteur en chef. Canadian Journal of Statistics = Revue canadienne de statistique, 33(1):1–2, March 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Yi:2005:MAR

[1249] Grace Y. Yi and Mary E. Thompson. Marginal and association regression models for longitudinal binary data with drop-outs: a likelihoodbased approach. Canadian Journal of Statistics = Revue canadienne de statistique, 33(1):3–20, March 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Bosq:2005:ESD

[1250] Denis Bosq. Estimation suroptimale de la densité par projection. (French) [Suroptimale estimate of the density by projection]. Canadian Journal of Statistics = Revue canadienne de statistique, 33(1):21– 37, March 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Braun:2005:LLD

[1251] John Braun, Thierry Duchesne, and James E. Stafford. Local likelihood density estimation for interval censored data. Canadian Journal of Statistics = Revue canadienne de statistique, 33(1):39–60, March 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2005:RAI

[1252] Zhigang Zhang, Liuquan Sun, Xingqiu Zhao, and Jianguo Sun. Regression analysis of interval-censored failure time data with linear transformation models. Canadian Journal of Statistics = Revue canadienne de statistique, 33(1):61–70, March 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Vandal:2005:CEL

[1253] Alain C. Vandal, Robert Gentleman, and Xuecheng Liu. Constrained estimation and likelihood intervals for censored data. Canadian Journal of Statistics = Revue canadienne de statistique, 33(1):71–83, March 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Sun:2005:SLT

[1254] Jianguo Sun and Liuquan Sun. Semiparametric linear transformation models for current status data. Canadian Journal of Statistics = Revue canadienne de statistique, 33(1):85–96, March 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Carvalho:2005:MNT

[1255] Alexandre X. Carvalho and Martin A. Tanner. Modeling nonlinear time series with local mixtures of generalized linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 33(1):97–113, March 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Alvo:2005:GSS

[1256] Mayer Alvo and Paul Cabilio. General scores statistics on ranks in the analysis of unbalanced designs. Canadian Journal of Statistics = Revue canadienne de statistique, 33(1):115–129, March 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Paul:2005:THS

[1257] Sudhir R. Paul and Xing Jiang. Testing the homogeneity of several twoparameter populations. Canadian Journal of Statistics = Revue canadienne de statistique, 33(1):131–143, March 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2005:AAR

[1258] Anonymous. Acknowledgements: Acknowledgement of referees’ services. remerciements aux membres des jurys. Canadian Journal of Statistics = Revue canadienne de statistique, 33(1):145, March 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2005:FPAa

[1259] Anonymous. Forthcoming papers/articles à paraître. Canadian Journal of Statistics = Revue canadienne de statistique, 33(1):146, March 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Haziza:2005:IDU

[1260] David Haziza and J. N. K. Rao. Inference for domains under imputation for missing survey data. Canadian Journal of Statistics = Revue canadienne de statistique, 33(2):149–161, June 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Goga:2005:RVS

[1261] Camelia Goga. Réduction de la variance dans les sondages en présence d’information auxiliaire: Une approche non paramétrique par splines de régression. (French) [Reduction of variance in polls in the presence auxiliary information: a nonparametric approach by regression splines]. Canadian Journal of Statistics = Revue canadienne de statistique, 33(2): 163–180, June 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Lombardi:2005:EFP

[1262] María-José Lombardí, Wenceslao González-Manteiga, and José-Manuel Prada-Sánchez. Estimation of a finite population distribution function based on a linear model with unknown heteroscedastic errors. Canadian Journal of Statistics = Revue canadienne de statistique, 33(2):181–200, June 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

MacKenzie:2005:UCM

[1263] Todd MacKenzie and Michal Abrahamowicz. Using categorical markers as auxiliary variables in log-rank tests and hazard ratio estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 33(2): 201–219, June 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Dette:2005:BMO

[1264] Holger Dette, Linda M. Haines, and Lorens A. Imhof. Bayesian and maximin optimal designs for heteroscedastic regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 33(2):221–241, June 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Asimit:2005:TOP

[1265] Jennifer Asimit and W. John Braun. Third order point process intensity estimation for reaction time experiment data. Canadian Journal of Statistics = Revue canadienne de statistique, 33(2):243–257, June 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Braun:2005:KSR

[1266] W. John Braun and Li-Shan Huang. Kernel spline regression. Canadian Journal of Statistics = Revue canadienne de statistique, 33(2):259–278, June 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Francisco-Fernandez:2005:SPS

[1267] Mario Francisco-Fernandez and Jean D. Opsomer. Smoothing parameter selection methods for nonparametric regression with spatially correlated errors. Canadian Journal of Statistics = Revue canadienne de statistique, 33(2):279–295, June 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Feng:2005:UVP

[1268] Zeny Z. Feng, Jiahua Chen, and Mary E. Thompson. The universal validity of the possible triangle constraint for affected sib pairs. Canadian Journal of Statistics = Revue canadienne de statistique, 33(2):297–310, June 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Genest:2005:PPI

[1269] Christian Genest. Preface/préface [International Conference on Dependence Modelling: Statistical Theory and Applications in Finance and Insurance (DeMoSTAFI), plus table of contents and photo of participants]. Canadian Journal of Statistics = Revue canadienne de statistique, 33 (3):313–316, September 2005. ISSN 0319-5724 (print), 1708-945X (electronic). Held in Québec City, QC, May 20–22, 2004, Dual English–French text.

Abdous:2005:EBB

[1270] Belkacem Abdous, Anne-Laure Fougères, and Kilani Ghoudi. Extreme behaviour for bivariate elliptical distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 33(3):317–334, September 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhao:2005:CLE

[1271] Yinshan Zhao and Harry Joe. Composite likelihood estimation in multivariate data analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 33(3):335–356, September 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Tsukahara:2005:SEC

[1272] Hideatsu Tsukahara. Semiparametric estimation in copula models. Canadian Journal of Statistics = Revue canadienne de statistique, 33(3):357– 375, September 2005. ISSN 0319-5724 (print), 1708-945X (electronic). See erratum [1553].

Vandenhende:2005:LDE

[1273] François Vandenhende and Philippe Lambert. Local dependence estimation using semiparametric Archimedean copulas. Canadian Journal of Statistics = Revue canadienne de statistique, 33(3):377–388, September 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2005:PLR

[1274] Xiaohong Chen and Yanqin Fan. Pseudo-likelihood ratio tests for semiparametric multivariate copula model selection. Canadian Journal of Statistics = Revue canadienne de statistique, 33(3):389–414, September 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Scaillet:2005:KST

[1275] Olivier Scaillet. A Kolmogorov-Smirnov type test for positive quadrant dependence. Canadian Journal of Statistics = Revue canadienne de statistique, 33(3):415–427, September 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Braekers:2005:CGE

[1276] Roel Braekers and Noël Veraverbeke. A copula-graphic estimator for the conditional survival function under dependent censoring. Canadian Journal of Statistics = Revue canadienne de statistique, 33(3):429–447, September 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Choi:2005:ALR

[1277] Yun-Hee Choi and David E. Matthews. Accelerated life regression modelling of dependent bivariate time-to-event data. Canadian Journal of Statistics = Revue canadienne de statistique, 33(3):449–464, September 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Oakes:2005:PCS

[1278] David Oakes. On the preservation of copula structure under truncation. Canadian Journal of Statistics = Revue canadienne de statistique, 33 (3):465–468, September 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2005:FPAb

[1279] Anonymous. Forthcoming papers/articles à paraître. Canadian Journal of Statistics = Revue canadienne de statistique, 33(3):469, September 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2005:VSRa

[1280] Anonymous. Volume 34 (2006): Subscription rates/frais d’abonnement. Canadian Journal of Statistics = Revue canadienne de statistique, 33(3): 470, September 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Tardif:2005:NPA

[1281] Serge Tardif, François Bellavance, and Constance Van Eeden. A nonparametric procedure for the analysis of balanced crossover designs. Canadian Journal of Statistics = Revue canadienne de statistique, 33(4):471–488, December 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Chacon:2005:CWD

[1282] José E. Chacón and Alberto Rodríguez-Casal. On the L_{1}-consistency of wavelet density estimates. Canadian Journal of Statistics = Revue canadienne de statistique, 33(4):489–496, December 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Karunamuni:2005:GRM

[1283] Rohana J. Karunamuni and Tom Alberts. A generalized reflection method of boundary correction in kernel density estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 33(4):497–509, December 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Bianco:2005:REL

[1284] Ana M. Bianco, Marta Garcia Ben, and Víctor J. Yohai. Robust estimation for linear regression with asymmetric errors. Canadian Journal of Statistics = Revue canadienne de statistique, 33(4):511–528, December 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Gao:2005:NTI

[1285] Xin Gao and Mayer Alvo. A nonparametric test for interaction in twoway layouts. Canadian Journal of Statistics = Revue canadienne de statistique, 33(4):529–543, December 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2005:FND

[1286] Lan Wang and Xiao-Hua Zhou. A fully nonparametric diagnostic test for homogeneity of variances. Canadian Journal of Statistics = Revue canadienne de statistique, 33(4):545–558, December 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Yin:2005:CRM

[1287] Guosheng Yin and Joseph G. Ibrahim. Cure rate models: a unified approach. Canadian Journal of Statistics = Revue canadienne de statistique, 33(4):559–570, December 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Iliopoulos:2005:BEK

[1288] George Iliopoulos, Dimitris Karlis, and Ioannis Ntzoufras. Bayesian estimation in Kibble’s bivariate gamma distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 33(4):571–589, December 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Sun:2005:NNP

[1289] Dongchu Sun and Paul L. Speckman. A note on nonexistence of posterior moments. Canadian Journal of Statistics = Revue canadienne de statistique, 33(4):591–601, December 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2005:VSRb

[1290] Anonymous. Volume 34 (2006): Subscription rates/Frais d’abnonnement. Canadian Journal of Statistics = Revue canadienne de statistique, 33(4): 602, December 2005. ISSN 0319-5724 (print), 1708-945X (electronic).

Wines:2006:ER

[1291] Douglas P. Wines. Editor’s report. Canadian Journal of Statistics = Revue canadienne de statistique, 34(1):1–2, March 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Wiens:2006:ERR

[1292] Douglas P. Wiens. Editor reports: Rapport du rédacteur en chef. Canadian Journal of Statistics = Revue canadienne de statistique, 34(1):2–3, March 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Canty:2006:BDR

[1293] Angelo J. Canty, Anthony C. Davison, David V. Hinkley, and Valérie Ventura. Bootstrap diagnostics and remedies. Canadian Journal of Statistics = Revue canadienne de statistique, 34(1):5–27, March 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Leger:2006:BCR

[1294] Christian Léger and Brenda MacGibbon. On the bootstrap in cube root asymptotics. Canadian Journal of Statistics = Revue canadienne de statistique, 34(1):29–44, March 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Tsao:2006:ELI

[1295] Min Tsao and Changbao Wu. Empirical likelihood inference for a common mean in the presence of heteroscedasticity. Canadian Journal of Statistics = Revue canadienne de statistique, 34(1):45–59, March 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Cao:2006:ELT

[1296] Ricardo Cao and Ingrid Van Keilegom. Empirical likelihood tests for twosample problems via nonparametric density estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 34(1):61–77, March 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

You:2006:BEL

[1297] Jinhong You, Gemai Chen, and Yong Zhou. Block empirical likelihood for longitudinal partially linear regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 34(1):79–96, March 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Lu:2006:CPL

[1298] Xuewen Lu, Gemai Chen, Radhey S. Singh, and Peter X.-K. Song. A class of partially linear single-index survival models. Canadian Journal of Statistics = Revue canadienne de statistique, 34(1):97–112, March 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Evans:2006:OCR

[1299] Michael J. Evans, Irwin Guttman, and Tim Swartz. Optimality and computations for relative surprise inferences. Canadian Journal of Statistics = Revue canadienne de statistique, 34(1):113–129, March 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Bibi:2006:PEP

[1300] Abdelouahab Bibi and Antony Gautier. Propriétés dans L^{2} et estimation des processus purement bilinéaires et strictement superdiagonaux à coefficients périodiques. (French) [Properties in L^{2} and estimation of purely-bilinear and strictly superdiagonal processes with periodic coefficients]. Canadian Journal of Statistics = Revue canadienne de statistique, 34(1):131–148, March 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Gonzalez-Manteiga:2006:GFT

[1301] Wenceslao González-Manteiga and Ana Peréz-González. Goodness-of-fit tests for linear regression models with missing response data. Canadian Journal of Statistics = Revue canadienne de statistique, 34(1):149–170, March 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Kim:2006:IUR

[1302] Jae Kwang Kim and Hyeonah Park. Imputation using response probability. Canadian Journal of Statistics = Revue canadienne de statistique, 34(1):171–182, March 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Rivest:2006:RCR

[1303] Louis-Paul Rivest and Ted Chang. Regression and correlation for 3 × 3 rotation matrices. Canadian Journal of Statistics = Revue canadienne de statistique, 34(2):187–202, June 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Boudreau:2006:SAB

[1304] Christian Boudreau and Jerald F. Lawless. Survival analysis based on the proportional hazards model and survey data. Canadian Journal of Statistics = Revue canadienne de statistique, 34(2):203–216, June 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Gombay:2006:CST

[1305] Edit Gombay and Abdulkadir Hussein. A class of sequential tests for twosample composite hypotheses. Canadian Journal of Statistics = Revue canadienne de statistique, 34(2):217–232, June 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

McLeish:2006:ERP

[1306] Donald L. McLeish and Cyntha A. Struthers. Estimation of regression parameters in missing data problems. Canadian Journal of Statistics = Revue canadienne de statistique, 34(2):233–259, June 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Sinha:2006:RIG

[1307] Sanjoy K. Sinha. Robust inference in generalized linear models for longitudinal data. Canadian Journal of Statistics = Revue canadienne de statistique, 34(2):261–278, June 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2006:ABI

[1308] Xiaogang Wang. Approximating Bayesian inference by weighted likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 34(2):279–298, June 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Puza:2006:IET

[1309] Borek Puza and Terence O’Neill. Interval estimation via tail functions. Canadian Journal of Statistics = Revue canadienne de statistique, 34(2): 299–310, June 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Rohani:2006:BSD

[1310] M. Farid Rohani, Khalil Shafie, and Siamak Noorbaloochi. A Bayesian signal detection procedure for scale-space random fields. Canadian Journal of Statistics = Revue canadienne de statistique, 34(2):311–325, June 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Viana:2006:CAO

[1311] Marios A. G. Viana and Hak-Myung Lee. Correlation analysis of ordered symmetrically dependent observations and their concomitants of order statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 34(2):327–340, June 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Mukherjee:2006:PLE

[1312] Kanchan Mukherjee. Pseudo-likelihood estimation in ARCH models. Canadian Journal of Statistics = Revue canadienne de statistique, 34 (2):341–356, June 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2006:FPAa

[1313] Anonymous. Forthcoming papers/Articles à paraître. Canadian Journal of Statistics = Revue canadienne de statistique, 34(2):357, June 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2006:OACa

[1314] Anonymous. Online access to The Canadian Journal of Statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 34(2):358, June 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:2006:PEL

[1315] Changbao Wu and J. N. K. Rao. Pseudo-empirical likelihood ratio confidence intervals for complex surveys. Canadian Journal of Statistics = Revue canadienne de statistique, 34(3):359–375, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Gustafson:2006:CPD

[1316] Paul Gustafson, Shahadut Hossain, and Ying C. MacNab. Conservative prior distributions for variance parameters in hierarchical models. Canadian Journal of Statistics = Revue canadienne de statistique, 34 (3):377–390, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

You:2006:CLP

[1317] Jinhong You, Yong Zhou, and Gemai Chen. Corrected local polynomial estimation in varying-coefficient models with measurement errors. Canadian Journal of Statistics = Revue canadienne de statistique, 34 (3):391–410, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Ferreira:2006:DMS

[1318] José T. A. S. Ferreira and Mark F. J. Steel. On describing multivariate skewed distributions: a directional approach. Canadian Journal of Statistics = Revue canadienne de statistique, 34(3):411–429, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Comte:2006:PCE

[1319] Fabienne Comte, Yves Rozenholc, and Marie-Luce Taupin. Penalized contrast estimator for adaptive density deconvolution. Canadian Journal of Statistics = Revue canadienne de statistique, 34(3):431–452, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Hill:2006:SOD

[1320] Jonathan B. Hill. Strong orthogonal decompositions and non-linear impulse response functions for infinite-variance processes. Canadian Journal of Statistics = Revue canadienne de statistique, 34(3):453–473, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Loubes:2006:RTD

[1321] Jean-Michel Loubes, Élie Maza, Marc Lavielle, and Luis Rodríguez. Road trafficking description and short term travel time forecasting, with a classification method. Canadian Journal of Statistics = Revue canadienne de statistique, 34(3):475–491, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2006:CSD

[1322] Anonymous. Case study in data analysis: Variables related to codling moth abundance and the efficacy of the Okanagan Sterile Insect Release Program. Canadian Journal of Statistics = Revue canadienne de statistique, 34(3):493, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Vernon:2006:IGA

[1323] Bob Vernon, Howard Thistlewood, Scott Smith, and Todd Kabaluk. Introduction: a GIS application to improve codling moth management in the Okanagan Valley of British Columbia. Canadian Journal of Statistics = Revue canadienne de statistique, 34(3):494–500, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Nathoo:2006:ACM

[1324] Farouk Nathoo, Laurie Ainsworth, Paramjit Gill, and Charmaine B. Dean. Analysis 1: Codling moth incidence in Okanagan orchards. Canadian Journal of Statistics = Revue canadienne de statistique, 34(3):500– 510, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Daigle:2006:AEI

[1325] Gaétan Daigle, Thierry Duchesne, Emmanuelle Reny-Nolin, and Louis-Paul Rivest. Analyse 2: Étude de l’influence de la topographie et des caractéristiques des vergers sur l’efficacité du programme d’épandage d’insectes steriles pour le carpocapse de la pomme (Laspeyresia pomonelle). (French) [Analysis 2: Study of the influence of topography and characteristics of orchards on the effectiveness of the sterile codling moth (Laspeyresia pomonelle) insect spraying program]. Canadian Journal of Statistics = Revue canadienne de statistique, 34(3):511– 521, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Esterby:2006:DAC

[1326] Sylvia R. Esterby, Howard Thistlewood, Bob Vernon, and Scott Smith. Discussion: Analysis of codling moth data from the Okanagan Sterile Insect Release Program. Canadian Journal of Statistics = Revue canadienne de statistique, 34(3):521–530, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2006:FPAb

[1327] Anonymous. Forthcoming papers/articles à paraítre. Canadian Journal of Statistics = Revue canadienne de statistique, 34(3):531, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2006:VSR

[1328] Anonymous. Volume 35 (2007): Subscription rates/frais d’abonnement. Canadian Journal of Statistics = Revue canadienne de statistique, 34(3): 532, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2006:OACb

[1329] Anonymous. Online access to The Canadian Journal of Statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 34(3):533, September 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Dette:2006:SMS

[1330] Holger Dette and Regine Scheder. Strictly monotone and smooth nonparametric regression for two or more variables. Canadian Journal of Statistics = Revue canadienne de statistique, 34(4):535–561, December 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

DaSilva:2006:KSM

[1331] Damião N. Da Silva and Jean D. Opsomer. A kernel smoothing method of adjusting for unit non-response in sample surveys. Canadian Journal of Statistics = Revue canadienne de statistique, 34(4):563–579, December 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Arellano-Valle:2006:UVS

[1332] Reinaldo B. Arellano-Valle, Márcia D. Branco, and Marc G. Genton. A unified view on skewed distributions arising from selections. Canadian Journal of Statistics = Revue canadienne de statistique, 34(4):581–601, December 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Biedermann:2006:SRD

[1333] Stefanie Biedermann, Holger Dette, and Andrey Pepelyshev. Some robust design strategies for percentile estimation in binary response models. Canadian Journal of Statistics = Revue canadienne de statistique, 34 (4):603–622, December 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Fang:2006:SRD

[1334] Zhide Fang. Some robust designs for polynomial regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 34 (4):623–638, December 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Dupuis:2006:RPE

[1335] Debbie J. Dupuis and Maria-Pia Victoria-Feser. A robust prediction error criterion for Pareto modelling of upper tails. Canadian Journal of Statistics = Revue canadienne de statistique, 34(4):639–658, December 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Barber:2006:MMP

[1336] Jarrett J. Barber, Alan E. Gelfand, and John A. Silander, Jr. Modelling map positional error to infer true feature location. Canadian Journal of Statistics = Revue canadienne de statistique, 34(4):659–676, December 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Schaubel:2006:MIM

[1337] Douglas E. Schaubel and Jianwen Cai. Multiple imputation methods for recurrent event data with missing event category. Canadian Journal of Statistics = Revue canadienne de statistique, 34(4):677–692, December 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Berrendero:2006:TMU

[1338] José R. Berrendero, Antonio Cuevas, and Francisco Vázquez-Grande. Testing multivariate uniformity: the distance-to-boundary method. Canadian Journal of Statistics = Revue canadienne de statistique, 34 (4):693–707, December 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Herbei:2006:CRO

[1339] Radu Herbei and Marten H. Wegkamp. Classification with reject option. Canadian Journal of Statistics = Revue canadienne de statistique, 34 (4):709–721, December 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2006:OAB

[1340] Anonymous. Online access to the Canadian Journal of Statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 34(4):731, December 2006. ISSN 0319-5724 (print), 1708-945X (electronic).

Gustafson:2007:E

[1341] Paul Gustafson. Editorial. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1):1–4, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2007:RPE

[1342] Anonymous. Report from the previous Editors. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1):5, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2007:ARS

[1343] Anonymous. Acknowledgement of referees’ services. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1):7, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Assuncao:2007:STC

[1344] Renato Assunção, Andréa Tavares, Thais Correa, and Martin Kulldorff. Space-time cluster identification in point processes. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1):9–25, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Ceyhan:2007:NFR

[1345] Elvan Ceyhan, Carey E. Priebe, and David J. Marchette. A new family of random graphs for testing spatial segregation. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1):27–50, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Labbe:2007:MTU

[1346] Aurélie Labbe and Mary E. Thompson. Multiple testing using the posterior probabilities of directional alternatives, with application to genomic studies. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1):51–68, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Lu:2007:MCG

[1347] Haolan Lu, James S. Hodges, and Bradley P. Carlin. Measuring the complexity of generalized linear hierarchical models. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1):69–87, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Delaigle:2007:NDE

[1348] Aurore Delaigle. Nonparametric density estimation from data with a mixture of Berkson and classical errors. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1):89–104, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Ilk:2007:MTR

[1349] Ozlem Ilk and Michael J. Daniels. Marginalized transition random effect models for multivariate longitudinal binary data. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1):105–123, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Lockhart:2007:CMS

[1350] Richard A. Lockhart, John J. Spinelli, and Michael A. Stephens. Cramérvon Mises statistics for discrete distributions with unknown parameters. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1): 125–133, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Fong:2007:MVA

[1351] P. W. Fong, W. K. Li, C. W. Yau, and C. S. Wong. On a mixture vector autoregressive model. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1):135–150, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Park:2007:CTS

[1352] Jung Wook Park, Marc G. Genton, and Sujit K. Ghosh. Censored time series analysis with autoregressive moving average models. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1):151–168, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Saidi:2007:CTN

[1353] Abdessamad Saidi. Consistent testing for non-correlation of two cointegrated ARMA time series. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1):169–188, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2007:M

[1354] Anonymous. Masthead. Canadian Journal of Statistics = Revue canadienne de statistique, 35(1):fmi, March 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Duchesne:2007:CTS

[1355] Pierre Duchesne. On consistent testing for serial correlation in seasonal time series models. Canadian Journal of Statistics = Revue canadienne de statistique, 35(2):193–213, June 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Nevalainen:2007:MSM

[1356] Jaakko Nevalainen, Denis Larocque, and Hannu Oja. On the multivariate spatial median for clustered data. Canadian Journal of Statistics = Revue canadienne de statistique, 35(2):215–231, June 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2007:UAE

[1357] Liqun Wang. A unified approach to estimation of nonlinear mixed effects and Berkson measurement error models. Canadian Journal of Statistics = Revue canadienne de statistique, 35(2):233–248, June 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Pardo-Fernandez:2007:GFT

[1358] Juan Carlos Pardo-Fernández, Ingrid Van Keilegom, and Wenceslao González-Manteiga. Goodness-of-fit tests for parametric models in censored regression. Canadian Journal of Statistics = Revue canadienne de statistique, 35(2):249–264, June 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2007:NEC

[1359] Song Xi Chen and Tzee-Ming Huang. Nonparametric estimation of copula functions for dependence modelling. Canadian Journal of Statistics = Revue canadienne de statistique, 35(2):265–282, June 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

DeOliveira:2007:OBA

[1360] Victor De Oliveira. Objective Bayesian analysis of spatial data with measurement error. Canadian Journal of Statistics = Revue canadienne de statistique, 35(2):283–301, June 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Garcia-Donato:2007:OPH

[1361] Gonzalo García-Donato and Dongchu Sun. Objective priors for hypothesis testing in one-way random effects models. Canadian Journal of Statistics = Revue canadienne de statistique, 35(2):303–320, June 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2007:LLT

[1362] Hui Li, Guosheng Yin, and Yong Zhou. Local likelihood with timevarying additive hazards model. Canadian Journal of Statistics = Revue canadienne de statistique, 35(2):321–337, June 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2007:FPA

[1363] Anonymous. Forthcoming papers/articles à paraître. Canadian Journal of Statistics = Revue canadienne de statistique, 35(2):338, June 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2007:COA

[1364] Anonymous. Complete online access to the Canadian Journal of Statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 35(2):339, June 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2007:VSR

[1365] Anonymous. Volume 35 (2007): Subscription rates/Frais d’abonnement. Canadian Journal of Statistics = Revue canadienne de statistique, 35(2): 340, June 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Andrews:2007:RLI

[1366] David F. Andrews. Robust likelihood inference for public policy. Canadian Journal of Statistics = Revue canadienne de statistique, 35(3):341– 350, September 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Feng:2007:APL

[1367] Zeny Z. Feng, Jiahua Chen, and Mary E. Thompson. Asymptotic properties of likelihood ratio test statistics in affected-sib-pair analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 35(3):351– 364, September 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Shin:2007:ODC

[1368] Hwashin H. Shin, Glen Takahara, and Duncan J. Murdoch. Optimal designs for calibration of orientations. Canadian Journal of Statistics = Revue canadienne de statistique, 35(3):365–380, September 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Jiang:2007:GLR

[1369] Jiancheng Jiang, Haibo Zhou, Xuejun Jiang, and Jianan Peng. Generalized likelihood ratio tests for the structure of semiparametric additive models. Canadian Journal of Statistics = Revue canadienne de statistique, 35(3):381–398, September 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2007:NEP

[1370] Runze Li and Lei Nie. A new estimation procedure for a partially nonlinear model via a mixed-effects approach. Canadian Journal of Statistics = Revue canadienne de statistique, 35(3):399–411, September 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Munk:2007:TNT

[1371] Axel Munk, Matthias Mielke, Guido Skipka, and Gudrun Freitag. Testing noninferiority in three-armed clinical trials based on likelihood ratio statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 35(3):413–431, September 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Jung:2007:TPT

[1372] Inkyung Jung and Martin Kulldorff. Theoretical properties of tests for spatial clustering of count data. Canadian Journal of Statistics = Revue canadienne de statistique, 35(3):433–446, September 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Lu:2007:SLR

[1373] QiQi Lu and Robert B. Lund. Simple linear regression with multiple level shifts. Canadian Journal of Statistics = Revue canadienne de statistique, 35(3):447–458, September 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Jorgensen:2007:SSS

[1374] Bent Jørgensen and Peter X.-K. Song. Stationary state space models for longitudinal data. Canadian Journal of Statistics = Revue canadienne de statistique, 35(4):461–483, December 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Buzkova:2007:LDA

[1375] Petra Bůžková and Thomas Lumley. Longitudinal data analysis for generalized linear models with follow-up dependent on outcome-related variables. Canadian Journal of Statistics = Revue canadienne de statistique, 35(4):485–500, December 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Kim:2007:NWA

[1376] Jae Kwang Kim and Jay J. Kim. Nonresponse weighting adjustment using estimated response probability. Canadian Journal of Statistics = Revue canadienne de statistique, 35(4):501–514, December 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

You:2007:ISP

[1377] Jinhong You and Gemai Chen. On inference for a semiparametric partially linear regression model with serially correlated errors. Canadian Journal of Statistics = Revue canadienne de statistique, 35(4):515–531, December 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

deLeon:2007:GMD

[1378] Alexander R. de Leon and K. C. Carrière. General mixed-data model: Extension of general location and grouped continuous models. Canadian Journal of Statistics = Revue canadienne de statistique, 35(4):533–548, December 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Carota:2007:FPD

[1379] Cinzia Carota. A family of power-divergence diagnostics for goodness-of-fit. Canadian Journal of Statistics = Revue canadienne de statistique, 35 (4):549–561, December 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Huang:2007:ELB

[1380] Chiung-Yu Huang, Jing Qin, and Fei Zou. Empirical likelihood-based inference for genetic mixture models. Canadian Journal of Statistics = Revue canadienne de statistique, 35(4):563–574, December 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Tong:2007:SRA

[1381] Xingwei Tong, Chao Zhu, and Jianguo Sun. Semiparametric regression analysis of two-sample current status data, with applications to tumorigenicity experiments. Canadian Journal of Statistics = Revue canadienne de statistique, 35(4):575–584, December 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Frey:2007:DFS

[1382] Jesse Frey. Distribution-free statistical intervals via ranked-set sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 35 (4):585–596, December 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Lian:2007:NFM

[1383] Heng Lian. Nonlinear functional models for functional responses in reproducing kernel Hilbert spaces. Canadian Journal of Statistics = Revue canadienne de statistique, 35(4):597–606, December 2007. ISSN 0319-5724 (print), 1708-945X (electronic).

Gustafson:2008:ER

[1384] Paul Gustafson. Editor’s report. Canadian Journal of Statistics = Revue canadienne de statistique, 36(1):1–2, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Rivest:2008:CRM

[1385] Louis-Paul Rivest. Capture-recapture models: Introduction. Canadian Journal of Statistics = Revue canadienne de statistique, 36(1):3–4, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Muthukumarana:2008:BAM

[1386] Saman Muthukumarana, Carl J. Schwarz, and Tim B. Swartz. Bayesian analysis of mark-recapture data with travel time-dependent survival probabilities. Canadian Journal of Statistics = Revue canadienne de statistique, 36(1):5–21, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Gimenez:2008:DTB

[1387] Olivier Gimenez. Discussion: Towards a Bayesian analysis template? Canadian Journal of Statistics = Revue canadienne de statistique, 36 (1):21–28, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Pradel:2008:ESS

[1388] Roger Pradel, Lory Maurin-Bernier, Olivier Gimenez, Meritxell Genovart, Rémi Choquet, and Daniel Oro. Estimation of sex-specific survival with uncertainty in sex assessment. Canadian Journal of Statistics = Revue canadienne de statistique, 36(1):29–42, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Choquet:2008:AGM

[1389] Rémi Choquet. Automatic generation of multistate capture-recapture models. Canadian Journal of Statistics = Revue canadienne de statistique, 36(1):43–57, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Gauthier:2008:ABR

[1390] Gilles Gauthier and Jean-Dominique Lebreton. Analysis of band-recovery data in a multistate capture-recapture framework. Canadian Journal of Statistics = Revue canadienne de statistique, 36(1):59–73, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Rivest:2008:WTE

[1391] Louis-Paul Rivest. Why a time effect often has a limited impact on capture-recapture estimates in closed populations. Canadian Journal of Statistics = Revue canadienne de statistique, 36(1):75–84, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Veran:2008:PIM

[1392] Sophie Véran and Jean-Dominique Lebreton. The potential of integrated modelling in conservation biology: a case study of the black-footed albatross (Phoebastria nigripes). Canadian Journal of Statistics = Revue canadienne de statistique, 36(1):85–98, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Mardia:2008:MMD

[1393] Kanti V. Mardia, Gareth Hughes, Charles C. Taylor, and Harshinder Singh. A multivariate von Mises distribution with applications to bioinformatics. Canadian Journal of Statistics = Revue canadienne de statistique, 36(1):99–109, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Pan:2008:PHM

[1394] Qing Pan and Douglas E. Schaubel. Proportional hazards models based on biased samples and estimated selection probabilities. Canadian Journal of Statistics = Revue canadienne de statistique, 36(1):111–127, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2008:THM

[1395] Jiahua Chen, Pengfei Li, and Yuejiao Fu. Testing homogeneity in a mixture of von Mises distributions with a structural parameter. Canadian Journal of Statistics = Revue canadienne de statistique, 36(1):129–142, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Alvo:2008:NTH

[1396] Mayer Alvo. Nonparametric tests of hypotheses for umbrella alternatives. Canadian Journal of Statistics = Revue canadienne de statistique, 36(1): 143–156, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Croux:2008:LDU

[1397] Christophe Croux, Gentiane Haesbroeck, and Kristel Joossens. Logistic discrimination using robust estimators: an influence function approach. Canadian Journal of Statistics = Revue canadienne de statistique, 36(1): 157–174, March 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2008:RSS

[1398] Haiying Chen, Elizabeth A. Stasny, and Douglas A. Wolfe. Ranked set sampling for ordered categorical variables. Canadian Journal of Statistics = Revue canadienne de statistique, 36(2):179–191, June 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Kedem:2008:FMR

[1399] Benjamin Kedem, Guanhua Lu, Rong Wei, and Paul D. Williams. Forecasting mortality rates via density ratio modeling. Canadian Journal of Statistics = Revue canadienne de statistique, 36(2):193–206, June 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Kokoszka:2008:TLD

[1400] Piotr Kokoszka, Inga Maslova, Jan Sojka, and Lie Zhu. Testing for lack of dependence in the functional linear model. Canadian Journal of Statistics = Revue canadienne de statistique, 36(2):207–222, June 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Sambucini:2008:LBA

[1401] Valeria Sambucini and Ludovico Piccinato. Likelihood and Bayesian approaches to inference for the stationary point of a quadratic response surface. Canadian Journal of Statistics = Revue canadienne de statistique, 36(2):223–238, June 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Sanso:2008:BST

[1402] Bruno Sansó, Alexandra M. Schmidt, and Aline A. Nobre. Bayesian spatio-temporal models based on discrete convolutions. Canadian Journal of Statistics = Revue canadienne de statistique, 36(2):239–258, June 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Shi:2008:LIM

[1403] Lei Shi and Gemai Chen. Local influence in multilevel models. Canadian Journal of Statistics = Revue canadienne de statistique, 36(2):259–275, June 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Sinha:2008:RMG

[1404] Sanjoy K. Sinha. Robust methods for generalized linear models with nonignorable missing covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 36(2):277–299, June 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Yang:2008:IGP

[1405] Zhenlin Yang, Eden Ka-Ho Wu, and Anthony F. Desmond. Inference for general parametric functions in Box–Cox-type transformation models. Canadian Journal of Statistics = Revue canadienne de statistique, 36 (2):301–319, June 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2008:NTC

[1406] Lan Wang. Nonparametric test for checking lack of fit of the quantile regression model under random censoring. Canadian Journal of Statistics = Revue canadienne de statistique, 36(2):321–336, June 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Field:2008:MBD

[1407] Christopher A. Field. Modeling biological data: Several vignettes. Canadian Journal of Statistics = Revue canadienne de statistique, 36(3):341– 354, September 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Ferreira:2008:BAE

[1408] Marco A. R. Ferreira and Marc A. Suchard. Bayesian analysis of elapsed times in continuous-time Markov chains. Canadian Journal of Statistics = Revue canadienne de statistique, 36(3):355–368, September 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Fils-Villetard:2008:PEP

[1409] Amélie Fils-Villetard, Armelle Guillou, and Johan Segers. Projection estimators of Pickands dependence functions. Canadian Journal of Statistics = Revue canadienne de statistique, 36(3):369–382, September 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Guillotte:2008:BED

[1410] Simon Guillotte and François Perron. A Bayesian estimator for the dependence function of a bivariate extreme-value distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 36(3):383–396, September 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Lopez-Fidalgo:2008:MRS

[1411] Jesús López-Fidalgo, Raul Martín-Martín, and Douglas P. Wiens. Marginally restricted sequential D-optimal designs. Canadian Journal of Statistics = Revue canadienne de statistique, 36(3):397–410, September 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Nelson:2008:PBM

[1412] Kerrie P. Nelson and Don Edwards. On population-based measures of agreement for binary classifications. Canadian Journal of Statistics = Revue canadienne de statistique, 36(3):411–426, September 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Nie:2008:BRN

[1413] Lei Nie, Haitao Chu, and Valeriy R. Korostyshevskiy. Bias reduction for nonparametric correlation coefficients under the bivariate normal copula assumption with known detection limits. Canadian Journal of Statistics = Revue canadienne de statistique, 36(3):427–442, September 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Plante:2008:NAL

[1414] Jean-François Plante. Nonparametric adaptive likelihood weights. Canadian Journal of Statistics = Revue canadienne de statistique, 36(3):443– 461, September 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhao:2008:ELI

[1415] Yichuan Zhao and Hongkun Wang. Empirical likelihood inference for the regression model of mean quality-adjusted lifetime with censored data. Canadian Journal of Statistics = Revue canadienne de statistique, 36 (3):463–478, September 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Bedard:2008:OSM

[1416] Mylène Bédard and Jeffrey S. Rosenthal. Optimal scaling of Metropolis algorithms: Heading toward general target distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 36(4):483–503, December 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Dannemann:2008:TTS

[1417] Jörn Dannemann and Hajo Holzmann. Testing for two states in a hidden Markov model. Canadian Journal of Statistics = Revue canadienne de statistique, 36(4):505–520, December 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Field:2008:BDM

[1418] Christopher A. Field, Zhen Pang, and Alan H. Welsh. Bootstrapping data with multiple levels of variation. Canadian Journal of Statistics = Revue canadienne de statistique, 36(4):521–539, December 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Hu:2008:DAB

[1419] Feifang Hu, Li-Xin Zhang, Siu H. Cheung, and Wai S. Chan. Doubly adaptive biased coin designs with delayed responses. Canadian Journal of Statistics = Revue canadienne de statistique, 36(4):541–559, December 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Kwok:2008:DCA

[1420] Sai Man Simon Kwok and Wai Keung Li. On diagnostic checking of the autoregressive conditional intensity model. Canadian Journal of Statistics = Revue canadienne de statistique, 36(4):561–576, December 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Ozturk:2008:IPR

[1421] Omer Ozturk. Inference in the presence of ranking error in ranked set sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 36(4):577–594, December 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Salibian-Barrera:2008:WQR

[1422] Matías Salibián-Barrera and Ying Wei. Weighted quantile regression with nonelliptically structured covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 36(4):595–611, December 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Staicu:2008:PMP

[1423] Ana-Maria Staicu and Nancy M. Reid. On probability matching priors. Canadian Journal of Statistics = Revue canadienne de statistique, 36 (4):613–622, December 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Willink:2008:SCI

[1424] Robin Willink. Shrinkage confidence intervals for the normal mean: Using a guess for greater efficiency. Canadian Journal of Statistics = Revue canadienne de statistique, 36(4):623–637, December 2008. ISSN 0319-5724 (print), 1708-945X (electronic).

Gustafson:2009:EER

[1425] Paul Gustafson. Editorial: Editor’s report. Canadian Journal of Statistics = Revue canadienne de statistique, 37(1):1–4, March 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Abd-Elfattah:2009:LRP

[1426] Ehab F. Abd-Elfattah and Ronald W. Butler. Log-rank permutation tests for trend: saddlepoint p-values and survival rate confidence intervals. Canadian Journal of Statistics = Revue canadienne de statistique, 37(1):5–16, March 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Bou-Hamad:2009:DTS

[1427] Imad Bou-Hamad, Denis Larocque, Hatem Ben-Ameur, Louise C. Mâsse, Frank Vitaro, and Richard E. Tremblay. Discrete-time survival trees. Canadian Journal of Statistics = Revue canadienne de statistique, 37 (1):17–32, March 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Brunner:2009:ITE

[1428] Jerry Brunner and Peter C. Austin. Inflation of Type I error rate in multiple regression when independent variables are measured with error. Canadian Journal of Statistics = Revue canadienne de statistique, 37(1): 33–46, March 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Frey:2009:EMT

[1429] Jesse Frey. An exact multinomial test for equivalence. Canadian Journal of Statistics = Revue canadienne de statistique, 37(1):47–59, March 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Hanson:2009:SIS

[1430] Timothy Hanson, Wesley Johnson, and Purushottam Laud. Semiparametric inference for survival models with step process covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 37(1):60– 79, March 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Mesfioui:2009:NGF

[1431] Mhamed Mesfioui, Jean-François Quessy, and Marie-Hélène Toupin. On a new goodness-of-fit process for families of copulas. Canadian Journal of Statistics = Revue canadienne de statistique, 37(1):80–101, March 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2009:NES

[1432] Xiao Wang. Nonparametric estimation of the shape function in a Gamma process for degradation data. Canadian Journal of Statistics = Revue canadienne de statistique, 37(1):102–118, March 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2009:NAB

[1433] Chunming Zhang, Yuan Jiang, and Zuofeng Shang. New aspects of Bregman divergence in regression and classification with parametric and nonparametric estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 37(1):119–139, March 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2009:ARS

[1434] Anonymous. Acknowledgement of referees’ services. Remerciements aux membres des jurys. Canadian Journal of Statistics = Revue canadienne de statistique, 37(1):140, March 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Swartz:2009:MSO

[1435] Tim B. Swartz, Paramjit S. Gill, and Saman Muthukumarana. Modelling and simulation for one-day cricket. Canadian Journal of Statistics = Revue canadienne de statistique, 37(2):143–160, June 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Fraser:2009:TEE

[1436] D. A. S. Fraser, A. Wong, and Y. Sun. Three enigmatic examples and inference from likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 37(2):161–181, June 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2009:LAJ

[1437] Baojiang Chen, Grace Y. Yi, and Richard J. Cook. Likelihood analysis of joint marginal and conditional models for longitudinal categorical data. Canadian Journal of Statistics = Revue canadienne de statistique, 37(2): 182–205, June 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Addona:2009:IPR

[1438] Vittorio Addona, Masoud Asgharian, and David B. Wolfson. On the incidence–prevalence relation and length-biased sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 37(2):206–218, June 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Sinha:2009:BTV

[1439] Sanjoy K. Sinha. Bootstrap tests for variance components in generalized linear mixed models. Canadian Journal of Statistics = Revue canadienne de statistique, 37(2):219–234, June 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Peng:2009:PMA

[1440] Liang Peng. A practical method for analysing heavy tailed data. Canadian Journal of Statistics = Revue canadienne de statistique, 37(2):235– 248, June 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Chacon:2009:DDC

[1441] José E. Chacón. Data-driven choice of the smoothing parametrization for kernel density estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 37(2):249–265, June 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2009:RTM

[1442] Peng Zhang, Zhenguo Qiu, Yuejiao Fu, and Peter X.-K. Song. Robust transformation mixed-effects models for longitudinal continuous proportional data. Canadian Journal of Statistics = Revue canadienne de statistique, 37(2):266–281, June 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Zheng:2009:THN

[1443] Xu Zheng. Testing heteroscedasticity in nonlinear and nonparametric regressions. Canadian Journal of Statistics = Revue canadienne de statistique, 37(2):282–300, June 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Sahu:2009:ENC

[1444] Sujit K. Sahu, Dipak K. Dey, and Márcia D. Branco. Erratum: “A new class of multivariate skew distributions with applications to Bayesian regression models” [MR2016224]. Canadian Journal of Statistics = Revue canadienne de statistique, 37(2):301–302, June 2009. ISSN 0319-5724 (print), 1708-945X (electronic). See [1190].

Ivanoff:2009:MAR

[1445] Gail Ivanoff. In memory of André Robert Dabrowski. Canadian Journal of Statistics = Revue canadienne de statistique, 37(3):305–306, September 2009. ISSN 0319-5724 (print), 1708-945X (electronic). Dual English-French text.

Dehling:2009:ADW

[1446] Herold Dehling. André Dabrowski’s work on limit theorems and weak dependence. Canadian Journal of Statistics = Revue canadienne de statistique, 37(3):307–326, September 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Dabrowski:2009:LDM

[1447] André Dabrowski, Jiyeon Lee, and David R. McDonald. Large deviations of multiclass M/G/1 queues. Canadian Journal of Statistics = Revue canadienne de statistique, 37(3):327–346, September 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Dabrowski:2009:SNP

[1448] André Dabrowski, Gail Ivanoff, and Rafał Kulik. Some notes on Poisson limits for empirical point processes. Canadian Journal of Statistics = Revue canadienne de statistique, 37(3):347–360, September 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Gottardo:2009:BRT

[1449] Raphael Gottardo and Adrian Raftery. Bayesian robust transformation and variable selection: a unified approach. Canadian Journal of Statistics = Revue canadienne de statistique, 37(3):361–380, September 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Sinha:2009:RSA

[1450] Sanjoy K. Sinha and J. N. K. Rao. Robust small area estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 37(3):381– 399, September 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Beaumont:2009:VEW

[1451] Jean-François Beaumont and Cynthia Bocci. Variance estimation when donor imputation is used to fill in missing values. Canadian Journal of Statistics = Revue canadienne de statistique, 37(3):400–416, September 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2009:DSP

[1452] Hongmei Zhang. Designing sampling plans to capture rare objects. Canadian Journal of Statistics = Revue canadienne de statistique, 37(3):417– 434, September 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:2009:LSC

[1453] Lang Wu, Wei Liu, and Juxin Liu. A longitudinal study of children’s aggressive behaviours based on multivariate mixed models with incomplete data. Canadian Journal of Statistics = Revue canadienne de statistique, 37(3):435–452, September 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Desmet:2009:LLF

[1454] Lieven Desmet and Irène Gijbels. Local linear fitting and improved estimation near peaks. Canadian Journal of Statistics = Revue canadienne de statistique, 37(3):453–475, September 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Lee:2009:FOB

[1455] Jaechoul Lee and Kyungduk Ko. First-order bias correction for fractionally integrated time series. Canadian Journal of Statistics = Revue canadienne de statistique, 37(3):476–493, September 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Atherton:2009:BOD

[1456] Juli Atherton, Benoit Charbonneau, David B. Wolfson, Lawrence Joseph, Xiaojie Zhou, and Alain C. Vandal. Bayesian optimal design for changepoint problems. Canadian Journal of Statistics = Revue canadienne de statistique, 37(4):495–513, December 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:2009:MHD

[1457] Jingjing Wu and Rohana J. Karunamuni. On minimum Hellinger distance estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 37(4):514–533, December 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Ghorbal:2009:GKR

[1458] Noomen Ben Ghorbal, Christian Genest, and Johanna Nešlehová. On the Ghoudi, Khoudraji, and Rivest test for extreme-value dependence. Canadian Journal of Statistics = Revue canadienne de statistique, 37 (4):534–552, December 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Bilder:2009:MMR

[1459] Christopher R. Bilder and Thomas M. Loughin. Modeling multipleresponse categorical data from complex surveys. Canadian Journal of Statistics = Revue canadienne de statistique, 37(4):553–570, December 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Horvath:2009:TSI

[1460] Lajos Horváth, Piotr Kokoszka, and Matthew Reimherr. Two sample inference in functional linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 37(4):571–591, December 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Sun:2009:EEP

[1461] Jianguo Sun and Junshan Shen. Efficient estimation for the proportional hazards model with competing risks and current status data. Canadian Journal of Statistics = Revue canadienne de statistique, 37(4):592–606, December 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2009:VSS

[1462] Haonan Wang and Jun Zhu. Variable selection in spatial regression via penalized least squares. Canadian Journal of Statistics = Revue canadienne de statistique, 37(4):607–624, December 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2009:IAV

[1463] Rui Wang and Stephen W. Lagakos. Inference after variable selection using restricted permutation methods. Canadian Journal of Statistics = Revue canadienne de statistique, 37(4):625–644, December 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Xi:2009:ECR

[1464] Liqun Xi, Ray Watson, Ji-Ping Wang, and Paul S. F. Yip. Estimation in capture–recapture models when covariates are subject to measurement errors and missing data. Canadian Journal of Statistics = Revue canadienne de statistique, 37(4):645–658, December 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Yin:2009:LSE

[1465] Guosheng Yin and Hui Li. Least squares estimation of varying-coefficient hazard regression with application to breast cancer dose-intensity data. Canadian Journal of Statistics = Revue canadienne de statistique, 37 (4):659–674, December 2009. ISSN 0319-5724 (print), 1708-945X (electronic).

Gustafson:2010:RPE

[1466] Paul Gustafson. Report from the previous editor. Canadian Journal of Statistics = Revue canadienne de statistique, 38(1):1–4, March 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2010:NNE

[1467] Jiahua Chen. Notes from the New Editor. Canadian Journal of Statistics = Revue canadienne de statistique, 38(1):5–7, March 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Hung:2010:EMT

[1468] Hung Hung and Chin-Tsang Chiang. Estimation methods for timedependent AUC models with survival data. Canadian Journal of Statistics = Revue canadienne de statistique, 38(1):8–26, March 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Yao:2010:NCA

[1469] Fang Yao, Radu V. Craiu, and Benjamin Reiser. Nonparametric covariate adjustment for receiver operating characteristic curves. Canadian Journal of Statistics = Revue canadienne de statistique, 38(1):27–46, March 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Bornn:2010:ECA

[1470] Luke Bornn, Arnaud Doucet, and Raphael Gottardo. An efficient computational approach for prior sensitivity analysis and cross-validation. Canadian Journal of Statistics = Revue canadienne de statistique, 38 (1):47–64, March 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Gombay:2010:CDL

[1471] Edit Gombay. Change detection in linear regression with time series errors. Canadian Journal of Statistics = Revue canadienne de statistique, 38(1):65–79, March 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Pycke:2010:STU

[1472] Jean-Renaud Pycke. Some tests for uniformity of circular distributions powerful against multimodal alternatives. Canadian Journal of Statistics = Revue canadienne de statistique, 38(1):80–96, March 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Davidov:2010:TOA

[1473] Ori Davidov and Amir Herman. Testing for order among K populations: theory and examples. Canadian Journal of Statistics = Revue canadienne de statistique, 38(1):97–115, March 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Moghtaderi:2010:UAL

[1474] Azadeh Moghtaderi, Glen Takahara, and David J. Thomson. Unaliasing of aliased line component frequencies. Canadian Journal of Statistics = Revue canadienne de statistique, 38(1):116–135, March 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Yu:2010:ASS

[1475] Qingzhao Yu, Bin Li, Zhide Fang, and Lu Peng. An adaptive sampling scheme guided by BART — with an application to predict processor performance. Canadian Journal of Statistics = Revue canadienne de statistique, 38(1):136–152, March 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

McNicholas:2010:MBC

[1476] Paul D. McNicholas and T. Brendan Murphy. Model-based clustering of longitudinal data. Canadian Journal of Statistics = Revue canadienne de statistique, 38(1):153–168, March 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2010:ARS

[1477] Anonymous. Acknowledgement of referees’ services. Remerciements aux membres des jurys. Canadian Journal of Statistics = Revue canadienne de statistique, 38(1):169, March 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Sutradhar:2010:PSI

[1478] Brajendra Sutradhar. Preface to the special issue on inferences in generalized linear longitudinal mixed models. Canadian Journal of Statistics = Revue canadienne de statistique, 38(2):171–173, June 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Sutradhar:2010:IGL

[1479] Brajendra C. Sutradhar. Inferences in generalized linear longitudinal mixed models. Canadian Journal of Statistics = Revue canadienne de statistique, 38(2):174–196, June 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Alfo:2010:TPR

[1480] Marco Alfò and Antonello Maruotti. Two-part regression models for longitudinal zero-inflated count data. Canadian Journal of Statistics = Revue canadienne de statistique, 38(2):197–216, June 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Sutradhar:2010:MWB

[1481] Brajendra C. Sutradhar and Taslim S. Mallick. Modified weights based generalized quasilikelihood inferences in incomplete longitudinal binary models. Canadian Journal of Statistics = Revue canadienne de statistique, 38(2):217–231, June 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Yi:2010:EFE

[1482] Grace Y. Yi, Richard J. Cook, and Baojiang Chen. Estimating functions for evaluating treatment effects in cluster-randomized longitudinal studies in the presence of drop-out and non-compliance. Canadian Journal of Statistics = Revue canadienne de statistique, 38(2):232–255, June 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Martinez:2010:LFP

[1483] Josue G. Martinez, Faming Liang, Lan Zhou, and Raymond J. Carroll. Longitudinal functional principal component modelling via stochastic approximation Monte Carlo. Canadian Journal of Statistics = Revue canadienne de statistique, 38(2):256–270, June 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Kang:2010:UTV

[1484] Emily L. Kang, Noel Cressie, and Tao Shi. Using temporal variability to improve spatial mapping with application to satellite data. Canadian Journal of Statistics = Revue canadienne de statistique, 38(2):271–289, June 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Sutradhar:2010:QLE

[1485] Brajendra Sutradhar, Alwell J. Oyet, and Veeresh G. Gadag. On quasilikelihood estimation for branching processes with immigration. Canadian Journal of Statistics = Revue canadienne de statistique, 38(2):290– 313, June 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Dai:2010:CNM

[1486] Hongying Dai and Richard Charnigo. Contaminated normal modeling with application to microarray data analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 38(3):315–332, September 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Song:2010:AHR

[1487] Xinyuan Song, Liuquan Sun, Xiaoyun Mu, and Gregg E. Dinse. Additive hazards regression with censoring indicators missing at random.

Canadian Journal of Statistics = Revue canadienne de statistique, 38 (3):333–351, September 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Hu:2010:VEC

[1488] Bo Hu, Jun Shao, and Mari Palta. Variability explained by covariates in linear mixed-effect models for longitudinal data. Canadian Journal of Statistics = Revue canadienne de statistique, 38(3):352–368, September 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Molina:2010:SAE

[1489] Isabel Molina and J. N. K. Rao. Small area estimation of poverty indicators. Canadian Journal of Statistics = Revue canadienne de statistique, 38(3):369–385, September 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Chiu:2010:BCR

[1490] Grace S. Chiu and Richard A. Lockhart. Bent-cable regression with autoregressive noise. Canadian Journal of Statistics = Revue canadienne de statistique, 38(3):386–407, September 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2010:DFT

[1491] Haiyan Wang, Siti Tolos, and Suojin Wang. A distribution free test to detect general dependence between a response variable and a covariate in the presence of heteroscedastic treatment effects. Canadian Journal of Statistics = Revue canadienne de statistique, 38(3):408–433, September 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Huang:2010:ELV

[1492] Zhensheng Huang and Riquan Zhang. Empirical likelihood for the varying-coefficient single-index model. Canadian Journal of Statistics = Revue canadienne de statistique, 38(3):434–452, September 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Molanes-lopez:2010:SEL

[1493] Elisa M. Molanes-lopez, Ricardo Cao, and Ingrid Van Keilegom. Smoothed empirical likelihood confidence intervals for the relative distribution with left-truncated and right-censored data. Canadian Journal of Statistics = Revue canadienne de statistique, 38(3):453–473, September 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Paige:2010:CFL

[1494] Robert Paige and Edward Allen. Closed-form likelihoods for stochastic differential equation growth models. Canadian Journal of Statistics = Revue canadienne de statistique, 38(3):474–487, September 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Lee:2010:SSM

[1495] Terry C. K. Lee, Min Tsao, and Francis W. Zwiers. State-space model for proxy-based millennial reconstruction. Canadian Journal of Statistics = Revue canadienne de statistique, 38(3):488–505, September 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Song:2010:BHM

[1496] Hae-Ryoung Song, Andrew B. Lawson, and Daniela Nitcheva. Bayesian hierarchical models for food frequency assessment. Canadian Journal of Statistics = Revue canadienne de statistique, 38(3):506–516, September 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Duchesne:2010:EAC

[1497] Pierre Duchesne and Simon Lalancette. Erratum: Authors’ corrigenda/ corrections des auteurs on testing for multivariate ARCH effects in vector time series models. Canadian Journal of Statistics = Revue canadienne de statistique, 38(3):517, September 2010. ISSN 0319-5724 (print), 1708-945X (electronic). See [1198].

Khalili:2010:NEF

[1498] Abbas Khalili. New estimation and feature selection methods in mixtureof-experts models. Canadian Journal of Statistics = Revue canadienne de statistique, 38(4):519–539, December 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Carrillo:2010:PGA

[1499] Iván A. Carrillo, Jiahua Chen, and Changbao Wu. The pseudo-GEE approach to the analysis of longitudinal surveys. Canadian Journal of Statistics = Revue canadienne de statistique, 38(4):540–554, December 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Gijbels:2010:PQD

[1500] Irène Gijbels, Marek Omelka, and Dominik Sznajder. Positive quadrant dependence tests for copulas. Canadian Journal of Statistics = Revue canadienne de statistique, 38(4):555–581, December 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2010:CIM

[1501] Hanfeng Chen, Jiahua Chen, and Shun-Yi Chen. Confidence intervals for the mean of a population containing many zero values under unequalprobability sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 38(4):582–597, December 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Torabi:2010:MSE

[1502] Mahmoud Torabi and Jon N. K. Rao. Mean squared error estimators of small area means using survey weights. Canadian Journal of Statistics = Revue canadienne de statistique, 38(4):598–608, December 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Tan:2010:NLD

[1503] Zhiqiang Tan. Nonparametric likelihood and doubly robust estimating equations for marginal and nested structural models. Canadian Journal of Statistics = Revue canadienne de statistique, 38(4):609–632, December 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Duan:2010:OES

[1504] Xiaogang Duan, Jing Qin, and Qihua Wang. Optimal estimation in surrogate outcome regression problems. Canadian Journal of Statistics = Revue canadienne de statistique, 38(4):633–646, December 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Frey:2010:TEV

[1505] Jesse Frey. Testing for equivalence of variances using Hartley’s ratio. Canadian Journal of Statistics = Revue canadienne de statistique, 38 (4):647–664, December 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Ma:2010:SMR

[1506] Yanyuan Ma and Guosheng Yin. Semiparametric median residual life model and inference. Canadian Journal of Statistics = Revue canadienne de statistique, 38(4):665–679, December 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Sinha:2010:ESA

[1507] Samiran Sinha. An estimated-score approach for dealing with missing covariate data in matched case–control studies. Canadian Journal of Statistics = Revue canadienne de statistique, 38(4):680–697, December 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Torabi:2010:STM

[1508] Mahmoud Torabi and Rhonda J. Rosychuk. Spatio-temporal modelling of disease mapping of rates. Canadian Journal of Statistics = Revue canadienne de statistique, 38(4):698–715, December 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Yan:2010:MBL

[1509] Guohua Yan, William J. Welch, and Ruben H. Zamar. Model-based linear clustering. Canadian Journal of Statistics = Revue canadienne de statistique, 38(4):716–737, December 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Gou:2010:SED

[1510] Tingting Gou and Duncan Murdoch. Simulation of extremes of diffusions. Canadian Journal of Statistics = Revue canadienne de statistique, 38(4):738–755, December 2010. ISSN 0319-5724 (print), 1708-945X (electronic).

Yang:2011:ELB

[1511] Baoying Yang, Gengsheng Qin, and Jing Qin. Empirical likelihood-based inferences for a low income proportion. Canadian Journal of Statistics = Revue canadienne de statistique, 39(1):1–16, March 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Qin:2011:NPI

[1512] Gengsheng Qin, Xiaoping Jin, and Xiao-Hua Zhou. Non-parametric interval estimation for the partial area under the ROC curve. Canadian Journal of Statistics = Revue canadienne de statistique, 39(1):17–33, March 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Yi:2011:RPL

[1513] Grace Y. Yi, Leilei Zeng, and Richard J. Cook. A robust pairwise likelihood method for incomplete longitudinal binary data arising in clusters. Canadian Journal of Statistics = Revue canadienne de statistique, 39(1): 34–51, March 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Kalogeropoulos:2011:LBI

[1514] Konstantinos Kalogeropoulos, Petros Dellaportas, and Gareth O. Roberts. Likelihood-based inference for correlated diffusions. Canadian Journal of Statistics = Revue canadienne de statistique, 39(1):52– 72, March 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Wen:2011:CRC

[1515] Chi-Chung Wen, Steve Y. H. Huang, and Yau-Hung Chen. Cox regression for current status data with mismeasured covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 39(1):73–88, March 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2011:TMC

[1516] Jianqiang C. Wang and Mary C. Meyer. Testing the monotonicity or convexity of a function using regression splines. Canadian Journal of Statistics = Revue canadienne de statistique, 39(1):89–107, March 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Song:2011:NTR

[1517] Weixing Song and Juan Du. A note on testing the regression functions via nonparametric smoothing. Canadian Journal of Statistics = Revue canadienne de statistique, 39(1):108–125, March 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Yanagihara:2011:SOB

[1518] Hirokazu Yanagihara, Ken-Ichi Kamo, and Tetsuji Tonda. Second-order bias-corrected AIC in multivariate normal linear models under nonnormality. Canadian Journal of Statistics = Revue canadienne de statistique, 39(1):126–146, March 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2011:DOR

[1519] Li-Chun Zhang and Nina Hagesæther. A domain outlier robust design and smooth estimation approach. Canadian Journal of Statistics = Revue canadienne de statistique, 39(1):147–164, March 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

MacEachern:2011:PSH

[1520] Steven N. MacEachern and Subharup Guha. Parametric and semiparametric hypotheses in the linear model. Canadian Journal of Statistics = Revue canadienne de statistique, 39(1):165–180, March 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Gibbs:2011:CSD

[1521] Alison L. Gibbs, Kevin J. Keen, and Liqun Wang. Case studies in data analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 39(2):181–217, June 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Niu:2011:THM

[1522] Xiaoqing Niu, Pengfei Li, and Peng Zhang. Testing homogeneity in a multivariate mixture model. Canadian Journal of Statistics = Revue canadienne de statistique, 39(2):218–238, June 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Min:2011:BMS

[1523] Aleksey Min and Claudia Czado. Bayesian model selection for D-vine pair-copula constructions. Canadian Journal of Statistics = Revue canadienne de statistique, 39(2):239–258, June 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Marra:2011:ESR

[1524] Giampiero Marra and Rosalba Radice. Estimation of a semiparametric recursive bivariate probit model in the presence of endogeneity. Canadian Journal of Statistics = Revue canadienne de statistique, 39(2):259–279, June 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Charnigo:2011:SCE

[1525] Richard Charnigo and Cidambi Srinivasan. Self-consistent estimation of mean response functions and their derivatives. Canadian Journal of Statistics = Revue canadienne de statistique, 39(2):280–299, June 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Park:2011:RPL

[1526] Seo Young Park and Yufeng Liu. Robust penalized logistic regression with truncated loss functions. Canadian Journal of Statistics = Revue canadienne de statistique, 39(2):300–323, June 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Lin:2011:SCB

[1527] Huazhen Lin, Ling Zhou, Heng Peng, and Xiao-Hua Zhou. Selection and combination of biomarkers using ROC method for disease classification and prediction. Canadian Journal of Statistics = Revue canadienne de statistique, 39(2):324–343, June 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Steiner:2011:PAM

[1528] Stefan H. Steiner, Nathaniel T. Stevens, Ryan Browne, and R. Jock MacKay. Planning and analysis of measurement reliability studies. Canadian Journal of Statistics = Revue canadienne de statistique, 39(2):344– 355, June 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Chenouri:2011:DDB

[1529] Shojaeddin Chenouri, Christopher G. Small, and Thomas J. Farrar. Data depth-based nonparametric scale tests. Canadian Journal of Statistics = Revue canadienne de statistique, 39(2):356–369, June 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2011:RCP

[1530] Minqiang Li, Liang Peng, and Yongcheng Qi. Reduce computation in profile empirical likelihood method. Canadian Journal of Statistics = Revue canadienne de statistique, 39(2):370–384, June 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Cook:2011:SIC

[1531] Richard J. Cook and Grace Y. Yi. A special issue of CJS in honour of Jack Kalbfleisch and Jerry Lawless. Canadian Journal of Statistics = Revue canadienne de statistique, 39(3):385–388, September 2011. ISSN 0319-5724 (print), 1708-945X (electronic). Dual English-French text.

Chen:2011:TET

[1532] Jiahua Chen and Pengfei Li. Tuning the EM-test for finite mixture models. Canadian Journal of Statistics = Revue canadienne de statistique, 39 (3):389–404, September 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Keiding:2011:APC

[1533] Niels Keiding. Age–period–cohort analysis in the 1870s: Diagrams, stereograms, and the basic differential equation. Canadian Journal of Statistics = Revue canadienne de statistique, 39(3):405–420, September 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Kim:2011:MSV

[1534] Daeyoung Kim and Bruce G. Lindsay. Modal simulation and visualization in finite mixture models. Canadian Journal of Statistics = Revue canadienne de statistique, 39(3):421–437, September 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Lee:2011:CIM

[1535] Terry C. K. Lee, Leilei Zeng, Darby J. S. Thompson, and C. B. Dean. Comparison of imputation methods for interval censored time-to-event data in joint modelling of tree growth and mortality. Canadian Journal of Statistics = Revue canadienne de statistique, 39(3):438–457, September 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2011:STM

[1536] Ni Li, Do-Hwan Park, Jianguo Sun, and KyungMann Kim. Semiparametric transformation models for multivariate panel count data with dependent observation process. Canadian Journal of Statistics = Revue canadienne de statistique, 39(3):458–474, September 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

McKeown:2011:CSO

[1537] Karen McKeown and Nicholas P. Jewell. Current status observation of a three-state counting process with application to simultaneous accurate and diluted HIV test data. Canadian Journal of Statistics = Revue canadienne de statistique, 39(3):475–487, September 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Neuhaus:2011:EMR

[1538] John M. Neuhaus and Charles E. McCulloch. The effect of misspecification of random effects distributions in clustered data settings with outcome-dependent sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 39(3):488–497, September 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Prentice:2011:MEM

[1539] Ross L. Prentice and Ying Huang. Measurement error modeling and nutritional epidemiology association analyses. Canadian Journal of Statistics = Revue canadienne de statistique, 39(3):498–509, September 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Qin:2011:OEF

[1540] Jing Qin and Biao Zhang. Optimal estimating functions in incomplete data and length biased sampling data problems. Canadian Journal of Statistics = Revue canadienne de statistique, 39(3):510–518, September 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Scott:2011:FRM

[1541] Alastair J. Scott and Chris J. Wild. Fitting regression models with response-biased samples. Canadian Journal of Statistics = Revue canadienne de statistique, 39(3):519–536, September 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Yang:2011:PIT

[1542] Yang Yang and Vijayan N. Nair. Parametric inference for time-to-failure in multi-state semi-Markov models: a comparison of marginal and process approaches. Canadian Journal of Statistics = Revue canadienne de statistique, 39(3):537–555, September 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2011:RRR

[1543] Zhiguo Li and Bin Nan. Relative risk regression for current status data in case-cohort studies. Canadian Journal of Statistics = Revue canadienne de statistique, 39(4):557–577, December 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Sun:2011:CMM

[1544] Liuquan Sun, Xingqiu Zhao, and Jie Zhou. A class of mixed models for recurrent event data. Canadian Journal of Statistics = Revue canadienne de statistique, 39(4):578–590, December 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Gu:2011:NRC

[1545] Chong Gu and Ping Ma. Nonparametric regression with cross-classified responses. Canadian Journal of Statistics = Revue canadienne de statistique, 39(4):591–609, December 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Bickel:2011:PAM

[1546] David R. Bickel. A predictive approach to measuring the strength of statistical evidence for single and multiple comparisons. Canadian Journal of Statistics = Revue canadienne de statistique, 39(4):610–631, December 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2011:AAN

[1547] Xiaoming Wang and Keumhee C. Carriere. Assessing additivity in nonparametric models — a kernel-based method. Canadian Journal of Statistics = Revue canadienne de statistique, 39(4):632–655, December 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Leng:2011:IVF

[1548] Chenlei Leng and Cheng Yong Tang. Improving variance function estimation in semiparametric longitudinal data analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 39(4):656–670, December 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Yu:2011:TSE

[1549] Jihnhee Yu, Albert Vexler, Seong-Eun Kim, and Alan D. Hutson. Twosample empirical likelihood ratio tests for medians in application to biomarker evaluations. Canadian Journal of Statistics = Revue canadienne de statistique, 39(4):671–689, December 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Zidan:2011:CIZ

[1550] Marwan Zidan, Jung-Chao Wang, and Magdalena Niewiadomska-Bugaj. Comparison of k independent, zero-heavy lognormal distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 39(4):690– 702, December 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Kojadinovic:2011:LST

[1551] Ivan Kojadinovic, Johan Segers, and Jun Yan. Large-sample tests of extreme-value dependence for multivariate copulas. Canadian Journal of Statistics = Revue canadienne de statistique, 39(4):703–720, December 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Rosenthal:2011:WCM

[1552] Jeffrey S. Rosenthal. Was the conservative majority predictable? Canadian Journal of Statistics = Revue canadienne de statistique, 39(4):721– 733, December 2011. ISSN 0319-5724 (print), 1708-945X (electronic).

Tsukahara:2011:ESE

[1553] Hideatsu Tsukahara. Erratum to “Semiparametric estimation in copula models”. Canadian Journal of Statistics = Revue canadienne de statistique, 39(4):734–735, December 2011. ISSN 0319-5724 (print), 1708-945X (electronic). See [1272].

Hajducek:2012:DAL

[1554] D. M. Hajducek and J. F. Lawless. Duration analysis in longitudinal studies with intermittent observation times and losses to followup. Canadian Journal of Statistics = Revue canadienne de statistique, 40(1):1–21, March 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Park:2012:CNM

[1555] Yongseok Park, John D. Kalbfleisch, and Jeremy M. G. Taylor. Constrained nonparametric maximum likelihood estimation of stochastically ordered survivor functions. Canadian Journal of Statistics = Revue canadienne de statistique, 40(1):22–39, March 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Yu:2012:BAM

[1556] Binbing Yu and Ram C. Tiwari. A Bayesian approach to mixture cure models with spatial frailties for population-based cancer relative survival data. Canadian Journal of Statistics = Revue canadienne de statistique, 40(1):40–54, March 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Hosseinkashi:2012:SGP

[1557] Yasaman Hosseinkashi, Shojaeddin Chenouri, Christopher G. Small, and Rob Deardon. A stochastic graph process for epidemic modelling. Canadian Journal of Statistics = Revue canadienne de statistique, 40(1):55– 67, March 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Brechmann:2012:TRV

[1558] E. C. Brechmann, C. Czado, and K. Aas. Truncated regular vines in high dimensions with application to financial data. Canadian Journal of Statistics = Revue canadienne de statistique, 40(1):68–85, March 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Bauer:2012:PCC

[1559] Alexander Bauer, Claudia Czado, and Thomas Klein. Pair-copula constructions for non-Gaussian DAG models. Canadian Journal of Statistics = Revue canadienne de statistique, 40(1):86–109, March 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Peng:2012:AJE

[1560] Liang Peng. Approximate jackknife empirical likelihood method for estimating equations. Canadian Journal of Statistics = Revue canadienne de statistique, 40(1):110–123, March 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Chauvet:2012:FEE

[1561] Guillaume Chauvet and David Haziza. Fully efficient estimation of coefficients of correlation in the presence of imputed survey data. Canadian Journal of Statistics = Revue canadienne de statistique, 40(1):124–149, March 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2012:RAE

[1562] Zilin Wang and Mary E. Thompson. A resampling approach to estimate variance components of multilevel models. Canadian Journal of Statistics = Revue canadienne de statistique, 40(1):150–171, March 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Bindele:2012:BIN

[1563] Huybrechts F. Bindele and Asheber Abebe. Bounded influence nonlinear signed-rank regression. Canadian Journal of Statistics = Revue canadienne de statistique, 40(1):172–189, March 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Meyer:2012:CPS

[1564] Mary C. Meyer. Constrained penalized splines. Canadian Journal of Statistics = Revue canadienne de statistique, 40(1):190–206, March 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Song:2012:NAJ

[1565] Hui Song, Yingwei Peng, and Dongsheng Tu. A new approach for joint modelling of longitudinal measurements and survival times with a cure fraction. Canadian Journal of Statistics = Revue canadienne de statistique, 40(2):207–224, June 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Papageorgiou:2012:RML

[1566] Georgios Papageorgiou. Restricted maximum likelihood estimation of joint mean-covariance models. Canadian Journal of Statistics = Revue canadienne de statistique, 40(2):225–242, June 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Davis:2012:TGL

[1567] Karelyn A. Davis, Chul G. Park, and Sanjoy K. Sinha. Testing for generalized linear mixed models with cluster correlated data under linear inequality constraints. Canadian Journal of Statistics = Revue canadienne de statistique, 40(2):243–258, June 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Haziza:2012:DRP

[1568] David Haziza and Frédéric Picard. Doubly robust point and variance estimation in the presence of imputed survey data. Canadian Journal of Statistics = Revue canadienne de statistique, 40(2):259–281, June 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Ding:2012:RAS

[1569] Jieli Ding, Yanyan Liu, David B. Peden, Steven R. Kleeberger, and Haibo Zhou. Regression analysis for a summed missing data problem under an outcome-dependent sampling scheme. Canadian Journal of Statistics = Revue canadienne de statistique, 40(2):282–303, June 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Kazianka:2012:OBA

[1570] Hannes Kazianka and Jürgen Pilz. Objective Bayesian analysis of spatial data with uncertain nugget and range parameters. Canadian Journal of Statistics = Revue canadienne de statistique, 40(2):304–327, June 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2012:NHB

[1571] Tingting Zhang and Jun S. Liu. Nonparametric hierarchical Bayes analysis of binomial data via Bernstein polynomial priors. Canadian Journal of Statistics = Revue canadienne de statistique, 40(2):328–344, June 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Hirose:2012:VSW

[1572] Kei Hirose and Sadanori Konishi. Variable selection via the weighted group lasso for factor analysis models. Canadian Journal of Statistics = Revue canadienne de statistique, 40(2):345–361, June 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhao:2012:SDN

[1573] Zhibiao Zhao and Weixin Yao. Sequential design for nonparametric inference. Canadian Journal of Statistics = Revue canadienne de statistique, 40(2):362–377, June 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Berrendero:2012:TUC

[1574] José R. Berrendero, Antonio Cuevas, and Beatriz Pateiro-López. Testing uniformity for the case of a planar unknown support. Canadian Journal of Statistics = Revue canadienne de statistique, 40(2):378–395, June 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2012:ARS

[1575] Anonymous. Acknowledgement of referees’ services. Remerciements aux membres des jurys. Canadian Journal of Statistics = Revue canadienne de statistique, 40(2):396–403, June 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Gel:2012:RLM

[1576] Yulia R. Gel and Bei Chen. Robust Lagrange multiplier test for detecting ARCH/GARCH effect using permutation and bootstrap. Canadian Journal of Statistics = Revue canadienne de statistique, 40(3):405–426, September 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Ketterer:2012:TIS

[1577] Florian Ketterer and Hajo Holzmann. Testing for intercept-scale switch in linear autoregression. Canadian Journal of Statistics = Revue canadienne de statistique, 40(3):427–446, September 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Duchesne:2012:TIB

[1578] Pierre Duchesne, Kilani Ghoudi, and Bruno Rémillard. On testing for independence between the innovations of several time series. Canadian Journal of Statistics = Revue canadienne de statistique, 40(3):447–479, September 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Kojadinovic:2012:GFT

[1579] Ivan Kojadinovic and Jun Yan. Goodness-of-fit testing based on a weighted bootstrap: a fast large-sample alternative to the parametric bootstrap. Canadian Journal of Statistics = Revue canadienne de statistique, 40(3):480–500, September 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Oller:2012:GFH

[1580] Ramon Oller and Guadalupe Gómez. A generalized Fleming and Harrington’s class of tests for interval-censored data. Canadian Journal of Statistics = Revue canadienne de statistique, 40(3):501–516, September 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Fok:2012:FMP

[1581] Carlotta Ching Ting Fok, James O. Ramsay, Michal Abrahamowicz, and Paul Fortin. A functional marked point process model for lupus data. Canadian Journal of Statistics = Revue canadienne de statistique, 40 (3):517–529, September 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Yi:2012:LBM

[1582] Grace Y. Yi and Jerald F. Lawless. Likelihood-based and marginal inference methods for recurrent event data with covariate measurement error. Canadian Journal of Statistics = Revue canadienne de statistique, 40 (3):530–549, September 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhu:2012:IAC

[1583] Hongjian Zhu and Feifang Hu. Interim analysis of clinical trials based on urn models. Canadian Journal of Statistics = Revue canadienne de statistique, 40(3):550–568, September 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Guan:2012:IBM

[1584] Zhong Guan, Jing Qin, and Biao Zhang. Information borrowing methods for covariate-adjusted ROC curve. Canadian Journal of Statistics = Revue canadienne de statistique, 40(3):569–587, September 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Jiang:2012:SAE

[1585] Jiming Jiang and Thuan Nguyen. Small area estimation via heteroscedastic nested-error regression. Canadian Journal of Statistics = Revue canadienne de statistique, 40(3):588–603, September 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Kim:2012:ISI

[1586] Jae Kwang Kim and Minki Hong. Imputation for statistical inference with coarse data. Canadian Journal of Statistics = Revue canadienne de statistique, 40(3):604–618, September 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:2012:ESI

[1587] Changbao Wu. Editorial: A special issue of CJS in honour of Mary Thompson. Canadian Journal of Statistics = Revue canadienne de statistique, 40(4):619–622, December 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Bellhouse:2012:EMT

[1588] David Bellhouse. Editorial: Mary Thompson. Canadian Journal of Statistics = Revue canadienne de statistique, 40(4):623–628, December 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Moodie:2012:LEO

[1589] Erica E. M. Moodie, Bibhas Chakraborty, and Michael S. Kramer. Qlearning for estimating optimal dynamic treatment rules from observational data. Canadian Journal of Statistics = Revue canadienne de statistique, 40(4):629–645, December 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2012:EVD

[1590] Mingyuan Zhang and Dylan S. Small. Effect of vitamin A deficiency on respiratory infection: Causal inference for a discretely observed continuous time non-stationary Markov process. Canadian Journal of Statistics = Revue canadienne de statistique, 40(4):646–662, December 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Chipman:2012:SDC

[1591] Hugh Chipman, Pritam Ranjan, and Weiwei Wang. Sequential design for computer experiments with a flexible Bayesian additive model. Canadian Journal of Statistics = Revue canadienne de statistique, 40(4):663–678, December 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Lohr:2012:BDE

[1592] Sharon L. Lohr and J. Michael Brick. Blending domain estimates from two victimization surveys with possible bias. Canadian Journal of Statistics = Revue canadienne de statistique, 40(4):679–696, December 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Skinner:2012:WRA

[1593] Chris Skinner and Ben Mason. Weighting in the regression analysis of survey data with a cross-national application. Canadian Journal of Statistics = Revue canadienne de statistique, 40(4):697–711, December 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Lindsay:2012:FIM

[1594] Bruce G. Lindsay and Weixin Yao. Fisher information matrix: a tool for dimension reduction, projection pursuit, independent component analysis, and more. Canadian Journal of Statistics = Revue canadienne de statistique, 40(4):712–730, December 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Reid:2012:LIC

[1595] N. Reid. Likelihood inference in complex settings. Canadian Journal of Statistics = Revue canadienne de statistique, 40(4):731–744, December 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2012:VSE

[1596] Zilin Li, Sijian Wang, and Xihong Lin. Variable selection and estimation in generalized linear models with the seamless L_{0} penalty. Canadian Journal of Statistics = Revue canadienne de statistique, 40(4):745–769, December 2012. ISSN 0319-5724 (print), 1708-945X (electronic).

Bagchi:2012:ENC

[1597] Parthasarathy Bagchi and Joseph B. Kadane. Erratum: Note of correction: “Laplace approximations to posterior moments and marginal distributions on circles, spheres, and cylinders”. Canadian Journal of Statistics = Revue canadienne de statistique, 40(4):770–771, December 2012. ISSN 0319-5724 (print), 1708-945X (electronic). See [618].

You:2013:SIM

[1598] Jinhong You, Yong Zhou, and Gemai Chen. Statistical inference for multivariate partially linear regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 41(1):1–22, March 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Dicker:2013:PUL

[1599] Lee Dicker and Xihong Lin. Parallelism, uniqueness, and large-sample asymptotics for the Dantzig selector. Canadian Journal of Statistics = Revue canadienne de statistique, 41(1):23–35, March 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Gijbels:2013:PQD

[1600] Irène Gijbels and Dominik Sznajder. Positive quadrant dependence testing and constrained copula estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 41(1):36–64, March 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Quessy:2013:MKT

[1601] Jean-François Quessy, Mériem Saïd, and Anne-Catherine Favre. Multivariate Kendall’s tau for change-point detection in copulas. Canadian Journal of Statistics = Revue canadienne de statistique, 41(1):65–82, March 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Min:2013:MPB

[1602] Xiaoyi Min and Dongchu Sun. A matching prior based on the modified profile likelihood in a generalized Weibull stress-strength model. Canadian Journal of Statistics = Revue canadienne de statistique, 41 (1):83–97, March 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Groeneboom:2013:CML

[1603] Piet Groeneboom, Geurt Jongbloed, and Stefanie Michael. Consistency of maximum likelihood estimators in a large class of deconvolution models. Canadian Journal of Statistics = Revue canadienne de statistique, 41 (1):98–110, March 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Sued:2013:RLE

[1604] Mariela Sued and Victor J. Yohai. Robust location estimation with missing data. Canadian Journal of Statistics = Revue canadienne de statistique, 41(1):111–132, March 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Jiang:2013:EPE

[1605] Wenyu Jiang and Bingshu E. Chen. Estimating prediction error in microarray classification: Modifications of the 0.632+ bootstrap when n < p. Canadian Journal of Statistics = Revue canadienne de statistique, 41(1):133–150, March 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Tan:2013:CSA

[1606] Zhiqiang Tan. A cluster-sample approach for Monte Carlo integration using multiple samplers. Canadian Journal of Statistics = Revue canadienne de statistique, 41(1):151–173, March 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhao:2013:APC

[1607] Hui Zhao, Yang Li, and Jianguo Sun. Analyzing panel count data with a dependent observation process and a terminal event. Canadian Journal of Statistics = Revue canadienne de statistique, 41(1):174–191, March 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2013:TLR

[1608] Zhiming Li, Tianfang Zhang, and Runchu Zhang. Three-level regular designs with general minimum lower-order confounding. Canadian Journal of Statistics = Revue canadienne de statistique, 41(1):192–210, March 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2013:ARS

[1609] Anonymous. Acknowledgement of referees’ services. Remerciements aux membres des jurys. Canadian Journal of Statistics = Revue canadienne de statistique, 41(1):211–215, March 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Maiboroda:2013:GEE

[1610] Rostyslav Maiboroda, Olena Sugakova, and Alexey Doronin. Generalized estimating equations for mixtures with varying concentrations. Canadian Journal of Statistics = Revue canadienne de statistique, 41(2):217–236, June 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhao:2013:ERC

[1611] Lihui Zhao and X. Joan Hu. Estimation with right-censored observations under a semi-Markov model. Canadian Journal of Statistics = Revue canadienne de statistique, 41(2):237–256, June 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Tsao:2013:EEL

[1612] Min Tsao. Extending the empirical likelihood by domain expansion. Canadian Journal of Statistics = Revue canadienne de statistique, 41 (2):257–274, June 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2013:LTM

[1613] Man-Hua Chen, Xingwei Tong, and Liang Zhu. A linear transformation model for multivariate interval-censored failure time data. Canadian Journal of Statistics = Revue canadienne de statistique, 41(2):275–290, June 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Schwarz:2013:ICB

[1614] Maik Schwarz, Geurt Jongbloed, and Ingrid Van Keilegom. On the identifiability of copulas in bivariate competing risks models. Canadian Journal of Statistics = Revue canadienne de statistique, 41(2):291–303, June 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Ozturk:2013:CMO

[1615] Omer Ozturk. Combining multi-observer information in partially rankordered judgment post-stratified and ranked set samples. Canadian Journal of Statistics = Revue canadienne de statistique, 41(2):304–324, June 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Lin:2013:OMF

[1616] Dennis K. J. Lin and Julie Zhou. D-optimal minimax fractional factorial designs. Canadian Journal of Statistics = Revue canadienne de statistique, 41(2):325–340, June 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Qian:2013:LRT

[1617] Meng Qian and Yongzhao Shao. A likelihood ratio test for goodness-of-fit of recessive and dominant models for case-control studies. Canadian Journal of Statistics = Revue canadienne de statistique, 41(2):341–352, June 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Zou:2013:MRS

[1618] Wei Zou and Jiahua Chen. A Markov regime-switching model for crudeoil markets: Comparison of composite likelihood and full likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 41 (2):353–367, June 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2013:VSE

[1619] Haixiang Zhang, Jianguo Sun, and Dehui Wang. Variable selection and estimation for multivariate panel count data via the seamless-L_{0} penalty. Canadian Journal of Statistics = Revue canadienne de statistique, 41(2): 368–385, June 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Owen:2013:SCE

[1620] Art B. Owen. Self-concordance for empirical likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 41(3):387–397, September 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2013:ELC

[1621] Binhuan Wang and Gengsheng Qin. Empirical likelihood confidence regions for the evaluation of continuous-scale diagnostic tests in the presence of verification bias. Canadian Journal of Statistics = Revue canadienne de statistique, 41(3):398–420, September 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Lin:2013:CMB

[1622] Wei Lin. On central matrix based methods in dimension reduction. Canadian Journal of Statistics = Revue canadienne de statistique, 41(3):421– 438, September 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Lakhal-Chaieb:2013:NEC

[1623] Lajmi Lakhal-Chaieb, Belkacem Abdous, and Thierry Duchesne. Nonparametric estimation of the conditional survival function for bivariate failure times. Canadian Journal of Statistics = Revue canadienne de statistique, 41(3):439–452, September 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2013:MOS

[1624] Tao Wang and Lang Wu. Multivariate one-sided tests for nonlinear mixed-effects models. Canadian Journal of Statistics = Revue canadienne de statistique, 41(3):453–465, September 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

AlLabadi:2013:BNG

[1625] Luai Al Labadi and Mahmoud Zarepour. A Bayesian nonparametric goodness of fit test for right censored data based on approximate samples from the beta-Stacy process. Canadian Journal of Statistics = Revue canadienne de statistique, 41(3):466–487, September 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Ren:2013:OBA

[1626] Cuirong Ren, Dongchu Sun, and Sujit K. Sahu. Objective Bayesian analysis of spatial models with separable correlation functions. Canadian Journal of Statistics = Revue canadienne de statistique, 41(3):488–507, September 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Escobar:2013:NRV

[1627] Emilio L. Escobar and Yves G. Berger. A new replicate variance estimator for unequal probability sampling without replacement. Canadian Journal of Statistics = Revue canadienne de statistique, 41(3):508–524, September 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhu:2013:ACT

[1628] Hongjian Zhu, Feifang Hu, and Hongyu Zhao. Adaptive clinical trial designs to detect interaction between treatment and a dichotomous biomarker. Canadian Journal of Statistics = Revue canadienne de statistique, 41(3):525–539, September 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhou:2013:FAE

[1629] Qi Zhou, Narayanaswamy Balakrishnan, and Runchu Zhang. The factor aliased effect number pattern and its application in experimental planning. Canadian Journal of Statistics = Revue canadienne de statistique, 41(3):540–555, September 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Lin:2013:NEM

[1630] Huazhen Lin and Jianxin Pan. Nonparametric estimation of mean and covariance structures for longitudinal data. Canadian Journal of Statistics = Revue canadienne de statistique, 41(4):557–574, December 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhao:2013:RED

[1631] Pu-Ying Zhao, Man-Lai Tang, and Nian-Sheng Tang. Robust estimation of distribution functions and quantiles with non-ignorable missing data. Canadian Journal of Statistics = Revue canadienne de statistique, 41 (4):575–595, December 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Du:2013:SFR

[1632] Yeting Du, Abbas Khalili, Johanna G. Nešlehová, and Russell J. Steele. Simultaneous fixed and random effects selection in finite mixture of linear mixed-effects models. Canadian Journal of Statistics = Revue canadienne de statistique, 41(4):596–616, December 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Chesneau:2013:FNE

[1633] Christophe Chesneau, Fabienne Comte, and Fabien Navarro. Fast nonparametric estimation for convolutions of densities. Canadian Journal of Statistics = Revue canadienne de statistique, 41(4):617–636, December 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Quessy:2013:GFS

[1634] Jean-François Quessy and Tarik Bahraoui. Graphical and formal statistical tools for assessing the symmetry of bivariate copulas. Canadian Journal of Statistics = Revue canadienne de statistique, 41(4):637–656, December 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Bedard:2013:EEL

[1635] Mylène Bédard and Matei Mireuta. On the empirical efficiency of local MCMC algorithms with pools of proposals. Canadian Journal of Statistics = Revue canadienne de statistique, 41(4):657–678, December 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Daemi:2013:TCR

[1636] Maryam Daemi and Douglas P. Wiens. Techniques for the construction of robust regression designs. Canadian Journal of Statistics = Revue canadienne de statistique, 41(4):679–695, December 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Lin:2013:RVE

[1637] C. Devon Lin, Wilson W. Lu, Keith Rust, and Randy R. Sitter. Replication variance estimation in unequal probability sampling without replacement: One-stage and two-stage. Canadian Journal of Statistics = Revue canadienne de statistique, 41(4):696–716, December 2013. ISSN 0319-5724 (print), 1708-945X (electronic).

Xu:2014:NCR

[1638] Jianfeng Xu and Yingwei Peng. Nonparametric cure rate estimation with covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 42(1):1–17, March 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Lin:2014:SLT

[1639] Huazhen Lin, Yi Li, Liang Jiang, and Gang Li. A semiparametric linear transformation model to estimate causal effects for survival data. Canadian Journal of Statistics = Revue canadienne de statistique, 42 (1):18–35, March 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Serban:2014:STM

[1640] Nicoleta Serban and Pengfei Li. A statistical test for mixture detection with application to component identification in multidimensional biomolecular NMR studies. Canadian Journal of Statistics = Revue canadienne de statistique, 42(1):36–60, March 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:2014:ATE

[1641] Yanyan Wu, Hélène Massam, and Augustine Wong. An accurate test for the equality of covariance matrices from decomposable graphical Gaussian models. Canadian Journal of Statistics = Revue canadienne de statistique, 42(1):61–77, March 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Brechmann:2014:HKC

[1642] Eike Christian Brechmann. Hierarchical Kendall copulas: Properties and inference. Canadian Journal of Statistics = Revue canadienne de statistique, 42(1):78–108, March 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Shi:2014:ADA

[1643] Haiwen Shi, Yu Cheng, and Jialiang Li. Assessing diagnostic accuracy improvement for survival or competing-risk censored outcomes. Canadian Journal of Statistics = Revue canadienne de statistique, 42(1):109–125, March 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Rao:2014:RSA

[1644] Jon N. K. Rao, Sanjoy K. Sinha, and Laura Dumitrescu. Robust small area estimation under semi-parametric mixed models. Canadian Journal of Statistics = Revue canadienne de statistique, 42(1):126–141, March 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Mashreghi:2014:BMI

[1645] Zeinab Mashreghi, Christian Léger, and David Haziza. Bootstrap methods for imputed data from regression, ratio and hot-deck imputation.

Canadian Journal of Statistics = Revue canadienne de statistique, 42(1): 142–167, March 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Tsai:2014:SPR

[1646] Shin-Fu Tsai and Chen-Tuo Liao. Selection of partial replication on two-level orthogonal arrays. Canadian Journal of Statistics = Revue canadienne de statistique, 42(1):168–183, March 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Bilodeau:2014:GLM

[1647] Martin Bilodeau. Graphical lassos for meta-elliptical distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 42(2):185– 203, June 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Avery:2014:RBF

[1648] Matthew Avery, Yichao Wu, Hao Helen Zhang, and Jiajia Zhang. RKHSbased functional nonparametric regression for sparse and irregular longitudinal data. Canadian Journal of Statistics = Revue canadienne de statistique, 42(2):204–216, June 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Hua:2014:SBS

[1649] Lei Hua, Ying Zhang, and Wanzhu Tu. A spline-based semiparametric sieve likelihood method for over-dispersed panel count data. Canadian Journal of Statistics = Revue canadienne de statistique, 42(2):217–245, June 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Xiang:2014:MPH

[1650] Sijia Xiang, Weixin Yao, and Jingjing Wu. Minimum profile Hellinger distance estimation for a semiparametric mixture model. Canadian Journal of Statistics = Revue canadienne de statistique, 42(2):246–267, June 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

McIsaac:2014:RDT

[1651] Michael A. McIsaac and Richard J. Cook. Response-dependent two-phase sampling designs for biomarker studies. Canadian Journal of Statistics = Revue canadienne de statistique, 42(2):268–284, June 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Hao:2014:REA

[1652] Meiling Hao, Xinyuan Song, and Liuquan Sun. Reweighting estimators for the additive hazards model with missing covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 42(2):285–307, June 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Abd-Elfattah:2014:SAR

[1653] Ehab Abd-Elfattah and Ronald Butler. Saddlepoint approximations for rank-invariant permutation tests and confidence intervals with intervalcensoring. Canadian Journal of Statistics = Revue canadienne de statistique, 42(2):308–324, June 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Shi:2014:AMR

[1654] Xiaoping Shi, Nancy Reid, and Yuehua Wu. Approximation to the moments of ratios of cumulative sums. Canadian Journal of Statistics = Revue canadienne de statistique, 42(2):325–336, June 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Miyawaki:2014:ETE

[1655] Koji Miyawaki. Estimating a treatment effect under uncertainty with application to a high-speed railway system. Canadian Journal of Statistics = Revue canadienne de statistique, 42(2):337–358, June 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2014:EAR

[1656] Anonymous. Editorial: Acknowledgement of referees’ services. Remerciements aux membres des jurys. Canadian Journal of Statistics = Revue canadienne de statistique, 42(2):359–363, June 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Schaubel:2014:SMS

[1657] Douglas E. Schaubel, Hui Zhang, John D. Kalbfleisch, and Xu Shu. Semiparametric methods for survival analysis of case-control data subject to dependent censoring. Canadian Journal of Statistics = Revue canadienne de statistique, 42(3):365–383, September 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Romdhani:2014:EKT

[1658] Hela Romdhani, Lajmi Lakhal-Chaieb, and Louis-Paul Rivest. An exchangeable Kendall’s tau for clustered data. Canadian Journal of Statistics = Revue canadienne de statistique, 42(3):384–403, September 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Han:2014:LDA

[1659] Peisong Han, Peter X.-K. Song, and Lu Wang. Longitudinal data analysis using the conditional empirical likelihood method. Canadian Journal of Statistics = Revue canadienne de statistique, 42(3):404–422, September 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Park:2014:HMA

[1660] Jin-Hong Park and S. Yaser Samadi. Heteroscedastic modelling via the autoregressive conditional variance subspace. Canadian Journal of Statistics = Revue canadienne de statistique, 42(3):423–435, September 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Xia:2014:BSA

[1661] Michelle Xia and Paul Gustafson. Bayesian sensitivity analyses for hidden sub-populations in weighted sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 42(3):436–450, September 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Frey:2014:EBG

[1662] Jesse Frey and Le Wang. EDF-based goodness-of-fit tests for ranked-set sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 42(3):451–469, September 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Jozani:2014:CDF

[1663] Mohammad Jafari Jozani, Alexandre Leblanc, and Éric Marchand. On continuous distribution functions, minimax and best invariant estimators, and integrated balanced loss functions. Canadian Journal of Statistics = Revue canadienne de statistique, 42(3):470–486, September 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Goegebeur:2014:LMT

[1664] Yuri Goegebeur, Armelle Guillou, and Michael Osmann. A local moment type estimator for the extreme value index in regression with random covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 42(3):487–507, September 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Rosenthal:2014:IS

[1665] Jeffrey S. Rosenthal. Interdisciplinary sojourns. Canadian Journal of Statistics = Revue canadienne de statistique, 42(4):509–524, December 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Pagui:2014:CCL

[1666] Euloge Clovis Kenne Pagui, Alessandra Salvan, and Nicola Sartori. Combined composite likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 42(4):525–543, December 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Ning:2014:RSN

[1667] Yang Ning, Kung-Yee Liang, and Nancy Reid. Reducing the sensitivity to nuisance parameters in pseudo-likelihood functions. Canadian Journal of Statistics = Revue canadienne de statistique, 42(4):544–562, December 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Baigorri:2014:MCO

[1668] Angel Rodolfo Baigorri, Cátia Regina Gonçalves, and Paulo Angelo Alves Resende. Markov chain order estimation based on the chi-square divergence. Canadian Journal of Statistics = Revue canadienne de statistique, 42(4):563–578, December 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Wichitaksorn:2014:GCS

[1669] Nuttanan Wichitaksorn, S. T. Boris Choy, and Richard Gerlach. A generalized class of skew distributions and associated robust quantile regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 42(4):579–596, December 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Fraiman:2014:DTT

[1670] Ricardo Fraiman, Ana Justel, Regina Liu, and Pamela Llop. Detecting trends in time series of functional data: a study of Antarctic climate change. Canadian Journal of Statistics = Revue canadienne de statistique, 42(4):597–609, December 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Fan:2014:SDM

[1671] Ruzong Fan, Bin Zhu, and Yuedong Wang. Stochastic dynamic models and Chebyshev splines. Canadian Journal of Statistics = Revue canadienne de statistique, 42(4):610–634, December 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Choi:2014:SIG

[1672] Sangbum Choi, Xuelin Huang, Janice N. Cormier, and Kjell A. Doksum. A semiparametric inverse-Gaussian model and inference for survival data with a cured proportion. Canadian Journal of Statistics = Revue canadienne de statistique, 42(4):635–649, December 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Haziza:2014:DRI

[1673] David Haziza, Christian-Olivier Nambeu, and Guillaume Chauvet. Doubly robust imputation procedures for finite population means in the presence of a large number of zeros. Canadian Journal of Statistics = Revue canadienne de statistique, 42(4):650–669, December 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Sinha:2014:AIL

[1674] Sanjoy K. Sinha and Abdus Sattar. Analysis of incomplete longitudinal data with informative drop-out and outliers. Canadian Journal of Statistics = Revue canadienne de statistique, 42(4):670–695, December 2014. ISSN 0319-5724 (print), 1708-945X (electronic).

Tan:2015:GPE

[1675] Zhiqiang Tan and Changbao Wu. Generalized pseudo empirical likelihood inferences for complex surveys. Canadian Journal of Statistics = Revue canadienne de statistique, 43(1):1–17, March 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Lunardon:2015:PCS

[1676] Nicola Lunardon. Prepivoting composite score statistics by weighted bootstrap iteration. Canadian Journal of Statistics = Revue canadienne de statistique, 43(1):18–41, March 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2015:RCA

[1677] Lei Wang, Jiahua Chen, and Xiaolong Pu. Resampling calibrated adjusted empirical likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 43(1):42–59, March 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Cote:2015:CBR

[1678] Marie-Pier Côté and Christian Genest. A copula-based risk aggregation model. Canadian Journal of Statistics = Revue canadienne de statistique, 43(1):60–81, March 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Torabi:2015:NPG

[1679] Mahmoud Torabi and Farhad Shokoohi. Non-parametric generalized linear mixed models in small area estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 43(1):82–96, March 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Ding:2015:PLS

[1680] Xiaobo Ding, Xiao-Hua Zhou, and Qihua Wang. A partially linear singleindex transformation model and its nonparametric estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 43(1):97–117, March 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Geraci:2015:ITB

[1681] Marco Geraci and M. C. Jones. Improved transformation-based quantile regression. Canadian Journal of Statistics = Revue canadienne de statistique, 43(1):118–132, March 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Xu:2015:RSS

[1682] Ximing Xu, Eva Cantoni, Joanna Mills Flemming, and Chris Field. Robust state space models for estimating fish stock maturities. Canadian Journal of Statistics = Revue canadienne de statistique, 43(1):133–150, March 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2015:ARS

[1683] Anonymous. Acknowledgement of referees’ services. Remerciements aux membres des jurys. Canadian Journal of Statistics = Revue canadienne de statistique, 43(1):151–155, March 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Nia:2015:SMB

[1684] Vahid Partovi Nia and Anthony C. Davison. A simple model-based approach to variable selection in classification and clustering. Canadian Journal of Statistics = Revue canadienne de statistique, 43(2):157–175, June 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Browne:2015:MGH

[1685] Ryan P. Browne and Paul D. McNicholas. A mixture of generalized hyperbolic distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 43(2):176–198, June 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Petralias:2015:VPB

[1686] Athanassios Petralias and Petros Dellaportas. Volatility prediction based on scheduled macroeconomic announcements. Canadian Journal of Statistics = Revue canadienne de statistique, 43(2):199–223, June 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Kapelner:2015:PMD

[1687] Adam Kapelner and Justin Bleich. Prediction with missing data via Bayesian Additive Regression Trees. Canadian Journal of Statistics = Revue canadienne de statistique, 43(2):224–239, June 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Zhao:2015:MIA

[1688] Jiwei Zhao, Richard J. Cook, and Changbao Wu. Multiple imputation for the analysis of incomplete compound variables. Canadian Journal of Statistics = Revue canadienne de statistique, 43(2):240–264, June 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Torkashvand:2015:PEB

[1689] Elaheh Torkashvand, Mohammad Jafari Jozani, and Mahmoud Torabi. Pseudo-empirical Bayes estimation of small area means based on James– Stein estimation in linear regression models with functional measurement error. Canadian Journal of Statistics = Revue canadienne de statistique, 43(2):265–287, June 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Yue:2015:VSI

[1690] Yu (Ryan) Yue and Ji Meng Loh. Variable selection for inhomogeneous spatial point process models. Canadian Journal of Statistics = Revue canadienne de statistique, 43(2):288–305, June 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Sujica:2015:ELS

[1691] Aleksandar Sujica and Ingrid Van Keilegom. Estimation of location and scale functionals in nonparametric regression under copula dependent censoring. Canadian Journal of Statistics = Revue canadienne de statistique, 43(2):306–335, June 2015. ISSN 0319-5724 (print), 1708-945X (electronic).

Bao:2015:BNM

[1692] Junshu Bao and Timothy E. Hanson. Bayesian nonparametric multivariate ordinal regression. Canadian Journal of Statistics = Revue canadienne de statistique, 43(3):337–357, September 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhou:2015:PHL

[1693] Qian M. Zhou, Peter X.-K. Song, and Mary E. Thompson. Profiling heteroscedasticity in linear regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 43(3):358–377, September 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2015:IBE

[1694] Min Chen, Changbao Wu, and Mary E. Thompson. An imputation based empirical likelihood approach to pretest–posttest studies. Canadian Journal of Statistics = Revue canadienne de statistique, 43(3):378– 402, September 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Gombay:2015:FRA

[1695] Edit Gombay, Abdulkadir A. Hussein, and Stefan H. Steiner. Flexible risk-adjusted surveillance procedures for autocorrelated binary series. Canadian Journal of Statistics = Revue canadienne de statistique, 43(3):403–419, September 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Choi:2015:ESM

[1696] Sangbum Choi, Xuelin Huang, and Janice N. Cormier. Efficient semiparametric mixture inferences on cure rate models for competing risks. Canadian Journal of Statistics = Revue canadienne de statistique, 43 (3):420–435, September 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yu:2015:SIA

[1697] Jichang Yu, Yanyan Liu, Dale P. Sandler, and Haibo Zhou. Statistical inference for the additive hazards model under outcome-dependent sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 43(3):436–453, September 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yan:2015:CPL

[1698] Ying Yan and Grace Y. Yi. A corrected profile likelihood method for survival data with covariate measurement error under the Cox model. Canadian Journal of Statistics = Revue canadienne de statistique, 43 (3):454–480, September 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2015:JPS

[1699] Jiahua Chen. The joy of proofs in statistical research. Canadian Journal of Statistics = Revue canadienne de statistique, 43(4):481–497, December 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yi:2015:VSI

[1700] Grace Y. Yi, Xianming Tan, and Runze Li. Variable selection and inference procedures for marginal analysis of longitudinal data with missing observations and covariate measurement error. Canadian Journal of Statistics = Revue canadienne de statistique, 43(4):498–518, December 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2015:REL

[1701] Kin yat Liu and Xingqiu Zhao. Robust estimation for longitudinal data with informative observation times. Canadian Journal of Statistics = Revue canadienne de statistique, 43(4):519–533, December 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhu:2015:CAR

[1702] Hongjian Zhu. Covariate-adjusted response adaptive designs incorporating covariates with and without treatment interactions. Canadian Journal of Statistics = Revue canadienne de statistique, 43(4):534–553, December 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhai:2015:RMB

[1703] Zhichun Zhai and Douglas P. Wiens. Robust model-based stratification sampling designs. Canadian Journal of Statistics = Revue canadienne de statistique, 43(4):554–577, December 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Mirakhmedov:2015:EET

[1704] Sherzod M. Mirakhmedov, Sreenivasa R. Jammalamadaka, and Magnus Ekström. Edgeworth expansions for two-stage sampling with applications to stratified and cluster sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 43(4):578–599, December 2015. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Charitidou:2015:BTF

[1705] Efstratia Charitidou, Dimitris Fouskakis, and Ioannis Ntzoufras. Bayesian transformation family selection: Moving toward a transformed Gaussian universe. Canadian Journal of Statistics = Revue canadienne de statistique, 43(4):600–623, December 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Hussami:2015:CL

[1706] Nadine Hussami and Robert J. Tibshirani. A component lasso. Canadian Journal of Statistics = Revue canadienne de statistique, 43(4):624–646, December 2015. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2016:IIEa

[1707] Anonymous. Issue information — Editorial Board. Canadian Journal of Statistics = Revue canadienne de statistique, 44(1):1, March 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2016:IIM

[1708] Anonymous. Issue information — masthead. Canadian Journal of Statistics = Revue canadienne de statistique, 44(1):2, March 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Thompson:2016:UBS

[1709] Mary E. Thompson, Lilia L. Ramirez Ramirez, Vyacheslav Lyubchich, and Yulia R. Gel. Using the bootstrap for statistical inference on random graphs. Canadian Journal of Statistics = Revue canadienne de statistique, 44(1):3–24, March 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Jiang:2016:SSP

[1710] Yiwei Jiang and Zehua Chen. A sequential scaled pairwise selection approach to edge detection in nonparanormal graphical models. Canadian Journal of Statistics = Revue canadienne de statistique, 44(1):25– 43, March 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Kurum:2016:SJM

[1711] Esra Kürüm, John Hughes, and Runze Li. A semivarying joint model for longitudinal binary and continuous outcomes. Canadian Journal of Statistics = Revue canadienne de statistique, 44(1):44–57, March 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wei:2016:SML

[1712] Wenhua Wei and Yong Zhou. Semiparametric maximum likelihood estimation for a two-sample density ratio model with right-censored data.

Canadian Journal of Statistics = Revue canadienne de statistique, 44(1): 58–81, March 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2016:SSC

[1713] Jiahua Chen, Pengfei Li, and Yukun Liu. Sample-size calculation for tests of homogeneity. Canadian Journal of Statistics = Revue canadienne de statistique, 44(1):82–101, March 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2016:JEL

[1714] Dongliang Wang and Yichuan Zhao. Jackknife empirical likelihood for comparing two Gini indices. Canadian Journal of Statistics = Revue canadienne de statistique, 44(1):102–119, March 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2016:ARS

[1715] Anonymous. Acknowledgement of referees’ services: Remerciements aux lecteurs critiques. Canadian Journal of Statistics = Revue canadienne de statistique, 44(1):120–124, March 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2016:IIEb

[1716] Anonymous. Issue information — Ed board and masthead. Canadian Journal of Statistics = Revue canadienne de statistique, 44(2):125–126, June 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2016:RPC

[1717] Qing Liu, Gong Tang, Joseph P. Costantino, and Chung-Chou H. Chang. Robust prediction of the cumulative incidence function under non-proportional subdistribution hazards. Canadian Journal of Statistics = Revue canadienne de statistique, 44(2):127–141, June 2016. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Long:2016:KDE

[1718] James P. Long, Noureddine El Karoui, and John A. Rice. Kernel density estimation with Berkson error. Canadian Journal of Statistics = Revue canadienne de statistique, 44(2):142–160, June 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Jin:2016:CTS

[1719] Baisuo Jin, Yuehua Wu, and Xiaoping Shi. Consistent two-stage multiple change-point detection in linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 44(2):161–179, June 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Tan:2016:BIH

[1720] Aixin Tan and Jian Huang. Bayesian inference for high-dimensional linear regression under mnet priors. Canadian Journal of Statistics = Revue canadienne de statistique, 44(2):180–197, June 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Xia:2016:BRM

[1721] Michelle Xia and Paul Gustafson. Bayesian regression models adjusting for unidirectional covariate misclassification. Canadian Journal of Statistics = Revue canadienne de statistique, 44(2):198–218, June 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

DeQueiroz:2016:SEK

[1722] Marina M. De Queiroz, Roger W. C. Silva, and Rosangela H. Loschi. Shannon entropy and Kullback–Leibler divergence in multivariate log fundamental skew-normal and related distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 44(2):219–237, June 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2016:CVNa

[1723] Anonymous. Cover — volume 44, number 2, June 2016. Canadian Journal of Statistics = Revue canadienne de statistique, 44(2):C1, June 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2016:IIEc

[1724] Anonymous. Issue information — Ed board and masthead. Canadian Journal of Statistics = Revue canadienne de statistique, 44(3):239–240, September 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Duchesne:2016:EME

[1725] Pierre Duchesne, Pierre Lafaye De Micheaux, and Joseph Tagne Tatsinkou. Estimating the mean and its effects on Neyman smooth tests of normality for ARMA models. Canadian Journal of Statistics = Revue canadienne de statistique, 44(3):241–270, September 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Ronchetti:2016:STQ

[1726] Elvezio Ronchetti and Radka Sabolová. Saddlepoint tests for quantile regression. Canadian Journal of Statistics = Revue canadienne de statistique, 44(3):271–299, September 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Sharipov:2016:SBB

[1727] Olimjon Sharipov, Johannes Tewes, and Martin Wendler. Sequential block bootstrap in a Hilbert space with application to change point analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 44(3):300–322, September 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Sang:2016:SGC

[1728] Yongli Sang, Xin Dang, and Hailin Sang. Symmetric Gini covariance and correlation. Canadian Journal of Statistics = Revue canadienne de statistique, 44(3):323–342, September 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2016:CSS

[1729] Peng Wang, Jianhui Zhou, and Annie Qu. Correlation structure selection for longitudinal data with diverging cluster size. Canadian Journal of Statistics = Revue canadienne de statistique, 44(3):343–360, September 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2016:FAM

[1730] Ruosha Li and Yu Cheng. Flexible association modelling and prediction with semi-competing risks data. Canadian Journal of Statistics = Revue canadienne de statistique, 44(3):361–374, September 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2016:VSF

[1731] Xu Liu, Xinyuan Song, Shangyu Xie, and Yong Zhou. Variable selection for frailty transformation models with application to diabetic complications. Canadian Journal of Statistics = Revue canadienne de statistique, 44(3):375–394, September 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2016:CVNb

[1732] Anonymous. Cover — volume 44, number 3, September 2016. Canadian Journal of Statistics = Revue canadienne de statistique, 44(3):C1, September 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2016:IIEd

[1733] Anonymous. Issue information — Ed board and masthead. Canadian Journal of Statistics = Revue canadienne de statistique, 44(4):395–396, December 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Rivest:2016:ULS

[1734] Louis-Paul Rivest, François Verret, and Sophie Baillargeon. Unit level small area estimation with copulas. Canadian Journal of Statistics = Revue canadienne de statistique, 44(4):397–415, December 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Hernandez-Stumpfhauser:2016:HBS

[1735] Daniel Hernandez-Stumpfhauser, F. Jay Breidt, and Jean D. Opsomer. Hierarchical Bayesian small area estimation for circular data. Canadian Journal of Statistics = Revue canadienne de statistique, 44(4):416–430, December 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:2016:SED

[1736] Jiwen Wu, Mary C. Meyer, and Jean D. Opsomer. Survey estimation of domain means that respect natural orderings. Canadian Journal of Statistics = Revue canadienne de statistique, 44(4):431–444, December 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2016:TAN

[1737] Yichi Zhang, Ana-Maria Staicu, and Arnab Maity. Testing for additivity in non-parametric regression. Canadian Journal of Statistics = Revue canadienne de statistique, 44(4):445–462, December 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Shepherd:2016:PSR

[1738] Bryan E. Shepherd, Chun Li, and Qi Liu. Probability-scale residuals for continuous, discrete, and censored data. Canadian Journal of Statistics = Revue canadienne de statistique, 44(4):463–479, December 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Pokharel:2016:GPE

[1739] Gyanendra Pokharel and Rob Deardon. Gaussian process emulators for spatial individual-level models of infectious disease. Canadian Journal of Statistics = Revue canadienne de statistique, 44(4):480–501, December 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2016:CVN

[1740] Anonymous. Cover — volume 44, number 4, December 2016. Canadian Journal of Statistics = Revue canadienne de statistique, 44(4):C1, December 2016. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2017:IIa

[1741] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 45(1):1–3, March 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Best:2017:NCC

[1742] Ana F. Best and David B. Wolfson. Nested case-control study designs for left-truncated survival data. Canadian Journal of Statistics = Revue canadienne de statistique, 45(1):4–28, March 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:2017:CRQ

[1743] Yuanshan Wu and Guosheng Yin. Cure rate quantile regression accommodating both finite and infinite survival times. Canadian Journal of Statistics = Revue canadienne de statistique, 45(1):29–43, March 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Mitra:2017:BMC

[1744] Riten Mitra, Peter Müller, and Yuan Ji. Bayesian multiplicity control for multiple graphs. Canadian Journal of Statistics = Revue canadienne de statistique, 45(1):44–61, March 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Purutcuoglu:2017:BAG

[1745] Vilda Purutçuoğlu, Melih Ağraz, and Ernst Wit. Bernstein approximations in glasso-based estimation of biological networks. Canadian Journal of Statistics = Revue canadienne de statistique, 45(1):62–76, March 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yu:2017:NMR

[1746] Chun Yu, Weixin Yao, and Kun Chen. A new method for robust mixture regression. Canadian Journal of Statistics = Revue canadienne de statistique, 45(1):77–94, March 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Bagnato:2017:MLN

[1747] Luca Bagnato, Antonio Punzo, and Maria G. Zoia. The multivariate leptokurtic-normal distribution and its application in model-based clustering. Canadian Journal of Statistics = Revue canadienne de statistique, 45(1):95–119, March 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2017:AAR

[1748] Anonymous. Acknowledgement of referees’ services: Remerciements aux lecteurs critiques. Canadian Journal of Statistics = Revue canadienne de statistique, 45(1):120–123, March 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2017:IIb

[1749] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 45(2):125–127, June 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Reid:2017:PSP

[1750] Stephen Reid, Jonathan Taylor, and Robert Tibshirani. Post-selection point and interval estimation of signal sizes in Gaussian samples. Canadian Journal of Statistics = Revue canadienne de statistique, 45(2):128– 148, June 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Xiao:2017:NAC

[1751] Xiao Xiao, Xiexin Liu, Xiaoling Lu, Xiangyu Chang, and Yufeng Liu. A new algorithm for computation of a regularization solution path for reinforced multicategory support vector machines. Canadian Journal of Statistics = Revue canadienne de statistique, 45(2):149–163, June 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Dutta:2017:DWR

[1752] Subhajit Dutta and Marc G. Genton. Depth-weighted robust multivariate regression with application to sparse data. Canadian Journal of Statistics = Revue canadienne de statistique, 45(2):164–184, June 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Xu:2017:SEM

[1753] Yuhang Xu, Jae Kwang Kim, and Yehua Li. Semiparametric estimation for measurement error models with validation data. Canadian Journal of Statistics = Revue canadienne de statistique, 45(2):185–201, June 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Ning:2017:ETE

[1754] Jing Ning, Chuan Hong, Liang Li, Xuelin Huang, and Yu Shen. Estimating treatment effects in observational studies with both prevalent and incident cohorts. Canadian Journal of Statistics = Revue canadienne de statistique, 45(2):202–219, June 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2017:NSE

[1755] Yixin Wang, Zhefang Zhou, Xiao-Hua Zhou, and Yong Zhou. Nonparametric and semiparametric estimation of quantile residual lifetime for length-biased and right-censored data. Canadian Journal of Statistics = Revue canadienne de statistique, 45(2):220–250, June 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2017:IIc

[1756] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 45(3):251–253, September 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Barcella:2017:RAC

[1757] William Barcella, Maria De Iorio, and Gianluca Baio. Review article: A comparative review of variable selection techniques for covariate dependent Dirichlet process mixture models. Canadian Journal of Statistics = Revue canadienne de statistique, 45(3):254–273, September 2017. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

DeOliveira:2017:BAD

[1758] Victor De Oliveira and Benjamin Kedem. Bayesian analysis of a density ratio model. Canadian Journal of Statistics = Revue canadienne de statistique, 45(3):274–289, September 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Susko:2017:BFB

[1759] Edward Susko. Bayes factor biases for non-nested models and corrections. Canadian Journal of Statistics = Revue canadienne de statistique, 45(3): 290–309, September 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Selvaratnam:2017:EGL

[1760] Selvakkadunko Selvaratnam, Alwell J. Oyet, Yanqing Yi, and Veeresh Gadag. Estimation of a generalized linear mixed model for responseadaptive designs in multi-centre clinical trials. Canadian Journal of Statistics = Revue canadienne de statistique, 45(3):310–325, September 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Frey:2017:TPR

[1761] Jesse Frey and Yimin Zhang. Testing perfect rankings in ranked-set sampling with binary data. Canadian Journal of Statistics = Revue canadienne de statistique, 45(3):326–339, September 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Jing:2017:TEL

[1762] Bing-Yi Jing, Min Tsao, and Wang Zhou. Transforming the empirical likelihood towards better accuracy. Canadian Journal of Statistics = Revue canadienne de statistique, 45(3):340–352, September 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2017:IId

[1763] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 45(4):353–355, December 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Khalili:2017:RSG

[1764] Abbas Khalili, Jiahua Chen, and David A. Stephens. Regularization and selection in Gaussian mixture of autoregressive models. Canadian Journal of Statistics = Revue canadienne de statistique, 45(4):356–374, December 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Schumacher:2017:CRM

[1765] Fernanda L. Schumacher, Victor H. Lachos, and Dipak K. Dey. Censored regression models with autoregressive errors: a likelihood-based perspective. Canadian Journal of Statistics = Revue canadienne de statistique, 45(4):375–392, December 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Morikawa:2017:SML

[1766] Kosuke Morikawa, Jae Kwang Kim, and Yutaka Kano. Semiparametric maximum likelihood estimation with data missing not at random. Canadian Journal of Statistics = Revue canadienne de statistique, 45 (4):393–409, December 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Hu:2017:RAB

[1767] Tao Hu, Qingning Zhou, and Jianguo Sun. Regression analysis of bivariate current status data under the proportional hazards model. Canadian Journal of Statistics = Revue canadienne de statistique, 45(4):410–424, December 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zambom:2017:NHT

[1768] Adriano Z. Zambom and Seonjin Kim. A nonparametric hypothesis test for heteroscedasticity in multiple regression. Canadian Journal of Statistics = Revue canadienne de statistique, 45(4):425–441, December 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

DeSouza:2017:SNR

[1769] Camila P. E. De Souza, Nancy E. Heckman, and Fan Xu. Switching nonparametric regression models for multi-curve data. Canadian Journal of Statistics = Revue canadienne de statistique, 45(4):442–460, December 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Holmquist:2017:TLD

[1770] Björn Holmquist and Peter Gustafsson. A two-level directional model for dependence in circular data. Canadian Journal of Statistics = Revue canadienne de statistique, 45(4):461–478, December 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Kim:2017:SIU

[1771] Jae Kwang Kim, Seunghwan Park, and Youngjo Lee. Statistical inference using generalized linear mixed models under informative cluster sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 45(4):479–497, December 2017. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2018:IIa

[1772] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 46(1):1–3, March 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cjs.11334.

Lockhart:2018:SIB

[1773] Richard Lockhart. Special issue on Big Data and the Statistical Sciences: Guest Editor’s Introduction. Canadian Journal of Statistics = Revue canadienne de statistique, 46(1):4–9, March 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cjs.11350.

Thompson:2018:DDS

[1774] Mary E. Thompson. Dynamic data science and official statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 46(1):10– 23, March 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
cjs.11322.

Bull:2018:SCH

[1775] Shelley B. Bull, Irene L. Andrulis, and Andrew D. Paterson. Statistical challenges in high-dimensional molecular and genetic epidemiology. Canadian Journal of Statistics = Revue canadienne de statistique, 46 (1):24–40, March 2018. CODEN ???? ISSN 0319-5724 (print), 1708-

945X (electronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1002/cjs.11342.

Taylor:2018:PSI

[1776] Jonathan Taylor and Robert Tibshirani. Post-selection inference for ℓ_{1}-penalized likelihood models. Canadian Journal of Statistics = Revue canadienne de statistique, 46(1):41–61, March 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https:/

/onlinelibrary.wiley.com/doi/abs/10.1002/cjs.11313.

Cook:2018:BDP

[1777] R. Dennis Cook and Liliana Forzani. Big data and partial least-squares prediction. Canadian Journal of Statistics = Revue canadienne de statistique, 46(1):62–78, March 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1002/cjs.11316.

Li:2018:LOB

[1778] Bing Li. Linear operator-based statistical analysis: a useful paradigm for big data. Canadian Journal of Statistics = Revue canadienne de statistique, 46(1):79–103, March 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1002/cjs.11329.

Nie:2018:MRA

[1779] Rui Nie, Douglas P. Wiens, and Zhichun Zhai. Minimax robust active learning for approximately specified regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 46(1):104–122, March 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
cjs.11327.

Wang:2018:OUM

[1780] Chun Wang, Ming-Hui Chen, Jing Wu, Jun Yan, Yuping Zhang, and Elizabeth Schifano. Online updating method with new variables for big data streams. Canadian Journal of Statistics = Revue canadienne de statistique, 46(1):123–146, March 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1002/cjs.11330.

Entezari:2018:LIS

[1781] Reihaneh Entezari, Radu V. Craiu, and Jeffrey S. Rosenthal. Likelihood inflating sampling algorithm. Canadian Journal of Statistics = Revue canadienne de statistique, 46(1):147–175, March 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cjs.11343.

Li:2018:ABE

[1782] Qiong Li, Xin Gao, and Hélène Massam. Approximate Bayesian estimation in large coloured graphical Gaussian models. Canadian Journal of Statistics = Revue canadienne de statistique, 46(1):176–203, March 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL

https://onlinelibrary.wiley.com/doi/abs/10.1002/cjs.11341.

Anonymous:2018:ARS

[1783] Anonymous. Acknowledgement of referees’ services: Remerciements aux lecteurs critiques. Canadian Journal of Statistics = Revue canadienne de statistique, 46(1):204–207, March 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1002/cjs.11353.

Anonymous:2018:IIb

[1784] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 46(2):209–211, June 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cjs.11335.

Cai:2018:ELI

[1785] Song Cai and Jiahua Chen. Empirical likelihood inference for multiple censored samples. Canadian Journal of Statistics = Revue canadienne de statistique, 46(2):212–232, June 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1002/cjs.11348.

Wang:2018:ACD

[1786] Qihua Wang and Xuan Wang. Analysis of censored data under heteroscedastic transformation regression models with unknown transformation function. Canadian Journal of Statistics = Revue canadienne de statistique, 46(2):233–245, June 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1002/cjs.11347.

Gijbels:2018:THE

[1787] Irène Gijbels, Mohammed A. Ibrahim, and Anneleen Verhasselt. Testing the heteroscedastic error structure in quantile varying coefficient models. Canadian Journal of Statistics = Revue canadienne de statistique, 46 (2):246–264, June 2018. CODEN ???? ISSN 0319-5724 (print), 1708-

945X (electronic). URL https://onlinelibrary.wiley.com/doi/abs/
10.1002/cjs.11346.

Chun:2018:RPM

[1788] Hyonho Chun, Myung Hee Lee, Sung-Ho Kim, and Jihwan Oh. Robust precision matrix estimation via weighted median regression with regularization. Canadian Journal of Statistics = Revue canadienne de statistique, 46(2):265–278, June 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1002/cjs.11356.

Kabaila:2018:MCP

[1789] Paul Kabaila. On the minimum coverage probability of model averaged tail area confidence intervals. Canadian Journal of Statistics = Revue canadienne de statistique, 46(2):279–297, June 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https://

onlinelibrary.wiley.com/doi/abs/10.1002/cjs.11349.

Plante:2018:CRC

[1790] Jean-François Plante and Jean-Baptiste Débordès. Combining ROC curves using MAMSE weighted distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 46(2):298–315, June 2018.

CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL

https://onlinelibrary.wiley.com/doi/abs/10.1002/cjs.11351.

Kawakubo:2018:CAI

[1791] Yuki Kawakubo, Shonosuke Sugasawa, and Tatsuya Kubokawa. Conditional Akaike information under covariate shift with application to small area estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 46(2):316–335, June 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1002/cjs.11354.

Zhai:2018:LOD

[1792] Yi Zhai and Zhide Fang. Locally optimal designs for binary doseresponse models. Canadian Journal of Statistics = Revue canadienne de statistique, 46(2):336–354, June 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). URL https://onlinelibrary.wiley.

com/doi/abs/10.1002/cjs.11355.

Abdelrazeq:2018:GFT

[1793] Ibrahim Abdelrazeq, B. Gail Ivanoff, and Rafal Kulik. Goodnessof-fit tests for Lévy-driven Ornstein–Uhlenbeck processes. Canadian Journal of Statistics = Revue canadienne de statistique, 46(2):355–376, June 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (elec-

tronic). URL https://onlinelibrary.wiley.com/doi/abs/10.1002/
cjs.11352.

Anonymous:2018:IIc

[1794] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 46(3):377–379, September 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Al-Labadi:2018:PBM

[1795] Luai Al-Labadi and Michael Evans. Prior-based model checking. Canadian Journal of Statistics = Revue canadienne de statistique, 46(3):380– 398, September 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Goh:2018:BME

[1796] Gyuhyeong Goh and Dipak K. Dey. Bayesian MAP estimation using Gaussian and diffused-gamma prior. Canadian Journal of Statistics = Revue canadienne de statistique, 46(3):399–415, September 2018. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhao:2018:VSR

[1797] Hui Zhao, Dayu Sun, Gang Li, and Jianguo Sun. Variable selection for recurrent event data with broken adaptive ridge regression. Canadian Journal of Statistics = Revue canadienne de statistique, 46(3):416–428, September 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Khalili:2018:HTF

[1798] Abbas Khalili and Anand N. Vidyashankar. Hypothesis testing in finite mixture of regressions: Sparsity and model selection uncertainty. Canadian Journal of Statistics = Revue canadienne de statistique, 46 (3):429–457, September 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yi:2018:RAD

[1799] Yanqing Yi and Xuan Li. Response adaptive designs with asymptotic optimality. Canadian Journal of Statistics = Revue canadienne de statistique, 46(3):458–469, September 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Hu:2018:RDE

[1800] Rui Hu. Robust design for the estimation of a threshold probability. Canadian Journal of Statistics = Revue canadienne de statistique, 46 (3):470–481, September 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Berg:2018:BSA

[1801] Emily Berg and Wayne A. Fuller. Benchmarked small area prediction. Canadian Journal of Statistics = Revue canadienne de statistique, 46 (3):482–500, September 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Bindele:2018:RBI

[1802] Huybrechts F. Bindele and Akim Adekpedjou. Rank-based inference with responses missing not at random. Canadian Journal of Statistics = Revue canadienne de statistique, 46(3):501–528, September 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2018:IId

[1803] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 46(4):529–531, December 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Joe:2018:PGD

[1804] Harry Joe. Parsimonious graphical dependence models constructed from vines. Canadian Journal of Statistics = Revue canadienne de statistique, 46(4):532–555, December 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2018:VAC

[1805] Hui Li, Zhiqiang Cao, and Guosheng Yin. Varying-association copula models for multivariate survival data. Canadian Journal of Statistics = Revue canadienne de statistique, 46(4):556–576, December 2018. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chatterjee:2018:NIL

[1806] Kiranmoy Chatterjee and Diganta Mukherjee. A new integrated likelihood for estimating population size in dependent dual-record system. Canadian Journal of Statistics = Revue canadienne de statistique, 46 (4):577–592, December 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Pandolfo:2018:DBD

[1807] Giuseppe Pandolfo, Davy Paindaveine, and Giovanni C. Porzio. Distance-based depths for directional data. Canadian Journal of Statistics = Revue canadienne de statistique, 46(4):593–609, December 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Adekpedjou:2018:CBQ

[1808] Akim Adekpedjou, Gayla R. Olbricht, and Gideon K. D. Zamba. Confidence bands for quantiles as a function of covariates in recurrent event models. Canadian Journal of Statistics = Revue canadienne de statistique, 46(4):610–634, December 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Maitra:2018:AIS

[1809] Trisha Maitra and Sourabh Bhattacharya. On asymptotic inference in stochastic differential equations with time-varying covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 46(4):635–655, December 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Bibbona:2018:EIP

[1810] Enrico Bibbona, Giovanni Pistone, and Mauro Gasparini. The empirical identity process: Asymptotics and applications. Canadian Journal of Statistics = Revue canadienne de statistique, 46(4):656–672, December 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Luo:2018:EPU

[1811] Yu Luo, David A. Stephens, and David L. Buckeridge. Estimating prevalence using indirect information and Bayesian evidence synthesis. Canadian Journal of Statistics = Revue canadienne de statistique, 46(4):673– 689, December 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Saleh:2018:RTA

[1812] A. K. Md. Ehsanes Saleh, Radim Navrátil, and Mina Norouzirad. Rank theory approach to ridge, LASSO, preliminary test and Stein-type estimators: a comparative study. Canadian Journal of Statistics = Revue canadienne de statistique, 46(4):690–704, December 2018. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2019:IIa

[1813] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 47(1):1–3, March 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2019:SIC

[1814] Anonymous. Special issue on Collaborative Research Team projects of the Canadian Statistical Sciences Institute: Guest Editor’s Introduction. Canadian Journal of Statistics = Revue canadienne de statistique, 47 (1):4–7, March 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Acar:2019:PAC

[1815] Elif F. Acar, Parisa Azimaee, and Md. Erfanul Hoque. Predictive assessment of copula models. Canadian Journal of Statistics = Revue canadienne de statistique, 47(1):8–26, March 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yin:2019:ISS

[1816] Yihao Yin, William H. Aeberhard, Stephen J. Smith, and Joanna Mills Flemming. Identifiable state-space models: a case study of the Bay of Fundy sea scallop fishery. Canadian Journal of Statistics = Revue canadienne de statistique, 47(1):27–45, March 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Albert-Green:2019:HPP

[1817] Alisha Albert-Green, W. John Braun, Charmaine B. Dean, and Craig Miller. A hierarchical point process with application to storm cell modelling. Canadian Journal of Statistics = Revue canadienne de statistique, 47(1):46–64, March 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Cardot:2019:ETE

[1818] Hervé Cardot, Anne De Moliner, and Camelia Goga. Estimation of total electricity consumption curves by sampling in a finite population when some trajectories are partially unobserved. Canadian Journal of Statistics = Revue canadienne de statistique, 47(1):65–89, March 2019. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Sun:2019:MAT

[1819] Jianping Sun, Karim Oualkacha, Celia M. T. Greenwood, and Lajmi Lakhal-Chaieb. Multivariate association test for rare variants controlling for cryptic and family relatedness. Canadian Journal of Statistics = Revue canadienne de statistique, 47(1):90–107, March 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Nathoo:2019:RSM

[1820] Farouk S. Nathoo, Linglong Kong, Hongtu Zhu, and for the Alzheimer’s Disease Neuroimaging Initiative. A review of statistical methods in imaging genetics. Canadian Journal of Statistics = Revue canadienne de statistique, 47(1):108–131, March 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2019:ARS

[1821] Anonymous. Acknowledgement of referees’ services: Remerciements aux lecteurs critiques. Canadian Journal of Statistics = Revue canadienne de statistique, 47(1):132–136, March 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2019:IIb

[1822] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 47(2):137–139, June 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wei:2019:CEL

[1823] Yizheng Wei, Yanyuan Ma, Tanya P. Garcia, and Samiran Sinha. A consistent estimator for logistic mixed effect models. Canadian Journal of Statistics = Revue canadienne de statistique, 47(2):140–156, June 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2019:LES

[1824] Jianxuan Liu and Yanyuan Ma. Locally efficient semiparametric estimators for a class of Poisson models with measurement error. Canadian Journal of Statistics = Revue canadienne de statistique, 47(2):157–181, June 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Su:2019:MHC

[1825] Chien-Lin Su, Johanna G. Neslehová, and Weijing Wang. Modelling hierarchical clustered censored data with the hierarchical Kendall copula. Canadian Journal of Statistics = Revue canadienne de statistique, 47(2): 182–203, June 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhou:2019:EFC

[1826] Ling Zhou, Haoqi Li, Huazhen Lin, and Peter X.-K. Song. Evaluating functional covariate-environment interactions in the Cox regression model. Canadian Journal of Statistics = Revue canadienne de statistique, 47(2):204–221, June 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Gustafson:2019:WES

[1827] Paul Gustafson and Mohammad Ehsanul Karim. When exposure is subject to nondifferential misclassification, are validation data helpful in testing for an exposure-disease association? Canadian Journal of Statistics = Revue canadienne de statistique, 47(2):222–237, June 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Jeganathan:2019:ESA

[1828] Pratheepa Jeganathan, Noroharivelo V. Randrianampy, Robert L. Paige, and A. Alexandre Trindade. An empirical saddlepoint approximation based method for smoothing survival functions under right censoring. Canadian Journal of Statistics = Revue canadienne de statistique, 47(2): 238–261, June 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2019:LMR

[1829] Xiang Li and Xianzheng Huang. Linear mode regression with covariate measurement error. Canadian Journal of Statistics = Revue canadienne de statistique, 47(2):262–280, June 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Cai:2019:ELC

[1830] Song Cai, Yongsong Qin, J. N. K. Rao, and Malgorzata Winiszewska. Empirical likelihood confidence intervals under imputation for missing survey data from stratified simple random sampling. Canadian Journal of Statistics = Revue canadienne de statistique, 47(2):281–301, June 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Shi:2019:DSS

[1831] Chenlu Shi and Boxin Tang. Design selection for strong orthogonal arrays. Canadian Journal of Statistics = Revue canadienne de statistique, 47(2):302–314, June 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Oliva-Aviles:2019:CVM

[1832] Cristian Oliva-Aviles, Mary C. Meyer, and Jean D. Opsomer. Checking validity of monotone domain mean estimators. Canadian Journal of Statistics = Revue canadienne de statistique, 47(2):315–331, June 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2019:IIc

[1833] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 47(3):333–335, September 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liao:2019:SMM

[1834] Jun Liao, Guohua Zou, and Yan Gao. Spatial Mallows model averaging for geostatistical models. Canadian Journal of Statistics = Revue canadienne de statistique, 47(3):336–351, September 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Sun:2019:AGS

[1835] Yanqing Sun, Li Qi, Fei Heng, and Peter B. Gilbert. Analysis of generalized semiparametric mixed varying-coefficients models for longitudinal data. Canadian Journal of Statistics = Revue canadienne de statistique, 47(3):352–373, September 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wong:2019:CBA

[1836] Weng Kee Wong and Julie Zhou. CVX-based algorithms for constructing various optimal regression designs. Canadian Journal of Statistics = Revue canadienne de statistique, 47(3):374–391, September 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

AlMohamad:2019:TSP

[1837] Diaa Al Mohamad and Michel Broniatowski. A two-step proximal-point algorithm for the calculus of divergence-based estimators in finite mixture models. Canadian Journal of Statistics = Revue canadienne de statistique, 47(3):392–408, September 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Lohr:2019:GFT

[1838] Sharon L. Lohr, Minsun K. Riddles, and J. Michael Brick. Goodness-of-fit tests for distributions estimated from complex survey data. Canadian Journal of Statistics = Revue canadienne de statistique, 47(3):409–425, September 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Torabi:2019:SGL

[1839] Mahmoud Torabi. Spatial generalized linear mixed models in small area estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 47(3):426–437, September 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2019:CBS

[1840] Fuyuan Li, Yanlin Tang, and Huixia Judy Wang. Copula-based semiparametric analysis for time series data with detection limits. Canadian Journal of Statistics = Revue canadienne de statistique, 47(3):438–454, September 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2019:BCE

[1841] Kun Chen, Ngai H. Chan, Man Wang, and Chun Y. Yau. On Bartlett correction of empirical likelihood for regularly spaced spatial data. Canadian Journal of Statistics = Revue canadienne de statistique, 47(3):455– 472, September 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Luo:2019:NPB

[1842] Hao Luo, Gabriela V. Cohen Freue, Xin Zhao, Alexandre Bouchard-Côté, Igor Burstyn, and Paul Gustafson. New perspective on the benefits of the gene-environment independence in case-control studies. Canadian Journal of Statistics = Revue canadienne de statistique, 47(3):473–486, September 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Lin:2019:GKE

[1843] Huazhen Lin, Baoying Yang, Ling Zhou, Paul S. F. Yip, Ying-Yeh Chen, and Hua Liang. Global kernel estimator and test of varying-coefficient autoregressive model. Canadian Journal of Statistics = Revue canadienne de statistique, 47(3):487–519, September 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2019:II

[1844] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 47(4):521–523, December 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Mostafaiy:2019:OEF

[1845] Behdad Mostafaiy, Mohammad Reza Faridrohani, and Shojaeddin Chenouri. Optimal estimation in functional linear regression for sparse noise-contaminated data. Canadian Journal of Statistics = Revue canadienne de statistique, 47(4):524–559, December 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Contreras-Cristan:2019:UPG

[1846] Alberto Contreras-Cristán, Richard A. Lockhart, Michael A. Stephens, and Shaun Z. Sun. On the use of priors in goodness-of-fit tests. Canadian Journal of Statistics = Revue canadienne de statistique, 47(4):560–579, December 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Gu:2019:SDM

[1847] Tian Gu, Jeremy M. G. Taylor, Wenting Cheng, and Bhramar Mukherjee. Synthetic data method to incorporate external information into a current study. Canadian Journal of Statistics = Revue canadienne de statistique, 47(4):580–603, December 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). See corrigenda [1985].

Ng:2019:EIF

[1848] Tin Lok J. Ng and Thomas B. Murphy. Estimation of the intensity function of an inhomogeneous Poisson process with a change-point. Canadian Journal of Statistics = Revue canadienne de statistique, 47(4):604–618, December 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Mccormack:2019:DLT

[1849] Andrew Mccormack, Nancy Reid, Nicola Sartori, and Sri-Amirthan Theivendran. A directional look at F -tests. Canadian Journal of Statistics = Revue canadienne de statistique, 47(4):619–627, December 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Kawamura:2019:ATL

[1850] Kenta Kawamura and Kanta Naito. Asymptotic theory for local estimators based on Bregman divergence. Canadian Journal of Statistics = Revue canadienne de statistique, 47(4):628–652, December 2019. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Guan:2019:IVE

[1851] Jing Guan, Hongjian Cheng, Kenneth A. Bollen, D. Roland Thomas, and Liqun Wang. Instrumental variable estimation in ordinal probit models with mismeasured predictors. Canadian Journal of Statistics = Revue canadienne de statistique, 47(4):653–667, December 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Samadi:2019:SAM

[1852] S. Yaser Samadi, Mahtab Hajebi, and Rahman Farnoosh. A semiparametric approach for modelling multivariate nonlinear time series. Canadian Journal of Statistics = Revue canadienne de statistique, 47(4):668– 687, December 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Song:2019:PMS

[1853] Yin Song, Farouk Nathoo, and Arif Babul. A Potts-mixture spatiotemporal joint model for combined magnetoencephalography and electroencephalography data. Canadian Journal of Statistics = Revue canadienne de statistique, 47(4):688–711, December 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Rivest:2019:REM

[1854] Louis-Paul Rivest and Shogo Kato. A random-effects model for clustered circular data. Canadian Journal of Statistics = Revue canadienne de statistique, 47(4):712–728, December 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Stephenson:2019:DSR

[1855] Matthew Stephenson, R. Ayesha Ali, Gerarda A. Darlington, and for the Alzheimer’s Disease Neuroimaging Initiative. Doubly sparse regression incorporating graphical structure among predictors. Canadian Journal of Statistics = Revue canadienne de statistique, 47(4):729–747, December 2019. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2020:IIa

[1856] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 48(1):1–3, March 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Hyndman:2020:SIS

[1857] Cody Hyndman. Special issue on stochastic models, statistics and finance: Guest Editor’s introduction. Canadian Journal of Statistics = Revue canadienne de statistique, 48(1):4–7, March 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Begin:2020:PBC

[1858] Jean-François Bégin and Geneviève Gauthier. Price bias and common practice in option pricing. Canadian Journal of Statistics = Revue canadienne de statistique, 48(1):8–35, March 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Fang:2020:CFS

[1859] Yizhou Fang, Martin Lysy, and Don Mcleish. Common-factor stochastic volatility modelling with observable proxy. Canadian Journal of Statistics = Revue canadienne de statistique, 48(1):36–61, March 2020. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2020:ICP

[1860] Fuqi Chen, Rogemar Mamon, and Sévérien Nkurunziza. Inference for a change-point problem under an OU setting with unequal and unknown volatilities. Canadian Journal of Statistics = Revue canadienne de statistique, 48(1):62–78, March 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Nasri:2020:GFR

[1861] Bouchra R. Nasri, Bruno N. Rémillard, and Mamadou Y. Thioub. Goodness-of-fit for regime-switching copula models with application to option pricing. Canadian Journal of Statistics = Revue canadienne de statistique, 48(1):79–96, March 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2020:EMT

[1862] Renjie Wang, Cody Hyndman, and Anastasis Kratsios. The entropic measure transform. Canadian Journal of Statistics = Revue canadienne de statistique, 48(1):97–129, March 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2020:ARS

[1863] Anonymous. Acknowledgement of referees’ services / remerciements aux lecteurs critiques. Canadian Journal of Statistics = Revue canadienne de statistique, 48(1):130–134, March 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2020:IIb

[1864] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 48(2):135–137, June 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

She:2020:VEA

[1865] Xichen She and Changbao Wu. Validity and efficiency in analyzing ordinal responses with missing observations. Canadian Journal of Statistics = Revue canadienne de statistique, 48(2):138–151, June 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chang:2020:POR

[1866] Lo-Bin Chang. Partial order relations for classification comparisons. Canadian Journal of Statistics = Revue canadienne de statistique, 48(2): 152–166, June 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2020:NDF

[1867] Su Chen. A new distribution-free k-sample test: Analysis of kernel density functionals. Canadian Journal of Statistics = Revue canadienne de statistique, 48(2):167–186, June 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2020:DED

[1868] Qingyang Zhang. Direct estimation of differential networks under highdimensional nonparanormal graphical models. Canadian Journal of Statistics = Revue canadienne de statistique, 48(2):187–203, June 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Holbrook:2020:EPE

[1869] Andrew Holbrook, Thomas Lumley, and Daniel Gillen. Estimating prediction error for complex samples. Canadian Journal of Statistics = Revue canadienne de statistique, 48(2):204–221, June 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhan:2020:SRM

[1870] Tianyu Zhan and Douglas E. Schaubel. Semiparametric regression methods for temporal processes subject to multiple sources of censoring. Canadian Journal of Statistics = Revue canadienne de statistique, 48(2):222– 237, June 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Jadhav:2020:FME

[1871] Sneha Jadhav and Shuangge Ma. Functional measurement error in functional regression. Canadian Journal of Statistics = Revue canadienne de statistique, 48(2):238–258, June 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Shu:2020:IMM

[1872] Heng Shu and Zhiqiang Tan. Improved methods for moment restriction models with data combination and an application to two-sample instrumental variable estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 48(2):259–284, June 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2020:ODC

[1873] Cai Li and Luo Xiao. Optimal design for classification of functional data. Canadian Journal of Statistics = Revue canadienne de statistique, 48(2): 285–307, June 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:2020:HDC

[1874] Yilei Wu, Yingli Qin, and Mu Zhu. High-dimensional covariance matrix estimation using a low-rank and diagonal decomposition. Canadian Journal of Statistics = Revue canadienne de statistique, 48(2):308–337, June 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2020:IIc

[1875] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 48(3):339–341, September 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2020:URS

[1876] Xinlei Wang, Mumu Wang, Johan Lim, and Soohyun Ahn. Using ranked set sampling with binary outcomes in cluster randomized designs. Canadian Journal of Statistics = Revue canadienne de statistique, 48(3):342– 365, September 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Shin:2020:BPC

[1877] Seung Jun Shin, Yichao Wu, and Ning Hao. A backward procedure for change-point detection with applications to copy number variation detection. Canadian Journal of Statistics = Revue canadienne de statistique, 48(3):366–385, September 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yang:2020:ELN

[1878] Zhihuang Yang and Niansheng Tang. Empirical likelihood for nonlinear regression models with nonignorable missing responses. Canadian Journal of Statistics = Revue canadienne de statistique, 48(3):386–416, September 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chenouri:2020:RMC

[1879] Shojaeddin Chenouri, Ahmad Mozaffari, and Gregory Rice. Robust multivariate change point analysis based on data depth. Canadian Journal of Statistics = Revue canadienne de statistique, 48(3):417–446, September 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Guan:2020:PMF

[1880] Leying Guan and Robert Tibshirani. Post model-fitting exploration via a “Next-Door’’ analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 48(3):447–470, September 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Ji:2020:SSM

[1881] Kexin Ji and Joel A. Dubin. A semiparametric stochastic mixed effects model for bivariate cyclic longitudinal data. Canadian Journal of Statistics = Revue canadienne de statistique, 48(3):471–498, September 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2020:EAH

[1882] Huiqiong Li, Han Zhang, Liang Zhu, Ni Li, and Jianguo Sun. Estimation of the additive hazards model with interval-censored data and missing covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 48(3):499–517, September 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Guo:2020:NCP

[1883] Lingzhe Guo and Reza Modarres. Nonparametric change point detection for periodic time series. Canadian Journal of Statistics = Revue canadienne de statistique, 48(3):518–534, September 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Shi:2020:PDE

[1884] Jianhong Shi, Xiuqin Bai, and Weixing Song. Partial deconvolution estimation in nonparametric regression. Canadian Journal of Statistics = Revue canadienne de statistique, 48(3):535–560, September 2020. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Dey:2020:RLB

[1885] Asim K. Dey, Cuneyt G. Akcora, Yulia R. Gel, and Murat Kantarcioglu. On the role of local blockchain network features in cryptocurrency price formation. Canadian Journal of Statistics = Revue canadienne de statistique, 48(3):561–581, September 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Bouezmarni:2020:NBK

[1886] Taoufik Bouezmarni, Sébastien Bellegem, and Yassir Rabhi. Nonparametric beta kernel estimator for long and short memory time series. Canadian Journal of Statistics = Revue canadienne de statistique, 48 (3):582–595, September 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Khodsiani:2020:OBB

[1887] Razieh Khodsiani and Saeid Pooladsaz. Optimal balanced block designs for correlated observations. Canadian Journal of Statistics = Revue canadienne de statistique, 48(3):596–604, September 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2020:IId

[1888] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 48(4):605–607, December 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Guo:2020:ECC

[1889] Lisha Guo, Yi Xiong, and X. Joan Hu. Estimation in the Cox cure model with covariates missing not at random, with application to disease screening/prediction. Canadian Journal of Statistics = Revue canadienne de statistique, 48(4):608–632, December 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

So:2020:CMB

[1890] Hon Yiu So, Mary E. Thompson, and Changbao Wu. Correlated and misclassified binary observations in complex surveys. Canadian Journal of Statistics = Revue canadienne de statistique, 48(4):633–654, December 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2020:IMM

[1891] Shijia Wang, Liangliang Wang, and Tim B. Swartz. Inference for misclassified multinomial data with covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 48(4):655–669, December 2020. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2020:HTU

[1892] Jiahua Chen, Pengfei Li, and Guanfu Liu. Homogeneity testing under finite location-scale mixtures. Canadian Journal of Statistics = Revue canadienne de statistique, 48(4):670–684, December 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Grover:2020:CBP

[1893] Kanika Grover, Elif F. Acar, and Mahmoud Torabi. Copula-based predictions in small area estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 48(4):685–711, December 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2020:SSC

[1894] Chengrui Li, Ying Hung, and Minge Xie. A sequential split-and-conquer approach for the analysis of big dependent data in computer experiments. Canadian Journal of Statistics = Revue canadienne de statistique, 48(4): 712–730, December 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Xia:2020:BME

[1895] Michelle Xia, P. Richard Hahn, and Paul Gustafson. A Bayesian mixture of experts approach to covariate misclassification. Canadian Journal of Statistics = Revue canadienne de statistique, 48(4):731–750, December 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2020:CTM

[1896] Shuo Shuo Liu and Bingshu E. Chen. Continuous threshold models with two-way interactions in survival analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 48(4):751–772, December 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Plante:2020:GAU

[1897] André Plante. A Gaussian alternative to using improper confidence intervals. Canadian Journal of Statistics = Revue canadienne de statistique, 48(4):773–801, December 2020. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2021:IIa

[1898] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):1–3, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Nathoo:2021:SIN

[1899] Farouk Nathoo, Linglong Kong, and Grace Y. Yi. Special issue on neuroimaging data analysis: Guest Editors’ introduction. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):4–9, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2021:SDM

[1900] Rongjie Liu, Hongtu Zhu, and for the Alzheimer’s Disease Neuroimaging Initiative. Statistical disease mapping for heterogeneous neuroimaging studies. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):10–34, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). See discussion [1901, 1902] and rejoinder [1903].

Wu:2021:DSD

[1901] Ben Wu and Jian Kang. Discussion of “Statistical disease mapping for heterogeneous neuroimaging studies”. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):35–38, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). See [1900, 1903].

Pietrosanu:2021:DSD

[1902] Matthew Pietrosanu and Bei Jiang. Discussion of “Statistical disease mapping for heterogeneous neuroimaging studies”. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):39–42, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). See [1900, 1903].

Liu:2021:RSD

[1903] Rongjie Liu and Hongtu Zhu. Rejoinder: “Statistical disease mapping for heterogeneous neuroimaging studies”. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):43–45, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). See [1900, 1901, 1902].

Dai:2021:BLS

[1904] Ning Dai, Hakmook Kang, Galin L. Jones, Mark B. Fiecas, and for the Alzheimer’s Disease Neuroimaging Initiative. A Bayesian latent spatial model for mapping the cortical signature of progression to Alzheimer’s disease. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):46–62, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2021:STF

[1905] Zeyi Wang, Haris I. Sair, Ciprian Crainiceanu, Martin Lindquist, Bennett A. Landman, Susan Resnick, Joshua T. Vogelstein, and Brian Caffo. On statistical tests of functional connectome fingerprinting. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):63–88, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Hart:2021:GBP

[1906] Brian Hart, Stephen Malone, and Mark Fiecas. A grouped beta process model for multivariate resting-state EEG microstate analysis on twins. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1): 89–106, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Mohammed:2021:SST

[1907] Shariq Mohammed and Dipak K. Dey. Scalable spatio-temporal Bayesian analysis of high-dimensional electroencephalography data. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):107–128, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Lan:2021:SBS

[1908] Zhou Lan, Brian J. Reich, and Dipankar Bandyopadhyay. A spatial Bayesian semiparametric mixture model for positive definite matrices with applications in diffusion tensor imaging. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):129–149, March 2021. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Ding:2021:MFR

[1909] Xiucai Ding, Dengdeng Yu, Zhengwu Zhang, and Dehan Kong. Multivariate functional response low-rank regression with an application to brain imaging data. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):150–181, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2021:SBW

[1910] Yanming Li, Bin Nan, Ji Zhu, and for the Alzheimer’s Disease Neuroimaging Initiative. A structured brain-wide and genome-wide association study using ADNI PET images. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):182–202, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Brzyski:2021:CIA

[1911] Damian Brzyski, Marta Karas, Beau M Ances, Mario Dzemidzic, Joaquín Goñi, Timothy W Randolph, and Jaroslaw Harezlak. Connectivityinformed adaptive regularization for generalized outcomes. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):203–227, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2021:ARS

[1912] Anonymous. Acknowledgement of referees’ services/remerciements aux lecteurs critiques. Canadian Journal of Statistics = Revue canadienne de statistique, 49(1):228–233, March 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2021:IIb

[1913] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):235–237, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Grenon-Godbout:2021:LPS

[1914] Nicolas Grenon-Godbout and Mylène Bédard. On-line partitioning of the sample space in the regional adaptive algorithm. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):238–261, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Hahn:2021:RMT

[1915] Eugene D. Hahn. Regression modelling with the tilted beta distribution: a Bayesian approach. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):262–282, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Kang:2021:VOC

[1916] Xiaoning Kang and Xinwei Deng. On variable ordination of Choleskybased estimation for a sparse covariance matrix. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):283–310, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Xu:2021:ENA

[1917] Guoying Xu and Yang Bai. Estimation of nonparametric additive models with high order spatial autoregressive errors. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):311–343, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Che:2021:ECL

[1918] Menglu Che, Jerald F. Lawless, and Peisong Han. Empirical and conditional likelihoods for two-phase studies. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):344–361, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:2021:REM

[1919] Ping Wu and Jiming Jiang. Robust estimation of mean squared prediction error in small-area estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):362–396, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2021:VSP

[1920] Baojiang Chen, Ao Yuan, and Grace Y. Yi. Variable selection for proportional hazards models with high-dimensional covariates subject to measurement error. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):397–420, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Newton:2021:WBB

[1921] Michael A. Newton, Nicholas G. Polson, and Jianeng Xu. Weighted Bayesian bootstrap for scalable posterior distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):421–437, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Li:2021:EAE

[1922] Yiou Li and Xinwei Deng. An efficient algorithm for Elastic I-optimal design of generalized linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):438–470, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Favre-Martinoz:2021:ENE

[1923] Cyril Favre-Martinoz, David Haziza, and Jean-François Beaumont. Efficient nonparametric estimation for skewed distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):471–496, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Casleton:2021:LSG

[1924] Emily M. Casleton, Daniel J. Nordman, and Mark S. Kaiser. Local structure graph models with higher-order dependence. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):497–513, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Betsch:2021:MDE

[1925] Steffen Betsch, Bruno Ebner, and Bernhard Klar. Minimum L^{q}-distance estimators for non-normalized parametric models. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):514–548, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2021:SJF

[1926] Yi Liu, Jinfeng Xu, and Gang Li. Sure joint feature screening in nonparametric transformation model for right censored data. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):549–565, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Queiroz:2021:BCZ

[1927] Francisco F. Queiroz and Artur J. Lemonte. A broad class of zero-or-one inflated regression models for rates and proportions. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):566–590, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Salamanca:2021:SMP

[1928] Juan J. Salamanca. Sets that maximize probability and a related variational problem. Canadian Journal of Statistics = Revue canadienne de statistique, 49(2):591–608, June 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2021:IIc

[1929] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):609–611, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2021:QAR

[1930] Ling-Wan Chen, Yu Cheng, Ying Ding, and Ruosha Li. Quantile association regression on bivariate survival data. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):612–636, September 2021. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Rabhi:2021:SRM

[1931] Yassir Rabhi and Masoud Asgharian. A semiparametric regression model under biased sampling and random censoring: a local pseudo-likelihood approach. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):637–658, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yuan:2021:SIR

[1932] Ao Yuan, Jin Piao, Jing Ning, and Jing Qin. Semiparametric isotonic regression modelling and estimation for group testing data. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):659–677, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Park:2021:ASP

[1933] Heewon Park, Rui Yamaguchi, Seiya Imoto, and Satoru Miyano. Automatic sparse principal component analysis. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):678–697, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Kobayashi:2021:FBQ

[1934] Genya Kobayashi, Taeyoung Roh, Jangwon Lee, and Taeryon Choi. Flexible Bayesian quantile curve fitting with shape restrictions under the Dirichlet process mixture of the generalized asymmetric Laplace distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):698–730, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Ding:2021:ECR

[1935] Maomao Ding, Jing Ning, and Ruosha Li. Evaluation of competing risks prediction models using polytomous discrimination index. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):731– 753, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhao:2021:SBA

[1936] Yanyan Zhao and Lei Sun. On set-based association tests: Insights from a regression using summary statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):754–770, September 2021. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Maiti:2021:UEF

[1937] Tapabrata Maiti, Abolfazl Safikhani, and Ping-Shou Zhong. On uncertainty estimation in functional linear mixed models. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):771–792, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Sang:2021:ABI

[1938] Hejian Sang and Jae Kwang Kim. An approximate Bayesian inference on propensity score estimation under unit nonresponse. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):793–807, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Xing:2021:LBC

[1939] Li Xing, Xuekui Zhang, Igor Burstyn, and Paul Gustafson. On logistic Box–Cox regression for flexibly estimating the shape and strength of exposure-disease relationships. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):808–825, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2021:VSS

[1940] Li Liu, Yanyan Liu, Feng Su, and Xingqiu Zhao. Variable selection and structure estimation for ultrahigh-dimensional additive hazards models. Canadian Journal of Statistics = Revue canadienne de statistique, 49 (3):826–852, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yi:2021:EHT

[1941] Grace Y. Yi and Ying Yan. Estimation and hypothesis testing with error-contaminated survival data under possibly misspecified measurement error models. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):853–874, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Hu:2021:PHD

[1942] Jie Hu, Yu Chen, Weiping Zhang, and Xiao Guo. Penalized highdimensional M-quantile regression: From L^{1} to L^{p} optimization. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):875– 905, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Qian:2021:ABC

[1943] Fang Qian, Weiping Zhang, and Yu Chen. Adaptive banding covariance estimation for high-dimensional multivariate longitudinal data. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):906–938, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yuan:2021:FCR

[1944] Yuze Yuan, Lihua Bai, and Jiancheng Jiang. Functional-coefficient regression models with GARCH errors. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):939–964, September 2021. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yuan:2021:SIY

[1945] Meng Yuan, Pengfei Li, and Changbao Wu. Semiparametric inference of the Youden index and the optimal cut-off point under density ratio models. Canadian Journal of Statistics = Revue canadienne de statistique, 49(3):965–986, September 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2021:IId

[1946] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):987–989, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Lawless:2021:CRC

[1947] Jerry Lawless. Covid-19-related content in The Canadian Journal of Statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):990–991, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Elliott:2021:ERR

[1948] Sean Elliott and Christian Gouriéroux. Estimated reproduction ratios in the SIR model. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):992–1017, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Parker:2021:URC

[1949] Matthew R. P. Parker, Yangming Li, Lloyd T. Elliott, Junling Ma, and Laura L. E. Cowen. Under-reporting of COVID-19 in the Northern Health Authority region of British Columbia. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1018–1038, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Fan:2021:OSL

[1950] Yan Fan, Yukun Liu, and Lixing Zhu. Optimal subsampling for linear quantile regression models. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1039–1057, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Hua:2021:ODU

[1951] Yi Hua and Min Yang. Optimal design under complete class with ancillary functions. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1058–1074, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yu:2021:IAS

[1952] Jun Yu and Dejian Lai. Interim analysis of sequential estimation-adjusted urn models with sample size re-estimation. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1075–1092, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2021:CDA

[1953] Qiang Zhang, Wenliang Pan, Chengwei Li, and Xueqin Wang. The conditional distance autocovariance function. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1093–1114, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Sang:2021:JEL

[1954] Yongli Sang, Xin Dang, and Yichuan Zhao. A Jackknife empirical likelihood approach for K-sample Tests. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1115–1135, December 2021. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Shu:2021:PBN

[1955] Di Shu and Wenqing He. Perturbation-based null hypothesis tests with an application to Clayton models. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1136–1151, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2021:QME

[1956] Tongwei Zhang, Dehui Wang, and Kai Yang. Quasi-maximum exponential likelihood estimation for double-threshold GARCH models. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1152– 1178, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Diao:2021:NSR

[1957] Guoqing Diao and Jing Qin. New semiparametric regression method with applications in selection-biased sampling and missing data problems. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1179–1195, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yoshida:2021:QFR

[1958] Takuma Yoshida. Quantile function regression and variable selection for sparse models. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1196–1221, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Tay:2021:PCG

[1959] Jingyi K. Tay, Jerome Friedman, and Robert Tibshirani. Principal component-guided sparse regression. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1222–1257, December 2021. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2021:WME

[1960] Yi Wang, Peng Wu, Xingwei Tong, and Jianguo Sun. A weighted method for the exclusive hypothesis test with application to typhoon data. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1258– 1272, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Brunel:2021:HRN

[1961] Elodie Brunel and Fabienne Comte. Hazard regression with noncompactly supported bases. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1273–1297, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Fang:2021:ILM

[1962] Junhan Fang and Grace Y. Yi. Imputation and likelihood methods for matrix-variate logistic regression with response misclassification. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1298– 1316, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2021:SIV

[1963] Zhengwei Liu, Qi Li, and Fukang Zhu. Semiparametric integer-valued autoregressive models on Z. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1317–1337, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Stefan:2021:EDB

[1964] Marius Stefan and Michael A. Hidiroglou. Estimation of design-based mean squared error of a small area mean model-based estimator under a nested error linear regression model. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1338–1363, December 2021. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2021:MLB

[1965] Fang Chen, Hong Wan, Hua Cai, and Guang Cheng. Machine learning in/for blockchain: Future and challenges. Canadian Journal of Statistics = Revue canadienne de statistique, 49(4):1364–1382, December 2021. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2022:IIa

[1966] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):1–4, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Neslehova:2022:E

[1967] Johanna G. Neslehová. Editorial. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):5–7, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Muller:2022:SIF

[1968] Hans-Georg Müller. Special issue on “Functional and object data analysis”: Guest Editor’s introduction. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):8–19, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2022:BSS

[1969] Xinlian Zhang, Gauri S. Datta, Ping Ma, and Wenxuan Zhong. Bayesian spline smoothing with ambiguous penalties. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):20–35, March 2022. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Lin:2022:DLL

[1970] Hongmei Lin, Tiejun Tong, Yuedong Wang, Wenchao Xu, and Riquan Zhang. Direct local linear estimation for Sharpe ratio function. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):36– 58, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2022:NSN

[1971] Yijia Liu, Lexin Li, and Xiao Wang. A nonlinear sparse neural ordinary differential equation model for multiple functional processes. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):59– 85, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Cheng:2022:AEF

[1972] Kun Cheng and Di-Rong Chen. Adaptive estimation for functional data: Using a framelet block-thresholding method. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):86–121, March 2022. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Nie:2022:RUT

[1973] Yunlong Nie, Yuping Yang, Liangliang Wang, and Jiguo Cao. Recovering the underlying trajectory from sparse and irregular longitudinal data. Canadian Journal of Statistics = Revue canadienne de statistique, 50 (1):122–141, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Ghosal:2022:VSN

[1974] Rahul Ghosal and Arnab Maity. Variable selection in nonparametric functional concurrent regression. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):142–161, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Cai:2022:REV

[1975] Xiong Cai, Liugen Xue, Jiguo Cao, and for the Alzheimer’s Disease Neuroimaging Initiative. Robust estimation and variable selection for function-on-scalar regression. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):162–179, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Mehrotra:2022:SVS

[1976] Suchit Mehrotra and Arnab Maity. Simultaneous variable selection, clustering, and smoothing in function-on-scalar regression. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):180–199, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhou:2022:CCC

[1977] Zhiyang Zhou and Peijun Sang. Continuum centroid classifier for functional data. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):200–220, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Jiang:2022:CAR

[1978] Jiakun Jiang, Huazhen Lin, Heng Peng, Gang-Zhi Fan, and Yi Li. Cluster analysis with regression of non-Gaussian functional data on covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1): 221–240, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2022:RKB

[1979] Meichen Liu, Matthew Pietrosanu, Peng Liu, Bei Jiang, Xingcai Zhou, Linglong Kong, and for the Alzheimer’s Disease Neuroimaging Initiative. Reproducing kernel-based functional linear expectile regression. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):241– 266, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhu:2022:EQE

[1980] Hanbing Zhu, Yehua Li, Baisen Liu, Weixin Yao, and Riquan Zhang. Extreme quantile estimation for partial functional linear regression models with heavy-tailed distributions. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):267–286, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Gervini:2022:DSM

[1981] Daniel Gervini. Doubly stochastic models for spatio-temporal covariation of replicated point processes. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):287–303, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Hormann:2022:TNS

[1982] Siegfried Hörmann, Piotr Kokoszka, and Thomas Kuenzer. Testing normality of spatially indexed functional data. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):304–326, March 2022. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Sidrow:2022:MMS

[1983] Evan Sidrow, Nancy Heckman, Sarah M. E. Fortune, Andrew W. Trites, Ian Murphy, and Marie Auger-Méthé. Modelling multi-scale, stateswitching functional data with hidden Markov models. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):327–356, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Shang:2022:ALM

[1984] Han Lin Shang. Not all long-memory estimators are born equal: the case of nonstationary functional time series. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):357–380, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Gu:2022:ACC

[1985] Tian Gu, Jeremy M. G. Taylor, Wenting Cheng, and Bhramar Mukherjee. Authors’ corrigenda/corrections des auteurs: On synthetic data method to incorporate external information into a current study. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):381– 384, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic). See [1847].

Gu:2022:ARS

[1986] Tian Gu, Jeremy M. G. Taylor, Wenting Cheng, and Bhramar Mukherjee. Acknowledgement of Referees’ Services Remerciements aux lecteurs critiques. Canadian Journal of Statistics = Revue canadienne de statistique, 50(1):385–390, March 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2022:IIb

[1987] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2):391–394, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Yi:2022:CCD

[1988] Grace Y. Yi, Pingbo Hu, and Wenqing He. Characterizing the COVID-19 dynamics with a new epidemic model: Susceptible-exposed-asymptomatic-symptomatic-active-removed. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2):395–416, June 2022. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Golchi:2022:EDO

[1989] Shirin Golchi. Estimating design operating characteristics in Bayesian adaptive clinical trials. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2):417–436, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Fraser:2022:LLL

[1990] D. A. S. Fraser and Mylène Bédard. The linear Lasso: a location model approach. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2):437–453, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wieczorek:2022:MSP

[1991] Jerzy Wieczorek and Jing Lei. Model selection properties of forward selection and sequential cross-validation for high-dimensional regression. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2): 454–470, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Luo:2022:RSS

[1992] Yiyun Luo, Yufeng Liu, and for the Alzheimer’s Disease Neuroimaging Initiative. Recovery of sums of sparse and dense signals by incorporating graphical structure among predictors. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2):471–490, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Egidi:2022:APD

[1993] Leonardo Egidi, Francesco Pauli, and Nicola Torelli. Avoiding prior-data conflict in regression models via mixture priors. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2):491–510, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Sun:2022:AHR

[1994] Xiaowei Sun, Xinyuan Song, and Liuquan Sun. Additive hazard regression of event history studies with intermittently measured covariates. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2): 511–532, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Moradian:2022:DER

[1995] Hoora Moradian, Weichi Yao, Denis Larocque, Jeffrey S. Simonoff, and Halina Frydman. Dynamic estimation with random forests for discretetime survival data. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2):533–548, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2022:SAF

[1996] Peng Liu, Shanshan Song, and Yong Zhou. Semiparametric additive frailty hazard model for clustered failure time data. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2):549–571, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Berger:2022:ACS

[1997] Moritz Berger and Matthias Schmid. Assessing the calibration of subdistribution hazard models in discrete time. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2):572–591, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Agniel:2022:SEC

[1998] Denis Agniel, Matthew Cefalu, and Bing Han. Synthetic estimation for the complier average causal effect. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2):592–611, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zheng:2022:MEM

[1999] Chaowen Zheng, Jingfang Huang, Ian A. Wood, and Yichao Wu. A modified expectation-maximization algorithm for latent Gaussian graphical model. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2):612–637, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Mynbaev:2022:NEU

[2000] Kairat Mynbaev, Carlos Martins-Filho, and Daniel J. Henderson. Nonparametric estimation of unrestricted distributions and their jumps. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2): 638–662, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Khalafi:2022:SCB

[2001] Mahnaz Khalafi, Ahmad Reza Soltani, Masoud Golalipour, Majid Azimmohseni, and Farzad Najafiamiri. On the spectral coherence between two periodically correlated processes. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2):663–680, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Taheriyoun:2022:EFD

[2002] Ali Reza Taheriyoun and Khalil Shafie. The estimation of the fractal dimension of a Gaussian field via Euler characteristic. Canadian Journal of Statistics = Revue canadienne de statistique, 50(2):681–707, June 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2022:IIc

[2003] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):709–712, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Douwes-Schultz:2022:EBE

[2004] Dirk Douwes-Schultz, Shuo Sun, Alexandra M. Schmidt, and Erica E. M. Moodie. Extended Bayesian endemic-epidemic models to incorporate mobility data into COVID-19 forecasting. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):713–733, September 2022. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2022:ESC

[2005] Liangliang Wang, Joosung Min, Renny Doig, Lloyd T. Elliott, and Caroline Colijn. Estimation of SARS-CoV-2 antibody prevalence through serological uncertainty and daily incidence. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):734–750, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Jiang:2022:MDS

[2006] Qiang Jiang, Yifan Xia, and Baosheng Liang. Matching distributions for survival data. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):751–775, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Diao:2022:NDR

[2007] Liqun Diao and Richard J. Cook. Nested doubly robust estimating equations for causal analysis with an incomplete effect modifier. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):776–794, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhong:2022:DET

[2008] Wei Zhong, Wei Zhou, Qingliang Fan, and Yang Gao. Dummy endogenous treatment effect estimation using high-dimensional instrumental variables. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):795–819, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2022:SQR

[2009] Liwen Zhang, Zhoufan Zhu, Xingdong Feng, and Yong He. Shrinkage quantile regression for panel data with multiple structural breaks. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):820– 851, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2022:GSD

[2010] Wu Wang and Zhongyi Zhu. Group structure detection for a highdimensional panel data model. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):852–866, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Duan:2022:SCP

[2011] Jiangtao Duan, Wei Gao, Hao Qu, and Hon Keung Tony Ng. Subspace clustering for panel data with interactive effects. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):867–887, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Ozturk:2022:TSC

[2012] Omer Ozturk, Olena Kravchuk, and Jennifer Brown. Two-stage cluster samples with judgment post-stratification. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):888–910, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Ozturk:2022:CRI

[2013] Omer Ozturk and Olena Kravchuk. Combining ranking information from different sources in ranked-set samples. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):911–927, September 2022. CO-DEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2022:SAC

[2014] Chunzhe Zhang, Curtis B. Storlie, and Thomas C. M. Lee. Sensitivity analysis in classification using Bayesian smoothing spline ANOVA probit regression. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):928–950, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2022:COD

[2015] Yanhong Liu, Haojie Ren, Xu Guo, Qin Zhou, and Changliang Zou. Cellwise outlier detection with false discovery rate control. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):951–971, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Jha:2022:BAD

[2016] Jayant Jha. Best approach direction for spherical random variables. Canadian Journal of Statistics = Revue canadienne de statistique, 50 (3):972–991, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Ebner:2022:TND

[2017] Bruno Ebner, Norbert Henze, and David Strieder. Testing normality in any dimension by Fourier methods in a multivariate Stein equation. Canadian Journal of Statistics = Revue canadienne de statistique, 50 (3):992–1033, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2022:THC

[2018] Guanfu Liu, Yuejiao Fu, Wenchen Liu, and Rongji Mu. Testing homogeneity in contaminated mixture models. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):1034–1046, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Al-Labadi:2022:TIB

[2019] Luai Al-Labadi, Forough Fazeli Asl, and Zahra Saberi. A test for independence via Bayesian nonparametric estimation of mutual information. Canadian Journal of Statistics = Revue canadienne de statistique, 50 (3):1047–1070, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2022:JRT

[2020] Qiang Chen, Wanzi Xu, and Yuting Gong. Jump-robust testing of volatility functions in continuous time models. Canadian Journal of Statistics = Revue canadienne de statistique, 50(3):1071–1095, September 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2022:IId

[2021] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1097–1100, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2022:ISI

[2022] Anonymous. Introduction to the special issue on the 50th anniversary of CJS. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1101–1103, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Reid:2022:DFS

[2023] Nancy Reid. D. A. S. Fraser: From structural inference to asymptotics. Canadian Journal of Statistics = Revue canadienne de statistique, 50 (4):1104–1115, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Csorgo:2022:RWT

[2024] Miklós Csörgő, Donald A. Dawson, Bouchra R. Nasri, and Bruno N. Rémillard. A random walk through Canadian contributions on empirical processes and their applications in probability and statistics. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1116–1142, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Mathai:2022:SGW

[2025] Arak M. Mathai and Serge B. Provost. On the singular gamma, Wishart, and beta matrix-variate density functions. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1143–1165, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chen:2022:PEL

[2026] Yilin Chen, Pengfei Li, J. N. K. Rao, and Changbao Wu. Pseudo empirical likelihood inference for nonprobability survey samples. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1166–1185, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Beaumont:2022:SIF

[2027] Jean-François Beaumont and David Haziza. Statistical inference from finite population samples: a critical review of frequentist and Bayesian approaches. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1186–1212, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Craiu:2022:RBI

[2028] Radu V. Craiu, Paul Gustafson, and Jeffrey S. Rosenthal. Reflections on Bayesian inference and Markov chain Monte Carlo. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1213–1227, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Chipman:2022:LPW

[2029] Hugh Chipman and Derek Bingham. Let’s practice what we preach: Planning and interpreting simulation studies with design and analysis of experiments. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1228–1249, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Andrews:2022:RR

[2030] David Andrews and Chris Field. Robust reflections. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1250–1253, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Xun:2022:SEH

[2031] Xiaolei Xun, Tianyu Guan, and Jiguo Cao. Sparse estimation of historical functional linear models with a nested group bridge approach. Canadian Journal of Statistics = Revue canadienne de statistique, 50 (4):1254–1269, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Cook:2022:LHA

[2032] Richard J. Cook and Jerald F. Lawless. Life history analysis with multistate models: a review and some current issues. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1270–1298, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Moodie:2022:CIC

[2033] Erica E. M. Moodie and David A. Stephens. Causal inference: Critical developments, past and future. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1299–1320, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2022:UGA

[2034] Lin Zhang and Lei Sun. Unifying genetic association tests via regression: Prospective and retrospective, parametric and nonparametric, and genotype- and allele-based tests. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1321–1338, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Susko:2022:CSM

[2035] Edward Susko. Complex statistical modelling for phylogenetic inference. Canadian Journal of Statistics = Revue canadienne de statistique, 50 (4):1339–1354, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Dean:2022:CCE

[2036] Charmaine B. Dean, Abdel H. El-Shaarawi, Sylvia R. Esterby, Joanna Mills Flemming, Richard D. Routledge, Stephen W. Taylor, Douglas G. Woolford, James V. Zidek, and Francis W. Zwiers. Canadian contributions to environmetrics. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1355–1386, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Thompson:2022:CSS

[2037] Mary Thompson, Nancy Reid, and Don Estep. The Canadian Statistical Sciences Institute 2003–2022. Canadian Journal of Statistics = Revue canadienne de statistique, 50(4):1387–1402, December 2022. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2023:IIa

[2038] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):1–4, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhao:2023:NTH

[2039] Fanrong Zhao, Nan Lin, and Baoxue Zhang. A new test for highdimensional regression coefficients in partially linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):5–18, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Miyawaki:2023:EVS

[2040] Koji Miyawaki and Steven N. MacEachern. Economic variable selection. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1): 19–37, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Cao:2023:DRW

[2041] Zhimiao Cao, Xiaoning Kang, and Mingqiu Wang. Doubly robust weighted composite quantile regression based on SCAD-L_{2}. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):38–76, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Burak:2023:NCR

[2042] Katherine L. Burak and Adam B. Kashlak. Nonparametric confidence regions via the analytic wild bootstrap. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):77–94, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Padilla:2023:GEN

[2043] Oscar Hernan Madrid Padilla and Yanzhen Chen. Graphon estimation via nearest-neighbour algorithm and two-dimensional fused-lasso denoising. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):95–110, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Luo:2023:MOR

[2044] Lan Luo and Peter X.-K. Song. Multivariate online regression analysis with heterogeneous streaming data. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):111–133, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Luo:2023:BCC

[2045] Yu Luo, David A. Stephens, and David L. Buckeridge. Bayesian clustering for continuous-time hidden Markov models. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):134–156, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Hu:2023:ZIP

[2046] Guanyu Hu, Hou-Cheng Yang, Yishu Xue, and Dipak K. Dey. Zeroinflated Poisson model with clustered regression coefficients: Application to heterogeneity learning of field goal attempts of professional basketball players. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):157–172, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zhang:2023:MAT

[2047] Xiaomeng Zhang, Xinyu Zhang, and Yanyuan Ma. A model-averaging treatment of multiple instruments in Poisson models with errors. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):173– 198, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Gavanji:2023:PLR

[2048] Parisa Gavanji, Wenyu Jiang, and Bingshu E. Chen. Penalized likelihood ratio test for a biomarker threshold effect in clinical trials based on generalized linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):199–215, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Feng:2023:RAM

[2049] Yanqin Feng, K. D. Prasangika, and Guoxin Zuo. Regression analysis of multivariate current status data under a varying coefficients additive hazards frailty model. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):216–234, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Sun:2023:ECC

[2050] Yanqing Sun, Fei Heng, Unkyung Lee, and Peter B. Gilbert. Estimation of conditional cumulative incidence functions under generalized semiparametric regression models with missing covariates, with application to analysis of biomarker correlates in vaccine trials. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):235–257, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Abousaleh:2023:MCO

[2051] Hanan Abousaleh and Julie Zhou. Minimax A-, c-, and I-optimal regression designs for models with heteroscedastic errors. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):258–274, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Krieger:2023:BED

[2052] Abba M. Krieger, David A. Azriel, and Adam Kapelner. Better experimental design by hybridizing binary matching with imbalance optimization. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):275–292, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2023:TDP

[2053] Sixu Liu, Yaping Wang, and Fasheng Sun. Two-dimensional projection uniformity for space-filling designs. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):293–311, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Erciulescu:2023:SDI

[2054] Andreea L. Erciulescu, Jean D. Opsomer, and Benjamin J. Schneider. Statistical data integration using multilevel models to predict employee compensation. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):312–326, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wu:2023:AAR

[2055] Yi Wu, Wei Yu, Wenzhi Yang, Saisai Ding, and Xuejun Wang. On asymptotic approximation of ratio models for weakly dependent sequences. Canadian Journal of Statistics = Revue canadienne de statistique, 51 (1):327–343, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2023:ARS

[2056] Anonymous. Acknowledgement of referees’ services / remerciements aux membres des jurys. Canadian Journal of Statistics = Revue canadienne de statistique, 51(1):344–349, March 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2023:IIb

[2057] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):351–354, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Han:2023:IIE

[2058] Peisong Han, Jeremy M. G. Taylor, and Bhramar Mukherjee. Integrating information from existing risk prediction models with no model details. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2): 355–374, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Liu:2023:DEE

[2059] Qianqian Liu and Zhouping Li. Distributed estimation with empirical likelihood. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):375–399, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Su:2023:DCA

[2060] Wen Su, Guosheng Yin, Jing Zhang, and Xingqiu Zhao. Divide and conquer for accelerated failure time model with massive time-to-event data. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):400–419, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Shao:2023:OSL

[2061] Li Shao, Shanshan Song, and Yong Zhou. Optimal subsampling for large-sample quantile regression with massive data. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):420–443, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Pashley:2023:CIM

[2062] Nicole E. Pashley and Marie-Abèle C. Bind. Causal inference for multiple treatments using fractional factorial designs. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):444–468, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Jiang:2023:DTR

[2063] Cong Jiang, Michael P. Wallace, and Mary E. Thompson. Dynamic treatment regimes with interference. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):469–502, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Manuel:2023:RBD

[2064] Christopher Manuel, Samiran Sinha, and Suojin Wang. Reducing bias due to misclassified exposures using instrumental variables. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):503–530, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Dai:2023:NTT

[2065] Maozhu Dai, Weining Shen, and Hal S. Stern. Nonparametric tests for treatment effect heterogeneity in observational studies. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):531–558, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Ma:2023:SAF

[2066] Haiqiang Ma, Chao Liu, Sheng Xu, and Jin Yang. Subgroup analysis for functional partial linear regression model. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):559–579, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Ma:2023:CGL

[2067] Haiqiang Ma and Jiming Jiang. Classified generalized linear mixed model prediction incorporating pseudo-prior information. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):580–595, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Wang:2023:EMC

[2068] Xianru Wang, Bin Liu, Xinsheng Zhang, Yufeng Liu, and for the Alzheimer’s Disease Neuroimaging Initiative. Efficient multiple change point detection for high-dimensional generalized linear models. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):596–629, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Zheng:2023:MDA

[2069] Siming Zheng, Alan T. K. Wan, and Yong Zhou. Missing data analysis with sufficient dimension reduction. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):630–651, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Cholaquidis:2023:WLD

[2070] Alejandro Cholaquidis, Ricardo Fraiman, Fabrice Gamboa, and Leonardo Moreno. Weighted lens depth: Some applications to supervised classification. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):652–673, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Williams:2023:EAG

[2071] Jonathan P. Williams, Yuying Xie, and Jan Hannig. The EAS approach for graphical selection consistency in vector autoregression models. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):674– 703, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Stupfler:2023:CBR

[2072] Gilles Stupfler and Antoine Usseglio-Carleve. Composite bias-reduced L^{p}-quantile-based estimators of extreme quantiles and expectiles. Canadian Journal of Statistics = Revue canadienne de statistique, 51(2):704– 742, June 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2023:IIc

[2073] Anonymous. Issue information. Canadian Journal of Statistics = Revue canadienne de statistique, 51(3):743–746, September 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Anonymous:2023:SIH

[2074] Anonymous. Special issue in honour of Nancy Reid: Guest Editors’ introduction. Canadian Journal of Statistics = Revue canadienne de statistique, 51(3):747–751, September 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Craiu:2023:CNR

[2075] Radu V. Craiu and Grace Y. Yi. A conversation with Nancy Reid. Canadian Journal of Statistics = Revue canadienne de statistique, 51 (3):752–759, September 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Battey:2023:IPS

[2076] Heather S. Battey. Inducement of population sparsity. Canadian Journal of Statistics = Revue canadienne de statistique, 51(3):760–768, September 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

McCullagh:2023:TTV

[2077] Peter McCullagh. A tale of two variances. Canadian Journal of Statistics = Revue canadienne de statistique, 51(3):769–779, September 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Elkantassi:2023:IIB

[2078] Soumaya Elkantassi, Ruggero Bellio, Alessandra R. Brazzale, and Anthony C. Davison. Improved inference for a boundary parameter. Canadian Journal of Statistics = Revue canadienne de statistique, 51(3):780– 799, September 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Carey:2023:SEW

[2079] Michelle Carey, Christian Genest, and James O. Ramsay. Sparse estimation within Pearson’s system, with an application to financial market risk. Canadian Journal of Statistics = Revue canadienne de statistique, 51(3):800–823, September 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Urban:2023:ONC

[2080] Konrad N. Urban, Heejong Bong, Josue Orellana, and Robert E. Kass. Oscillating neural circuits: Phase, amplitude, and the complex normal distribution. Canadian Journal of Statistics = Revue canadienne de statistique, 51(3):824–851, September 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

McCormack:2023:TLH

[2081] Andrew McCormack and Peter D. Hoff. Tests of linear hypotheses using indirect information. Canadian Journal of Statistics = Revue canadienne de statistique, 51(3):852–876, September 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Pace:2023:CSC

[2082] Luigi Pace, Alessandra Salvan, and Nicola Sartori. Confidence sequences with composite likelihoods. Canadian Journal of Statistics = Revue canadienne de statistique, 51(3):877–896, September 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Kalbfleisch:2023:ROM

[2083] John D. Kalbfleisch and Zhenzhen Xu. Rerandomization and optimal matching. Canadian Journal of Statistics = Revue canadienne de statistique, 51(3):897–913, September 2023. CODEN ???? ISSN 0319-5724 (print), 1708-945X (electronic).

Ogawa:1974:STA

[2084] Junjiro Ogawa. Statistical theory of the analysis of experimental designs. Marcel Dekker, Inc., New York, NY, USA, 1974. vi + 465 pp. Statistics: Textbooks and Monographs, Vol. 8.

Fraser:1976:PST

[2085] D. A. S. (Donald Alexander Stuart) Fraser. Probability and Statistics: Theory and Applications. Duxbury Press, North Scituate, MA, USA, 1976. ISBN 0-87872-099-5. xi + 623 pp. LCCN QA273 .F82.

Anderson:1977:BMS

[2086] Theodore Wilbur Anderson, Somesh Das Gupta, and George Peter Hansbenno Styan. A Bibliography of Multivariate Statistical Analysis. Robert E. Krieger Publishing Company, Huntington, NY, USA, 1977. ISBN 0-88275-477-7. x + 642 pp. LCCN Z6654.M8 A53 1977 QA278.

Cassel:1977:FIS

[2087] Claes Cassel, Carl-Erik Särndal, and Jan Håkan Wretman. Foundations of Inference in Survey Sampling. Wiley series in probability and mathematical statistics. John Wiley, New York, NY, USA, 1977. ISBN 0-471-02563-1. xi + 192 pp. LCCN QA276.6 .C34.

Chambers:1977:CMD

[2088] John M. Chambers. Computational Methods for Data Analysis. Wiley series in probability and mathematical statistics. John Wiley, New York, NY, USA, 1977. ISBN 0-471-02772-3. xi + 268 pp. LCCN QA276.4 .C48 1977.

Dixon:1977:BBC

[2089] Wilfrid Joseph Dixon and M. B. (Morton B.) Brown, editors. BMDP-77, Biomedical Computer Programs, P-series. University of California Press, Berkeley, CA, USA, 1977. ISBN 0-520-03569-0. xiii + 880 pp. LCCN RA409.5 .C33 1977.

Fienberg:1977:ACC

[2090] Stephen E. Fienberg. The Analysis of Cross-Classified Categorical Data. MIT Press, Cambridge, MA, USA, 1977. ISBN 0-262-06063-9. x + 151 pp. LCCN QA278 .F53.

Giri:1977:MSI

[2091] Narayan C. Giri. Multivariate Statistical Inference. Probability and mathematical statistics. Academic Press, New York, USA, 1977. ISBN 0-12-285650-3. xvi + 319 pp. LCCN QA278 .G56.

Krishnan:1977:MMS

[2092] P. (Parameswara) Krishnan, editor. Mathematical Models of Sociology, volume 24 of Sociological review monograph. University of Keele, Keele, UK, 1977. ISBN 0-904425-03-7. 229 pp. LCCN HM15 .S545 no. 24 HM24.

Mathai:1977:CNP

[2093] A. M. Mathai and Giorgio Pederzoli. Characterizations of the Normal Probability Law. John Wiley, New York, NY, USA, 1977. ISBN 0-470-99322-7. x + 149 pp. LCCN QA273.6 .M37.

Mathai:1977:PS

[2094] A. M. Mathai. Probability and statistics. MacMillan Publishing Company, New York, NY, USA, 1977. ISBN 0-333-90115-0. xv + 406

pp. LCCN QA273 .M38. URL http://catalog.hathitrust.org/api/
volumes/oclc/4276431.html.

Odeh:1977:PBS

[2095] Robert E. Odeh et al. Pocket Book of Statistical Tables. Statistics, textbooks and monographs; v. 22. Marcel Dekker, Inc., New York, NY, USA, 1977. ISBN 0-8247-6515-X. x + 166 pp. LCCN QA276.25 .P6.

Owen:1977:STM

[2096] D. B. (Donald Bruce) Owen and R. Odeh. Selected Tables in Mathematical Statistics, volume 5 of Selected tables in mathematical statistics. American Mathematical Society, Providence, RI, USA, 1977. ISBN 0-8218-1905-4. 263 pp. LCCN QA276.25 .S43 vol. 5.

Wonnacott:1977:IS

[2097] Thomas H. Wonnacott and Ronald J. Wonnacott. Introductory Statistics. Wiley series in probability and mathematical statistics. John Wiley, New York, NY, USA, 3d edition, 1977. ISBN 0-471-95982-0. xxii + 650 pp. LCCN HA29 .W62 1977.

Dubuc:1978:POC

[2098] Serge Dubuc. Problèmes d’optimisation en calcul des probabilités, volume 62 of Séminaire de mathématiques supérieures [Seminar on Higher Mathematics]. Presses de l’Université de Montréal, Montréal, QC, Canada, 1978. ISBN 0-8405-0413-6. 137 pp. LCCN QA273 .D784.

Freedman:1978:IMS

[2099] David Freedman, Robert Pisani, and Robert Purves. Instructor’s Manual for Statistics. W. W. Norton & Co., New York, NY, USA, 1978. ISBN 0-393-09041-8. 135 pp. LCCN QA276 .F74 1978 Suppl.

Freedman:1978:S

[2100] David Freedman, Robert Pisani, and Roger Purves. Statistics. W. W. Norton & Co., New York, NY, USA, 1978. ISBN 0-393-09076-0. xv + 506 + 83 pp. LCCN QA276 .F683 1978.

Krotki:1978:DDS

[2101] Karol Jósef Krótki, editor. Developments in Dual System Estimation of Population Size and Growth. University of Alberta Press, Edmonton, AB, Canada, 1978. ISBN 0-88864-017-X. xviii + 260 pp. LCCN HB886 .D48.

Kruskal:1978:IES

[2102] William Kruskal and Judith M. Tanur, editors. International Encyclopedia of Statistics. Free Press, New York, NY, USA, 1978. ISBN 0-02-917960-2 (paperback). xxi + 1350 (2 volumes) pp. LCCN HA17 .I63.

Letac:1978:CMP

[2103] Gérard Letac. Chaînes de Markov sur les permutations, volume 63 of Séminaire de mathématiques supérieures [Seminar on Higher Mathematics]. Presses de l’Université de Montréal, Montréal, QC, Canada, 1978. ISBN 0-8405-0412-8. 117 pp. LCCN QA274.7 .L47.

Malliavin:1978:GDS

[2104] Paul Malliavin. Géométrie différentielle stochastique, volume 64 of Séminaire de mathématiques supérieures [Seminar on Higher Mathematics]. Presses de l’Université de Montréal, Montréal, QC, Canada, 1978. ISBN 0-8405-0425-X. 144 pp. LCCN ???? Notes prepared by Danièle Dehen and Dominique Michel.

Mathai:1978:FAS

[2105] A. M. Mathai and Rajendra Kumar Saxena. The H-Function with Applications in Statistics and Other Disciplines. John Wiley, New York, NY, USA, 1978. ISBN 0-470-26380-6. xii + 192 pp. LCCN QA353.H9 M37.

Nash:1978:CNM

[2106] John C. Nash. Compact Numerical Methods for Computers: Linear Algebra and Function Minimisation. Adam Hilger Ltd., Bristol, UK, 1978. ISBN 0-85274-330-0. 227 pp. LCCN QA184.N37 1978.

Neter:1978:AS

[2107] John Neter, William Wasserman, G. A. Whitmore, Kenneth Schneider, Randall Byers, and Daniel Brick. Applied Statistics. Allyn and Bacon, Needham Heights, MA, USA, 1978. ISBN 0-205-05982-1, 0-205-05983-X (test manual), 0-205-06038-2 (solutions manual). xix + 743, 4 + 130 (test manual), iii + 143 (solutions manual) pp. LCCN QA276; HA29 .N437.

Sankoff:1978:LVM

[2108] David Sankoff, editor. Linguistic Variation: Models and Methods. Academic Press, New York, USA, 1978. ISBN 0-12-618850-5. xvii + 296 pp. LCCN P120.V37 L52.

Walpole:1978:PSE

[2109] Ronald E. Walpole and Raymond H. Myers. Probability and Statistics for Engineers and Scientists. MacMillan Publishing Company, New York, NY, USA, second edition, 1978. ISBN 0-02-424110-5. xii + 580 pp. LCCN TA340 .W35 1978.

Whitmore:1978:SDA

[2110] G. A. Whitmore and M. Chapman Findlay, editors. Stochastic Dominance: an Approach to Decision-Making Under Risk. Lexington Books, Lexington, MA, USA, 1978. ISBN 0-669-01310-2. xiv + 398 pp. LCCN HG174 .S76.

Williams:1978:SS

[2111] Bill Williams. A Sampler on Sampling. Wiley series in probability and mathematical statistics. John Wiley, New York, NY, USA, 1978. ISBN 0-471-03036-8. xv + 254 pp. LCCN QA276.6 .W54 1978.

Blake:1979:IAP

[2112] Ian F. Blake. An Introduction to Applied Probability. John Wiley, New York, NY, USA, 1979. ISBN 0-471-03210-7. 528 pp. LCCN QA273 .B586.

Brillinger:1979:MS

[2113] David R. Brillinger et al. Mathematical Statistics, volume 5 of Selecta Statistica Canadiana; Selecta Statistica Canadiana; Mathematical statistics. Selecta Statistica Canadiana, Department of Mathematics, McMaster University, Hamilton, ON, Canada, 1979. ISBN 0-920788-04-1. 150 pp. LCCN ????

Das:1979:DAE

[2114] M. N. (Manindra Nath) Das and Narayan C. Giri. Design and Analysis of Experiments. John Wiley, New York, NY, USA, 1979. ISBN 0-470-26861-1. vii + 295 pp. LCCN QA279 .D35 1979.

Flaherty:1979:PGD

[2115] David H. Flaherty. Privacy and Government Data Banks: an International Perspective. Mansell, London, UK, 1979. ISBN 0-7201-0930-2. 353 pp. LCCN K3560 .F53.

Geramita:1979:ODQ

[2116] A. V. Geramita and Jennifer Seberry. Orthogonal Designs: Quadratic Forms and Hadamard Matrices, volume 45 of Lecture notes in pure and applied mathematics. Marcel Dekker, Inc., New York, NY, USA, 1979. ISBN 0-8247-6774-8. x + 460 pp. LCCN QA166.4 .G47.

Goodman:1979:MAC

[2117] Leo A. Goodman and William Kruskal. Measures of Association for Cross Classifications, volume 1 of Springer series in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1979. ISBN 0-387-90443-3. x + 146 pp. LCCN HM24 .G627. With a foreword by Stephen E. Fienberg.

Kocherlakota:1979:PP

[2118] K. Kocherlakota. A Primer in Probability. Statistics, textbooks and monographs; v. 28. Marcel Dekker, Inc., New York, NY, USA, 1979. ISBN 0-8247-6836-1. viii + 326 pp. LCCN QA273 .S766.

Marshall:1979:ITM

[2119] Albert W. Marshall and Ingram Olkin. Inequalities: Theory of Majorization and its Applications, volume 143 of Mathematics in science and engineering. Academic Press, New York, USA, 1979. ISBN 0-12-473750-1. xx + 569 pp. LCCN QA295 .M42. URL

http://www.loc.gov/catdir/description/els031/79050218.html;
 http://www.loc.gov/catdir/toc/els031/79050218.html.

Mathai:1979:PS

[2120] Arakaparampil M. Mathai and Pushpa N. Rathie. Probability and Statistics. MacMillan Publishing Company, New York, NY, USA, 1979. ISBN 0-333-19056-4. xv + 406 pp. LCCN ????

Mohanty:1979:LPC

[2121] Gopal Mohanty. Lattice Path Counting and Applications. Probability and mathematical statistics. Academic Press, New York, USA, 1979. ISBN 0-12-504050-4. xi + 185 pp. LCCN QA273.45 .M63.

Narayana:1979:LPC

[2122] T. V. (Tadepalli Venkata) Narayana. Lattice Path Combinatorics, with Statistical Applications, volume 23 of Mathematical expositions. University of Toronto Press, Toronto, ON, Canada, 1979. ISBN 0-8020-5405-6. xii + 106 pp. LCCN QA276 .N33.

Nash:1979:CNM

[2123] John C. Nash. Compact numerical methods for computers: linear algebra and function minimisation. John Wiley, New York, NY, USA, 1979. ISBN 0-470-26559-0. ix + 227 pp. A Halsted Press Book.

Srivastava:1979:IMS

[2124] M. S. Srivastava and C. G. Khatri. An Introduction to Multivariate Statistics. North-Holland Publishing Co., Amsterdam, The Netherlands, 1979. ISBN 0-444-00302-9. xvi + 350 pp. LCCN QA278 .S69.

Adler:1980:PCI

[2125] Hans J. Adler and David Allan Brusegard. Perspectives Canada III. Ministre des Approvisionnements et Services Canada, Ottawa, ON, Canada, 1980. ISBN 0-660-90202-8. 312 pp. LCCN HA742 .C355 1980.

Afriat:1980:DFS

[2126] S. N. Afriat. Demand Functions and the Slutsky Matrix, volume 7 of Princeton studies in mathematical economics. Princeton University Press, Princeton, NJ, USA, 1980. ISBN 0-691-04222-5. xii +

269 pp. LCCN HB801 .A33. URL http://www.loc.gov/catdir/
enhancements/fy1411/79083973-d.html.

Hubert:1980:BQ

[2127] J. J. (John Joseph) Hubert and E. M. (Edward M.) Carter. Biostatistics, 1064 Questions. Kendall/Hunt Pub., Dubuque, IA, USA, preliminary edition, 1980. ISBN 0-8403-2287-9 (paperback). 155 pp. LCCN QH323.5 .H82 1980.

Kalbfleisch:1980:SAF

[2128] J. D. Kalbfleisch and Ross L. Prentice. The Statistical Analysis of Failure Time Data. Wiley series in probability and mathematical statistics. John Wiley, New York, NY, USA, 1980. ISBN 0-471-05519-0. xi + 321 pp. LCCN QA270 .K34. URL http://www.

loc.gov/catdir/description/wiley037/79021889.html;   http:/
/www.loc.gov/catdir/enhancements/fy0607/79021889-b.html;   http:
//www.loc.gov/catdir/toc/onix02/79021889.html.

Neff:1980:DSM

[2129] Norman D. Neff and Joseph I. Naus. The distribution of the size of the maximum cluster of points on a line, volume 6 of Selected Tables in Mathematical Statistics. American Mathematical Society, Providence, RI, USA, 1980. ISBN 0-8218-1906-2. vii + 207 pp. LCCN ????

Nishisato:1980:ACD

[2130] Shizuhiko Nishisato. Analysis of Categorical Data: Dual Scaling and Its Applications, volume 24 of Mathematical expositions. University of Toronto Press, Toronto, ON, Canada, 1980. ISBN 0-8020-5489-7. ISSN 0076-5333. xiii + 276 pp. LCCN QA278 .N57.

Odeh:1980:TNT

[2131] Robert E. Odeh and D. B. (Donald Bruce) Owen. Tables for Normal Tolerance Limits, Sampling Plans, and Screening. Statistics, textbooks and monographs; v. 32. Marcel Dekker, Inc., New York, NY, USA, 1980. ISBN 0-8247-6944-9. ix + 316 pp. LCCN TS156 .O26.

Sinha:1980:LTR

[2132] S. K. (Snehesh Kumar) Sinha and B. K. (Balvant Keshav) Kale. Life Testing and Reliability Estimation. John Wiley, New York, NY, USA, 1980. ISBN 0-470-26911-1, 0-85226-820-3. vi + 196 pp. LCCN QA276

.S53. URL http://www.gbv.de/dms/hbz/toc/ht002578317.pdf.

Brillinger:1981:TSD

[2133] David R. Brillinger. Time Series: Data Analysis and Theory. Holden-Day series in time series analysis. Holden-Day, San Francisco, CA, USA, expanded edition, 1981. ISBN 0-8162-1150-7. xii + 540 pp. LCCN QA280 .B74 1981.

Freedman:1981:DMM

[2134] H. I. Freedman. Deterministic Mathematical Models in Population Ecology, volume 57 of Pure and applied Mathematics. Marcel Dekker, Inc., New York, NY, USA, 1981. ISBN 0-8247-6653-9. x + 254 pp. LCCN ????

Grassmann:1981:SSM

[2135] Winfried K. Grassmann. Stochastic Systems for Management. North-Holland Publishing Co., Amsterdam, The Netherlands, 1981. ISBN 0-444-00449-1. xii + 358 pp. LCCN T57.6 .G72 1981.

Kanasewich:1981:TSA

[2136] E. R. Kanasewich. Time Sequence Analysis in Geophysics. University of Alberta Press, Edmonton, AB, Canada, third edition, 1981. ISBN 0-88864-074-9. xiii + 480 pp. LCCN QC806 .K36 1981.

Meeker:1981:PTN

[2137] William Q. Meeker, Larry W. Cornwell, and Leo Avedis Aroian. The Product of Two Normally Distributed Random Variables, volume 7 of Selected tables in mathematical statistics. American Mathematical Society, Providence, RI, USA, 1981. ISBN 0-8218-1907-0. ISSN 0094-8837. vii + 256 pp. LCCN QA276.25 .S43 vol. 7 QA274.

Mikhail:1981:AAS

[2138] Edward M. Mikhail and Gordon Gracie. Analysis and Adjustment of Survey Measurements. Van Nostrand Reinhold Co., New York, NY, USA, 1981. ISBN 0-442-25369-9. xii + 340 pp. LCCN TA549 .M537.

Wonnacott:1981:RSC

[2139] Thomas H. Wonnacott and Ronald J. Wonnacott. Regression, a Second Course in Statistics. Wiley series in probability and mathematical statistics. John Wiley, New York, NY, USA, 1981. ISBN 0-471-95974-X. xix + 556 pp. LCCN QA278.2 .W66.

Guttman:1982:IES

[2140] Irwin Guttman, S. S. (Samuel Stanley) Wilks, and J. Stuart Hunter. Introductory Engineering Statistics. Wiley series in probability and mathematical statistics. John Wiley, New York, NY, USA, third edition, 1982. ISBN 0-471-07859-X. xi + 580 pp. LCCN TA340 .G8 1982.

Lawless:1982:SMM

[2141] Jerald F. Lawless. Statistical Models and Methods for Lifetime Data. Wiley series in probability and mathematical statistics. Applied probability and statistics. John Wiley, New York, NY, USA, 1982. ISBN 0-471-08544-8. ISSN 0271-6356. xi + 580 pp. LCCN QA276

.L328. URL http://www.loc.gov/catdir/description/wiley034/
81011446.html;   http://www.loc.gov/catdir/enhancements/fy0607/
81011446-b.html;   http://www.loc.gov/catdir/toc/onix05/81011446.
html.

Wonnacott:1982:SDP

[2142] Ronald J. Wonnacott and Thomas H. Wonnacott. Statistics: Discovering its Power. Wiley series in probability and mathematical statistics: Applied probability and statistics section. John Wiley, New York, NY, USA, 1982. ISBN 0-471-01412-5. ISSN 0271-6356. xviii + 378 pp. LCCN QA276.12 .W66.

Abraham:1983:SMF

[2143] Bovas Abraham and Johannes Ledolter. Statistical Methods for Forecasting. Wiley series in probability and mathematical statistics, Applied probability and statistics. John Wiley, New York, NY, USA, 1983. ISBN 0-471-86764-0. ISSN 0271-6336. xv + 445 pp. LCCN QA279.2 .A27 1983.

URL http://www.loc.gov/catdir/bios/wiley043/83007006.html;
 http://www.loc.gov/catdir/description/wiley035/83007006.html;
 http://www.loc.gov/catdir/toc/onix06/83007006.html.

Basilevsky:1983:MAS

[2144] Alexander Basilevsky. Matrix Algebra in the Statistical Sciences. North-Holland Publishing Co., Amsterdam, The Netherlands, 1983. ISBN 0-444-00756-3. xiii + 389 pp. LCCN ????

Freedman:1983:ACM

[2145] David Freedman. Approximating Countable Markov Chains. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1983. ISBN 0-387-90804-8. x + 140 pp. LCCN QA274.7 .F73 1983.

Freedman:1983:BMD

[2146] David Freedman. Brownian Motion and Diffusion. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-387-90805-6. xii + 231 pp. LCCN QA274.75 .F74 1983.

Freedman:1983:MC

[2147] David Freedman. Markov Chains. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1983. ISBN 0-387-90808-0. xiv + 382 pp. LCCN QA274.7 .F74 1983. Corrected reprint of the 1971 original.

Sankoff:1983:TWS

[2148] David Sankoff and Joseph B. Kruskal, editors. Time Warps, String Edits, and Macromolecules: the Theory and Practice of Sequence Comparison. Addison-Wesley Pub. Co., Advanced Book Program, Reading, MA, USA, 1983. ISBN 0-201-07809-0. xii + 382 pp. LCCN QA292 .T55 1983.

Srivastava:1983:IAM

[2149] M. S. Srivastava and E. M. (Edward M.) Carter. An Introduction to Applied Multivariate Statistics. North-Holland Publishing Co., Amsterdam, The Netherlands, 1983. ISBN 0-444-00621-4. x + 394 pp. LCCN QA278 .S687 1983.

Greenwood:1985:CSI

[2150] P. E. Greenwood and Al§bert Nikolaevich Shirëiìaev. Contiguity and the Statistical Invariance Principle, volume 1 of Stochastics monographs, theory and applications of stochastic processes. Gordon and Breach, New York, NY, USA, 1985. ISBN 2-88124-013-5 (Switzerland). ISSN 0275-5785. viii + 236 pp. LCCN QA274 .G74 1985.

Keyfitz:1985:AMD

[2151] Nathan Keyfitz. Applied Mathematical Demography. Springer texts in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 1985. ISBN 0-387-96155-0. xxi + 441 pp. LCCN HB849.51 .K49 1985.

Miettinen:1985:TEP

[2152] Olli S. (Olli Sakari) Miettinen. Theoretical Epidemiology: Principles of Occurrence Research in Medicine. A Wiley medical publication. John Wiley, New York, NY, USA, 1985. ISBN 0-471-86769-1. xxii + 359 pp. LCCN RA652.4 .M54 1985.

Manoukian:1986:MCT

[2153] Edward B. Manoukian. Modern Concepts and Theorems of Mathematical Statistics. Springer series in statistics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1986. ISBN 0-387-96186-0. xvi + 156 pp. LCCN QA276 .M333 1986.

Constantine:1987:CTS

[2154] Gregory M. Constantine. Combinatorial theory and statistical design. Wiley Series in Probability and Mathematical Statistics: Probability and Mathematical Statistics. John Wiley, New York, NY, USA, 1987. ISBN 0-471-84097-1. xvi + 472 pp.

Spanier:1987:AF

[2155] Jerome Spanier and Keith B. Oldham. An Atlas of Functions. Hemisphere Publishing Corp., Washington, DC, USA, 1987. ISBN 0-89116-573-8, 3-540-17395-1. ix + 700 pp. LCCN QA331 .S685 1987. See also the second edition [2161].

Turner:1987:TTC

[2156] Roland Turner, editor. Thinkers of the twentieth century. St. James Press, Chicago, IL, USA, second edition, 1987. ISBN 0-912289-83-X. viii + 977 pp.

Das:1988:DAE

[2157] M. N. Das and N. C. Giri. Design and analysis of experiments. John Wiley, New York, NY, USA, second edition, 1988. ISBN 0-470-20714-0. xii + 488 pp.

Nash:1990:CNM

[2158] J. C. Nash. Compact numerical methods for computers. Adam Hilger Ltd., Bristol, UK, second edition, 1990. ISBN 0-85274-318-1; 0-85274-319-X. xii + 278 pp. Linear algebra and function minimisation.

Ismail:1999:DRA

[2159] Mourad E. H. Ismail and Dennis Stanton, editors. Dedicated to Richard A. Askey on the occasion of his 65th birthday. Part I. International Press, Somerville, MA, 1999. ISSN 1073-2772. 1–127 pp. In Methods Appl. Anal. 6 (1999), no. 1.

Das:2006:DAE

[2160] M. N. (Manindra Nath) Das and Narayan C. Giri. Design and Analysis of Experiments. New Age International, New Delhi, India, 2 edition, 2006. ISBN 0-470-26861-1, 0-85226-158-6, 0-85226-914-5. 488 pp. LCCN QA279 .D35 1979.

Oldham:2009:AFE

[2161] Keith B. Oldham, Jan Myland, and Jerome Spanier, editors. An Atlas of Functions: With Equator, the Atlas Function Calculator. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., second edition, 2009. ISBN 0-387-48807-3 (softcover), 0-387-48806-5 (hardcover). xi + 748 pp. LCCN QA331 .S685 2009. US$129.95.

Freedman:2012:MC

[2162] David Freedman. Markov Chains. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2012. ISBN 1-4612-5502-3. 382 pp. LCCN QA274.7 .F74 2012.

Joffe:1978:BP

[2163] Anatole Joffe and Peter Ney, editors. Branching Processes, volume 5 of Advances in Probability and Related Topics. Marcel Dekker, Inc., New York, NY, USA, 1978. ISBN 0-8247-6800-0. LCCN QA273.A1 A4 vol. 5 QA274.76.

Gupta:1980:MSA

[2164] R. P. Gupta, editor. Multivariate Statistical Analysis: Proceedings of the Second Conference on Multivariate Statistical Inference held at Dalhousie University, Halifax, NS, October 5–7, 1979. North-Holland Publishing Co., Amsterdam, The Netherlands, 1980. ISBN 0-444-86019-3. LCCN ????

Krewski:1981:CTS

[2165] D. Krewski, Richard Platek, J. N. K. Rao, et al., editors. Current Topics in Survey Sampling. Academic Press, New York, USA, 1981. ISBN 0-12-426280-5. LCCN HN29 .I62 1980.